A Bibliography of Papers in Lecture Notes in Computer Science (1991)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 February 2014 Version 1.21

Title word cross-reference

(2^{3539} + 1)/3 [10]. 1 [1867]. 2 [1358, 777]. 2^{e} [1583]. 3 [1706, 580, 237, 602, 569, 576, 585, 601, 574, 715]. [16, 7, 6] [453]. ^{∗} [1852]. ^{2} [487]. ^{3} [1456]. _{n} [772]. AC^{0} [1057]. C^{n} [1374]. d [1391]. ℓ [26]. ϵ_{0} [683]. f [1924]. GF(2^{m}) [479]. ≫ [42]. K [1853, 141, 1924, 1887, 1483, 937, 193]. καππα [2105]. L ≥ 2 [27]. ≪ [42]. M [193, 457, 745, 459, 1391]. GF(3)^{†} [474]. GraphEd [1161]. N [409, 1566, 475]. NF^{2} [1880]. q [459]. R^{d} [1839]. S^{+} [938]. V_{k}(F_{2}) [1391]. X + Y [146]. Z_{n} [2191].

* [706, 685, 702, 185, 683, 679, 704, 709].

-ary [457, 459]. -bisimulations [772]. -Coloring [141]. -Coteries [1887]. -Cube [409]. -D [1706, 569, 777, 574]. -Dependent [1924]. -Dimensional [237, 1867, 937, 715]. -fold [26]. -Functions [1391]. -Hash [1566]. -homogeneous [639]. -Machine [336]. -Matchings [1924]. -Means-Type [1853]. -Mutual [1887]. -nd [1358]. -out [193]. -regular [639]. -Sequences [459, 1391]. -tree [1852]. -Trees [938]. -Variable [1483]. -Watchmen [745].

1 [1776, 2041, 1736]. 1-Bounded [2041]. 1/ lambda [639]. 1990s [1762, 1990].

2 [1896, 1535, 1517, 1085, 516, 1530, 1366, 480, 1243, 1874, 1851, 1564, 1563]. 2-Category [1243]. 2-Manifolds [1851].

2-Monotonic [1896]. 2-Structures [1874]. 2-Way [1085, 516]. 2000s [1992].

3 [2145]. 3-Sorted [2145]. 3.0 [2098]. 370 [449].

’89 [1568].

90s [1965]. ’91 [1614]. 9X [238, 2121].

= [633, 519].

Abduction [1620, 667, 1442, 1684, 1260]. Abductive [1690, 666, 2086]. ABEL [1828]. abnormal [592]. absence [659]. Absolute [454]. Abstract [1830, 2107, 887, 1319, 1300, 1231, 1296, 2202, 6, 1312, 1304, 1, 2229, 41, 765, 1159, 1076, 2, 1308, 2228, 1584, 8, 1041, 2201, 1620, 947, 898, 1002, 1341, 7, 2215, 812, 18, 1239, 2223, 2204, 1040, 1329, 1612, 2219, 2211, 5, 9, 978, 603, 2193, 1691]. Abstraction [948, 955, 206, 227, 370]. Abstractions [1039, 2048, 904, 1700]. AC [1930]. AC-Equation [1930]. Academia [1286, 1290]. Acceptable [1085]. Acceptance [366, 1877]. Accepting [553]. acceptors [777]. Access [2192, 722, 929, 917, 2034, 2076, 932]. accessibility [789]. Accomplishments [1732]. According [1490, 806]. Account [1537]. Accounts [673]. accuracy [1446]. ACE [2111]. Achieving [1318]. Ackermannian [1373]. ACP [986]. acquire [1443]. Acquisition [426, 1441, 1687]. Action [542, 2151, 981, 1000, 868, 969, 2144]. Action/Event [2144]. Actions [1302, 1182, 866, 609, 868, 1404, 969, 1682, 1703]. Active [1647, 1712]. Activity [245]. Actor [1188]. Actors [368]. Ada [253, 251, 239, 243, 258, 241, 236, 249, 237, 259, 252, 244, 238, 2121, 255, 1877, 250, 240, 246, 247, 242]. Adaptability [445]. Adaptation [125, 2019]. Adaption [107]. Adaptive [119, 361, 137, 428, 1089, 1223, 1853, 128, 113, 1097, 194, 1412, 656, 600, 1445, 1452, 332]. Adaptively [415]. Adders [1497]. Adding [542, 692, 1017, 489, 945]. Addition [1340, 20]. Address [1269]. Addressable [358]. addressing [1688]. Adequacy [959]. Adiga [48]. ADIS [332]. ADT [1230]. Advance [1633]. Advances [1767, 431]. Advantages [2012]. Advice [230]. AFADA [240]. affected [600]. Affine [352]. Affix [1516, 1510]. AFIT [1277]. Against [1082, 1324, 2227, 1308, 42]. Agency [113]. Agenda [1762, 1956]. Agent [1401]. Agents [861, 839, 772, 1425]. Aggregate [279]. Agna [896]. Agonistic [849]. AI [1712, 1565, 862, 1252]. Aid [858]. Aided [2078, 1815, 1140]. AIFresco [1704]. Aims [1205]. Album [1270]. ALECSI [209]. ALF [1052, 2095]. Algebra [542, 2140, 987, 285, 2188, 1115, 993, 315, 996, 1002, 2186, 2141, 1153, 1005, 2160, 1239, 471, 2174, 212, 1013, 2051, 1352, 794]. Algebraic [468, 1165, 692, 1231, 262, 263, 264, 265, 266, 267, 980, 1955, 1776, 138, 286, 1356, 1176, 1735, 1233, 1361, 991, 1179, 1158, 1235, 1349, 814, 1087, 677, 470, 1368, 1936, 1370, 1374, 1845, 464, 1379, 1749, 1748, 1389, 1390, 1145, 709, 905, 769]. Algebras [1208, 539, 981, 1109, 1154, 950, 2182, 756, 795]. Algol [1101]. Algol-like [1101]. Algorithm [410, 1025, 1835, 749, 1836, 181, 1172, 1357, 164, 1305, 2152, 1359, 1569, 1212, 89, 536, 560, 428, 77, 405, 2024, 751, 111, 72, 1897, 1916, 535, 1084, 46, 226, 1920, 1365, 1898, 470, 1089, 1918, 178, 48, 1092, 93, 736, 182, 937, 434, 1946, 413, 346, 186, 1317, 433, 392, 1876, 1853, 1855, 82, 83, 1902, 192, 1298, 92, 770, 601, 1703, 797, 652]. Algorithmic [474, 1215, 142]. Algorithmics [531]. Algorithms [469, 68, 1895, 2066, 743, 545, 843, 87, 1058, 1926, 1355, 1348, 170, 2176, 1626, 88, 75, 1142, 177, 73, 1924, 2012, 361, 69, 160, 747, 1915, 115, 424, 90, 1907, 1867, 139, 553, 1060, 2030, 71, 131, 78, 79, 1923, 2165, 80, 753, 1061, 721, 800, 561, 169, 2007, 1347, 391, 1368, 527, 1886, 1351, 81, 748, 2023, 916, 1093, 1063, 180, 408, 1064, 1847, 455, 94, 345, 84, 415, 167, 394, 735, 130, 1439, 52, 779, 2212]. Allocation [1032, 193, 351, 1502]. Allows [77]. Almost [560, 4]. Along [1843]. alpha [65]. Alphabet [1760, 1937]. Alphabets [70]. Alternating [1081, 777, 789]. Alternation [2152]. Alternation-Free [2152]. Alternative [72, 1537]. Alternatives [238]. Amalgamated [1455]. Ambiguities [923]. Ambiguity [547]. Ambiguous [821]. among [1712]. AMTS [664]. Analog [2012, 1601]. analogical [1726]. Analogies [1187]. Analysis [409, 816, 355, 962, 2156, 1120, 623, 73, 249, 1644, 1837, 69, 507, 382, 1908, 1511, 1060, 891, 892, 1897, 886, 814, 298, 81, 1091, 1582, 1503, 748, 1129, 1616, 1093, 1775, 1747, 1929, 1745, 1671, 958, 924, 624, 2124, 1628, 580, 1695, 903, 58, 593, 61, 594, 1436, 1473, 1250]. Analysis/Formal [1775]. Analyze [2204]. Analyzing [1148, 243, 1164]. Anaphora [1541, 1540]. anatomical [573, 572, 576]. anatomical-based [576]. Ancestor [178]. Ancestors [716]. And-Parallellism [2074]. And/Or [2111, 1311]. And/Or-parallel [2111]. Andorra [375, 2105, 2106]. Andorra-I [2106]. Angle [1953]. Angle-Sweep [1953]. Animation [1050]. Annealing [100, 102, 103, 130, 581]. Annotated [262]. Answers [320]. Antagonistic [849]. Any [1084]. Applicable [877]. Application [685, 329, 1173, 1735, 669, 1606, 1078, 111, 95, 1488, 1740, 496, 1222, 1842, 852, 1192, 826, 97, 461, 317, 1575, 140, 850, 1663, 1796, 603, 802, 1671, 82, 1889, 592, 1715]. Applications [746, 1160, 964, 267, 1734, 1806, 482, 1738, 1322, 1117, 309, 259, 729, 493, 1339, 1982, 618, 2063, 1240, 121, 1585, 255, 368, 1323, 135, 246, 1891, 1298]. Applicative [879]. Applied [2129, 565]. Approach [201, 277, 1636, 2157, 474, 1779, 278, 1170, 1782, 1283, 1348, 1176, 241, 1733, 1158, 160, 1315, 2155, 45, 819, 315, 419, 1770, 288, 1651, 111, 1187, 1488, 847, 2131, 1109, 336, 1032, 1347, 1739, 1410, 1655, 379, 626, 1108, 2133, 396, 1663, 279, 1664, 1407, 1243, 1405, 1598, 425, 851, 1674, 1411, 1894, 232, 1143, 845, 741, 439, 2190, 1708, 1693, 1422, 765, 769, 2231, 1725, 1259, 1266]. Approaches [251, 1867, 222, 1604]. Appropriate [1275, 1762]. Approximate [1683, 430, 1648, 1720, 779]. Approximated [563]. Approximates [147]. Approximating [1835, 1901, 112]. Approximation [743, 181, 139, 1061, 391, 1945, 1589]. Approximations [28]. Approximative [926]. Arbiter [1308]. Arbitrary [1838]. Arbitrated [1308]. Arbitration [2009]. Arc [717, 1460, 670]. Architectural [1023, 894, 1731, 2011, 486]. Architecture [165, 856, 861, 359, 1991, 616, 2105, 299, 343, 420, 304, 183, 1413, 1714, 62]. Architectures [1980, 371, 1360, 479, 1962, 163, 500, 1598, 874, 700, 1706, 1253]. Arcs [1001, 1754]. Area [1835, 350]. Argument [1557]. Arguments [1]. Arising [1985, 1129]. Arithmetic [1354, 1371, 1224, 455, 187, 242]. Arithmetics [1211]. ARQ [662]. Arrangement [714]. Arrangements [1864, 1907]. Array [174, 475, 1095]. Arrays [1949, 1950, 418, 1386, 349, 351]. Artificial [132, 196, 197, 105, 309, 1740, 1268, 1871, 334, 85, 1709, 1415]. ary [457, 459]. Asexual [445]. ASPARAGOS [2024]. Aspect [1514]. Aspects [295, 1588, 508, 1546, 802, 1066, 594]. Assembly [1407, 449]. Assertion [1737]. Assertional [1638]. Assessing [806]. Assessment [1641, 249]. Assets [1640]. Assignment [701, 1916]. Assignments [879, 1889]. Assimilation [667]. Assistance [1150, 24]. Assistant [2162, 2173]. Associating [862]. Association [1231, 1701]. Associative [2012, 290, 381, 1709, 784, 1263]. associative-commutative [784]. Assumption [669, 1723, 1438]. Assumption-Based [669, 1723]. Assumptions [1, 1313, 1666, 1716]. ASTRAL [1737]. Asymmetric [456, 1385]. Asymptotically [470, 642, 630]. Asynchronous [725, 2140, 2184, 610, 2130, 1133, 2138, 988]. Asynchronously [2139]. Ates [1805]. ATINF [1249]. Atlant [327]. atlas [589, 576]. ATMOSPHERE [484]. ATMS [671, 670]. Atomic [767, 191]. Atomized [1086]. ATPG [1127]. Attachment [846]. Attack [2205, 1581, 1297, 1082, 2226, 2203, 1134, 2206, 29]. Attacks [2227, 1595, 1308, 42, 27]. Attempt [2138, 1787]. attentive [1717]. Attribute [1500, 1505, 1498, 1514, 1515, 1511, 1517, 1506, 1512, 1499, 1501, 1508, 1509, 833, 1036, 1513, 1507, 1035, 227, 1519, 1504, 1502, 1447, 1448]. attribute-based [1448]. Attributes [1038, 969]. Attributive [1549]. Augmentation [739, 1919]. Aurora [384, 2104]. Authentication [957, 26, 2194, 1308, 1370, 40, 1605, 25, 27, 2196, 654, 2195]. Authenticators [2228]. Author [1346, 2230]. Auto [1125]. Auto/ Autograph [1125]. Autoepistemic [1486, 1226, 1227, 1679, 1462, 1463]. Automata [366, 1206, 1071, 2122, 1076, 1077, 283, 2005, 1085, 516, 1193, 1154, 1938, 1582, 1007, 1094, 537, 1145, 2190, 766, 773, 777, 792]. Automated [685, 1642, 1601, 1130, 2169, 1263, 1228, 712, 1453, 580, 582]. Automatic [277, 278, 579, 1441, 1735, 1780, 1119, 1136, 257, 1087, 2153, 302, 878]. Automatically [1155]. Automating [1681, 578, 223, 2181]. Automaton [1070, 1939]. automorphism [629, 628]. Autonomous [861, 1968]. Availability [163]. Avalanche [1336]. Average [1207, 551, 1897, 147, 1091].

Average-Case [1091]. Averaging [859]. Avionic [256, 255]. AVL [145]. Avoid [1133]. Avoiding [1118, 2163, 1044]. Axiomatizations [539]. Axioms [1861].

B [1830, 1820]. B-Tool [1820]. Back

[295, 664, 404, 1238, 106, 107].

Back-and-Forth [1238]. Back-End [295]. Back-Propagation [106, 107]. Backend [161]. Background [1550, 592]. Backtracking [665, 149, 1012]. Backup [1957]. Balance [145]. Balanced [461, 635]. Balancing [550]. Ballots [1875]. Balls [1881]. Bananas [876]. Bandwidth [139, 172]. Barbed [876]. Base [211, 626, 323]. Based [1636, 2157, 1750, 1554, 848, 1548, 6, 1172, 1148, 2207, 1626, 2047, 858, 713, 1595, 2151, 120, 669, 1234, 1158, 1157, 89, 1514, 1183, 1051, 44, 45, 983, 819, 1599, 357, 1923, 1189, 330, 463, 1192, 2191, 222, 480, 1130, 1396, 248, 1280, 2068, 1370, 427, 857, 1397, 2075, 1613, 859, 1337, 212, 140, 959, 1021, 832, 395, 425, 114, 1853, 1202, 1669, 1840, 3, 1751, 1340, 28, 1672, 449, 670, 316, 1065, 2214, 421, 1053, 1411, 161, 2061, 4, 1338, 1419, 1706, 591, 1723, 1695, 53, 1694, 981, 1443, 1416, 1691, 568, 2111, 588, 1448, 576]. based [271, 2216, 646, 794, 911, 396, 1663, 1433, 1450, 65, 914, 1710, 1677, 1428, 565, 1212]. Bases [478, 920, 290, 2068, 2063, 477, 476, 1420, 2051]. Basic [263, 1976, 2059]. Basis [1980, 1353, 1084, 1559, 481, 475, 1064, 1452]. Bayesian [573, 570, 809]. BBN [2108]. BCH [1358, 633, 1383]. BCI [972]. BCI-Lambda-Terms [972]. BCK [1106]. BCK-Formulas [1106]. BDDs [2166]. Be [39, 144, 1308, 8, 147, 806, 1568, 1558, 496, 1660, 978]. beam [567]. Beenker [1379]. Behavior [368]. Behavioral [982]. Behaviour [2118, 2136, 2173, 2139, 49]. Behavioural [1783, 512, 2151, 2164, 1241]. Behaviours [2143]. Belief [808, 823, 1626, 803, 669, 1629, 668, 673, 1486, 804, 809, 839, 1634, 1221]. Beliefs [799, 806, 1680]. Belt [201]. Benchmarking [241]. Bendix [687]. Bends [741]. Benefits [135]. Bent [461, 13]. Bermuda [1539]. Bernstein [1845]. Bernstein-Bezier [1845]. betaeta [976]. Better [545, 893, 1888]. Betting [1009]. Between [1539, 1626, 1139, 1034, 2056, 143, 1747, 466, 1442, 1179, 1003, 483]. Beyond [722, 820, 1965, 1438]. Bezier [1845]. Bi [1586, 433]. Bi-Partitioning [433]. Bi-Proof [1586]. Bibliography [262, 915]. Biconnect [1919]. Biconnected [529]. Bidirectional [1399]. Bifinite [1100]. Binary [1069, 658, 1583, 1910, 165, 1075, 1358, 925, 1579, 1497, 143, 11, 1091, 163, 182, 1378, 1383, 416, 634, 630, 635, 633, 234]. Binary-Search [1075]. Binding [2114, 891]. Binding-Time [891]. Bintree [140]. Biological [443]. Biology [2021]. Biomedical [848, 597, 589, 596]. Bipartite [561]. Bipartitions [732]. Birkhoff [704]. bisection [776]. Bisimilarity [378, 2148, 541]. Bisimulation [1155, 2149, 790]. bisimulations [772]. Bit [1299, 1583, 1326, 1602, 1325]. Bit-by-Bit [1583]. Bitcommitment [4]. Black [144]. Blackboard [2112]. Blast [196]. Block [1334, 35, 462, 82, 394, 635]. block-codes [635]. Block-Complete [394]. Body [830, 806]. Boltzmann [129, 99, 101]. Boltzmann- [129]. Boltzmann-Learned [99]. Boolean [1896, 631, 2188, 1214, 559, 1083, 1577, 14, 1942, 1096]. Borderline [1034]. Both [854, 1766]. Bottom [1911]. Bottom-Up [1911]. Bound [1068, 146, 1911, 661, 642]. Boundary [1838, 586]. Bounded [992, 467, 547, 1081, 2165, 2209, 1224, 906, 789, 2041]. Bounded-memory [2165]. Bounds [1312, 138, 16, 273, 1220, 1080, 1225, 1062, 476, 25, 789]. Box [1595]. Boxes [1597, 1336]. BP [1943]. brain [579, 577]. Branching [509, 1144, 2184, 2148, 2149, 1062, 569, 789]. Brave [2103]. Breadth [1071]. Breadth-Depth [1071]. Break [2219]. bridges [57]. Brief [2091, 1627]. Bristol [384]. Broadcast [1615, 1572, 2123, 1361, 437, 167]. Broadcasting [1313]. Broken [1568]. browsing [65]. BST [1091]. Buchberger [1359]. Buchi [1070]. Buddy [932]. Buddy-Tree [932]. Buffer [2153]. Build [813, 1700]. Building [1565, 132, 1121, 485, 2075, 1593, 1261, 270, 1712, 1256]. Bulk [319]. Burst [456]. Business [1768, 221]. Butterfly [2108, 1888]. Byte [1385]. Bytecode [2080]. Byzantine [2126].

C [603]. Cache [894, 2155, 1149].

Cache-level [894]. Caches [604]. Caching [340]. CAD [1185, 1736]. CAD-Package [1736]. CAD-Systems [1185]. Cadiz [1815]. calcifications [582]. Calculating [1495, 1376]. Calculation [458]. calculations [565]. calculi [770]. Calculus [2192, 2017, 692, 2152, 525, 2118, 1117, 610, 1130, 985, 1662, 281, 1411, 967, 1825, 1442, 771, 883, 783]. Call [190, 771]. call-by-value [771]. Calls [173]. Can [39, 144, 1568, 1308, 8, 1020, 734, 147, 1660, 217]. Canonical [523, 1374, 2050]. Can’t [1558]. capable [1444]. Capsules [606]. Capture [221]. capturing [1724]. Card [1332, 1331]. cardiac [580]. Cardinal [818]. cardinality [792]. Cards [2201, 1603, 957]. Carlo [1060, 565]. Carry [1340]. CAS [487]. Cascading [413]. Case [1100, 474, 1207, 75, 713, 1911, 551, 467, 220, 1897, 147, 1091, 1490, 1493, 1775, 597, 763, 1865, 787, 570, 1418, 490, 499, 1766, 485, 216, 1763, 496, 1764, 1767, 486, 229, 1765, 217, 231, 1290]. Cases [744]. Cash [2218]. CASTLE [1635]. CAT [670]. CAT-Correct [670]. Categorial [466, 1529]. Categorical [965, 1109, 312, 774]. Categories [402, 1102, 1191]. Categorization [116].

Category [1243]. Causal [980, 198, 1660, 1699, 1429]. Causality [983, 998]. causation [1464]. Cayley [176, 482]. CCS [990, 2166, 772, 1006, 790, 519, 2172]. Cell [1559, 1169]. Cellular [2005, 1938, 1582, 766]. Cellwork [1183, 1196]. Center [736, 494]. Centers [743, 740]. Certain [1581, 789]. Certified [1607]. Chain [1171, 69]. Chaining [1210, 1416]. Chains [20]. Challenge [307]. Challenges [1522, 1730, 1961, 1990]. Challenging [448]. chamfer [588]. Change [207, 635]. Changes [1742]. Channel [1916, 463, 1600, 1890, 434, 1894, 167, 659, 661]. Channel-Routing [434]. Channels [1392]. Chaos [550]. Chaotic [849, 1614, 380, 1576, 2020]. Chaotic-Map [1614]. character [643]. Characterisation [1669]. Characteristic [1306]. Characteristics [14]. Characterization [1069, 1180, 53, 1255]. Characterizations [2041, 2195]. Characters [459]. Chart [1525]. Chart-Parsing [1525]. Check [2172]. Checked [2154]. Checker [1819, 608]. Checkers [515, 1333]. Checking [2150, 1927, 1127, 1403, 1118, 2152, 1773, 2171, 2166, 1041, 2168, 997, 1144, 2184, 1123, 2167, 942, 2177, 958, 151, 511]. Chemical [197, 444]. Chernoff [661]. Chess [1307]. Chinese [465]. Chip [558]. Choice [992, 1086, 58, 906]. Chordal [169, 1891]. Chosen [2227, 2226]. Chromatic [391]. CICS [1801]. CIM [493]. Cipher [33, 1344]. Ciphers [1581, 1334, 1954, 1567, 1604, 18, 1340]. Ciphertext [2227, 2226]. Circle [170]. Circuit [1895, 2225, 1078, 78, 1497, 433]. Circuits [1122, 1127, 57, 2180, 2184, 1079, 1364, 1080, 1153, 1339, 1947, 1472]. Circular [717]. Circular-Arc [717]. Circumscription [1222]. Citizens [899]. Clarity [1004]. Class [625, 1304, 466, 139, 742, 1909, 376, 899, 434, 1811, 1943, 649, 63].

Classes [440, 1213, 1841, 1938, 870, 715, 270, 1290, 774]. Classical [2088, 861]. Classification [474, 402, 1637, 1838, 1410, 1655, 1745, 1439, 597, 587, 1449]. Classifications [2050]. Classifier [117, 118, 109, 1698]. Classifiers [406]. Clausal [691, 1240]. Clause [687, 1213, 710]. Clauses [686, 523, 675, 2096, 999, 1256, 1468]. Clean [373]. Client [187]. Client-Server [187]. Clock [45, 2126]. Clock-Controlled [45]. CLOS [329]. Closed [2057, 1664]. Closely [1963]. Closest [1922]. Closure [286, 1077, 2204, 153]. CLPS [1056]. Cluster [847]. Clustering [928, 1853]. CM [2213]. CM-Curves [2213]. CMS [666]. CMS/ATMS [666]. Coarse [1049]. Coarse-grained [1049]. Code [1171, 2135, 1360, 1033, 387, 621, 1747, 897, 874, 652]. Coded [653, 463, 462, 455, 1388, 661]. Codes [457, 253, 548, 1358, 466, 1611, 1364, 1349, 1367, 1369, 1351, 1371, 414, 1372, 1223, 1378, 1379, 1381, 1383, 456, 1385, 462, 25, 453, 27, 1387, 2196, 1392, 634, 629, 630, 635, 658, 657, 631, 637, 628, 627, 653, 651, 640, 636, 638, 786, 643, 633, 641, 2195, 642]. codeword [657]. Coding [1830, 659, 77, 78, 1370, 464, 413, 656, 660, 654]. Coercive [872]. Cognition [132]. Cognitive [1401, 2016, 1466, 1253]. Coherence [1107, 2155, 974]. Coherent [638]. Coin [1299, 1313]. coincidence [564]. COLD [1776, 1785, 1736]. COLD-1 [1776, 1736]. COLD-K [1785]. Collaboration [1290]. Collage [1186]. Collapsed [965]. Collapsing [533, 1103]. Collection [894, 2102, 1212, 346, 345]. Collective [1313, 2022]. Collector [1495]. collimator [564]. Collision [30, 1856]. Color [392]. Color-Exchange [392]. Colored [1904, 368, 903, 911]. Coloring [141, 392]. Coloured [61, 60]. colposcopic [597]. Combination [917, 222, 801]. Combinator [381]. Combinatorial [857, 2023, 1387, 2195, 1906]. Combinatorics [1907]. combinatory [794]. Combined [458, 696]. Combiners [18]. Combining [1420, 1742, 1774, 485, 1435]. Come [1522]. Command [441]. Commands [360]. Commercial [2063, 494]. Commitment [1299, 1668]. Common [178]. Communicating [985, 2158, 1778, 1143, 2139]. Communication [2140, 1300, 1231, 2123, 2112, 610, 422, 172, 1128, 299, 214, 1951, 2210, 167, 370, 988, 662, 660, 64]. Communications [134]. Commutations [520, 796]. Commutative [521, 1606, 54, 784, 1263]. Compact [171, 1217]. Compaction [1867]. Comparability [170]. Comparative [251, 130, 52]. Comparing [2189]. Comparison [1913, 713, 45, 1509, 1785, 1613, 1337, 932, 660, 2044, 989, 1338, 593]. Comparison-Efficient [1913]. Comparisons [803]. Compatibility [705]. Compete [2089]. Competitive [2110, 735]. Competitiveness [1915]. Compilation [179, 1039, 241, 152, 2082, 1722, 1697, 914]. Compiler [282, 1821, 1027, 1512, 873, 1811, 1053, 1248, 448]. Compiler-Controlled [873]. Compiling [697, 2064]. complement [784]. Complementarity [1347, 2002]. Complete [2066, 1904, 539, 951, 868, 1942, 2193, 394, 546, 77]. Completely [466, 1381]. Completeness [693, 1005, 1470]. Completion [686, 1230, 687, 689, 690, 1236, 524, 563, 1243]. Complex [288, 934, 2002, 1807, 1660, 1931, 2042, 619, 130, 649, 768, 1721]. Complexes [936]. Complexity [468, 355, 1904, 1207, 1072, 1935, 1334, 713, 16, 1580, 1076, 520, 2179, 350, 283, 273, 478, 1940, 1078, 551, 1144, 963, 1578, 971, 191, 1088, 126, 1938, 162, 389, 1374, 1575, 1223, 15, 916, 353, 1380, 2021, 1594, 1937, 2210, 1065, 1096, 871, 902, 647, 781]. Component [1019, 1560]. Components [498, 256, 529, 1491, 611]. Composite [44]. Composition [1752, 1233, 2163, 510, 210, 796, 64, 905, 913]. Compositional [2150, 2117, 992, 280, 995, 1137, 1925, 2131, 951, 1128]. Compound [419]. comprehensibility [1446]. Comprehensions [896]. Compressible [1345]. Compression [1588, 1862, 602, 1410, 413]. Computability [2043]. Computation [722, 2069, 138, 1304, 1301, 185, 1217, 533, 1123, 2167, 1188, 1365, 2056, 1341, 841, 2223, 1268, 1374, 1846, 1380, 1384, 153, 337, 52, 628]. Computational [508, 557, 1076, 261, 917, 389, 2018, 1573, 353, 1124, 450, 1066, 1702, 798]. Computations [1353, 543, 1059, 2128, 341, 275, 1390, 1254]. computed [567]. Computer [1150, 1174, 952, 1866, 1908, 909, 1815, 1280, 279, 1873, 134, 1875, 1140, 1672, 1882, 1708, 2232]. Computer-Aided [1140]. Computerized [2224]. Computers [2006, 411]. Computing [1279, 2222, 554, 749, 1072, 1356, 512, 528, 719, 1205, 438, 1952, 559, 751, 742, 289, 339, 1939, 2149, 470, 1974, 1654, 1195, 1848, 937, 1968, 1589, 135, 1992, 338, 187]. Concatenated [627, 652]. Concatenation [462]. Concept [2005, 45, 2094, 2067, 86, 2068, 1696, 1685, 1682, 1257]. Concept-Based [2068]. Concepts [265, 1976, 1029, 1732, 1509, 332, 1533]. Conceptual [816, 234, 2064, 1620, 226, 199]. conceptualizations [1434]. Concerning [1334, 1590, 777]. Concise [2070]. Conclog [2231]. Concrete [2094]. Concurrency [365, 2161, 678, 1009, 1010, 1787, 759, 793]. Concurrent [697, 355, 2156, 1861, 2151, 367, 983, 1364, 1190, 2081, 616, 1769, 2143, 373, 611, 411, 2231]. Condition [706]. Conditional [1623, 688, 693, 674, 990, 1641, 810, 677, 678, 695, 710, 802]. Conditions [1070, 1154]. Cone [567]. Confidence [855]. Configurable [501]. Configuration [1730, 379, 1203, 1692]. configuring [776]. Confirmation [30]. Confirmatory [594]. Conflict [167]. Conflict-Free [167]. Conflicting [803]. Confluence [976, 695, 682]. Confluences [1403]. Confluent [525, 520, 1375]. Conformational [81]. Confront [2138]. Congruence [544]. conjunction [761]. conjunctive [1686]. Connected [1949, 174]. connection [1261]. Connectionist [2013, 857, 2075]. Connectionist-Like [2075]. Connections [863]. Connectivity [816, 738, 781]. Conquer [1909, 2007]. cons [1466]. Consciousness [2001]. Consecutive [2056]. consequences [631]. Consequential [1961]. Considerations [1210, 2011, 1590]. Considered [1246]. Consistency [1067, 1966, 688, 1403, 1773, 1819, 1406, 608, 2099, 408, 2039, 834, 1460]. Consistency-Checking [1403]. consisting [1420]. Constant [1941]. constants [1466]. Constrained [75, 707, 1007]. Constraint [1025, 104, 2084, 2078, 2102, 1022, 2092, 535, 1396, 1397, 1394, 1663, 151, 1048, 1460, 1471]. Constraint-Based [1396, 1397, 1663]. Constraints [1068, 1049, 2046, 2047, 926, 1393, 297, 804, 961, 1890, 1241, 1395, 1411, 1686, 1684, 1258, 1688]. Constraints-Based [1411]. Construct [12]. Constructing [1836, 256, 1306, 748, 1902, 1454, 1563]. Construction [457, 277, 278, 321, 1515, 17, 1512, 1336, 1367, 1921, 636, 1154, 106, 1379, 2004, 1602, 27, 1053, 1169, 2212]. Constructions [692, 1361, 1363, 13, 1112, 456, 1352, 780]. Constructor [175]. Constructs [1020, 407]. Containment [701]. Content [196, 918, 358, 157, 1540]. Content-Addressable [358]. Content-Oriented [918]. Contents [1832, 1823, 1422]. Context [1553, 1177, 1180, 1181, 2148, 948, 1139, 1632, 1194, 1986, 1098, 157, 1563, 600, 780, 1703]. Context-Free [1180, 1181, 1098, 1177, 1194, 780, 1703]. Contextual [664, 2086]. Continuations [881, 966]. Continuing [1277]. Continuous [1661, 1429, 571]. contours [579]. Contracting [1948]. Contraction [166, 1093]. Contractor [254]. Contradiction [1400]. contrast [600]. Control [2192, 457, 1548, 498, 664, 1147, 1652, 2130, 252, 1394, 1104, 489, 2127, 1385, 1311, 83, 420]. Controlled [2027, 1272, 45, 320, 873, 909]. Controller [1646, 257]. Controllers [1148]. Controlling [2103]. Conventional [1488]. Convergence [69, 107]. Convertible [1309]. Convex [52, 1903, 1921, 936, 1893]. Convolutional [1364, 653, 651, 652]. Conway [410]. Cooperating [225, 1254]. Cooperative [405, 1711]. Coordination [1537, 1677]. Coping [1969]. Coprocessor [244]. copulative [1727]. Copy [1795]. Copying [2111]. Copying-based [2111]. Corporate [1652]. Correct [435, 1014, 670]. Correcting [1358, 867, 1223, 456, 1385]. Correction [1613, 1390, 400, 1362, 566]. Correctly [806]. Correctness [963, 1934, 2173]. correlated [661]. Correlation [2197, 19, 1581, 461, 18, 460, 645]. Correlation-Immune [2197]. CORSAIR [1332, 1331]. COSMOS [2175]. Cost [527, 1929]. Costly [144]. Coteries [1887]. Counterpoint [254]. Counting [1061, 346]. Coupled [363, 1963, 700, 1469]. Couples [849]. Coupling [224]. Course [448]. Courses [1270, 1277, 1273]. Cover [1945]. Covering [451, 2163, 632, 1369, 1387, 658]. Covers [745, 412]. Cpo [1784]. Created [197, 108]. creation [1681]. Creativeness [1941]. Creativity [403]. CREW [1923]. Criteria [1595, 852, 527, 1577]. Criterion [1336, 834]. Critical [188, 1082, 1243, 246, 2124]. Critique [226]. Crop [854]. Cross [460]. Cross-Correlation [460]. Crossed [684].

Crossover [73, 1451]. crplus [1943]. Cryptanalysis [1614, 1566, 1296, 2202, 2207, 1601, 45, 1570, 1567, 28, 1298]. Cryptanalytic [1613]. Crypto [1568]. Cryptographic [1, 1322, 43, 21, 2224, 2213, 1339, 1575, 1323, 37, 1340, 2193, 483, 650, 645]. Cryptographically [2229]. Cryptography [23, 32, 1333, 1608, 1327, 1591, 5]. Cryptology [1058, 1606]. Cryptosystem [1614, 2207, 1335, 1611, 1297, 1576, 32, 1570, 1612, 28]. Cryptosystems [1592, 1296, 1306, 34, 2204, 1337, 2211, 1332, 1331, 1338, 2212]. CSP [1783, 989]. Cube [409, 770]. Cubes [390]. Cubic [530, 431, 1204]. Cumulative [1474, 1480, 1678]. Cumulativity [1475, 1668]. Current [251, 1870, 593]. Curricula [1279]. Curriculum [1278]. Curve [1837, 1306]. Curved [1382]. Curves [1354, 1592, 472, 2213, 2191, 1593, 640]. Customer [254]. Customization [376]. Cut [721]. Cuts [530, 738]. Cutting [1216]. Cycle [179, 1171]. Cycles [536, 1660]. Cyclic [1360, 1593, 1689].

d [936, 1706, 580, 602, 1652, 569, 777, 576, 585, 601, 574, 475]. D-volume [602]. D.E.S. [1334]. Dactl [1184]. Dark [2022]. Darwin [129]. Darwin- [129]. Data [2069, 1231, 1588, 2170, 918, 2047, 893, 2083, 1735, 1644, 1863, 321, 309, 929, 287, 380, 315, 360, 920, 1033, 727, 752, 1410, 312, 934, 1090, 1944, 1126, 212, 127, 2076, 1956, 1853, 1671, 958, 2033, 2113, 924, 970, 1247, 845, 157, 921, 938, 1712, 592, 579, 1707, 658, 765, 578, 1716, 572, 576, 1259, 1266, 1453, 575]. Data-Driven [2083, 380, 1259]. Data-Flow [1033]. Database [324, 1554, 218, 1028, 2060, 2046, 2038, 1643, 154, 608, 1182, 424, 311, 226, 917, 1868, 2056, 317, 159, 1796, 935, 323, 328, 316, 2044, 155, 292, 1880, 844, 327, 1713, 1422, 1423]. Databases [325, 314, 1647, 2037, 1857, 930, 357, 2058, 331, 2034, 2035, 2040, 2049, 426, 1673, 2045, 1674, 213, 2061, 1424, 1426, 1427, 1428]. Dataflow [359, 951, 2134, 1116, 874]. Datalog [325, 2041, 1685]. Date [1487]. DB [1554]. DBE [218]. DBMS [933, 939, 250]. deadline [1469]. deadline-coupled [1469]. Deadlock [361, 2178, 1143]. Deadlock-Free [361]. Deadlocks [243]. Deal [2201]. Dealing [1403, 856, 213]. Debugger [1802]. Debugging [1409, 2101, 362, 1050]. Decade [235]. Decentralized [123]. Decidability [520, 1098]. Decidable [2148, 538, 787]. Deciding [1213, 982, 2187, 270, 541]. Decimation [1201, 1391]. Decision [689, 1211, 858, 853, 1077, 559, 419, 357, 852, 86, 1092, 1884, 797]. Decision-Making [1884]. Decisions [854]. Declarative [2089, 2080, 1409, 2038, 307, 269, 2036, 2076, 2077, 2061]. decoder [663]. Decoding [651, 640, 1351, 641, 1388, 662]. Decomposability [548]. Decomposition [469, 429, 2131, 931, 1193, 1891]. Decrementing [2012]. Dedicated [2062]. Deduction [707, 678, 1480, 712, 1431, 1263]. Deductions [1228]. Deductive [1492, 2046, 1620, 2037, 2034, 2035, 1546, 1673, 2045, 2044, 844, 1424, 1427]. Default [201, 822, 1624, 1638, 1630, 863, 1654, 1658, 807, 1661, 1668, 1669, 1480, 1675, 1698, 1716, 1466, 1431]. Defaults [1653, 1676, 1684, 1466]. Defeasible [1643, 1458, 1256]. Deficient [429]. Defined [1173, 288]. Definition [1044, 611, 945, 944]. Definitions [2080]. deformable [591]. Degree [548, 776]. Degrees [533, 476, 1094]. Delaunay [556, 718]. Delay [409, 2183, 2180, 547, 1153]. Delay-Insensitive [1153]. Delayed [1838]. Delays [900, 1403]. Delegation [957]. Delft [1812]. Delta [1190, 633]. Delta-Grammars [1190]. Demand [885]. Dempster [800, 1657]. Dempster-Shafer [800, 1657]. Dendrograms [414].

Denotational [372, 1108]. Dense [536, 560]. Density [1569, 1082]. Dependable [1982, 1974]. Dependence [1625, 959, 1221]. Dependence-Based [959]. Dependencies [2047, 2045, 1689]. Dependency [151, 1695, 1705]. dependency-recording [1695]. Dependent [1924, 2184, 496, 1139, 1246]. Depictions [1559]. Depth [1071, 1079, 1096]. Dequeue [1071]. Derandomization [729]. Derivability [1478]. Derivation [1493, 345]. Derivations [1197]. Deriving [2084, 2193, 1792, 351, 1422]. DES-Like [1296]. Describing [1167]. Description [2091, 1752, 1117, 185, 2093, 1768, 1549, 453, 703, 767]. Descriptions [1119, 1497, 2077]. Design [2027, 1830, 614, 355, 1741, 189, 432, 1284, 698, 2194, 218, 1734, 208, 225, 1731, 76, 1595, 608, 249, 361, 1234, 1772, 983, 1185, 110, 406, 1740, 342, 2058, 1285, 2131, 1815, 852, 1739, 1128, 1289, 461, 1604, 2120, 2105, 1286, 985, 1287, 1775, 1873, 418, 1063, 443, 37, 217, 1743, 316, 2159, 1247, 155, 1248, 909, 61, 1426, 1711, 1721, 564]. Designed [449, 633]. Designer [614, 1711]. Designing [407, 159, 1951]. desire [1414]. detecting [637]. Detection [881, 1042, 1364, 1043, 389, 186, 345, 1390, 1143, 579, 582, 585]. Determinate [1011]. Determination [1357, 1842, 868]. Determined [1660]. Determines [1660]. Determining [553]. Determinism [1043]. Deterministic [1071, 283, 1949, 1062]. DEVA [1792]. Develop [189]. Developing [1291, 259, 1161, 1292]. Development [1750, 1496, 262, 263, 264, 265, 266, 267, 499, 492, 1492, 491, 329, 623, 1774, 1828, 1528, 1831, 245, 497, 1488, 1000, 1497, 493, 1829, 1489, 222, 487, 1805, 1778, 2137, 1878, 1794, 494, 1986, 1014, 2144, 447, 624, 1795, 1721, 911]. Developmental [1169]. Developments [1790, 342, 1490]. Diagnose [421]. Diagnosis [1637, 711, 2158, 1695, 1694]. diagnostic [1699]. Diagram [719, 1853].

Diagrams [1121, 1850]. Diameter

[1899, 1848]. Dickson [1610]. dictionaries [788]. Dictionary [625]. Difference [1645, 2056, 13]. Differences [131]. Different [124, 259, 798, 1694]. Differential [1566, 1296, 2202, 1074, 1595, 1363, 1567, 477, 1389, 583]. Differential-Like [1595]. Digest [46, 1317]. Digital [1122, 1310, 1603, 1831, 42, 419, 1328, 2127, 1714, 582, 587, 660]. Dilemma [124]. Dimension [478, 233]. Dimensional [390, 2026, 1105, 237, 1867, 1916, 937, 715]. Dimensions [723, 1838, 936]. Diminishing [1215]. Diophantine [769, 1589]. dipole [575]. DIPSY [1716]. DIPSY-E [1716]. Direct [1712, 1111, 1573, 629]. Directed [1210, 1027, 536, 753, 1225, 670, 1247]. Directory [423]. DIS [42]. disambiguation [1701]. Discipline [943]. discovery [1453]. Discrete [6, 1304, 44, 1147, 95, 1599, 1341, 826, 1337, 281, 1589, 1594, 1345, 1338, 568, 1258]. Discrete-log [1345]. discriminants [599]. Discrimination [834, 1725]. Discussion [1960, 1967, 1971, 1975, 1983, 1988, 1997, 1160]. Dishes [557]. Disjoint [752, 1920, 695]. Disjunctions [2071]. Disjunctive [2035, 1471, 1456, 1255]. Disks [1881]. Disorder [1865]. Dispersion [744]. displays [1417]. dissolution [1467]. Distance [466, 737, 853, 393, 1381, 1450, 634, 630, 588, 633]. Distances [528]. Distinguishing [2149]. Distributable [908]. Distributed [2192, 124, 188, 436, 189, 2071, 1023, 251, 1966, 383, 1734, 856, 1976, 438, 1029, 1887, 280, 995, 173, 1003, 1006, 1982, 618, 1974, 252, 1932, 2009, 344, 10, 1999, 1870, 711, 1871, 1585, 346, 186, 193, 2015, 180, 98, 1956, 2137, 1996, 250, 94, 345, 122, 194, 192, 50, 1423, 59, 596, 1729]. Distributed-Memory [383]. Distributing [1786]. Distribution [1985, 1315, 1314, 630]. Distributions [1632, 592, 773]. disturbed [648]. Divertible [1]. Divide [1909, 2007].

Divide-and-Conquer [1909, 2007]. DNA [81]. Do [2089, 1880, 806, 1784]. Document [402, 1328, 156, 161]. Does [1866, 941]. Domain [1642, 1541, 2099, 1811, 1699]. Domains [1100, 860, 138, 853, 1393, 2094, 505, 1394, 946, 1959, 1048]. Domesticating [1020]. Dominance [1923]. DOMINO [495]. Door [1511]. Dot [274, 2051]. Double [1358, 1158, 1918, 1196]. Double-Error [1358]. Double-Wall [1196]. Doubtful [839]. Down [623, 2155, 624, 906]. Doyle [1227]. Drawing [1736]. Driven [2083, 1406, 380, 1147, 487, 443, 1745, 1248, 2172, 1259]. Driver [664]. DSP56000 [21]. Dual [1358, 481, 578]. dual-echo [578]. Duality [1099, 483]. Duals [1383, 633]. Duplicates [303]. Duration [2116]. Dynamic [1905, 718, 716, 747, 529, 405, 2067, 298, 1463, 928, 1043, 2048, 175, 1242, 1850, 107, 1922, 83, 1746, 2053, 765, 593, 790, 1418]. Dynamical [2017, 1645]. Dynamically [1741, 367, 288, 604, 750]. Dynamically-Typed [604]. Dynamics [124, 2028, 120, 2020, 889].

E-Unification [710]. Easy [1030]. echo [578]. echocardiogram [595]. Ecological [124]. Ecosystem [851]. EcrinsDesign [219]. Edge [1904, 600, 1920, 738, 1204, 585]. Edge-affected [600]. Edge-Colored [1904]. Edge-Disjoint [1920]. Editor [219, 1819, 1163, 1880, 1249, 1163]. EDS [292]. Education [1279, 1277]. Effect [1070]. Effective [77, 2009, 2043]. Effects [445]. Efficiency [949, 2084, 1210, 47, 1060]. Efficient [432, 1913, 1722, 2225, 181, 2114, 1172, 1045, 2176, 1626, 1142, 1361, 341, 883, 1603, 536, 359, 2178, 1219, 1033, 1424, 2095, 2087, 891, 930, 1909, 971, 780, 721, 2056, 527, 435, 530, 2034, 159, 2212, 22, 1011, 1407, 186, 897, 151, 343, 1035, 1053, 194, 1254, 183, 2061, 351, 337, 1425, 1703, 797]. Efficiently [354, 1948, 148]. Effort [1015]. EILENBERG [549]. Election [755, 438, 59]. Elections [2224]. ELECTRE [1738]. Electrical [1715]. electromagnetic [575]. Electronic [2218, 247]. Electronics [464, 85]. Elementary [1182, 1616, 1676, 900]. Elements [1496, 202, 476, 1383]. Elicitation [1631, 1295]. Elimination [1173, 756]. ElipSys [2102]. Ellipsoids [1881]. Elliptic [1354, 1592, 1306, 2191, 1593]. Embedded [259, 260]. Embedding [506, 1910, 1095, 731, 989]. Embeddings [348]. Embryological [85]. Embryos [85]. Emergence [132, 127]. emission [567]. Emphasis [1965]. Empirical [828, 200, 748, 1453]. Encapsulation [1993]. Enclosing [1881]. Encoders [463, 454]. Encoding [1360, 938, 1688]. Encryption [35, 1370, 1330, 29]. End [295, 1812, 207]. End-User [207]. Engine [1548, 1547]. Engineering [1553, 1291, 1274, 1281, 1279, 239, 1552, 1282, 209, 1752, 1759, 1760, 1294, 1761, 1757, 1275, 1739, 1280, 1289, 1277, 484, 1762, 1293, 1273, 1295, 1290]. Engineers [1276]. Engines [386]. Enhance [449]. Enhanced [188]. Enhancement [230, 600]. Enhancing [1588]. Ensembles [1578]. entailment [1683]. Entailments [1670]. Entities [862, 1838, 1242, 425]. Entity [1182, 316, 1674]. Entity-Relationship [1182, 316, 1674]. Entropy [859, 1666, 655]. Enumerating [738, 1596]. Enumeration [724]. Envelopes [876]. Environment [613, 436, 189, 1120, 1552, 329, 1744, 1738, 1528, 216, 1512, 493, 1198, 229, 1796, 1847, 215, 1715, 1440, 1162]. Environmental [1961, 919, 1671]. Environments [1288, 1750, 1955, 492, 1752, 248, 484, 1753]. Epistemic [799]. EPOS [225]. Equality [686, 947, 1091, 710, 1438]. Equation [1930]. Equational [706, 685, 688, 707, 1236, 2053, 704, 1261]. Equations [1168, 2047, 1645, 352, 1017, 1389, 709, 817, 769, 1453]. Equilibria [88]. EQUIP [300]. equipment [1715].

Equivalence [701, 1174, 1139, 2169, 1124, 1929, 49, 902]. Equivalences [2176, 2164, 982, 1238, 679]. Equivalent [1611]. Error [457, 1358, 1220, 1364, 1613, 1372, 1223, 2158, 102, 1385, 455, 1390, 1362, 637, 628, 636]. Error-Coded [455]. Error-Correcting [1223]. Error-Correction [1613, 1362]. error-detecting [637]. Error-Free [102]. Errors [456]. ESIGN [1603]. Essential [228]. Essentially [26]. Establishment [1275]. ESTEREL [2138]. Estimates [291]. estimation [586]. Euler [1918]. Eulerian [1904]. Eureka [1732]. EUROCRYPT [1614]. Evaluating [843, 1883]. Evaluation [1500, 1505, 1046, 1492, 2038, 160, 399, 336, 1506, 2059, 826, 340, 299, 1617, 108, 2042, 1035, 381, 878, 1504, 1424]. Evaluator [877]. Evaluators [1501, 1502]. Event [1147, 998, 1972, 1464]. Event-Triggered [1972]. Events [1641, 810]. Every [1660]. EVES [1791]. Evidence [803, 811, 806]. Evidential [836, 1655]. Evolution [2025, 104, 76, 131, 96, 97, 126, 1286, 443, 98, 2022, 113, 2016, 130]. Evolutionarily [108]. Evolutionary [2027, 2029, 114, 122]. Evolutionsstrategie [95]. Evolving [315]. Exact [1093, 392]. examination [570]. Example [1492, 960, 135, 1883]. Examples [1230, 842, 1447]. Exceptions [618, 1682]. Exchange [2, 2201, 1888, 392, 9]. Exclusion [1887]. Exclusive [1947]. Executable [1803]. Execution [860, 1738, 2102, 2111, 1256, 224, 1050, 1151, 2033, 343, 306]. Execution-Led [860]. Exhaustive [885]. Existence [1227]. Expanded [1595]. Expansions [1486, 1226]. Expectations [1621]. Expected [556, 718]. Expediency [1762]. Experience [259, 289, 1754, 1711]. Experiences [1801, 500, 93, 1878, 1053]. Experiment [294]. Experimental [23, 1184, 1846, 447]. Experimentation [220, 1702]. Experiments [293, 1012]. Expert [218, 197, 209, 1641, 810, 856, 855, 399, 1631, 1187, 1656, 403, 1663, 1666, 300, 838, 596]. Expertise [1291, 1294, 1293, 1292, 1295]. Explainable [1994]. Explaining [1155]. Explanation [1443]. Explanation-based [1443]. Explanations [198]. Explicit [1402, 90, 793]. Exploiting [2221, 258, 939, 83]. Exploring [2099, 113]. Explosion [1118, 2163, 1133, 1134]. Exponential [452, 467]. Exponentiating [20]. Exponentiation [31, 481]. Exponents [1345]. Expressed [1631]. Expression [1159]. Expressions [890, 274, 321, 699, 1555]. Expressive [518, 955]. ExSpect [1814]. Extended [1319, 1300, 1296, 2202, 179, 6, 1312, 1304, 1, 1211, 2229, 41, 243, 2, 1308, 2228, 1584, 1182, 8, 2201, 1629, 920, 999, 676, 1341, 7, 18, 1239, 2204, 40, 1329, 1612, 1400, 2219, 2211, 1245, 37, 5, 9, 670, 2193]. Extending [933, 895, 1700, 1050, 1260, 867, 793, 316]. Extensible [2060, 939, 2098]. Extension [2107, 1152, 2019, 1783, 977, 2168, 1648, 1368, 1681]. Extensional [506, 794]. Extensionality [976]. Extensionally [965]. Extensions [1305, 1119, 1654, 1534, 149, 415, 1437]. External [1047]. Externally [749]. Externals [1517]. Extraction [592, 526, 975, 1665, 1743, 1018, 1426]. Extremes [556]. EZ [1780].

F [977]. Face [110, 1876]. Facets [619]. Facility [326, 744]. factor [593, 594]. Factoring [6, 7, 1589, 4]. Factorization [1337, 1338]. Factorizations [521]. Factors [473]. Factory [1732]. Facts [2017, 1673]. Fading [1392]. FADS [593]. Failure [822, 1746, 988, 1684, 1267]. Failures [1933, 421, 988]. Fair [1301, 369]. Fairness [1777, 1010]. faithfulness [758]. families [642]. Family [1270, 1344, 630, 1679]. Fare [557]. Farthest [723]. Fast [1980, 1054, 1581, 177, 1212, 534, 729, 800, 1333, 1949, 470, 1351, 358, 1330, 1944, 1092, 38, 1946]. Faster [1920, 2217, 20]. Fault [465, 1979, 1980, 2157, 176, 195, 1978, 165, 2117, 2145, 1976, 2128, 1987, 2129, 2127, 1977, 2126, 184, 439, 1432]. Fault-Masking [2127]. Fault-Tolerance [176, 1978, 1977, 439]. Fault-Tolerant [465, 1980, 195, 165, 2117, 2145, 1976, 2128, 2129, 1432]. Faulty [1910]. FBP [568]. FEAL [1297, 1344, 2203, 1566]. FEAL-4 [2203]. FEAL-6 [2203]. FEAL-8 [1297]. Feasible [458]. Feature [702, 106]. Features [327, 1440]. FEC [662]. FEC-ARQ [662]. Federated [229]. Feedback [1412, 1145]. Feedforward [112]. Fence [1766]. Fewness [1073]. FFTs [649]. FGCS [956]. FI [918]. FI-Quadtree [918]. Fiat [40, 38]. fiberings [1262]. Fibrations [1116]. FIDO [2099]. Field [474, 479, 470, 1660, 581]. Fields [1553, 1061, 1341, 1303, 481, 1380, 9]. Fifth [497]. Figures [719]. File [1958, 1890, 423, 1957, 302, 760]. Find [1869]. Finding [2026, 713, 716, 1924, 536, 727, 467, 544, 732, 1343, 1386, 1919]. Findings [502]. Fine [339]. Fine-Grain [339]. Fingers [1880]. Finitary [281]. Finite [542, 1070, 138, 479, 1393, 1137, 680, 1085, 1061, 1303, 544, 2169, 1394, 2099, 481, 2158, 1380, 475, 1013, 270, 1048, 1143, 817, 777, 792]. Finite-graph [1013]. Finiteness [892, 1154, 2172]. FINSIEL [1756]. First [686, 1806, 376, 899, 970, 118, 1882, 1722, 1420, 1260]. First-Class [376]. First-Order [686, 1260]. Five [1573]. Fixed [702]. Fixed-Point [702]. Fixpoints [1227, 204]. Flang [2097, 2079]. flattening [1257]. Flexible [384, 1528, 404, 1849, 215, 2106, 1723, 1257, 1432]. Flight [2127]. Flight-Control [2127]. Flip [202]. Flip-Flops [202]. Flirt [314]. Flops [202]. Flow [893, 360, 1033, 1503, 958, 1436]. Flowgraph [1194]. Flowshop [2010]. FLR [1054]. Flsim [976]. fluctuations [659].

Fluid [120]. fly [2164, 2165]. FNC [1517]. FNC-2 [1517]. Focus [674, 207]. fold [26]. Folding [142, 1044]. foldings [912]. FOLI [334]. Folklore [722]. Forbidden [1206, 544]. Forgery [39]. FORK [356]. Form [886, 1784]. Formal [979, 1789, 499, 1209, 1122, 2145, 1828, 1733, 377, 2155, 185, 2184, 133, 1491, 471, 1253, 1771, 1598, 2127, 2137, 2175, 1794, 1799, 1795, 1430, 758]. Formalisation [368]. Formalism [1523, 1244, 1435]. Formalisms [1549]. Formalization [1648, 1777, 1527, 2053]. Formalizing [827, 1772, 1650]. Formally [257]. Formation [402, 86, 1698]. Formed [1883]. forms [778]. Formula [1376]. Formulae [1083]. Formulas [1214, 514, 2149, 1106]. Formulations [991]. Formulator [1562]. Forth [1238]. Forward [1210, 2096, 1416]. Foundation [952, 710]. Foundations [2222, 263, 671, 712]. Four [1187, 1343, 135]. FP [694]. FPL [694]. Fractal [415]. fractional [655]. Framework [366, 205, 1028, 2151, 926, 863, 999, 2082, 501, 1804, 2137, 1995, 1140, 203, 791, 591, 1690, 1458, 1251, 1258]. FRATS [1024]. Free [2152, 361, 521, 1180, 1181, 992, 2148, 30, 1092, 102, 1856, 1098, 167, 1177, 58, 906, 780, 1194, 1703]. Freedom [2178]. Frege [1216, 945]. Fresco [606]. Freshman [1287]. Front [1812]. FSS [1687]. FSS-WASTL [1687]. FTAM [1131]. Full [1067, 1910, 948]. Fully [718, 752, 175, 1469, 351]. Fully-Pipelined [175]. Function [1618, 1568, 1935, 1583, 461, 343, 122, 1424]. Functional [895, 440, 1496, 2080, 2060, 208, 2047, 2168, 1020, 2095, 336, 899, 694, 1043, 162, 2079, 2097, 876, 486, 212, 437, 897, 2210, 2045, 1018, 318, 573, 572]. Functionality [2077]. Functions [887, 813, 1896, 1072, 138, 2197, 1173, 19, 1626, 803, 1099, 272, 273, 1940, 1301, 559, 112, 1482, 804, 11, 1917, 2215, 1330, 30, 13, 1376, 12, 1577, 14, 1602, 37, 1494, 1096, 343, 1609, 1391, 1316, 351, 631]. Fundamental [1926]. Funds [247]. Furnaces [196]. Fuss [1858]. Future [1979, 2062, 1730, 1993, 1961, 1974, 927, 307, 1984, 1245, 1879]. Futurebus [340]. Fuzzy [816, 1834, 831, 848, 843, 918, 1645, 1646, 819, 1652, 814, 852, 1396, 1659, 859, 815, 850, 832, 851, 118, 598].

G [336, 344, 2036]. G-Line [344]. G-Log [2036]. GALATEA [338]. Galois [1389]. game [591, 1417]. game-theoretic [591]. game-tree [1417]. Games [124, 99]. GAML [875]. GANNET [110]. Gap [1229, 1529]. GAP-Handling [1529]. Gap-Language-Technique [1229]. Garbage [1495, 894, 2102, 1212, 346, 345]. GauBeta [1084]. Gaussian [659]. GCD [470]. GCLA [2090]. General [954, 836, 1168, 831, 1356, 514, 1301, 2011, 289, 2224, 2082, 34, 527, 2134, 204]. generality [1446]. Generalization [1222, 1418, 1854, 1411]. Generalized [1358, 360, 2059, 326, 1921, 1379, 462, 2126, 184, 818]. Generated [1618, 1582, 912]. Generating [766, 2166, 1181, 1033, 553, 198, 442, 753, 2020, 148, 1851, 897, 415]. Generation [1978, 1138, 1027, 550, 1862, 2228, 202, 1991, 534, 497, 515, 435, 1449, 1560, 38, 320, 2033, 1753, 1480, 874, 66]. Generative [1514]. Generator [1579, 427, 1372, 458, 1811]. Generators [2215, 1326, 1602, 2216]. Generic [2189, 1247]. Genetic [68, 87, 88, 75, 73, 69, 89, 115, 424, 77, 2024, 90, 110, 111, 72, 2030, 71, 112, 131, 78, 79, 80, 86, 91, 81, 2023, 93, 117, 94, 82, 83, 84, 85, 92, 1439, 1451]. Genetics [2024, 2023]. Gentle [1053]. Geo [924]. Geo-SAL [924]. GeoBench [1847]. Geodesic [555]. Geographic [929, 939, 917, 934, 940, 922]. Geographical [923, 935]. Geometric [728, 19, 737, 1907, 998, 452, 1349, 389, 1844, 1846, 916, 1378, 1847, 1840, 1311, 731, 1709, 643, 642]. Geometrical [1068]. Geometry [1895, 557, 261, 917, 52, 2232]. Gestalt [1558]. Getting [1759]. Gilbert [642]. Girths [1609]. GIS [941, 921]. Given [8, 742, 1084, 833]. Giving [42]. Glass [2026]. GLEAM [74]. Gleaned [119]. Glimpse [308]. Global [2026, 69, 341, 98]. Goal [1210]. Goal-Directed [1210]. Goals [706, 1686]. Golden [2148]. Good [2213, 427, 460, 630, 642]. Goppa [1611, 642]. Gottler [1167]. Gourmet [557]. Government [1274]. Graded [1458, 1630]. gradient [584]. Graduate [1277]. Graduate-Level [1277]. Grain [380, 339]. grained [1049]. Grammar [1517, 1503, 1513]. Grammars [1498, 1160, 1167, 1170, 1071, 1172, 1173, 1175, 1177, 1179, 1178, 1158, 1157, 1181, 1514, 1185, 1515, 1186, 1511, 1161, 1188, 1190, 1512, 1499, 1536, 1529, 1192, 1508, 1516, 1510, 1034, 1509, 1198, 1036, 1202, 1525, 1527, 1507, 1203, 1035, 1164, 1705, 1253]. Grammatical [1189]. Grandmaster [1307]. Granularity [213]. GRAPES [495]. Graph [432, 1160, 1165, 1895, 1167, 1170, 1172, 1173, 2047, 1115, 1175, 1159, 1179, 1205, 1178, 1158, 1180, 1157, 1184, 1185, 382, 1161, 1103, 739, 1191, 1192, 1088, 1885, 1195, 530, 344, 1198, 1946, 1200, 696, 563, 1063, 1201, 392, 1163, 1202, 1203, 1919, 750, 1204, 1098, 1164, 741, 700, 92, 789, 62, 1013]. graph-accessibility [789]. Graph-Based [2047]. Graph-Reducible [1200]. Graph-Theoretic [741]. Graphic [1174, 1187, 1249]. Graphical [219, 923, 158, 1183, 2036, 215, 1430]. Graphics [1866, 140, 2232]. Graphoids [1199]. Graphs [1166, 348, 1904, 545, 1625, 170, 466, 482, 1115, 1176, 528, 2064, 2163, 1948, 536, 529, 1078, 1135, 1188, 753, 1920, 169, 1886, 389, 148, 1225, 1381, 1855, 184, 717, 1891, 2065, 731, 1609, 394, 776, 910, 1431]. Greedy [147]. grey [584]. grey-level [584]. Grid [1903]. Grids [1893]. Gröbner [1353, 476, 475]. Groping [2022]. Group [1302, 1587, 1756, 32, 616, 1389, 629, 1711]. Groups [1304, 1926, 1059, 524, 628]. Grow [927]. GSBL [1232]. Guarantees [187]. Guard [745]. Guarded [360]. Guards [996]. Guidance [257]. Guide [618, 2177, 217]. guided [575].

Habits [1860]. Habituation [113]. Haeckel [129]. Haeckel-Strategies [129]. Hamiltonian [390, 1904, 536]. Hamming [1358]. Handle [1177, 921]. Handle-rewriting [1177]. handled [1468]. handles [57]. Handling [2102, 1647, 297, 1653, 812, 1672, 1418, 1265, 1529]. Hard [188, 436, 514, 1739, 2120, 532, 1269, 552, 1973, 1941]. Hardest [1076]. Hardware [2221, 2062, 1119, 22, 1124, 2181, 421]. Hash [1566, 1568, 1940, 30, 1602, 37, 1609, 1316]. Hashing [2196, 1889]. Having [1106]. Head [1538, 578]. Heap [354]. Heaps [959]. Heapsort [1911]. Hedge [200]. Help [253, 1866, 860]. helps [1443]. Herbrand [1438]. Heterogeneous [485, 1968, 192]. Heuristic [1032, 156, 1398, 1399, 395, 1883]. Heuristic-Based [395]. Heuristics [744]. hidden [1454]. Hiding [1319]. Hierarchic [1402]. Hierarchical [860, 406, 340, 1658, 40, 788, 911]. Hierarchies [821, 60, 619, 1422, 1698]. Hierarchy [951, 989, 1257, 760]. High [1708, 258, 1233, 1178, 356, 339, 1962, 244, 316, 1781, 1707, 50, 658, 54, 905, 61]. High-Level [1233, 356, 316, 1708, 1707, 905]. High-Performance [1962]. High-Speed [339]. Higher [723, 1105, 2135, 1039, 891, 376, 885, 1149, 1239, 2174, 708, 1577, 1244, 1507, 1035, 762]. Higher-Order [1039, 891, 376, 885, 1149, 708, 1244, 762]. Highly [75, 153]. Hill [1558]. Hints [119, 1844]. Historical [1912]. History [541]. Hitting [1903]. HK [429]. HMS [2185]. HOL [2173, 2181, 2175]. homogeneous [639, 566]. Homogenity [1353]. Homotopy [742]. Honest [1324]. Hood [907, 248]. Horn [687, 523, 675, 2096, 999, 1468, 710]. Hospital [442]. HP [1800, 1816]. HP-SL [1800]. HP-ST [1816]. Huffman [657, 413]. Human [1998, 1600, 1873]. Human-Computer [1873]. Hybrid [359, 1444, 1651, 2146, 616, 839, 1163, 1688, 662, 1685, 1691, 1435]. Hypercube [181, 172]. Hypercubes [348, 171, 1910]. Hyperedge [1156, 1189, 1193]. Hyperelliptic [1594, 640, 2212]. Hyperentailment [1676]. Hyperfinite [1151]. Hypergraph [1177, 1181]. Hypergraphs [1168, 825]. Hypermap [928]. Hypermedia [1879]. Hyperplane [714, 1864]. Hypersurface [467]. Hypertext [330, 65]. HYTECH [330].

IBM [1801, 1818, 449]. ICC [608]. ICsup [185]. ID [2214]. ID-Based [2214]. ID/LP [1527]. Ideal [1365, 1151, 1373]. Ideals [507, 1606, 477]. Ideas [465]. Identification [1307, 6, 47, 1935, 44, 80, 1600, 644]. Identify [1666]. Identifying [1896]. identities [634]. Identity [44, 3]. Identity-Based [44, 3]. Ignore [941]. II [2090, 2202, 1335, 478, 473, 1368]. III [1209, 113]. Image [848, 550, 1862, 835, 920, 2076, 103, 109, 183, 599, 596, 574, 1729, 594]. image-understanding [1729]. Images [1655, 1706, 595, 597, 589, 598, 573, 584, 571, 586, 577, 581, 565, 575]. Imaginary [2025]. imaging [592, 580, 572, 601, 564]. Immobilizing [733]. Immoral [1301]. Immortality [1970]. Immune [119, 2197, 115]. Impact [1757, 1758, 1746]. Imperative [1020]. Imperatively [1404]. Implement [288, 250]. Implementation [614, 1230, 295, 2080, 1208, 2221, 1172, 1026, 225, 883, 1603, 360, 2095, 2087, 226, 1501, 2007, 1237, 622, 2009, 2097, 875, 48, 1240, 22, 1124, 279, 102, 9, 300, 327, 4, 301, 1689, 914]. Implementational [617].

Implementations [1031, 415]. Implemented [147, 932]. Implementing [890, 1355, 89, 1682, 407, 696, 1847, 1428]. Implications [1279, 1286]. Implicit [1772, 518]. Implicitization [472]. Importance [2, 500, 919]. Impossibility [1327]. Impossible [1487]. Imprecise [1623, 842, 833, 1659, 1663]. Imprecision [1664]. Impredicative [1113]. Impressions [1806]. Improve [1136, 1004, 1439]. Improved [543, 1312, 1357, 1569, 1950, 107, 1064, 151, 1298, 572]. Improvements [112]. Improving [2084, 53, 1446, 1082, 532, 1380]. IMSE [220]. In-Place [1089, 1944, 754]. Inclusion [2171]. Incommensurability [2028]. Incomplete [1857, 357, 842, 1486, 1884]. incompletely [763]. Incompleteness [844]. Inconsistencies [1647, 1265]. inconsistency [1251]. Incorporation [572]. Increasing [172]. Incremental [888, 1505, 1025, 625, 1511, 1534, 601, 1035, 155, 1902, 1162]. Independence [802]. Independent [152, 734, 1885]. Index [1346, 2230, 424, 927, 936]. Indexing [928, 1957, 1717]. Indirect [346]. Indiscernibility [814]. individual [1436]. individual-flow [1436]. Induction [1230, 86, 1657, 2174, 1696]. Inductive [276, 274, 272, 273, 1482, 842, 275, 1236, 945, 1484, 1485, 1439, 1461, 1448]. Industrial [1757, 1276, 421, 1715]. Industry [1274, 1286, 941, 1290, 207]. Industry-Academia [1290]. Inequality [473]. Inequivalence [1155]. inf [549]. inf-Languages [549]. Inference [2066, 2088, 1636, 1548, 1547, 1474, 825, 272, 273, 1477, 1621, 943, 891, 1482, 1189, 1483, 1396, 1751, 710, 1480, 1484, 1485, 1726]. Inference-Based [1751]. Infinitary [281]. Infinite [1070, 1059, 522, 1940, 680, 778, 1936, 2147, 681, 649, 792]. Informatics [1859]. Information [2028, 831, 239, 1312, 1618, 923, 397, 208, 446, 810, 1595, 1102, 835, 1857, 939, 2059, 326, 230, 222, 801, 830, 1917, 940, 162, 2048, 805, 214, 1884, 2200, 919, 832, 332, 1533, 1065, 207, 1421, 1445, 570, 1432, 1710, 1415]. Information-Based [1065]. Information-Theoretic [2200]. Informational [401, 2014, 1883]. Informatization [335]. Inheritance [2052, 612, 1402, 821, 1650, 312, 1749, 157, 318, 1682, 1436, 1473]. Inhibitor [1001]. INLEN [1426]. Inline [604]. inner [652]. Innovation [1877]. Input [255, 1452]. Input/Output [255]. Insecure [1600]. Insensitive [2180, 1153]. Insights [1711]. instance [1450]. instance-based [1450]. Instances [1319]. Instantiation [712]. Institution [1242]. Institutions [827]. Instruction [1022, 1294]. Integer [2156, 1068, 1359, 473, 1366, 1337, 1340, 1098, 1338]. Integers [1589]. Integral [622]. Integrate [1799]. Integrated [498, 216, 939, 1805, 117, 1796, 1162, 1753, 921, 333, 1708, 911]. Integrating [1695, 154, 1733, 861, 1416, 260, 1560, 1036]. Integration [430, 1045, 498, 428, 501, 694, 224, 486, 489, 921, 1694]. Integrity [608, 297, 2055, 615, 1617]. Intelligence [309, 1740, 1268, 1871, 334, 1415]. Intelligent [1958, 400, 1702, 1677]. Intensional [320]. Interacting [73, 994]. Interaction [1993, 1873, 2211, 1712]. Interactions [127]. Interactive [2222, 6, 1572, 1322, 1586, 1220, 1161, 442, 1324, 2209, 1320, 1321, 574, 2077, 1413, 1687, 1608, 601, 2200, 2226, 1323]. Interconnection [195, 1909, 1591]. Interdistance [724]. Interface [1824, 1515, 331, 1878, 1423]. Interfaces [324, 1731, 922, 1677]. Interfacing [386]. Interference [463, 1104, 1101, 1388]. Interior [1347]. Interleaving [1067, 519]. Intermittent [396]. Internals [1517]. Interoperability [1055]. Interpolation [140, 431, 588]. Interpolation-Based [140].

Interpretation [887, 848, 1041, 1183, 821, 947, 1040, 1671, 1931, 944, 1727, 595, 1691, 596, 1704]. Interpretations [683]. Interpreters [2083]. Intersection [973, 975, 974, 1856, 761]. Intersymbol [463, 1388]. Interval [2118, 819, 2115, 727, 563, 1855, 187, 1433]. Interval-Based [819, 1433]. Intervals [1905, 141, 727, 1405, 2072, 1435]. Introducing [1713, 248, 1521]. Introduction [1498, 67, 1622, 1158, 1157, 2067, 1124]. Introspection [1486]. Introspective [1222]. Intuitionistic [1111, 1267]. Intuitive [74]. Invariance [1042]. Invariant [1001]. Invariants [365, 1110, 54, 583]. inverse [1442, 1700]. Inversion [2210]. Inverting [31]. Investigating [1442]. Investigation [281]. Invitations [317]. Invited [2192, 1037, 1016, 1296, 517, 557, 504, 551, 2215, 537, 531, 686, 689, 676, 678, 1311]. Invocation [1772]. Involving [1661, 1942]. IP [330]. IPSE [485]. IPSEN [1162]. IPSEN-Environment [1162]. IQ [1433]. Irish [1826]. Iron [196]. Irreducible [473, 414, 1380]. ISL [2115]. ISO [42, 1131]. ISO/IEC [42]. Isomorphism [466, 872]. Isotopy [715]. Issue [43]. Issues [614, 1278, 1282, 491, 1935, 1334, 1985, 1976, 1773, 2102, 1760, 1987, 1129, 1984]. ISTSPN [333]. Iterated [124, 430]. Iterating [1576]. Iterative [1613]. ITL [2070, 2093]. IV [1802, 1803].

Jacobi [1376]. Jacobian [1352]. JAM [2107]. Jersey [1016]. Join [291, 151]. Joint [655, 1370]. JTMS [1474]. Judgements [1641]. Jump [15]. Junction [1179]. Justifications [669]. Justified [1970, 1464].

Kanban [909]. Kantor [1355]. Karp [1152]. Keep [1487]. Kernel [201, 2154, 375, 2105, 1244]. Key [1592, 2207, 2227, 2, 2201, 1315, 1611, 1576, 1570, 2191, 34, 1608, 1327, 1337, 5, 9, 3, 28, 2214, 1332, 1314, 1331, 1338, 29]. Key-Updating [3]. Keynote [1269]. Keys [1607, 1327]. Keystream [15]. Khafre [2202]. KIDS [596]. Kind [1040]. Kinds [1548, 947, 317]. KIWIS [211]. KL [1724]. Kleene [522]. Klein [641]. Knapsack [728, 1568]. Knapsacks [2207, 36]. Knee [1894]. Knock [1894]. Knock-Knee [1894]. Knowledge [201, 808, 211, 1419, 1553, 2066, 1636, 1319, 1554, 848, 1548, 2080, 1552, 1539, 1, 1697, 41, 1644, 1561, 1535, 1514, 1222, 667, 1426, 2082, 1318, 1550, 1320, 1321, 290, 2068, 857, 626, 271, 2063, 1573, 1329, 1544, 2226, 1665, 840, 869, 1672, 2193, 2051, 231, 1545, 1677, 1295, 2065, 1719, 1713, 1420, 1699, 1693, 1457, 1458, 1694, 1443, 1687, 2208, 1430, 1721, 1434, 1264, 1435]. Knowledge-Based [848, 1672, 271, 1677]. Knowledge-bases [2051]. knowledge-systems [1721]. Known [755, 532, 2203, 29]. Known-Plaintext [29]. Knuth [1490, 687]. Kripke [1139, 1004]. Kripke-Models [1004]. KS [1520].

L [2091, 1183, 547]. L- [1183]. L.0 [1129]. Labeled [1874]. Labelled [1620, 1128, 1942, 1015]. Laboratory [1054]. Lack [854]. LaCoS [1806]. Lagarias [1082]. Lambda [699, 972, 525, 770, 771, 883, 794]. lambda-calculi [770]. lambda-Calculus [525, 771, 883]. lambda-terms [794]. LAN [1711]. land [660]. Landscape [2025, 128]. Language [1933, 2090, 322, 282, 607, 895, 440, 1539, 2060, 923, 441, 1232, 258, 236, 2171, 1523, 1524, 1029, 1629, 1827, 1737, 1184, 865, 1824, 1515, 356, 2094, 2095, 553, 968, 376, 2032, 899, 2059, 1320, 618, 1237, 934, 1397, 2079, 2097, 221, 1807, 605, 620, 1126, 1560, 233, 2036, 2134, 1544, 332, 2138, 1202, 924, 449, 846, 1229, 1247, 327, 334, 989, 374, 1443, 61, 1802]. Languages [1031, 1941, 706, 979, 284, 1209, 371, 1171, 2084, 1, 1076, 1117, 1180, 2179, 522, 1511, 604, 311, 1085, 516, 1190, 1516, 1510, 1483, 885, 1194, 694, 1043, 307, 271, 269, 955, 974, 873, 897, 1246, 879, 1101, 1204, 967, 1098, 549, 2061, 1696, 767, 768, 1685, 773, 780, 782]. Larch [1770, 1824]. Larch/C [1824]. Large [1830, 1565, 319, 1166, 920, 139, 357, 852, 1303, 7, 260, 1593, 2063, 1343, 1765, 1956, 1751, 1096, 1609, 921]. Large-Scale [1956]. Last [2205, 2206]. Lattice [1148, 1084, 427, 1064, 1264]. Lattice-Based [1148]. Lattices [817]. LAURE [2078]. Law [2020]. Layered [1283, 995, 2131]. Layers [1921, 1169]. Layout [1170, 1867, 1849]. Lazard [521]. Laziness [880]. Lazy [1046, 1026, 885, 875, 897, 1248, 771]. LCL [1824]. Leader [755, 438]. Leakage [1618]. Learn [734, 1699]. Learned [99, 1293, 495]. Learning [1635, 2019, 625, 597, 1637, 74, 1440, 842, 395, 107, 114, 1066, 1412, 1443, 1698, 1447, 1448, 1259, 1450, 1452]. Lecture [1037, 1016, 517, 557, 504, 551, 1826, 537, 531]. Led [860]. left [590]. Lemma [1078, 948]. Lempel [1575]. Length [548, 742, 413, 657, 773]. Lengths [741]. Lenient [873]. Lenses [876]. Less [898]. Lessons [1285, 495]. Let [1880]. Letter [1085]. LEU [1535, 1530, 1564, 1563]. LEU/2 [1535, 1530, 1564, 1563]. leucocyte [592]. Level [258, 716, 1233, 1178, 388, 356, 1982, 1277, 1126, 1531, 2175, 316, 439, 1781, 1708, 1707, 894, 54, 905, 584, 61]. Level-Ancestors [716]. Levin [430]. Lexical [1535, 1725, 1533, 426, 1540]. Lexicology [1530]. lexicon [1687]. LFSR [648, 458]. Library [21, 256, 2092, 1811]. LIFE [1037, 310, 2071]. Like [962, 348, 1296, 1595, 2075, 38, 1802, 1101]. LILOG [1554, 1548, 1547, 1552, 1555, 1536, 1550, 1551, 1543, 1544, 1521]. LILOG-DB [1554]. limit [772]. Line [749, 1953, 1836, 1915, 930, 174, 1898, 344, 423, 455, 777, 561]. Line-Sweep [1953]. Linear [599, 687, 749, 1583, 2152, 54, 16, 1580, 690, 647, 1393, 560, 380, 559, 1144, 1916, 2186, 1350, 191, 1347, 1303, 147, 1918, 404, 1372, 15, 966, 1104, 281, 868, 149, 1095, 1855, 1395, 1389, 1097, 726, 1298, 709, 351, 817, 769, 1424, 778, 643]. Linear-Time [2152, 560, 726]. Linearity [880]. linearly [53]. Lines [1903, 715]. Linguistic [1528, 2020, 1551, 845]. Linguistically [1631]. Linguistics [1539]. Link [737, 1901, 1848, 736, 1709, 585]. Links [1650]. Liouvillean [2028]. Lisp [362]. LISPO2 [607]. List [727, 896, 1097, 735]. Lists [508, 143]. Literal [675]. Literate [1493]. Live [992, 1020, 906]. Livelocks [243]. Liveness [1141, 1148, 2180, 913]. LL [1071]. Local [469, 101, 1932, 1557, 127, 1101, 84, 834, 585, 1710]. localities [764]. Locality [928, 313, 2002]. Localization [1556]. Locally [1939]. Locating [740]. Location [1922, 1746]. Loci [1384]. Lof [964]. Log [1038, 2036, 1345]. Logarithm [1599, 1594]. Logarithmic [1073, 170, 718, 417]. Logarithms [6, 1304, 44, 1341, 1337, 1589, 1338]. Logging [296]. Logic [685, 1046, 1206, 822, 2088, 2080, 828, 1047, 1474, 1049, 2084, 2100, 1210, 2135, 314, 1039, 1118, 2145, 1175, 1643, 1475, 1646, 1627, 1180, 1022, 341, 1629, 956, 504, 1038, 2115, 257, 998, 2111, 1144, 2184, 2095, 2129, 1476, 1123, 2167, 505, 387, 971, 2081, 2073, 2086, 826, 694, 1149, 1043, 1050, 1658, 1509, 2079, 2097, 678, 2099, 2119, 605, 1126, 1659, 2174, 966, 1104, 1036, 281, 1021, 1044, 2035, 1665, 1513, 969, 851, 1546, 1669, 1670, 710, 1227, 2113, 970, 1048, 385, 1053, 1675, 2016, 2053, 1018, 203, 2190, 1719, 1455, 1456, 1459, 765]. logic [1414, 1461, 2231, 1462, 783, 1463, 1258, 1260, 793, 1470, 1471, 1267]. Logic-Based [1021]. Logical [201, 2074, 2052, 508, 446, 1176, 1159, 2151, 671, 821, 886, 1785, 1262, 1251, 1258]. Logically [512]. Logicians [978]. Logics [827, 1624, 981, 1630, 1219, 1651, 1932, 1479, 984, 2004, 1668, 537, 704, 1221, 1679, 1458, 1467, 1252]. LOGIDATA [322]. LOKI [2202]. Long [314]. Long-Term [314]. Look [1981]. Loop [1412]. Loops [1228]. Loose [1803]. Loosely [363, 700]. Loosely-Coupled [363]. Loosening [994]. LOQIS [328]. Loss [835]. Lotos [1152, 2157]. Low [390, 2026, 1300, 1569, 1708, 660]. Low-Density [1569]. Low-Dimensional [2026]. low-level [1708]. low-rate [660]. Lower [138, 146, 16, 1911, 1080, 1225, 1062, 25, 789]. Lowest [178]. LP [961]. LPF [1785]. Lsub [1543, 1544]. Lucifer [2202]. LZW [413].

M [895]. M-Structures [895]. MAA [1798]. Machine [697, 2088, 132, 2080, 2185, 2167, 342, 336, 2169, 161, 700, 2107]. Machines [286, 1121, 898, 1081, 101, 2189, 841, 1151, 978, 1143, 337, 798, 1254]. Machining [261]. Made [1583]. Magic [1040, 2043]. magnetic [581]. Magnitude [1481]. Maintaining [529, 1899, 788, 750]. Maintenance [1272, 669, 668, 223, 672, 1715, 1720, 1462]. Majority [1301]. Making [853, 1251, 898, 357, 852, 730, 1884]. mammograms [582]. Management [211, 849, 2114, 491, 1283, 923, 854, 1730, 287, 622, 934, 940, 867, 626, 403, 1959, 846, 232, 190, 1795]. Managing [354, 1535, 245, 1671]. Manifolds [1851]. Manipulating [938]. Manipulation [1880]. Manipulator [2160]. Manipulators [76]. Manufacturing [1849, 1432]. Many [685, 1322, 1323]. Many-Prover [1322, 1323]. Many-Valued [685]. Map [1614, 1576, 867, 1169]. Mapping [195, 463, 123, 93, 349]. Mapping-Problem [93]. Mappings [474]. Maps [226, 1714]. Marker [1183, 575]. markings [51]. Markov [69, 2020, 1567, 1660]. Marriage [314]. Marriages [561]. Martin [964]. Martin-Lof [964]. Mas [1019]. Masking [2127]. Masks [1736]. Mass [1632, 135]. Massive [287]. Massively [371, 342, 135, 415]. Match [1234]. Matching [1172, 177, 561, 562, 1490, 770, 779, 782]. Matchings [1924]. Mathematical [1208, 397, 1831, 1457]. Matrices [543, 1559, 471, 1352]. Matrix [1305, 429, 427, 1372]. Maturity [502]. Max [826, 1258]. Maxima [1090]. Maximal [1072, 675, 1924, 530, 1405, 2072]. Maximal-Literal [675]. Maximum [723, 2183, 1578, 1885, 859, 1666, 584]. May [806]. Maze [837]. Mbits [663]. Mbits/s [663]. McDermott [1227]. McEliece [1611, 1570]. McNaughton [1936]. MD4 [2205, 46, 1317, 2206]. Mean [160, 581]. Meaning [1037, 540, 200]. Meanings [1532, 2015, 1441]. Means [816, 299, 98, 1853, 281]. Measure [853]. Measurement [1671, 240]. Measurements [855]. Measures [1865, 837]. Measuring [833]. Mechanical [1777, 1897, 2126]. Mechanically [2154, 2180]. Mechanics [2004]. Mechanism [812, 1794, 250]. Mechanisms [1772, 1529, 945]. Mechanized [2135]. Mechanizing [2174]. media [566, 1677]. Medical [1637, 858, 399, 570]. medicine [565]. Medium [1271, 64]. Membership [813, 40]. Membranes [236]. Memories [451]. Memory [725, 722, 1023, 1966, 383, 1142, 2, 287, 1217, 1981, 339, 1998, 1024, 18, 385, 2085, 2165, 788, 1418]. Mental [34]. Merge [1742]. Merging [350]. Meristems [1196]. Mesh [1908, 1949, 174]. Mesh-Connected [1949, 174]. Message [725, 42, 289, 46, 1317, 439]. Message-Passing [289, 439]. Meta [2083, 2103, 1802, 684, 1829, 1489, 1681, 490, 1802, 1803]. Meta- [1829]. Meta-Brave [2103]. Meta-CASE [490]. Meta-Interpreters [2083]. Meta-IV [1803]. Meta-IV-like [1802].

Meta-language [1802]. Meta-rule [684]. meta-theory [1681]. MetaEdit [215]. Metaknowledge [1548, 919]. Metalevel [1726]. Metamodels [1488]. Metamorphic [1771]. Metaphor [120]. Metaphors [121]. Metaproof [1322, 1323]. Method [1830, 457, 1741, 253, 179, 813, 1776, 1491, 1350, 1499, 1336, 501, 487, 1921, 928, 404, 212, 1014, 1395, 704, 2172, 1439, 577, 1261]. methodological [2231]. Methodologies [361]. Methodology [1829, 2175, 215, 911]. Methods [1500, 1742, 499, 1862, 154, 803, 1733, 929, 1136, 1831, 133, 847, 2032, 1506, 1087, 917, 156, 435, 2034, 620, 1063, 932, 1840, 1758, 1390, 593]. Metric [1745, 364]. Metric-Driven [1745]. Metrics [1747, 1748, 1450]. Microinstruction [421]. Microprocessors [1366]. Middle [390]. Might [806]. Miller [1152]. Million [1343]. MIMD [289]. Mimicking [2020]. Min [826, 1258]. Min-Max [826, 1258]. Mind [1860, 830]. Mind-Body [830]. Mini [2093]. Mini-Description [2093]. Minimal [176, 1138, 2145, 2163, 1940, 544, 1386, 475, 552, 1937, 52]. minimax [1417]. Minimization [1137, 139, 1464]. Minimum [2026, 1835, 635, 657, 2183, 412, 747, 835, 1901, 742, 721, 126, 748, 1869, 634]. Minimum-change [635]. Minors [1206, 544]. Minsumcut [720]. Mixed [1892]. Mixed-Searching [1892]. ML [1016, 441, 947, 889, 875, 1245, 871]. Mobile [984, 660]. Modal [2017, 822, 2152, 2145, 863, 1479, 984, 969, 1628, 1250, 1659]. Model [511, 2069, 1895, 322, 1023, 1067, 2026, 2157, 849, 1138, 2114, 591, 1776, 506, 1127, 397, 1118, 2152, 858, 234, 228, 2102, 237, 2166, 2168, 309, 359, 995, 185, 1217, 997, 920, 1144, 2184, 1123, 2167, 226, 1271, 101, 1632, 312, 882, 857, 1091, 1238, 221, 2177, 2127, 1634, 603, 958, 2033, 316, 2144, 1746, 1675, 519, 232, 157, 318, 337, 1699, 1695, 1694, 598, 1724, 1702, 582, 785, 1466, 1427, 1432].

Model-based [591, 1695, 1694]. Model-Checking [2152, 511]. Model-Oriented [2144, 1776]. Model-Theoretic [1238]. Modeling [189, 210, 1651, 198, 500, 1840, 603, 1723, 1693]. Modelization [2138]. Modelled [1203]. Modelling [236, 2118, 2129, 1401, 248, 815, 851, 215, 1674, 231, 122, 318, 242, 585]. Models [722, 432, 1742, 264, 980, 1122, 2047, 2013, 367, 983, 315, 503, 965, 1103, 1113, 951, 1004, 275, 2073, 2132, 678, 759, 1010, 1400, 946, 1226, 1273, 1013, 270, 2144, 167, 838, 590, 1713, 595, 59, 1256, 758, 62, 1454]. modification [1444]. Modifications [607, 798]. Modified [114]. Modular [2207, 1234, 1342, 1029, 510, 1315, 1631, 997, 1108, 605, 1140, 2217, 903]. Modularity [950]. Modulation [463, 462, 1392, 659, 653, 661, 660]. modulation-coding [660]. Module [1735, 1179]. Modules [1384, 2077]. Moduli [1590]. Modulo [44, 1593, 1583]. Mohan [48]. Mohan-Adiga [48]. Molecular [2021]. Monad [703]. Monadic [1206, 1180, 524, 1659, 1375]. Monien [1897]. Monoid [1092]. Monoids [521]. Monotone [1950, 944]. Monotonic [1482, 1227, 791, 1455, 822, 824, 1473, 1896]. Monte [1060, 565]. Moore [1227]. Morphisms [547, 1191, 703, 55]. Morphological [1531]. Morphology [1839]. Morris [1490]. Most [2181]. Motion [1557, 1850, 935, 590]. Motivations [118]. Motorola [21]. Movement [1538, 1203]. Moves [1944]. Moving [1843]. MPLomega [1785]. MR [580, 598, 578, 572]. MRI [579]. MTMM [867]. Mu [2152]. Mu-Calculus [2152]. Muller [631, 1369]. Multi [1966, 339, 852, 1320, 620, 434, 874, 167, 1693, 1690, 1417, 577]. Multi-Channel [167]. Multi-Criteria [852]. Multi-Language [1320]. Multi-Methods [620]. multi-modeling [1693]. multi-player [1417]. Multi-Protocol [1966]. multi-resolution [577]. Multi-Terminal [434]. multi-theory [1690]. Multi-Thread [339, 874]. Multicomputers [361, 439]. Multicriteria [850]. Multicycles [2119]. Multidimensional [1376, 1853, 2049]. Multigroups [40]. Multilanguage [1055]. Multilective [350]. Multilevel [1423, 1392, 1725]. Multimedia [2000, 1999]. multimodality [576]. Multiparadigm [324]. Multiparty [2225, 2201]. Multiple [827, 828, 558, 1219, 1368, 462, 619, 806, 591, 1696, 1467]. Multiple-Valued [827, 828, 1219, 1467]. Multiplication [479, 1342, 1079, 1366, 2217]. Multiplicative [643]. Multiplier [481]. Multiprocessing [1965]. Multiprocessor [383, 380, 347, 1149, 244, 2009, 1399, 183]. Multiprocessors [2007, 153, 2085]. Multipurpose [2006]. Multisets [754]. multistrategy [1452]. Multithreading [873]. Multitype [838]. Multivariate [1377]. Munin [1966]. Mural [1790, 1810]. MUSE [2108]. Musicians [1187]. Mutation [71, 445]. Mutual [1887]. Mutually [24].

n [1944]. Narrowing [1025, 1026, 1051, 2087, 1248]. Narrowing-Based [1051]. Natural [1539, 401, 1523, 1515, 1516, 1560, 121, 846, 1395, 1443, 583]. Nature [130]. Navigating [1898]. Navigation [120, 123, 1710]. Navigational [2032]. NC [170, 536]. NCE [1180]. nd [1358, 1386]. Nearest [713, 1914]. Neat [1760]. Necessary [1246, 806]. Needs [235]. Negation [822, 1402, 1684, 1267, 1456, 756, 1470]. Negative [540]. Neighborhood [731]. Neighbors [723, 1914]. Neighbourhood [412]. NERV [2011]. Nervous [85]. Nested [303]. Net [2157, 110, 2125, 900, 55, 905, 59, 912, 65]. Nets [378, 1120, 1231, 1107, 901, 243, 1233, 992, 112, 1001, 1128, 434, 950, 368, 333, 1781, 49, 50, 51, 903, 52, 53, 54, 904, 905, 56, 57, 906, 908, 775, 60, 910, 61, 911, 915, 64, 913, 914, 907]. Network [409, 743, 195, 401, 197, 858, 105, 824, 862, 202, 444, 341, 855, 245, 2011, 357, 857, 1660, 353, 279, 168, 108, 109, 1689, 1718]. Network-Based [357]. Networks [808, 1069, 1858, 176, 119, 196, 854, 1361, 2012, 1621, 111, 1841, 852, 1918, 809, 2075, 1591, 1875, 114, 1116, 1143, 199, 1430]. Neural [104, 1858, 196, 197, 858, 105, 236, 2012, 202, 444, 855, 1621, 110, 2011, 111, 357, 1410, 857, 108, 114, 338, 109, 1706, 1724, 1718]. Neurobiological [446]. Neuron [859]. neurosurgery [576]. News [202]. Newton [634]. Next [1991, 1989, 235, 1325]. NFA [552]. NFAs [1937]. Nineties [1999]. NLC [1157]. Node [1157]. Noise [1388]. Noisy [45]. Non [590, 822, 2088, 895, 1354, 1592, 1583, 208, 824, 1322, 1606, 899, 1321, 2002, 1608, 2050, 2200, 2226, 1323, 1227, 1101, 874, 791, 1455, 778, 566, 1468, 643]. Non-Canonical [2050]. Non-Classical [2088]. Non-Commutative [1606]. Non-Functional [208]. non-homogeneous [566]. non-Horn [1468]. Non-Interactive [1322, 1321, 1608, 2200, 2226, 1323]. Non-Interference [1101]. Non-Linear [1583, 778]. Non-locality [2002]. Non-Monotonic [1227, 822, 824, 791, 1455]. Non-rigid [590]. Non-Strict [899, 874, 895]. Non-Supersingular [1354]. Nonclassical [2073]. Nondegenerate [1596]. Nondeterminism [2037, 1094]. Nondeterministic [1062, 1254]. Noninterleaving [1067]. Nonlinear [1080, 1339, 1597]. Nonmonotonic [1621, 1476, 1478, 1482, 1464, 1479, 1676, 203, 1265]. Nonrecursive [2031]. Nonregular [516]. Nonterminals [553]. Nonterminating [682]. Nontrivial [1225]. Nor [1313]. Norm [1084]. Normal [1475, 1456, 778, 1431]. Normalization [692, 964]. Notation [1776, 1492]. Notations [538]. Note [410, 887, 1156, 174, 1483, 1201, 1472, 628]. Notes [1643, 1079, 442, 1974, 1826, 1678]. Notion [823, 1237]. Notions [1777, 1478, 500]. noun [1441]. Novel [339, 418]. Novelty [1465]. NP [1941, 1, 77, 2208, 1225]. NP-Complete [77]. NP-hard [1941]. NP-Problems [1225]. NU [1017]. NU-Prolog [1017]. Nuclear [464, 565]. Number [1070, 1058, 44, 1579, 553, 1080, 470, 476, 1096]. Numbers [819, 28, 1395, 818]. Numerical [428, 1060, 1390, 1697, 1695].

O [1944]. Object [613, 979, 2069, 607, 2060, 2052, 1955, 329, 1558, 2046, 314, 623, 1752, 525, 612, 154, 234, 608, 1866, 309, 1958, 315, 1235, 920, 1511, 968, 1488, 896, 604, 610, 2055, 1842, 1559, 1508, 931, 618, 934, 2034, 1804, 616, 615, 605, 1871, 233, 1749, 425, 2040, 2049, 1995, 1996, 603, 250, 2033, 328, 2044, 2042, 967, 1880, 232, 157, 624, 1422]. Object-Based [425]. Object-Orientation [1508]. Object-Oriented [979, 2069, 607, 2060, 329, 2046, 623, 1752, 154, 1866, 309, 315, 1235, 920, 1511, 968, 1488, 604, 2055, 1559, 618, 2034, 1804, 616, 605, 1871, 233, 2040, 2049, 1995, 1996, 250, 2033, 328, 2044, 967, 232, 157, 624]. Objects [1558, 854, 288, 312, 313, 1769, 945, 1852, 1203, 2053, 619, 591]. Oblivious [2220, 1574, 4]. OBSCURE [1237]. Observation [993, 1002]. Observations [2142]. Observing [764, 2121]. Odd [412]. Odlyzko [1082]. of- [193]. Off [174]. Off-Line [174]. Office [1834, 425]. Offline [747]. OISEAU [230]. OL [1169]. OL-Systems [1169]. Old [557]. omega [789, 195]. omega-branching [789]. On-Line [1836, 1915, 1898, 423, 455, 561, 777]. On-the-fly [2165]. Once [795]. One [1302, 1085, 1917, 2215, 1602, 806, 1316, 2208, 1469]. One-Letter [1085]. one-round [2208]. One-Way [1302, 1917, 2215, 1316]. Online [1899, 878]. onto [697, 349]. Ontology [1551]. OO [234]. OO-Binary [234]. OODBMS [614]. OODBs [2031]. OOZE [613]. Open [317, 488, 1968, 1065, 1485]. Openness [1384]. Operand [2217]. Operating [1979, 1978, 1991, 1993, 1961, 1964, 1962, 1974, 1970, 1977, 1996, 305, 1994, 1965]. Operation [718, 1234, 194]. Operational [1167, 1779, 1209, 2074, 1047, 375, 2142, 1931, 1048]. Operations [933, 285, 1477, 1852, 242, 1416]. Operator [990, 82]. Operator-Oriented [82]. Operators [117, 1480, 583, 1263, 1451]. Optical [358, 109]. Optimal [1913, 1835, 164, 1910, 1900, 1686, 1022, 424, 751, 71, 1916, 910, 1135, 1083, 1921, 1089, 178, 461, 732, 736, 182, 353, 434, 1095, 1485]. Optimisation [79, 826, 1258]. Optimization [625, 129, 1174, 95, 2032, 101, 96, 2031, 2023, 868, 98, 2021, 1387, 94, 2104, 1906, 2042, 122, 1419, 1697]. Optimizing [100, 604, 834, 130]. Optimum [745, 147]. Option [1086]. Optionality [1557]. OPTRAN [1518, 1519]. OR-Parallel [386, 385, 2114, 2104]. OR-Parallelism [384, 2110]. OR-Parallellism [1023]. orbits [567]. Order [1206, 686, 2135, 1039, 550, 1180, 891, 376, 1578, 1939, 1538, 885, 1149, 2174, 708, 870, 1577, 2177, 1133, 1044, 1244, 27, 703, 1507, 1906, 970, 1035, 705, 709, 762, 1722, 1420, 50, 770, 782, 1260, 795]. Order-Sorted [870, 703, 705, 782, 795]. Ordered [1038, 1653, 859]. Ordering [160, 563, 818]. orderings [757]. Orders [1136, 2178]. Ordinal [683, 538]. Organisation [235]. Organisms [105, 445]. Organization [2014, 339, 2018, 423]. Organizational [230]. Organizing [143, 1714]. Orientation [314, 1508]. Oriented [613, 979, 2069, 607, 2060, 1955, 329, 2046, 623, 918, 2057, 1752, 612, 154, 234, 608, 1866, 309, 315, 1235, 920, 1511, 968, 1488, 604, 32, 2055, 1501, 1559, 618, 2034, 1804, 616, 605, 1871, 233, 1749, 2040, 1163, 2049, 1995, 1996, 250, 2033, 328, 82, 2144, 2044, 967, 232, 157, 624, 1776, 1249]. Orphans [1369]. Orthogonal [1351, 460]. Other [317, 494, 648]. Our [1768]. Outline [884]. Outlines [1530]. Output [1579, 652]. Overall [1205]. Overhead [1300]. Overheads [2081]. Overlap [727]. Overlay [937]. Overloading [870, 871]. Overview [310, 1800, 286, 1791, 1627, 1056, 1518, 2182].

P [52, 53, 546, 1943]. P-complete [546]. P-semiflows [52]. P/T [52, 53]. Package [1736]. Packages [255]. Packet [1545]. Packets [1545]. Packing [91]. Pair [1243, 1922]. Pairing [525]. Pairs [180]. Pam [2160]. Pancake [168]. Pandora [2107]. Panel [1278, 1160, 1755, 941, 2089]. paper [686, 689, 676, 678]. PAPETRI [1120]. paraconsistent [1265]. Paradigm [1859, 498, 86, 307, 443, 1531, 988]. Paradigms [2089, 365]. Paradise [308]. Paradox [1314]. Parallel [410, 722, 366, 355, 979, 284, 895, 543, 179, 371, 286, 164, 354, 177, 1924, 89, 388, 1948, 2101, 185, 2024, 297, 534, 173, 289, 896, 342, 387, 1909, 362, 1923, 1000, 336, 1506, 535, 407, 169, 1086, 101, 91, 480, 1024, 290, 1918, 178, 97, 358, 379, 162, 875, 123, 2001, 2105, 2023, 182, 121, 168, 868, 1063, 102, 408, 873, 103, 135, 300, 1014, 302, 153, 531, 420, 447, 415, 381, 292, 306, 337, 700, 242, 92, 1460, 386, 385, 2114, 2111, 2104]. Parallel-Choice [1086]. Parallelism [293, 2221, 90, 2113, 384, 2110]. Parallelizing [2085]. Parallellism [2074, 1023]. Parameter [420, 415]. Parameterisation [1244]. Parameters [649, 1349]. Parametric [740, 472]. Parametricity [965, 884]. Paramodulation [687]. Parentheses [177].

Parity [1339]. Parlog [2074, 1030]. PARSAC [1366, 480]. PARSAC-2 [1366, 480]. Parser [1528]. Parsing [1071, 1393, 1705, 1398, 1525, 1527, 1504, 1703]. Part [2074, 1215, 467, 619]. Part-Whole [619]. Partial [762, 1647, 1136, 2178, 1109, 1191, 877, 156, 2177, 1133, 1388, 878, 1797, 1425, 794]. Partial-Order [2177, 1133]. Partially [521, 1653]. Participants [2193]. Partition [1078, 1946]. Partitioning [746, 1895, 78, 162, 433, 92]. Partitions [816, 2020, 638]. Parts [2181]. Party [2194, 24, 1612]. Pascal [962]. Passing [725, 289, 439]. Past [1730, 307, 1969, 1245, 1879]. Path [2047, 1952, 467, 136, 527, 730, 123, 1126, 180, 1892, 775, 1467]. pathological [598]. pathology [1417]. Paths [1070, 150, 528, 1901, 742, 1920, 1856, 741, 584]. Pathwidth [545]. patient [574]. patient-image [574]. Pattern [782, 1483, 562, 1490, 1248, 770, 1444]. Patterns [1482]. Payment [2219]. PC [449]. PCF [506]. PCPL [2092]. PCTE [489]. [508]. ATMS [666]. Autograph [1125]. Bags [508]. C [1824]. DB [306]. Event [2144]. Formal [1775]. frame-based [1691]. IEC [42]. lambda [639]. LP [1527]. or [1311]. Or-parallel [2111]. Output [255]. T [52, 53]. Percept [1434]. Perfect [1312, 1940, 1597]. Performance [2108, 1837, 160, 2110, 1865, 399, 2000, 1740, 298, 931, 1962, 244, 291, 443, 411, 408, 108, 932, 187, 1439]. Periodic [347, 1575, 454]. Periodically [1580]. Permutation [1926, 174, 1885, 2216]. Permutations [195, 1075, 534, 417, 1596, 12, 646]. Permuted [17]. Perplexed [2177]. Persistence [1649, 1464]. Persistent [2054, 607, 896, 752, 2048]. Personal [1757, 1872]. perspectives [1698]. Pertinence [1650]. PET [572]. Petri [2157, 1120, 901, 49, 50, 51, 243, 56, 992, 59, 775, 60, 61, 1128, 911, 915, 368, 65, 914, 333, 1781]. Petri-net-based [65]. Phase [436]. Phenomena [1556]. phenomenon [1257]. Phoneme [398]. photon [567]. Phrase [1537]. Physical [198]. Physically [1183]. Physically-Based [1183]. Physics [2021, 1269]. pi [967]. pi-Calculus [967]. Pictorial [938]. Picture [1171]. Piecewise [404, 431, 571]. Pilot [256]. Pipelined [175]. Place [1089, 1944, 754]. Placed [367]. Placement [1170, 1382]. places [913]. Plaintext [2203, 29]. Plan [79, 2033]. Planar [528, 719, 1948, 529, 1078, 1899, 739, 1920, 1844]. Plane [1216, 751, 1841]. Planes [1836]. Planned [1877]. Planner [860, 862]. Planning [1406, 861, 863, 865, 2058, 224, 123, 850, 1407, 1850, 199, 1469, 1433]. Plans [864, 868, 1404, 869, 1446]. Plant [1196, 1169]. plate [589]. Platform [293]. Plausible [828, 1728]. player [1417]. PLEXUS [1164]. Plurality [1543]. Plus [694]. Pocket [261]. Point [1834, 862, 702, 727, 1899, 1347, 1551, 1922, 254]. Point-Counterpoint [254]. Point-Set [1899]. pointer [798]. Points [1841, 1903, 732, 1435]. Poker [34]. Policy [221, 1959]. Polish [1492]. Polka [1030]. polycontextural [1262]. Polygon [749, 1952, 142, 1903, 1848, 736]. Polygonal [1852]. Polygons [745, 555, 1901, 732, 1382]. Polyhedra [1842, 1876]. Polyhedron [150]. Polymorphic [887, 943, 604]. Polymorphism [888, 965, 708, 1114]. Polynomial [962, 1068, 1896, 1367, 1483, 470, 1194, 1386, 475, 1373]. Polynomial-Time [1367, 1483]. Polynomials [543, 1363, 473, 1365, 1380]. Polyphase [460]. Polytope [733]. Polytopes [1839]. POMP [342]. POMSET [2179]. Pomsets [993]. POOL [295, 301]. POOL-X [295, 301]. POOMA [304, 305]. Poor [1934]. Population [73, 2024, 90, 2023]. Populations [97]. Port [1890]. Portable [295]. Posit [1774].

Posit-and-Prove [1774]. Position

[1836, 1558, 1557]. Positive [1896, 1222]. Possibilistic [1627, 825, 669, 1258, 826, 832]. Possibility [1648, 1628, 1250]. Post [1834, 927]. Postcondition [1793]. Potential [42, 927]. Power [468, 513, 558, 1181, 152, 518, 955, 1573, 798]. Powerful [416]. PPP [216]. Practical [2220, 278, 1588, 2227, 137, 169, 1337, 1064, 1338]. Practice [1757, 1869, 1799]. Pragmatic [837, 1413]. PRAM [1923]. PRAMs [356, 1947]. Pratt [1490]. Pre [1793, 1973]. Pre-and [1793]. Pre-Scheduling [1973]. Precautions [42]. Precedence [535]. Precedences [1031]. Precision [854]. Preclassification [1838]. Precondition [2159]. Predicate [1039, 1130, 783]. Predicates [1659, 1112, 944, 1722]. Predicting [196, 444, 1746]. prediction [1257]. Predictions [1856]. Preemption [2120, 2121]. Preface [1822, 1833]. Preference [815]. Preferential [1675]. preferred [1256]. Prefix [414]. Preliminary [225, 1643, 1650, 570, 1323]. Premises [540]. Preorder [949, 63]. Preorders [2164, 1103]. Preprocessing [562, 109, 1571]. Presburger [1211]. Presence [1301, 710, 1388, 709, 659]. Present [1245]. Presentation [1809, 1837, 1755, 2070]. Presenting [524]. preservation [913]. Preserve [78]. Preserved [692]. Preserving [1004, 431, 541, 49]. PRESS [1656]. Presupposition [1540]. Primality [10, 1610]. Prime [1341]. Primes [7, 1593, 1343]. Primitive [1380, 1389]. Primitives [1617, 323, 483]. Principal [972, 1946, 327]. Principle [2100, 126]. Principles [1613, 1426]. prior [570]. Priorities [2161, 1195]. Prioritized [2141, 1258]. priors [573]. PRISMA [294, 293, 306, 296]. PRISMA/DB [306]. Prisoner [124]. Private [1327]. PRIZ [268]. Probabilistic [755, 511, 2176, 1643, 283, 1649, 1085, 516, 470, 1656, 1660, 807, 1661, 1616, 1666, 1670, 1673, 1943, 1427]. Probabilities [1623, 803]. Probability [1626, 71, 515, 1091, 628]. probe [564]. Problem [1834, 119, 1307, 1105, 164, 1118, 720, 664, 1887, 560, 77, 405, 1857, 139, 136, 178, 830, 93, 121, 2010, 1133, 193, 180, 1594, 1922, 82, 846, 84, 1143, 735, 589, 776, 781, 784, 787, 797]. Problems [723, 728, 1904, 1928, 129, 87, 1211, 75, 2183, 546, 747, 1077, 865, 452, 95, 1923, 752, 1082, 739, 1061, 721, 1347, 1767, 826, 527, 730, 175, 1849, 744, 563, 1968, 1064, 1877, 1878, 1942, 2104, 552, 1065, 2008, 1098, 1485, 130, 631, 1258, 789, 1225]. Procedural [2089]. Procedure [1152, 1219, 173, 666, 524, 421, 190]. Procedures [689, 1047, 854]. Process [542, 1933, 2140, 987, 2170, 1282, 981, 1117, 377, 1408, 996, 503, 245, 1740, 1763, 502, 1002, 2186, 2141, 1005, 1764, 222, 1767, 248, 2160, 224, 1154, 1430, 2174, 2182, 1013, 767, 912]. Processes [949, 980, 1049, 2176, 2188, 1861, 2123, 367, 994, 259, 2187, 2148, 1769, 984, 363, 1008, 1011, 443, 1012, 122, 1146, 2139, 2172]. Processing [2080, 1539, 854, 446, 202, 89, 2092, 388, 1555, 930, 298, 931, 1868, 2086, 290, 159, 212, 2076, 1531, 846, 1097, 447, 183, 2061, 598, 1471, 1266]. Processor [341, 1949, 174, 358, 344, 563, 103, 343]. Processors [372, 1080, 535, 2116]. produced [1439]. Product [1145]. Production [2013, 1756, 388, 79, 135, 909]. Productivity [1758]. Professional [1277, 207]. Profile [15]. profiles [1421]. PROFIT [1036]. Program [1496, 1927, 1492, 2135, 1150, 526, 510, 1042, 975, 963, 515, 971, 1275, 1489, 1151, 1493, 1513, 411, 1929, 369, 1746, 606, 1018]. Programmable [341]. Programmed [1197, 1163]. Programming [1031, 2090, 1905, 319, 282, 2156, 1166, 284, 2088, 1068, 1496, 508, 1955, 1474, 2084, 314, 1039, 1359, 1175, 154, 1022, 1866, 956, 960, 133, 968, 173, 505, 311, 387, 1510, 86, 2086, 961, 694, 1050, 379, 616, 1807, 876, 2099, 268, 955, 620, 175, 1036, 1771, 1749, 1847, 1751, 710, 323, 328, 2113, 1048, 385, 1053, 967, 203, 327, 85, 374, 1455, 1461, 2231, 1451, 1718]. Programming-by-Example [960]. Programs [954, 962, 687, 1742, 1047, 1210, 1475, 526, 504, 280, 2101, 2115, 2111, 1033, 1476, 1934, 362, 1109, 336, 2081, 1034, 2073, 162, 1509, 679, 966, 985, 959, 1044, 2035, 1062, 2043, 1494, 1014, 874, 1883, 2041, 204, 1485, 1456, 1459, 650, 1425, 789, 1470, 1471, 1266]. Progress [442, 1163, 1202]. PROGRESS-Editor [1163]. progressing [790]. Project [1270, 329, 1283, 1272, 956, 245, 1801, 330, 1271, 1846, 332, 1162, 1273, 1753, 338, 448, 334]. Projection [882, 1942]. projections [602]. Projective [478, 2232]. Projects [251]. Proliferation [1670]. PROLOG [1212, 2092, 954, 1028, 1209, 1208, 1045, 1027, 2112, 2062, 1029, 375, 694, 2098, 2105, 2063, 1017, 386, 2104, 2085]. Proof [1789, 1319, 688, 1107, 1148, 1111, 1586, 1215, 280, 1216, 1219, 1220, 1087, 677, 1320, 1936, 884, 2169, 2162, 2174, 2226, 945, 203, 629, 1249, 1252]. Proof-Theoretic [203]. Proofs [2154, 1, 1572, 1106, 1324, 2209, 1321, 1573, 1329]. Propagating [804]. Propagation [1625, 2057, 404, 106, 1577, 14, 107]. Proper [1892, 1441]. Proper-Path-Width [1892]. Properties [964, 171, 1141, 1558, 1121, 1159, 1142, 2151, 1077, 929, 2185, 2178, 2180, 965, 2187, 2213, 11, 461, 18, 437, 184, 1316, 901, 905]. Property [1402, 538, 1657, 460, 1436]. property-flow [1436]. Proposal [35, 1689]. Proposed [1568]. Propositional [2017, 1662, 1665, 1411, 1442, 1260, 1470]. Propositions [1639]. pros [1466]. PROSPECTRA [1808, 1829]. Prosthesis [1998]. Protean [1131]. Protected [2157, 1308]. Protecting [4]. Protection [43, 1372, 1984, 650, 636]. Protocol [1966, 24, 2131, 1131, 1149, 437, 9, 2126, 2125, 59, 2208]. Protocols [2225, 1231, 2194, 1299, 2, 1599, 422, 34, 1128, 1951, 279, 134, 2193, 420]. PROTOS [2091]. PROTOS-L [2091]. Prototyping [1780, 1813, 1740, 1633, 1513]. Provable [1324]. Provably [526, 435, 33, 2219, 37, 3]. Provably-Secure [33, 37]. Prove [1774, 2173]. Proved [39]. Prover [1322, 1051, 1323, 2208]. Provers [1220, 2209, 1585]. Provide [1290]. Providing [2048, 2076]. Proving [954, 901, 960, 1897, 963, 10, 2173, 1249]. Proving-by-Example [960]. Pruning [1398]. Pseudo [31, 1353, 2215, 1582, 460]. Pseudo-Homogenity [1353]. Pseudo-Polyphase [460]. Pseudo-random [2215]. Pseudoprimes [1610]. Pseudorandom [646, 12, 1602, 2216]. PSF [2162]. Public [1592, 2207, 2227, 1611, 1607, 1570, 2191, 34, 1608, 1337, 2211, 5, 28, 1332, 1331, 1338]. Public-Key [2207, 1570, 2191, 1608, 1337, 5, 28]. Public-Randomness [5]. Publicly [1321]. Punctuated [88]. Purely [1936]. Purpose [2006, 2011, 289]. Pushdown [1094]. Pushouts [1167, 1158]. Putting [664]. Puzzle [82].

QPR [1388]. QPR-TCM [1388]. Quadratic [9]. Quadrature [1388]. Quadriphase [461]. Quadtree [918]. Qualitative [1641, 272, 802, 1481, 1412]. Quality [1758, 599]. Quantified [1214]. Quantifiers [1623, 2100]. Quantitative [1132, 1440, 593]. Quantum [2017, 23, 2220, 1299, 2004]. Quartic [641]. Quasi [176, 351]. Quasi-Linear [351]. Quasi-Minimal [176]. Queries [843, 158, 930, 2031, 1852, 1097, 303]. Query [2060, 923, 152, 2032, 2059, 326, 298, 931, 1868, 159, 2036, 924, 306, 1419]. Query-Independent [152]. Querying [2058, 2068, 1397]. Question [214]. Queue [2122]. Queues [1502]. Quickest [136, 180]. Quiescence [186]. Quotient [539].

R [1652]. Race [1617]. Radical [1365]. radio [660]. radiolabelled [592]. Radius [451, 1369, 632]. RAISE [1806, 1827, 1817]. RAM [1217]. Ramsey [1111, 1224]. Random [517, 1580, 2201, 534, 417, 1326, 1327, 1582, 1664, 148, 1343, 1385, 1889, 2215]. Randomization [2225]. Randomized [1908, 1032, 33, 1605]. Randomizers [1616]. Randomizing [1928]. Randomly [1618]. Randomness [1218, 1329, 5]. Range [721, 1466]. Ranges [1479]. Rank [429, 1092]. Ranking [156, 161]. Rate [1312, 107, 660]. Rational [980, 1734, 472, 1401, 28, 773]. Rationality [1475]. Raw [196]. Ray [555, 715]. Ray-shooting [715]. Rayleigh [661, 1392]. RDBMS [258]. Reachability [1135, 910]. Reactions [444]. Reactive [120, 861, 237, 1446]. Reactor [197]. Read [1947]. Reads [558]. Real [188, 436, 511, 189, 1288, 2140, 987, 1231, 2122, 2135, 2117, 2145, 258, 2183, 2118, 2123, 2136, 2185, 1737, 259, 2115, 2187, 1740, 372, 1925, 1005, 1739, 1972, 2132, 1974, 260, 1289, 461, 238, 2120, 2143, 2133, 2134, 9, 2137, 2144, 1973, 420, 519, 194, 247, 192, 2139, 337, 1699, 785]. Real-Time [188, 436, 189, 2122, 2135, 2117, 2145, 258, 2183, 2118, 2123, 2136, 2185, 1737, 259, 2115, 1740, 372, 1925, 1739, 1972, 2132, 1974, 260, 1289, 238, 2120, 2133, 2134, 2137, 2144, 1973, 194, 247, 192, 2139, 511, 420, 785]. Real-Timed [2143]. Real-Valued [461]. Realistic [1217]. Reality [2138]. Realizability [944]. Really [2089]. Reals [546]. Reaping [135]. Reason [840, 1787].

Reasoning [201, 685, 822, 836, 1790, 828, 1639, 824, 1642, 926, 925, 1788, 1132, 1648, 1650, 1476, 2096, 1632, 2086, 885, 1932, 1655, 1656, 1844, 807, 1661, 1663, 1405, 2072, 206, 1673, 2077, 1676, 1481, 203, 838, 791, 1455, 1678, 1728, 1692, 1690, 1720, 1251, 1464, 1466, 1431, 1265]. Recasting [1875]. Reciprocal [427]. Recognition [171, 110, 1842, 396, 591, 598, 1444]. Recognizability [1070]. Recognizable [768, 792]. Recognize [516, 1918, 1855]. Recognizing [1194, 404, 1204]. Recombination [445]. Recomposition [1054]. Reconfigurable [383, 252]. Reconfiguration [513]. Reconstruct [182]. Reconstruction [419, 569, 1088, 103, 726, 1706, 573, 568, 571, 570]. recording [1695]. Recovery [42, 2127, 302, 296]. Rectangles [1835, 91]. Rectangular [1922]. Rectilinear [435, 1848, 736, 741]. Recurrence [505, 352]. Recurrences [349]. Recursion [962, 1784, 1110]. Recursions [1583, 1424]. Recursive [1168, 158, 321, 526, 272, 273, 864, 1042, 1336, 323, 944, 343]. Red [144]. Red-Black [144]. REDOC [2202, 1335]. REDOC-II [2202]. Reduce [1135]. Reduced [2080, 1015, 66]. Reducibility [1218]. Reducible [1200]. Reducing [2081, 2173]. Reduction [1165, 1207, 525, 387, 1084, 1024, 344, 696, 1064, 976, 565]. Reductions [1928, 1004]. Redundancy [1173, 657]. Redundant [76, 77]. Reed [629, 631, 1369, 663]. reentrant [903]. Reevaluation [1519]. Reference [346]. Refineable [2143]. Refinement [542, 1933, 1150, 1000, 1929, 1796, 369, 1825, 370, 767]. refinements [49]. Reflection [617]. Reflective [616]. Regarding [2027]. Regions [1649]. Register [45, 1032, 1126]. registered [573]. Registration [577, 576, 574, 575]. Regression [404]. Regular [890, 1836, 466, 56, 1144, 2184, 553, 467, 2187, 639, 780].

Regularity [547, 1085]. regularization [568]. Regulated [841]. Regulation [849]. Reification [1792]. Relabelling [1195]. Related [723, 1553, 1334, 1773, 1351, 631]. Relating [1113, 1008]. Relation [1595, 1645, 1139, 1659, 1747]. Relational [843, 285, 1176, 1857, 2055, 226, 1197, 159, 212, 1252, 292, 817, 1428]. Relations [1474, 2057, 512, 2171, 925, 1149, 815, 1255, 1473]. Relationship [1626, 234, 1182, 143, 488, 316, 1674]. Relationships [134, 1429, 1442]. relative [1414]. Relativised [1938]. Relativity [829]. Release [1966]. Reliability [188, 1639, 2129, 811]. Reliable [2125, 781]. Remaindering [465]. Remark [39, 47]. Remarks [2008, 648]. Remote [190]. Remotely [921]. Removal [1400]. renal [593]. rendering [601]. Rendezvous [364]. Reorganisation [1097]. Repair [1408, 2219]. Repeated [1580]. Replaceable [1741]. Replaced [978]. Replacement [1156, 1178, 1189, 1193, 1204]. Report [1650, 1813, 1801, 1617, 449]. Reporting [164]. Repositories [1958, 488, 1956]. Repository [494]. Representation [808, 2066, 1539, 1693, 1457, 208, 105, 2146, 1559, 1845, 811, 1544, 1664, 1404, 1852, 416, 2051, 619, 1719, 1713, 1429, 1726, 1430, 1434, 1264, 1435]. Representational [926]. Representations [680, 959, 2015, 2016, 1709]. Representing [1558, 1542, 1629, 865, 1532, 738, 233]. Requirement [2133]. Requirements [499, 209, 208, 1759, 1760, 929, 210, 1761, 1762, 2124]. Research [1553, 1652, 1764, 1970, 1762, 134, 851, 941, 1956]. Researches [1757]. Reservation [2000]. Residue [16, 34, 649]. Resolution [205, 687, 2100, 2078, 1213, 514, 1467, 870, 270, 670, 1442, 1700, 577, 1472, 1047]. resonance [581]. Resource [1148, 2000, 2059, 214, 194]. Resources [2, 1215, 326, 193]. Respect [1008]. respectable [1251]. respecting [55]. Response [104, 1915, 1388]. Response-Time [104]. Response-Trellis [1388]. Rest [1549]. Restraining [1670]. Restraints [1218]. Restricted [687, 2100, 745, 719, 434]. Restriction [1541]. Restructuring [179, 335]. Resultants [1074]. Results [251, 520, 1205, 2179, 1801, 142, 1371, 1372, 1375, 1114, 1065, 118, 777, 2216]. RETE [1172]. Retriangulating [1900]. Retrieval [1707, 918, 161, 1421, 1445, 1437]. Retrieving [358]. Reusability [1490]. Reusable [611, 255]. Reuse [253, 1491, 621, 254]. Reusing [210]. Reverse [1492]. Reversibility [1574]. Revision [823, 669, 668, 673, 1662]. Revisited [829, 378, 1229, 155, 1914, 704, 1465]. Rewrite [697, 698, 684, 757, 679, 705]. REwriting [1163, 697, 692, 276, 674, 1115, 1157, 1184, 382, 2087, 683, 1191, 676, 524, 678, 695, 2173, 691, 1240, 1375, 1200, 696, 682, 700, 1177, 778]. Rewritings [681]. RIGAL [282]. Right [1759]. Rigid [2019, 590]. Ring [1606, 2191, 1365, 59, 1254]. Rings [755, 16, 1363, 649, 653]. RIPE [1617]. RISC [1964]. Risky [1640]. RLucid [2134]. RMS [206]. RNC [738]. RNS [418]. Roadmaps [1356]. Robot [120, 123, 1710]. Robotic [860]. Robots [421]. Robust [1217]. Robustly [1318]. Robustness [1840]. Role [854, 1863, 956, 2000, 927]. Roles [73, 1473, 1415]. Root [1910]. Rooted [1194]. Roots [1886]. ROSE [253]. ROSE-Ada [253]. round [2208]. Rounds [2205, 1220, 1573, 2206]. Routes [745, 1902]. Routing [361, 1916, 174, 434, 168, 1894, 1893]. RSA [1305, 8, 1590, 22]. RTL [2136, 1130, 2119]. RUEGEN [1520]. Rule [697, 1750, 2078, 1234, 1192, 222, 221, 2061, 1416, 1691, 684, 914]. Rule-Aided [2078].

Rule-Based

[1750, 1234, 1192, 2061, 1416, 914]. rule/ frame [1691]. rule/frame-based [1691]. Rules [124, 698, 1637, 1639, 311, 427, 1657, 615, 2173, 1021, 832, 1480, 1673, 705, 1439, 597, 1449]. Run [43, 17]. Run-Time [43]. Running [516].

s [713, 663, 1047, 1595, 393, 1336, 1597]. S-Box [1595]. S-Boxes [1597, 1336]. S-Distance [393]. S-SLD-Resolution [1047]. S5 [1659]. S5-Modal [1659]. SAC [480]. SAC-2 [480]. Safe [621]. Safety [509, 1141, 2178, 2180, 246, 2124]. Safety-Critical [2124]. SAL [924]. Salesman [87, 2010, 84]. Sample [275]. Sand [27]. Satellite [1655, 2076]. Satisfaction [1025, 2150, 1460]. Satisfiability [538, 2119]. Satisfying [1336, 1577]. Save [1015]. Scalable [153]. Scale [319, 584, 583, 1956]. Scale-space [583]. Scaling [1750, 2217]. scatter [566]. Scenario [1961, 207]. Scheduler [384, 2106]. Schedulers [386, 1146]. Scheduling [188, 436, 384, 2109, 181, 2122, 1022, 2000, 79, 535, 2081, 347, 2132, 162, 238, 1890, 563, 1973, 194, 192]. Schema [614, 607, 2039, 1422]. Schemata [1558]. Scheme [436, 39, 6, 44, 42, 729, 78, 2224, 298, 1370, 40, 38, 726, 1571, 645, 2216, 646]. Schemes [1312, 47, 2198, 1173, 1584, 1837, 1586, 972, 24, 2191, 313, 159, 3, 1311, 345, 644, 660, 652, 1434]. Schnorr [1571]. School [1826]. SCI [2155]. Science [952, 134, 450]. scopings [1704]. Scramblers [1601]. Scriptic [374]. Scruffy [1761]. SE [1292]. Search [1069, 1928, 918, 1075, 424, 740, 388, 77, 2103, 143, 427, 1398, 1399, 416, 84, 122, 1686]. Searching [1913, 737, 358, 1912, 1892]. Season [314]. Seat [664]. Second [1206, 1180, 709, 770]. Second-Order [1180, 709, 770]. Secrecy [1588, 4]. Secret [1615, 1312, 2198, 2199, 2201, 1576, 24, 2200, 1875, 1311]. Secure [2222, 26, 2229, 1310, 2227, 1308, 33, 2223, 1951, 2200, 2219, 37, 3, 1794]. Security [1300, 1985, 2, 1987, 1324, 1984, 1598, 1986, 1571]. Seen [1766]. Segment [749, 164, 930, 932]. Segmentation [835, 581, 578, 584, 571]. Selected [1849]. Selecting [743]. Selection [124, 197, 146, 1735, 424, 1590, 1949, 1950, 724, 1702]. Selectionist [116]. Selections [2042]. Selective [1042]. Self [2014, 1714, 1607, 143, 877, 1604, 2018, 1391, 1444]. Self-Applicable [877]. Self-Certified [1607]. Self-Decimation [1391]. self-modification [1444]. Self-Organization [2014, 2018]. Self-Organizing [143, 1714]. Self-Synchronizing [1604]. Selforganization [2029]. Semantic [219, 1727, 854, 824, 1511, 673, 826, 2015, 1437, 2172, 1419, 1255, 1687, 1703, 1704, 1473]. Semantics [1933, 1779, 1209, 2074, 1539, 509, 371, 1047, 2052, 1782, 2117, 1027, 1402, 905, 1475, 702, 504, 315, 375, 968, 372, 311, 999, 1001, 971, 1086, 269, 1108, 2142, 1400, 942, 953, 1133, 2035, 1048, 1675, 967, 672, 204, 364, 791, 1456, 1459, 774, 1725, 1427]. Semantics-Directed [1027]. Semi [689, 1356, 520, 448, 787, 796, 797]. Semi-Algebraic [1356]. Semi-Challenging [448]. Semi-Commutations-Decidability [520]. Semi-Decision [689]. semi-unification [787, 797]. Semialgebraic [1348, 467]. semiflows [52]. Semirings [730]. Sensed [921]. Sensitization [113]. sensor [1710]. sensor-based [1710]. sentences [1727]. Separations [1057]. Separators [1080]. SEPIA [2098]. Sequence [45, 1579, 1578, 460]. Sequences [1583, 19, 16, 1580, 1579, 680, 17, 2020, 461, 1582, 1575, 15, 458, 459, 454, 1391, 1169, 902, 648, 647, 602, 594]. Sequential [1049, 1127, 1121, 99, 2167, 1364, 2173, 1124, 180, 662, 577]. Serial [1795]. Series [468, 1918]. Series-Parallel [1918]. Server [187]. Services [1752]. SESADA [1744]. Set [1835, 2057, 1595, 321, 982, 24, 752, 1899, 1885, 1374, 1945, 1114, 1386, 475, 1097, 913]. Set-Oriented [2057]. Set-Theoretic [1114]. Sets [1941, 1834, 277, 831, 278, 508, 1356, 1211, 1348, 523, 1924, 1213, 1940, 467, 13, 1664, 460, 817, 592, 579, 780, 792, 66]. Sets/ Bags/ [508]. Setup [447]. Seven [1285]. Several [1777]. Sexual [445]. Shadow [1736]. Shafer [800, 1657]. Shaft [454]. Shallow [724, 709]. Shamir [40, 38]. Shamir-Like [38]. Shannon [654]. Shannon-theoretic [654]. Shape [1348, 588, 431]. Shape-based [588]. shapes [1706]. shaping [1399]. Shared [1966, 2228, 372, 24, 1024, 1311, 1787, 1014, 2085, 2116]. Shared-Memory [2085]. Shared-State [1787, 1014]. Shares [2198]. Sharing [725, 1312, 2198, 2199, 2200, 969, 2214, 1432]. Shear [1908]. Shell [1656]. Shells [231]. Shift [45]. Shooting [555, 715]. Short [1056, 1893]. Shortest [554, 749, 150, 528, 1952, 1841, 1902]. Should [496]. Shrinking [179]. Shuffle [348, 1888]. Shuffle-Exchange [1888]. Shuffle-Like [348]. Side [1761]. Sides [1766]. Siemens [1652]. Sign [1357]. Signal [125, 419]. Signals [404, 1388]. Signature [39, 1584, 1603, 1586, 42, 38, 161]. Signature-Based [161]. Signatures [1309, 2229, 1587, 1619, 1310, 41, 2228, 8, 1934, 417, 1585]. Signer [2229]. Significance [1340]. Silence [2148]. Silica [617]. Silicon [196]. SIMD [371, 242]. Similarities [131]. Simple [1933, 191, 142, 885, 178, 416, 1855]. Simplical [936]. Simplices [428]. Simplification [1045, 2055]. Simplifier [699]. Simulated [74, 100, 102, 103, 130]. Simulating [113]. Simulation [2171, 2011, 1149, 809, 299, 279, 122, 62]. Simulations [1135, 1003, 910, 1254]. simulator [50]. Simultaneous [866]. Single [1158, 881, 452, 467, 775, 12, 563, 167, 337, 567]. Single-Channel [167]. Single-path [775]. Single-Processor [563]. Single-Threading [881]. Singleton [975]. Singular [429]. situations [1680]. Sixth [1978]. Sixth-Generation [1978]. Size [755, 2198, 73, 1271, 349]. Skeptical [672]. Sketch [2083]. Skip [727]. SKOLEM [1494]. SLD [1047]. Sliding [82]. Small [342, 1096, 349]. Smallest [553, 1919, 1881]. Smalltalk [606]. Smart [957, 1603, 1332, 1331]. Smart-cards [957]. Snapshots [191]. Snefru [2202]. SNePS [1401]. Society [335]. Sociopathicity [1657]. Soft [1269]. Software [1830, 1741, 189, 1291, 1270, 1274, 1281, 1750, 1279, 1742, 499, 492, 1282, 491, 1283, 1272, 623, 498, 1744, 1752, 43, 1730, 1732, 1756, 1294, 259, 503, 245, 497, 1763, 502, 1757, 223, 1275, 493, 1764, 1767, 248, 252, 224, 1276, 1289, 1277, 1330, 500, 1287, 1805, 611, 1293, 1273, 1748, 1743, 254, 914, 1295, 1290, 187, 333, 624, 783, 911]. Solomon [629, 663]. Solution [1168, 191, 193, 784]. Solutions [175, 532, 1389, 817]. Solvable [452, 270]. Solve [178]. Solver [664]. Solvers [380]. Solving [706, 119, 1307, 87, 769, 99, 1393, 1350, 1303, 1394, 1930, 1845, 93, 121, 1064, 2104, 1395]. Some [465, 1678, 964, 1904, 843, 1312, 1355, 648, 1619, 1985, 1643, 177, 8, 1363, 1079, 1590, 777, 1923, 847, 1191, 275, 269, 30, 1372, 471, 1989, 1375, 1384, 368, 1968, 1990, 1013, 1066, 2008, 634, 631, 539, 1213, 999, 1466, 1252, 1225]. Sort [1908, 173, 703]. Sortal [1533]. Sorted [146, 1038, 870, 703, 705, 782, 795, 2145]. Sorting [410, 1913, 137, 1865, 1089, 1912, 1944, 754, 353, 168, 167]. Soul [1995]. Sounds [396]. Source [1409, 811, 1747, 656]. Sources [801, 1694]. Soviet [335]. Space [746, 2140, 987, 1073, 2005, 387, 1081, 2209, 1398, 1399, 715, 2004, 420, 1696, 587, 583, 66].

Space-Bounded [1081]. Spaces [125, 1772, 376, 937, 2147, 1203, 1448]. SPADE [217]. Spanning [723, 747, 753, 435, 748, 1869]. Sparse [1350, 1303]. Spatial [933, 926, 993, 929, 1555, 930, 917, 931, 1868, 928, 123, 1852, 924, 936, 921, 1719, 1716]. SpecBox [1810]. Special [1288, 2006, 744]. Specialization [1744, 1042]. Specialized [2084, 524]. specific [570]. Specification [1830, 1789, 262, 263, 264, 265, 266, 267, 1955, 2038, 1232, 2145, 158, 1179, 377, 2136, 1827, 2155, 2115, 1770, 1491, 1925, 1499, 422, 1019, 1803, 1237, 1807, 605, 1198, 2133, 233, 1798, 1771, 1021, 279, 437, 1778, 2127, 425, 1244, 1796, 1246, 1797, 1736, 783]. Specifications [1031, 1152, 2154, 540, 693, 2052, 1776, 1786, 1735, 1809, 1780, 1179, 1788, 2185, 1235, 1003, 1087, 677, 1236, 518, 1238, 221, 271, 2048, 2174, 1241, 1951, 2144, 1799, 1146, 2116, 762]. Specificity [820]. Specified [1910, 1372, 763]. Specify [1028]. Specifying [366, 1231, 1737, 257, 1190, 2130, 1959]. Speckenmeyer [1897]. SPECT [566]. Speculations [1989]. Speculative [2109]. Speech [1601, 1724]. Speed [2184, 339, 1008, 658, 1697, 1254]. Speed-Dependent [2184]. speed-up [1254]. Spelling [400]. Spherical [1367, 1378, 627]. Spin [2026]. Spin-Glass [2026]. Spirit [450]. Spline [568]. Spline-based [568]. splines [589]. split [772]. Splitting [25]. Spohn [799]. Spoofing [27]. SQL [303]. Square [1886]. ST [1816]. Stability [1954, 1222]. Stable [1100, 1980, 506, 1099, 1981, 1083, 561, 1944, 1427, 1400, 302]. Stacks [1502]. Staircase [1854]. Stamp [160, 1328]. Stamps [191]. Standard [1016, 478, 35, 477, 1798, 488]. Standardization [492]. Standing [76]. Star [184]. State [895, 1118, 1121, 981, 2163, 1137, 968, 2189, 2169, 1399, 1133, 2147, 1787, 1014, 1134, 1143, 63, 66]. State-based [981]. Statecharts [2146, 953]. Statement [1807]. Statements [1086]. States [553, 2056]. Static [1782, 1460, 893, 249, 382, 1511, 779]. Static-Semantic [1511]. Statically [620]. Statically-Typed [620]. Statistical [1297, 1326]. Statistics [517, 1906, 1695]. Status [1813, 1617]. Steiner [2030, 435]. Step [953, 1469]. Steps [1030]. Stepwise [1778, 2144]. Stereo [405, 1707]. still [1970]. Stochastic [773, 809, 433, 56, 582]. Stock [1366]. Stone [1099]. Storage [1980, 940, 1502]. Storages [1076]. Storing [150]. STOSS [809]. Straight [1903]. Strand [2101]. Strategic [214]. Strategies [2027, 698, 2109, 129, 286, 131, 96, 2029, 98]. Strategy [675, 1024, 97, 114, 2022, 130, 1721, 1433]. Stratification [1374]. Stream [1581, 1954, 1604, 18]. Street [1843]. Strict [686, 899, 1336, 874, 895, 786]. Strictness [507, 886]. String [524, 562, 1375, 779]. String-Rewriting [524, 1375]. Strong [761, 1057, 378, 692, 2229, 533, 1479, 1459]. Strongly [506, 877, 33, 318]. Strongly-Randomized [33]. Strongly-Typed [877]. Structural [1779, 375, 142, 1154, 531, 1316, 1266]. Structure [2170, 918, 446, 1542, 1940, 727, 78, 1747, 1937, 1537, 1724, 1688, 1454]. Structured [432, 1733, 1188, 312, 1775, 423, 1404]. Structures [219, 895, 2170, 1176, 1561, 90, 998, 752, 1139, 1090, 1197, 127, 945, 1942, 1545, 1712, 1707, 569, 593, 1874]. Structuring [231]. Stubborn [1134, 66]. Studies [1645, 593]. Study [251, 1208, 200, 713, 803, 220, 1490, 1493, 130, 763, 52, 1451]. Studying [447]. STUF [1523, 1524, 1526, 1525]. Sturm [1377]. Style [1787]. Styles [1281]. Subclasses [1214]. Subdividing [415]. Subdivisions [1901]. Subfunctions [1096]. Subgoals [869]. Subgraph [1098]. Sublinear [1080].

Subresultants [1074]. Subset [1821, 1569, 560, 1082, 1803, 1064]. Subset-Sum [560]. Subsets [1092]. Subsorts [708]. Substitution [459, 1891]. subsume [1680]. Subsumption [2069, 2065]. Subterm [538]. Subtree [163]. Subtyping [977, 976, 1749]. Subway [1735]. Succeeds [1084]. Successful [1070]. Such [1660]. Sufficient [693, 553]. Suffix [734]. Suggested [1614]. Sum [1569, 560, 1082, 391, 1064]. Summarization [845]. Summary [1960, 1967, 1971, 1975, 1983, 1988, 1997, 287, 37]. Summation [236]. sums [643]. Superposition [686]. Supersingular [1354, 1592]. Supervisory [1147]. Supply [740]. Support [1979, 1554, 1790, 266, 893, 210, 1763, 1764, 1982, 1767, 1962, 1765, 1162, 1995, 1996, 1751, 1753, 385, 194]. supported [909]. Supporting [1744, 312, 1873, 1852, 333]. Supracompact [1477]. Surface [714, 602, 585]. Surfaces [472]. Surgical [564]. Surjective [525]. Survey [262, 683, 1506, 1371, 1199, 1066, 1485, 49, 644]. Swedish [235]. Sweep [1953, 1836]. Switching [2204, 1797]. symbol [628]. Symbolic [1122, 2128, 2166, 2168, 2184, 2167, 1268, 1151, 1709, 1439, 1440]. symbols [1424]. Symmetric [1376, 1537]. Symmetries [1842]. Symmetry [1917, 389]. Symplectic [474]. Symptoms [711]. Synchronization [2126, 1973, 2008]. Synchronized [1081, 900]. Synchronizing [1049, 1604, 609]. Synchronous [2138]. Synchrony [994]. Syndrome [1298]. Synergetic [841]. Synergy [722]. Syntactic [881, 2041, 1701]. Syntactical [275]. Syntactics [204]. Syntax [791]. Synthesis [276, 274, 992, 971, 275, 684, 2182, 58, 906, 1266, 326]. Synthesizing [1425, 1146]. System [211, 1565, 2066, 2157, 1168, 1054, 2091, 262, 263, 264, 265, 266, 267, 848, 74, 540, 2060, 1122, 401, 197, 209, 208, 856, 1402, 120, 608, 228, 1780, 669, 1234, 115, 855, 380, 280, 2093, 1052, 2095, 2011, 1488, 896, 1808, 442, 80, 1517, 1550, 1739, 677, 1829, 1982, 1520, 1396, 1962, 1056, 252, 260, 1518, 1656, 809, 2098, 268, 605, 1560, 233, 403, 1660, 1036, 2219, 1805, 186, 2076, 839, 1957, 117, 935, 1996, 710, 300, 1986, 328, 302, 449, 2104, 2214, 118, 1097, 292, 305, 1994, 1564, 400, 921, 439, 2106, 1795, 1706, 1681, 977]. system [662, 1423, 1685, 1444, 1445, 1448, 61, 1462, 1716, 1426, 1711, 576, 1717, 62, 596, 1729, 564, 65, 1704]. System-Level [439]. Systematic [2194, 1491]. Systems [2192, 725, 188, 366, 511, 355, 189, 2054, 2156, 1979, 1023, 1636, 276, 251, 1319, 2019, 1554, 849, 823, 26, 402, 1978, 239, 2003, 923, 2117, 2145, 1641, 810, 258, 2183, 1976, 241, 154, 2014, 2227, 2151, 1322, 249, 1645, 1102, 1178, 438, 237, 702, 256, 89, 1183, 1586, 388, 1620, 2185, 399, 1737, 1991, 1958, 259, 668, 1631, 1216, 1147, 995, 983, 1831, 1137, 939, 382, 2129, 2000, 1220, 1187, 198, 289, 683, 1993, 1740, 1925, 387, 2067, 1000, 1003, 538, 1961, 1350, 1190, 1964, 230, 684, 676, 804, 1972, 1192, 917, 1868, 222, 1320]. Systems [1303, 1280, 1974, 928, 1195, 1932, 2029, 940, 2002, 1196, 1534, 857, 524, 2009, 1397, 317, 486, 695, 2120, 352, 1999, 1870, 1845, 2018, 2133, 1970, 1984, 1849, 484, 1375, 1605, 1200, 696, 488, 1663, 1133, 919, 2158, 1242, 1666, 1765, 1778, 2127, 1323, 1963, 1977, 2137, 3, 1996, 369, 2153, 1990, 386, 83, 1992, 1973, 2044, 1015, 109, 385, 519, 682, 1965, 1797, 192, 85, 2124, 838, 1712, 900, 1413, 769, 1698, 1416, 58, 1691, 776, 778, 785, 757, 1721, 1252, 1262, 63, 1432, 912, 1264, 914, 1415, 1428, 1163, 1185, 332, 1169]. Systolic [1360, 479, 1342, 1077, 418, 351].

Table [1823, 75]. Tableau [1219, 1669]. Tableau-Based [1669]. Tactical [1786]. Tactics [99]. Tail [1110]. Taken [42]. Talk [1296, 1311]. Talking [401]. TAM [2137].

Taming [2147]. Tarski [1211]. Task [436, 1147, 2121, 1452]. task-adaptive [1452]. Task-Driven [1147]. Tasking [243, 244]. Tasks [188, 1407, 1692]. TAXON [2094]. Taxonomic [1692, 1397]. TC2000 [2108]. TCM [1388]. TCSP [366, 1067]. Teach [1285]. Teaching [1281, 1284, 1282, 1283, 1289, 1287, 450, 449]. Technical [711, 1753]. Technique [746, 1148, 894, 1042, 198, 1793, 2059, 2169, 1229]. Techniques [979, 1955, 1118, 258, 931, 622, 927, 2099, 850, 418, 327, 1797, 1713, 651, 565]. Technology [490, 497, 330, 1762, 332]. teleological [1693]. Template [396]. Template-based [396]. Temporal [1118, 1121, 1142, 236, 926, 1542, 1132, 1144, 2184, 2129, 1813, 1716, 2133, 1126, 233, 868, 1405, 1404, 2153, 206, 2159, 213, 1146, 2190, 901, 765, 1256, 1464, 793, 1433]. Temporally [711]. Ten [2027]. Tense [829, 1932, 1702]. Term [276, 1231, 314, 539, 1181, 382, 965, 676, 2075, 695, 2173, 946, 1200, 696, 2051, 778, 756]. Terminal [525, 434]. Termination [706, 954, 683, 696, 345, 757]. Terminological [2066, 1651]. Terms [680, 2173, 681, 794, 972]. Ternary [1122]. Terrain [1898]. tessellation [777]. Test [277, 278, 1735, 1326, 2204, 1325, 780]. Testable [1939]. Testing [986, 245, 260, 1610, 1009, 1873, 2043, 682, 902]. Tests [1856]. Text [1522, 1860, 1561, 1534, 562, 1163, 1753, 1564, 161, 1437]. Text-Oriented [1163]. Texts [1542, 1443, 779]. Textual [2146]. Their [1358, 1606, 965, 2146, 916, 2016, 818, 1121, 1339, 1372, 760, 798, 1434, 1254]. Theme [1271]. Theorem [1111, 1864, 1051, 522, 1897, 1936, 1945, 1224, 549, 1249, 654]. Theorems [1107]. Theoretic [1058, 1334, 1238, 2200, 1114, 741, 203, 2190, 591, 654]. Theoretical [277, 1588, 1063, 269]. Theories [2017, 688, 1638, 1113, 1654, 1226, 703, 1420, 763, 784, 1261, 760]. Theory [986, 1165, 836, 964, 1207, 2135, 2047, 952, 546, 1595, 1954, 991, 272, 1408, 1648, 70, 866, 419, 365, 800, 1370, 1657, 884, 1844, 464, 1199, 1377, 1124, 832, 799, 1743, 316, 1066, 712, 1628, 1681, 1690, 1461, 654, 1710, 1250]. Theory-Based [316]. Theses [2027]. Thesis [1484]. Thin [589]. Thin-plate [589]. Thoughts [1990]. Thread [339, 874]. Threading [881]. Threads [480]. Three [436, 1837, 1916, 1093, 1546, 1679, 783, 1267]. Three-Dimensional [1916]. Three-Phase [436]. Three-Valued [1546, 1679, 783, 1267]. Threshold [1079, 1612]. TICKLE [1752]. Tight [146, 1911]. Tilings [537]. TIM [2088]. Time [1941, 188, 436, 189, 104, 829, 1288, 749, 1896, 509, 2122, 2135, 1403, 170, 2152, 2117, 2145, 258, 75, 2183, 2118, 2123, 718, 43, 862, 160, 1915, 2136, 2185, 560, 1737, 259, 2115, 1328, 1144, 2184, 452, 891, 1740, 372, 1925, 1916, 2186, 516, 191, 1005, 463, 1739, 1972, 2132, 1367, 1483, 1974, 147, 1194, 1918, 260, 1289, 867, 238, 1007, 2120, 2133, 2134, 281, 2072, 1380, 1929, 2137, 2144, 1973, 1204, 519, 194, 213, 247, 844, 726, 192, 2139, 511, 785, 1469, 1435, 420, 519]. Time-Constrained [1007]. Time-Stamp [160, 1328]. Time-Stamps [191]. Time-Table [75]. Time-Triggered [1972]. Time-Varying [463]. Timed [2187, 2141, 2146, 2143, 1009, 2142, 2182, 1002]. Timeliness [2124]. Timing [101]. Timings [352]. TMS [1653, 665]. Tokens [1307]. Tokio [1126]. Tolerance [1979, 176, 1978, 1987, 1977, 1840, 184, 439]. Tolerance-Based [1840]. Tolerant [465, 1980, 195, 165, 2117, 2145, 1976, 2128, 2129, 2126, 1432]. tomographic [575]. tomography [567]. Tool [432, 1789, 253, 1790, 218, 982, 210, 220, 1161, 1818, 422, 501, 487, 1820, 2075, 2120, 240, 333, 989, 1712, 1723, 650, 1430, 1814]. Toolbox [922]. Toolkit [362]. Tools [266, 1786, 154, 485, 497, 330, 2181, 217, 494, 1164, 1290]. Toolset [1816, 1817]. Top [623, 906, 2155, 932, 624, 976]. Top-Down [623, 624, 906]. Topics [1288]. Topological [926, 925, 847, 852, 587]. Topologies [1909, 108]. Topology [83]. Toroidal [1095]. Tossing [1299, 1313]. Totally [1950, 1014]. Toulouse [2088]. tournaments [639]. tower [760]. TR [1852]. Trace [1565, 991, 522, 768]. Tracking [1228, 590]. Tracks [1885]. Trail [1918]. Trainable [595]. Training [1274, 197, 1275, 1276, 1715]. Transaction [1794, 250]. Transactions [225, 190]. Transfer [2220, 1574, 2130, 1126, 121, 247, 4]. Transferable [1634]. Transfers [1890]. Transformation [1774, 2083, 1205, 1236, 631, 63]. Transformational [1497, 1829, 1489, 1490]. Transformations [1496, 430, 994, 1188, 2159]. transformed [647]. Transforming [2096, 1682]. Transient [2127]. Transient-Recovery [2127]. Transition [540, 369, 1015, 902, 1430]. Transitive [286, 1381, 153, 817]. Transitively [2057]. Transitivity [1792]. Translating [2141]. Translation [303]. Translations [1942]. Translator [398, 1520]. transmission [658]. Transport [850]. Transputer [89, 776]. Transversals [554]. Trapdoor [36]. Trapdoors [1611]. Travelling [2025, 87, 2010, 84]. Traversals [182]. Traverses [1876]. Treatment [1543]. Tree [890, 1206, 1070, 165, 747, 1077, 1952, 1123, 2167, 166, 86, 1193, 435, 163, 748, 1869, 182, 1093, 932, 1417, 1852]. Trees [723, 138, 145, 144, 517, 1910, 1075, 412, 1117, 716, 1180, 740, 559, 139, 2030, 734, 393, 753, 143, 927, 1050, 1091, 1872, 416, 938, 792]. Treewidth [545, 717]. Trellis [377, 463, 1388, 661, 65]. Trends [1870, 207]. Triangle [1539]. Triangulated [142]. Triangulation [556, 718, 147]. Triangulations [555, 1900, 1851]. Tries [734]. Triggered [1972]. Trilinear [474]. TRIMM [103]. Triple [29]. Triplets [429]. Trivalued [1665]. True [1009, 1010]. Truly [234]. Trust [1951]. Trusted [24, 1612]. Truth [669, 1720, 668, 1004, 1667, 672, 1462]. Truth-Preserving [1004]. truths [1680]. Tseitin [514]. Tuple [376]. Tuple-Spaces [376]. Turing [1218, 1081]. tutor [1702]. Tutorial [1158, 1827, 1824, 1826, 1825]. Tutoring [400]. Two [1835, 2194, 2205, 1118, 546, 1030, 1488, 779, 783, 7, 1092, 1114, 1531, 2175, 483, 2206, 29, 567, 259, 535, 2208, 796, 660]. Two-Key [29]. Two-Level [1531, 2175]. Two-Party [2194]. two-prover [2208]. two-valued [783]. Typability [871]. Type [701, 964, 1209, 952, 1773, 702, 1041, 943, 891, 972, 683, 1113, 2048, 269, 870, 621, 942, 1853, 872, 454, 242, 774]. Type-Safe [621]. Type-Schemes [972]. Type-theoretical [269]. Typed [701, 1038, 604, 877, 620, 318, 770]. Types [319, 973, 1231, 975, 882, 317, 1239, 974, 1246, 1931, 970, 1247, 606, 761, 765, 760]. Typicality [1728, 1466]. Typing [2047, 313].

U [1038]. U-Log [1038]. Ultimate [132]. ultrasound [586, 601]. Unambiguity [1073]. Unary [1937]. Unbounded [191]. Unboxed [899]. Uncertain [823, 1637, 397, 1642, 669, 853, 1657, 805, 1663, 840, 597, 51, 1458]. Uncertainty [1636, 1689, 2003, 1625, 1640, 856, 1644, 1631, 1651, 841, 837, 812, 833, 1664, 1671, 1672, 846, 838, 585]. Unconditionally [2229, 1310, 1308, 4]. Undecidability [1375]. Undeniable [1309, 2229, 1619, 41, 1584, 1586, 1585]. Undergraduate [1273]. Underlying [862]. Understanding [1522, 1863, 1534, 1560, 2022, 1564, 1729, 1452]. Undirected [753]. Unequal [1372, 636]. Unequal-Error-Protection [1372]. Unfolding [1044].

Unfolding-Definition-Folding [1044]. Unidirectional [457, 637, 1385, 59]. Unification [763, 2075, 708, 870, 710, 709, 787, 797, 1527]. Unification-ID [1527]. Unification-ID/ LP [1527]. Unified [2047, 1347, 1664, 1143]. Uniform [507, 1829, 1947, 349, 797]. Unifying [1075]. Unimodular [1389]. Unintrusive [1799]. Union [973, 975, 752, 695]. Unique [1106, 703]. Unique-Sort [703]. Unit [675, 230]. Universal [706, 1110, 1326, 1239, 2218, 2196]. Universality [1325]. universe [1438]. University [941]. Unknown [1898, 1447]. Unnecessary [1044]. Unordered [166]. Unreliable [1639]. Unsafe [1666]. Unstructured [730]. Untraceable [2219]. Untyped [699]. Updatable [2040]. Update [1028, 735]. Updates [2057, 2038, 2039, 1438]. Updating [325, 805, 3, 2045]. Upgrading [1793]. Upon [859]. Upper [789]. Use [1160, 1496, 1699, 1127, 903, 237, 1631, 1185, 1801, 934, 464, 1591, 2211, 1693, 2232, 570]. Usefulness [1953]. User [1421, 1249, 1731, 1394, 1873, 1878, 922, 207, 1723, 1677]. User-oriented [1249]. Uses [1730]. Using [1439, 2157, 1701, 2225, 1966, 1790, 1776, 1122, 430, 1806, 441, 555, 243, 2145, 236, 1643, 249, 2171, 1022, 2136, 2201, 1514, 881, 259, 1136, 2178, 1515, 2184, 1511, 173, 220, 1364, 78, 1135, 80, 1004, 86, 222, 544, 1131, 1128, 1149, 81, 162, 404, 1804, 615, 585, 481, 1431, 2173, 1126, 40, 1376, 1133, 1666, 1513, 9, 1959, 250, 94, 270, 2051, 2053, 1412, 1571, 1502, 591, 567, 662, 1686, 573, 584, 588, 2175, 652, 581, 1710]. Utility [1667]. Utilization [208]. Utilize [1329]. Utilizing [497, 463]. Utterance [2013]. uUing [2135].

v [336]. Vacillatory [1935]. Vague [810]. Vagueness [1674]. Validation [503, 279, 1748, 783]. Valley [1558].

Valuation [1640, 1428]. Value [713, 160, 429, 771]. Valued [827, 685, 828, 1219, 461, 1546, 1679, 783, 1467, 1267]. Values [899, 833, 1447]. Variable [205, 548, 380, 1483, 413]. Variable-Grain [380]. Variable-Length [548]. Variable-Resolution [205]. Variables [1661, 1044, 1101, 1416]. variance [565]. Variant [1213, 96, 652]. Variations [665, 1271]. Various [42]. Varshamov [642]. Varying [463]. VDM [1790, 1821, 1782, 1786, 1774, 1773, 1819, 1733, 1809, 1788, 1813, 1826, 1804, 1785, 1798, 1812, 1811, 1787, 1792]. VDM-Reification [1792]. VDM-SL [1782, 1819, 1812]. Vector [96, 1386]. Vectorized [1123, 2167]. Vectors [404, 754]. Vehicles [420]. ventricular [590]. Verb [1538]. Verbs [1557]. Verifiable [1321, 2200]. Verification [2157, 1122, 2135, 2176, 1142, 2199, 1119, 2164, 2136, 510, 1136, 2178, 503, 997, 2184, 2167, 1925, 2165, 422, 1130, 1128, 1149, 1151, 353, 1805, 1011, 2127, 2175, 2126, 1140, 2153, 1015, 2159, 606, 2125]. Verifier [1126]. Verifiers [1324]. Verify [1131]. Verifying [366, 1141, 1121, 2151, 2185, 2180, 1875]. Version [1696, 2092, 622, 1323, 726, 1417]. Version-space [1696]. Versioning [2049]. Versions [2084]. Versus [201, 283, 1972, 1895, 1216, 1083]. Vertex [1945]. Vertical [164]. Very [1166, 200]. VHL [1202]. VHL-Language [1202]. Via [1916, 634, 533, 2086, 870, 1589, 64, 913]. Vicinity [55]. VICTOR [22]. View [1648, 1757, 1551, 1872, 1749, 128, 1462, 1436]. Viewed [1494]. Viewpoint [1764]. Viewpoints [1680]. Views [1028, 2040]. Virtual [1860, 70, 385]. Visibility [164, 1952, 751, 1021, 1854, 750]. Visible [749]. VISICOM [331]. Vision [2054, 205, 235, 1672, 1708, 1709, 1707]. Visit [1501]. Visit-Oriented [1501]. Visual [1733, 1809, 331, 2029, 1717]. Visualization [1863, 411]. Viterbi [652]. VLSI [432, 88, 2012, 418, 183]. Vocabulary [1534]. voice [1413]. Volume [1876, 921, 601, 602]. Voluntary [2120]. Voronoi [719, 1850, 1853]. Voxel [565]. vs [1488, 224, 254, 1436]. VSF [488].

Walking [1880]. Wall [1196, 590]. WAM [1208, 1212]. WAM-Based [1212]. Ware [1269]. WASTL [1687]. Watchman [1902]. Watchmen [745]. Wave [1399]. Wave-shaping [1399]. Way [1302, 1514, 837, 1917, 2215, 1602, 1316, 1085, 516]. Ways [1799]. Weak [1952]. Weakly [749, 1327, 1375]. Weaknesses [1619, 1584]. Weight [1358, 811]. Weighted [753, 561, 859, 1853, 1889, 735, 652]. weighted-output [652]. Weights [1383, 633]. well [635, 1468]. WF [1456]. Where [39]. Which [749, 8, 635, 226]. Whitney [1374]. Whole [619]. Width [1899, 1892]. Wiedemann [1350]. Will [978]. WIMs [1362]. Window [1643]. Wire [876, 1893]. within [861, 1702, 1725, 1804, 1704]. Without [1121, 1313, 1660, 24, 1612, 564]. Withstanding [27]. Word [1072, 1105, 398, 1532, 2015, 1701]. Words [1936, 766]. Work [2109]. Workbench [1526, 2161]. Working [819]. Works [1784]. Workshop [1280]. Workspaces [225]. World [1020, 1549]. Worlds [1670]. Worst [1865, 1911]. Worst-case [1865]. Wrapper [968]. Write [1913]. Write-Optimal [1913]. Writing [282, 1779, 451, 1520, 1753, 65]. Wrong [1759, 133]. WUMs [1362].

X [295, 301]. XOR [1583]. XRA [303].

XYZ [1846, 1847].

Y-Tree [1077]. Year [1287, 207].

Z [613, 1779, 612, 1780, 2136, 1801, 1818, 1815, 1781]. Zero [1319, 1, 41, 478, 1318, 1320, 1321, 1573, 1329, 2226, 2208]. Zero-Knowledge [1319, 1, 41, 1318, 1320, 1321, 1329, 2226, 2208]. Zipf [2020]. Ziv [1575]. Zone [714, 1864].

References

Burmester:1991:ALN

[1] Mike V. D. Burmester and Yvo Desmedt. All languages in NP have divertible zero-knowledge proofs and arguments under cryptographic assumptions (extended abstract). Lecture Notes in Computer Science, 473:1– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730001.
pdf.

Davida:1991:IMR

[2] George I. Davida, Yvo Desmedt, and René Peralta. On the importance of memory resources in the security of key exchange protocols (extended abstract). Lecture Notes in Computer Science, 473:11– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730011.
pdf.

Shinozaki:1991:PSK

[3] S. Shinozaki, Toshiya Itoh, Atsushi Fujioka, and Shigeo Tsujii. Provably secure key-updating schemes in identity-based systems. Lecture Notes in Computer Science, 473:16– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730016.
pdf.

denBoer:1991:OTP

[4] Bert den Boer. Oblivious transfer protecting secrecy — an implementation for oblivious transfer protecting secrecy almost unconditionally and a bitcommitment based on factoring protecting secrecy unconditionally. Lecture Notes in Computer Science, 473:31– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730031.
pdf.

Santis:1991:PRP

[5] Alfredo De Santis and Giuseppe Persiano. Public-randomness in public-key cryptography (extended abstract). Lecture Notes in Computer Science, 473: 46–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730046.
pdf.

Brickell:1991:IIS

[6] Ernest F. Brickell and Kevin S. Mc-Curley. An interactive identification scheme based on discrete logarithms and factoring (extended abstract). Lecture Notes in Computer Science, 473: 63–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730063.
pdf.

Lenstra:1991:FTL

[7] Arjen K. Lenstra and Mark S. Manasse. Factoring with two large primes (extended abstract). Lecture Notes in Computer Science, 473:72– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730072.
pdf.

Evertse:1991:WNR

[8] Jan-Hendrik Evertse and Eugène van Heyst. Which new RSA signatures can be from some given RSA signatures? (extended abstract). Lecture Notes in Computer Science, 473:83– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730083.
pdf.

Scheidler:1991:IKE

[9] Renate Scheidler, Johannes A. Buchmann, and Hugh C. Williams. Implementation of a key exchange protocol using real quadratic fields (extended abstract). Lecture Notes in Computer Science, 473:98–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730098.
                                        pdf.

Morain:1991:DPP

[10] François Morain. Distributed primality proving and the primality of (2^{3539} + 1)/3. Lecture Notes in Computer Science, 473:110–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730110.
                                        pdf.

Lloyd:1991:PBF

[11] Sheelagh Lloyd. Properties of binary functions. Lecture Notes in Computer Science, 473:124–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730124.
                                        pdf.

Pieprzyk:1991:HCP

[12] Józef P. Pieprzyk. How to construct pseudorandom permutations from single pseudorandom functions. Lecture Notes in Computer Science, 473:140– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730140.
pdf.

Nyberg:1991:CBF

[13] Kaisa Nyberg. Constructions of bent functions and difference sets. Lecture Notes in Computer Science, 473:151– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730151.
pdf.

Preneel:1991:PCB

[14] Bart Preneel, Werner Van Leekwijck, Luc Van Linden, René Govaerts, and Joos Vandewalle. Propagation characteristics of Boolean functions. Lecture Notes in Computer Science, 473:161– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730161.
pdf.

Niederreiter:1991:LCP

[15] Harald Niederreiter. The linear complexity profile and the jump complexity of keystream sequences. Lecture Notes in Computer Science, 473:174– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730174.
                                        pdf.

Dai:1991:LBL

[16] Zong-Duo Dai, Thomas Beth, and Dieter Gollmann. Lower bounds for the linear complexity of sequences over residue rings. Lecture Notes in Computer Science, 473:189–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730189.
                                        pdf.

Jansen:1991:CRP

[17] Cees J. A. Jansen. On the construction of run permuted sequences. Lecture Notes in Computer Science, 473: 196–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730196.
                                        pdf.

Meier:1991:CPC

[18] Willi Meier and Othmar Staffelbach. Correlation properties of combiners with memory in stream ciphers (extended abstract). Lecture Notes in Computer Science, 473:204– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730204.
pdf.

Chan:1991:CFG

[19] Agnes Hui Chan, Mark Goresky, and Andrew Klapper. Correlation functions of geometric sequences. Lecture Notes in Computer Science, 473:214– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730214.
pdf.

Yacobi:1991:EFA

[20] Yacov Yacobi. Exponentiating faster with addition chains. Lecture Notes in Computer Science, 473:222–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730222.
pdf.

Dusse:1991:CLM

[21] Stephen R. Dusse and Burt S. Kaliski Jr. A cryptographic library for the Motorola DSP56000. Lecture Notes in Computer Science, 473:230– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730230.
pdf.

Orup:1991:VER

[22] Holger Orup, Erik Svendsen, and Erik Andreasen. VICTOR — an efficient RSA hardware implementation. Lecture Notes in Computer Science, 473: 245–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730245.
pdf.

Bennett:1991:EQC

[23] Charles H. Bennett, François Bessette, Gilles Brassard, Louis Salvail, and John Smolin. Experimental quantum cryptography. Lecture Notes in Computer Science, 473:253– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730253.
pdf.

Ingemarsson:1991:PSS

[24] Ingemar Ingemarsson and Gustavus J. Simmons. A protocol to set up shared secret schemes without the assistance of mutually trusted party. Lecture Notes in Computer Science, 473:266– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730266.
pdf.

Sgarro:1991:LBA

[25] Andrea Sgarro. Lower bounds for authentication codes with splitting. Lecture Notes in Computer Science, 473: 283–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730283.
pdf.

Beutelspacher:1991:EFS

[26] Albrecht Beutelspacher and Ute Rosenbaum. Essentially ℓ-fold secure authentication systems. Lecture Notes in Computer Science, 473:294–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730294.
pdf.

Smeets:1991:CAC

[27] Ben Smeets, Peter Vanroose, and Zhe xian Wan. On the construction of authentication codes with sand codes withstanding spoofing attacks of order L ≥ 2. Lecture Notes in Computer Science, 473:306– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730306.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730306.
pdf.

Stern:1991:CPK

[28] Jacques Stern and Philippe Toffin. Cryptanalysis of a public-key cryptosystem based on approximations by rational numbers. Lecture Notes in Computer Science, 473:313–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730313.
pdf.

vanOorschot:1991:KPA

[29] Paul C. van Oorschot and Michael J. Wiener. A known-plaintext attack on two-key triple encryption. Lecture Notes in Computer Science, 473:318– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730318.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/0473/04730318.
pdf.

Miyaguchi:1991:CSH

[30] Shoji Miyaguchi, Kazuo Ohta, and Masahiko Iwata. Confirmation that some hash functions are not collision free. Lecture Notes in Computer Science, 473:326–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0473/04730326.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0473/04730326.pdf.

Bauspiess:1991:IPE

[31] Fritz Bauspieß, Hans-Joachim Knobloch, and Peer Wichmann. Inverting the pseudo exponentiation. Lecture Notes in Computer Science, 473:344–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730344.
pdf.

Hwang:1991:CGO

[32] Tzonelih Hwang. Cryptosystem for group oriented cryptography. Lecture Notes in Computer Science, 473:352– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730352.
pdf.

Maurer:1991:PSS

[33] Ueli M. Maurer. A provably-secure strongly-randomized cipher. Lecture Notes in Computer Science, 473:361– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730361.
                                        pdf.

Kurosawa:1991:GPK

[34] Kaoru Kurosawa, Yutaka Katayama, Wakaha Ogata, and Shigeo Tsujii. General public key residue cryptosystems and mental poker protocols. Lecture Notes in Computer Science, 473: 374–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730374.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730374.
                                        pdf.

Lai:1991:PNB

[35] Xuejia Lai and James L. Massey. A proposal for a new block encryption standard. Lecture Notes in Computer Science, 473:389–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730389.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730389.
                                        pdf.

Niemi:1991:NTK

[36] Valtteri Niemi. A new trapdoor in knapsacks. Lecture Notes in Computer Science, 473:405–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730405.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730405.
pdf.

Santis:1991:DPS

[37] Alfredo De Santis and Moti Yung. On the design of provably-secure cryptographic hash functions (extended summary). Lecture Notes in Computer Science, 473:412–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730412.
pdf.

Ong:1991:FSG

[38] H. Ong and Claus P. Schnorr. Fast signature generation with a Fiat Shamir-like scheme. Lecture Notes in Computer Science, 473:432–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730432.
pdf.

Bleumer:1991:RSS

[39] Gerrit Bleumer, Birgit Pfitzmann, and Michael Waidner. A remark on signature scheme where forgery can be proved. Lecture Notes in Computer Science, 473:441–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730441.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730441.
pdf.

Ohta:1991:MAH

[40] Kazuo Ohta, Tatsuaki Okamoto, and Kenji Koyama. Membership authentication for hierarchical multigroups using the extended Fiat– Shamir scheme. Lecture Notes in Computer Science, 473:446–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730446.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730446.
pdf.

Chaum:1991:ZKU

[41] David Chaum. Zero-knowledge undeniable signatures (extended abstract). Lecture Notes in Computer Science, 473:458–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0473/04730458.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0473/04730458.
pdf.

Guillou:1991:PTA

[42] Louis C. Guillou, Jean-Jacques Quisquater, Mike Walker, Peter Landrock, and Caroline Shaer. Precautions taken against various potential attacks in ISO/IEC DIS 9796 ≪Digital Signature Scheme Giving Message Recovery≫. Lecture Notes in Computer Science, 473:465– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730465.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730465.
pdf.

Domingo-Ferrer:1991:SRT

[43] Josep Domingo-Ferrer. Software runtime protection: A cryptographic issue. Lecture Notes in Computer Science, 473:474–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0473/04730474.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0473/04730474.
pdf.

Girault:1991:IBI

[44] Marc Girault. An identity-based identification scheme based on discrete logarithms modulo a composite number. Lecture Notes in Computer Science, 473:481–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730481.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/0473/04730481.
                                        pdf.

Golic:1991:NCC

[45] Jovan Dj. Golic and Miodrag J. Mihaljevic. A noisy clock-controlled shift register cryptanalysis concept based on sequence comparison approach. Lecture Notes in Computer Science, 473: 487–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730487.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730487.
                                        pdf.

Kaliski:1991:MMD

[46] Burt S. Kaliski Jr. The MD4 message digest algorithm. Lecture Notes in Computer Science, 473:492–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730492.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0473/04730492.
                                        pdf.

Burmester:1991:REI

[47] Mike V. D. Burmester. A remark on the efficiency of identification schemes. Lecture Notes in Computer Science, 473:493–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0473/04730493.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0473/04730493.
pdf.

Meister:1991:IMA

[48] Gisela Meister. On an implementation of the Mohan-Adiga algorithm. Lecture Notes in Computer Science, 473: 496–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0473/04730496.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0473/04730496.
pdf.

Brauer:1991:SBE

[49] W. Brauer, R. Gold, and W. Vogler. A survey of behaviour and equivalence preserving refinements of Petri nets. Lecture Notes in Computer Science, 483:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1991:DSH

[50] B. Butler, R. Esser, and R. Mattmann. A distributed simulator for high order Petri nets. Lecture Notes in Computer Science, 483:47–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardoso:1991:PNU

[51] J. Cardoso, R. Valette, and D. Dubois. Petri nets with uncertain markings. Lecture Notes in Computer Science, 483:64–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colom:1991:CGS

[52] J. M. Colom and M. Silva. Convex geometry and semiflows in P/T nets. A comparative study of algorithms for computation of minimal Psemiflows. Lecture Notes in Computer Science, 483:79–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colom:1991:ILB

[53] J. M. Colom and M. Silva. Improving the linearly based characterization of P/T nets. Lecture Notes in Computer Science, 483:113–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couvreur:1991:LIC

[54] J. M. Couvreur and J. Martinez. Linear invariants in commutative high level nets. Lecture Notes in Computer Science, 483:146–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desel:1991:VRN

[55] J. Desel and A. Merceron. Vicinity respecting net morphisms. Lecture Notes in Computer Science, 483:165– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dutheillet:1991:RSP

[56] C. Dutheillet and S. Haddad. Regular stochastic Petri nets. Lecture Notes in Computer Science, 483:186–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1991:CHB

[57] J. Esparza and M. Silva. Circuits, handles, bridges and nets. Lecture Notes in Computer Science, 483:210–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1991:ASF

[58] J. Esparza and M. Silva. On the analysis and synthesis of free choice systems. Lecture Notes in Computer Science, 483:243–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Florin:1991:PNM

[59] G. Florin, C. Kaiser, and S. Natkin. Petri net models of a distributed election protocol on a unidirectional ring. Lecture Notes in Computer Science, 483:287–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huber:1991:HCP

[60] P. Huber, K. Jensen, and R. M. Shapiro. Hierarchies in coloured Petri nets. Lecture Notes in Computer Science, 483:313–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1991:CPN

[61] K. Jensen. Coloured Petri nets: A high level language for system design and analysis. Lecture Notes in Computer Science, 483:342–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nutt:1991:SSA

[62] G. J. Nutt. A simulation system architecture for graph models. Lecture Notes in Computer Science, 483:417– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pomello:1991:STP

[63] L. Pomello and C. Simone. A state transformation preorder over a class of EN systems. Lecture Notes in Computer Science, 483:436–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Souissi:1991:CNC

[64] Y. Souissi and G. Memmi. Composition of nets via a communication medium. Lecture Notes in Computer Science, 483:457–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stotts:1991:ATS

[65] P. D. Stotts and R. Furuta. alpha Trellis: A system for writing and browsing Petri-net-based hypertext. Lecture Notes in Computer Science, 483:471– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valmari:1991:SSR

[66] A. Valmari. Stubborn sets for reduced state space generation. Lecture Notes in Computer Science, 483:491– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:Ia

[67] Anonymous. Introduction. Lecture Notes in Computer Science, 496:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:GA

[68] Anonymous. Genetic algorithms. Lecture Notes in Computer Science, 496:3, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiben:1991:GCG

[69] A. E. Eiben, E. H. L. Aarts, and K. M. Van Hee. Global convergence of genetic algorithms: A Markov chain analysis. Lecture Notes in Computer Science, 496:4–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1991:TVA

[70] D. E. Goldberg. The theory of virtual alphabets. Lecture Notes in Computer Science, 496:13–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesser:1991:TOM

[71] J. Hesser and R. Manner. Towards an optimal mutation probability for genetic algorithms. Lecture Notes in Computer Science, 496:23–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesser:1991:AGA

[72] J. Hesser and R. Manner. An alternative genetic algorithm. Lecture Notes in Computer Science, 496:33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeJong:1991:AIR

[73] K. A. De Jong and W. M. Spears. An analysis of the interacting roles of population size and crossover in genetic algorithms. Lecture Notes in Computer Science, 496:38–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blume:1991:GSS

[74] C. Blume. GLEAM — A system for simulated intuitive learning. Lecture Notes in Computer Science, 496:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colorni:1991:GAH

[75] A. Colorni, M. Dorigo, and V. Maniezzo. Genetic algorithms and highly constrained problems: The time-table case. Lecture Notes in Computer Science, 496:55–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davidor:1991:ESD

[76] Y. Davidor and Y. Goldberg. An evolution standing on the design of redundant manipulators. Lecture Notes in Computer Science, 496:60–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerrits:1991:RCN

[77] M. Gerrits and P. Hogeweg. Redundant coding of an NP-Complete problem allows effective genetic algorithm search. Lecture Notes in Computer Science, 496:70–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hulin:1991:CPG

[78] M. Hulin. Circuit partitioning with genetic algorithms using a coding scheme to preserve the structure of a circuit. Lecture Notes in Computer Science, 496:75–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Husbands:1991:GAP

[79] P. Husbands, F. Mill, and S. Warrington. Genetic algorithms, production plan optimisation, and scheduling. Lecture Notes in Computer Science, 496: 80–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:1991:SIU

[80] T. Johnson and P. Husbands. System identification using genetic algorithms. Lecture Notes in Computer Science, 496:85–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucasius:1991:CAD

[81] C. B. Lucasius, S. Werten, A. H. J. M. van Aert, G. Kateman, and M. J. J. Blommers. Conformational analysis of DNA using genetic algorithms. Lecture Notes in Computer Science, 496:90–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suh:1991:OOG

[82] J.-Y. Suh and C.-D. Lee. Operatororiented genetic algorithm and its application to the sliding block puzzle problem. Lecture Notes in Computer Science, 496:98–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thierens:1991:TEG

[83] D. Thierens and L. Vercauteren. A topology exploiting genetic algorithm to control dynamic systems. Lecture Notes in Computer Science, 496:104– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ulder:1991:GLS

[84] N. L. J. Ulder, E. H. L. Aarts, H.- J. Bandelt, P. J. M. van Laarhoven, and E. Pesch. Genetic local search algorithms for the travelling salesman problem. Lecture Notes in Computer Science, 496:109–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deGaris:1991:GPA

[85] H. de Garis. Genetic programming: Artificial nervous systems, artificial embryos and embryological electronics. Lecture Notes in Computer Science, 496:117–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koza:1991:CFD

[86] J. R. Koza. Concept formation and decision tree induction using the genetic programming paradigm. Lecture Notes in Computer Science, 496:124– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braun:1991:STS

[87] H. C. Braun. On solving travelling salesman problems by genetic algorithms. Lecture Notes in Computer Science, 496:129–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohoon:1991:GAP

[88] J. P. Cohoon, W. N. Martin, and D. S. Richards. Genetic algorithms and punctuated equilibria in VLSI. Lecture Notes in Computer Science, 496:134– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogarty:1991:IGA

[89] T. C. Fogarty and R. Huang. Implementing the genetic algorithm on Transputer based parallel processing systems. Lecture Notes in Computer Science, 496:145–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorges-Schleuter:1991:EPG

[90] M. Gorges-Schleuter. Explicit parallelism of genetic algorithms through population structures. Lecture Notes in Computer Science, 496:150–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroger:1991:PGP

[91] B. Kroger, P. Schwenderling, and O. Vornberger. Parallel genetic packing of rectangles. Lecture Notes in Computer Science, 496:160–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonLaszewski:1991:PGP

[92] G. von Laszewski and H. Muhlenbein. Partitioning a graph with a parallel genetic algorithm. Lecture Notes in Computer Science, 496:165–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neuhaus:1991:SMP

[93] P. Neuhaus. Solving the mappingproblem — experiences with a genetic algorithm. Lecture Notes in Computer Science, 496:170–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Starkweather:1991:OUD

[94] T. Starkweather, D. Whitley, and K. Mathias. Optimization using distributed genetic algorithms. Lecture Notes in Computer Science, 496:176– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herdy:1991:AED

[95] M. Herdy. Application of the Evolutionsstrategie to discrete optimization problems. Lecture Notes in Computer Science, 496:188–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kursawe:1991:VES

[96] F. Kursawe. A variant of evolution strategies for vector optimization. Lecture Notes in Computer Science, 496: 193–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lohmann:1991:AES

[97] R. Lohmann. Application of evolution strategy in parallel populations. Lecture Notes in Computer Science, 496: 198–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rudolph:1991:GOM

[98] G. Rudolph. Global optimization by means of distributed evolution strategies. Lecture Notes in Computer Science, 496:209–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Exman:1991:SSG

[99] I. Exman. Solving sequential games with Boltzmann-learned tactics. Lecture Notes in Computer Science, 496: 216–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1991:OSA

[100] K. H. Hoffmann, M. Christoph, and M. Hanf. Optimizing simulated annealing. Lecture Notes in Computer Science, 496:221–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kreuzer:1991:LTM

[101] L. Kreuzer and K. Goser. A local timing model for parallel optimization with Boltzmann machines. Lecture Notes in Computer Science, 496:226– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roussel-Ragot:1991:EFP

[102] P. Roussel-Ragot, N. Kouicem, and G. Dreyfus. Error-free parallel implementation of simulated annealing. Lecture Notes in Computer Science, 496: 231–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shoemaker:1991:TPP

[103] R. L. Shoemaker, H. H. Barrett, and R. H. Seacat. TRIMM: A parallel processor for image reconstruction by simulated annealing. Lecture Notes in Computer Science, 496:242–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrew:1991:RTC

[104] A. M. Andrew. The response-time constraint in neural evolution. Lecture Notes in Computer Science, 496:254– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1991:ANN

[105] R. J. Collins and D. R. Jefferson. An artificial neural network representation for artificial organisms. Lecture Notes in Computer Science, 496:259– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piramuthu:1991:FCB

[106] S. Piramuthu. Feature construction for back-propagation. Lecture Notes in Computer Science, 496:264–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomon:1991:ICR

[107] R. Salomon. Improved convergence rate of back-propagation with dynamic adaption of the learning rate. Lecture Notes in Computer Science, 496:269– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiffmann:1991:PEE

[108] W. Schiffmann, M. Joost, and R. Werner. Performance evaluation of evolutionarily created neural network topologies. Lecture Notes in Computer Science, 496:274–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vedenov:1991:OIP

[109] A. A. Vedenov, M. V. Zurin, E. B. Levchenko, G. D. Myl’nikov, A. N. Timashev, and A. Yu. Turygin. Optical image preprocessing for neural network classifier systems. Lecture Notes in Computer Science, 496:284–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hancock:1991:GGD

[110] P. J. B. Hancock and L. S. Smith. GANNET: Genetic design of a neural net for face recognition. Lecture Notes in Computer Science, 496:292– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heistermann:1991:AGA

[111] J. Heistermann. The application of a genetic approach as an algorithm for neural networks. Lecture Notes in Computer Science, 496:297–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffgen:1991:GIF

[112] K.-U. Hoffgen, H. P. Siemon, and A. Ultsch. Genetic improvements of feedforward nets for approximating functions. Lecture Notes in Computer Science, 496:302–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Todd:1991:EAA

[113] P. M. Todd and G. F. Miller. Exploring adaptive agency III: Simulating the evolution of habituation and sensitization. Lecture Notes in Computer Science, 496:307–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholz:1991:LSN

[114] M. Scholz. A learning strategy for neural networks based on a modified evolutionary strategy. Lecture Notes in Computer Science, 496:314–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forrest:1991:GAI

[115] S. Forrest and A. S. Perelson. Genetic algorithms and the immune system. Lecture Notes in Computer Science, 496:320–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manderick:1991:SC

[116] B. Manderick. Selectionist categorization. Lecture Notes in Computer Science, 496:326–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schachtner:1991:CSI

[117] A. Schachtner. A classifier system with integrated genetic operators. Lecture Notes in Computer Science, 496:331– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valenzuela-Rendon:1991:FCS

[118] M. Valenzuela-Rendon. The fuzzy classifier system: Motivations and first results. Lecture Notes in Computer Science, 496:338–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bersini:1991:HAP

[119] H. Bersini and F. J. Varela. Hints for adaptive problem solving gleaned from immune networks. Lecture Notes in Computer Science, 496:343–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decuyper:1991:RRN

[120] J. Decuyper and D. Keymeulen. A reactive robot navigation system based on a fluid dynamics metaphor. Lecture Notes in Computer Science, 496: 356–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paton:1991:TNM

[121] R. C. Paton, H. S. Nwana, M. J. R. Shave, T. J. M. Bench-Capon, and S. Hughes. Transfer of natural metaphors to parallel problem solving applications. Lecture Notes in Computer Science, 496:363–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voigt:1991:MSD

[122] H.-M. Voigt, J. Born, and I. Santibanez-Koref. Modelling and simulation of distributed evolutionary search processes for function optimization. Lecture Notes in Computer Science, 496: 373–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mataric:1991:PDS

[123] M. J. Mataric. Parallel, decentralized spatial mapping for robot navigation and path planning. Lecture Notes in Computer Science, 496:381–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adachi:1991:EDU

[124] N. Adachi and K. Matsuo. Ecological dynamics under different selection rules in distributed and iterated prisoner’s dilemma games. Lecture Notes in Computer Science, 496:388–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnard:1991:ASS

[125] S. T. Barnard and A. Bergman. Adaptation in signal spaces. Lecture Notes in Computer Science, 496:395–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1991:PMC

[126] L. R. Lopez and H. J. Caulfield. A principle of minimum complexity in evolution. Lecture Notes in Computer Science, 496:405–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paredis:1991:EDS

[127] J. Paredis. The emergence of data structures from local interactions. Lecture Notes in Computer Science, 496: 410–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schull:1991:VAL

[128] J. Schull. The view from the adaptive landscape. Lecture Notes in Computer Science, 496:415–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boseniuk:1991:BDH

[129] T. Boseniuk and W. Ebeling. Boltzmann- , Darwin-, and Haeckel-strategies in optimization problems. Lecture Notes in Computer Science, 496:430–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deGroot:1991:OCP

[130] C. de Groot, D. Wurtz, and K. H. Hoffmann. Optimizing complex problems by Nature’s algorithms: Simulated annealing and evolution strategy — A comparative study. Lecture Notes in Computer Science, 496:445–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmeister:1991:GAE

[131] F. Hoffmeister and T. Back. Genetic algorithms and evolution strategies — similarities and differences. Lecture Notes in Computer Science, 496:455– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bock:1991:BUM

[132] P. Bock. Building the ultimate machine: The emergence of artificial cognition. Lecture Notes in Computer Science, 496:472–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hehner:1991:WWF

[133] E. C. R. Hehner. What’s wrong with formal programming methods? Lecture Notes in Computer Science, 497:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Probert:1991:CSC

[134] R. L. Probert. Computer science and communications protocols: Research relationships. Lecture Notes in Computer Science, 497:24–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1991:RBM

[135] S. J. Smith. Reaping the benefits of mass production with massively parallel computing — four example applications. Lecture Notes in Computer Science, 497:33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hung:1991:QPP

[136] Y.-C. Hung and G.-H. Chen. On the quickest path problem. Lecture Notes in Computer Science, 497:44–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estivill-Castro:1991:PAS

[137] V. Estivill-Castro and D. Wood. Practical adaptive sorting. Lecture Notes in Computer Science, 497:47–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bshouty:1991:LBA

[138] N. H. Bshouty. Lower bounds for algebraic computation trees of functions with finite domains. Lecture Notes in Computer Science, 497:55–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haralambides:1991:AAB

[139] J. Haralambides and F. Makedon. Approximation algorithms for the bandwidth minimization problem for a large class of trees. Lecture Notes in Computer Science, 497:66–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ouksel:1991:IBB

[140] M. A. Ouksel and A. Yaagoub. The interpolation-based bintree and its application to graphics. Lecture Notes in Computer Science, 497:78–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlisle:1991:CI

[141] M. C. Carlisle and E. L. Lloyd. On the k-coloring of intervals. Lecture Notes in Computer Science, 497:90–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kooshesh:1991:FTS

[142] A. A. Kooshesh and B. M. E. Moret. Folding a triangulated simple polygon: Structural and algorithmic results. Lecture Notes in Computer Science, 497:102–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1991:RBS

[143] T. W. Lai and D. Wood. A relationship between self-organizing lists and binary search trees. Lecture Notes in Computer Science, 497:111–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cameron:1991:HCC

[144] H. Cameron and D. Wood. How costly can red-black trees be? Lecture Notes in Computer Science, 497: 117–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cameron:1991:BAT

[145] H. Cameron and D. Wood. Balance in AVL trees. Lecture Notes in Computer Science, 497:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosnard:1991:TLB

[146] M. Cosnard and A. G. Ferreira. A tight lower bound for selection in sorted X + Y. Lecture Notes in Computer Science, 497:134–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levcopoulos:1991:GTA

[147] C. Levcopoulos and A. Lingas. Greedy triangulation approximates the optimum and can be implemented in linear time in the average case. Lecture Notes in Computer Science, 497:139– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raman:1991:GRG

[148] R. Raman. Generating random graphs efficiently. Lecture Notes in Computer Science, 497:149–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rival:1991:LEB

[149] I. Rival and S. Zhan. Linear extensions with backtracking. Lecture Notes in Computer Science, 497:161–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1991:SSP

[150] J. Chen and Y. Han. Storing shortest paths for a polyhedron. Lecture Notes in Computer Science, 497:169– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1991:IJD

[151] K. P. Tan and T. C. Tan. An improved join dependency for efficient constraint checking. Lecture Notes in Computer Science, 497:182–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1991:PQI

[152] J. Han. On the power of queryindependent compilation. Lecture Notes in Computer Science, 497:185– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toptsis:1991:PTC

[153] A. A. Toptsis. Parallel transitive closure computation in highly scalable multiprocessors. Lecture Notes in Computer Science, 497:197–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czejdo:1991:MTI

[154] B. Czejdo and M. Taylor. Methods and tools for integrating database systems and object-oriented programming systems. Lecture Notes in Computer Science, 497:207–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1991:IDD

[155] K. Wang and L. Y. Yuan. Incremental database design revisited. Lecture Notes in Computer Science, 497:219– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1991:PDR

[156] D. L. Lee and W. Y. P. Wong. Partial document ranking by heuristic methods. Lecture Notes in Computer Science, 497:231–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1991:CIC

[157] L. Yu and S. L. Osborn. Context inheritance and content inheritance in an object-oriented data model. Lecture Notes in Computer Science, 497:240– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czejdo:1991:GSR

[158] B. D. Czejdo, R. P. Tucci, and D. W. Embley. Graphical specification of recursive queries. Lecture Notes in Computer Science, 497:252–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakanishi:1991:DRD

[159] M. Nakanishi, Y. Katsuyama, M. Ito, and A. Hashimoto. On designing relational database schemes for efficient query processing. Lecture Notes in Computer Science, 497:264–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Toweissy:1991:MVA

[160] M. El-Toweissy, N. El-Makky, M. Abougabal, and S. Fouad. The mean value approach to performance evaluation of time-stamp ordering algorithms. Lecture Notes in Computer Science, 497: 276–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuwono:1991:BTR

[161] B. Yuwono and D. L. Lee. A backend text retrieval machine for signaturebased document ranking. Lecture Notes in Computer Science, 497:288– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maheshwari:1991:PSP

[162] P. Maheshwari. Partitioning and scheduling of parallel functional programs using complexity information. Lecture Notes in Computer Science, 497:300–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mittal:1991:SAB

[163] R. Mittal, B. N. Jain, and R. K. Patney. Subtree availability in binary tree architectures. Lecture Notes in Computer Science, 497:312–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1991:OPA

[164] I.-W. Chan and D. K. Friesen. An optimal parallel algorithm for the vertical segment visibility reporting problem. Lecture Notes in Computer Science, 497:323–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chau:1991:FTB

[165] S.-C. Chau and A. L. Liestman. A fault-tolerant binary tree architecture. Lecture Notes in Computer Science, 497:335–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1991:UTC

[166] L.-H. Hsu and J. J. J. Wang. Unordered tree contraction. Lecture Notes in Computer Science, 497:347–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1991:CFS

[167] C.-B. Yang, R. C. T. Lee, and W.-T. Chen. Conflict-free sorting algorithms under single-channel and multi-channel broadcast communication models. Lecture Notes in Computer Science, 497: 350–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qiu:1991:PRS

[168] K. Qiu, H. Meijer, and S. G. Akl. Parallel routing and sorting on the pancake network. Lecture Notes in Computer Science, 497:360–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirsch:1991:PPA

[169] E. S. Kirsch and J. R. S. Blair. Practical parallel algorithms for chordal graphs. Lecture Notes in Computer Science, 497:372–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1991:LTN

[170] L. Chen. Logarithmic time NC algorithms for comparability graphs and circle graphs. Lecture Notes in Computer Science, 497:383–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boals:1991:CHP

[171] A. J. Boals, A. K. Gupta, J. A. Hashmi, and N. A. Sherwani. Compact hypercubes: Properties and recognition. Lecture Notes in Computer Science, 497:395–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1991:ICB

[172] Z. Liu and J.-H. You. Increasing communication bandwidth on hypercube. Lecture Notes in Computer Science, 497:403–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herman:1991:PPS

[173] D. Herman and L. Trilling. Programming a parallel sort using distributed procedure calls. Lecture Notes in Computer Science, 497:415–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krizanc:1991:NLP

[174] D. Krizanc. A note on off-line permutation routing on a mesh-connected processor array. Lecture Notes in Computer Science, 497:418–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myoupo:1991:FPS

[175] J. F. Myoupo. A fully-pipelined solutions constructor for dynamic programming problems. Lecture Notes in Computer Science, 497:421–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumslag:1991:FTQ

[176] M. Baumslag. On the fault-tolerance of quasi-minimal Cayley networks. Lecture Notes in Computer Science, 497: 431–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1991:SFP

[177] S. K. Das, C. C.-Y. Chen, G. Lewis, and S. Prasad. Some fast parallel algorithms for parentheses matching. Lecture Notes in Computer Science, 497:443–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1991:SOP

[178] R. Lin and S. Olariu. A simple optimal parallel algorithm to solve the lowest common ancestor problem. Lecture Notes in Computer Science, 497:455– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biswas:1991:ECS

[179] S. Biswas, A. Sanyal, and L. M. Tewari. Extended cycle shrinking: A restructuring method for parallel compilation. Lecture Notes in Computer Science, 497:462–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosen:1991:SDA

[180] J. B. Rosen and G.-L. Xue. Sequential and distributed algorithms for the all pairs quickest path problem. Lecture Notes in Computer Science, 497: 471–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boals:1991:EAA

[181] A. Boals, A. Gupta, J. Hashmi, and N. Sherwani. An efficient approximation algorithm for hypercube scheduling. Lecture Notes in Computer Science, 497:474–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olariu:1991:OPA

[182] S. Olariu, M. Overstreet, and Z. Wen. An optimal parallel algorithm to reconstruct a binary tree from its traversals. Lecture Notes in Computer Science, 497:484–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Youn:1991:EMA

[183] H. Y. Youn. An efficient multiprocessor architecture for image processing in VLSI. Lecture Notes in Computer Science, 497:496–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srimani:1991:GFT

[184] P. K. Srimani. Generalized fault tolerance properties of star graphs. Lecture Notes in Computer Science, 497: 510–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gopinath:1991:FDI

[185] B. Gopinath, A. K. Singh, and P. Uppaluru. A formal description of the ICsup * model parallel computation. Lecture Notes in Computer Science, 497: 520–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rana:1991:EAQ

[186] S. P. Rana. An efficient algorithm for quiescence detection in a distributed system. Lecture Notes in Computer Science, 497:532–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woodside:1991:IAC

[187] C. M. Woodside, S. Majumdar, and J. E. Neilson. Interval arithmetic for computing performance guarantees in client-server software. Lecture Notes in Computer Science, 497:535–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alijani:1991:ERS

[188] G. S. Alijani and H. F. Wedde. Enhanced reliability in scheduling critical tasks for hard real-time distributed systems. Lecture Notes in Computer Science, 497:547–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ammar:1991:DME

[189] R. Ammar and P. Zhang. A design and modeling environment to develop real-time, distributed software systems. Lecture Notes in Computer Science, 497:559–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1991:MRP

[190] W. Zhou and B. Molinari. On the management of remote procedure call transactions. Lecture Notes in Computer Science, 497:571–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirousis:1991:SAS

[191] L. M. Kirousis, P. Spirakis, and P. Tsigas. Simple atomic snapshots: A linear complexity solution with unbounded time-stamps. Lecture Notes in Computer Science, 497:582–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ZeinElDine:1991:DSA

[192] O. ZeinElDine, M. El-Toweissy, and R. Mukkamala. A distributed scheduling algorithm for heterogeneous realtime systems. Lecture Notes in Computer Science, 497:588–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raynal:1991:DSR

[193] M. Raynal. A distributed solution to the k-out of-M resources allocation problem. Lecture Notes in Computer Science, 497:599–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wedde:1991:EDR

[194] H. F. Wedde, D. C. Daniels, and D. Huizinga. Efficient distributed resource scheduling for adaptive realtime operation support. Lecture Notes in Computer Science, 497:610– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhattacharya:1991:MFT

[195] U. Bhattacharya and S. Bhattacharya. Mapping of fault-tolerant permutations in Omega interconnection network. Lecture Notes in Computer Science, 497:628–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulsari:1991:ANN

[196] A. B. Bulsari and H. Saxen. Artificial neural networks for predicting silicon content in raw iron from blast furnaces. Lecture Notes in Computer Science, 497:642–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulsari:1991:CRS

[197] A. B. Bulsari, B. Saxen, and H. Saxen. A chemical reactor selection expert system created by training an artificial neural network. Lecture Notes in Computer Science, 497:645–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hewett:1991:MTG

[198] R. Hewett. A modeling technique for generating causal explanations of physical systems. Lecture Notes in Computer Science, 497:657–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1991:PCN

[199] L. Yang and S. Szpakowicz. Planning in conceptual networks. Lecture Notes in Computer Science, 497:669–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerrito:1991:ESM

[200] P. Cerrito, W. Karwowski, and Krzysztof Ostaszewski. Empirical study of the meaning of the hedge “VERY”. Lecture Notes in Computer Science, 497:672–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdallah:1991:KKV

[201] M. A. N. Abdallah. Kernel knowledge versus belt knowledge in default reasoning: a logical approach. Lecture Notes in Computer Science, 497:675– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberbach:1991:NNP

[202] E. Eberbach. Neural network processing elements as a news generation of “flip-flops”. Lecture Notes in Computer Science, 497:687–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1991:PTF

[203] L. Y. Yuan, J.-H. You, and C. Wang. A proof-theoretic framework for nonmonotonic reasoning and logic programming. Lecture Notes in Computer Science, 497:699–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1991:SSF

[204] L. Y. Yuan. On semantics, syntactics and fixpoints of general programs. Lecture Notes in Computer Science, 497:709–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basu:1991:FVR

[205] A. Basu and X. Li. A framework for variable-resolution vision. Lecture Notes in Computer Science, 497:721– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sundararajan:1991:RTR

[206] V. Sundararajan, H. N. Dholakia, and N. Parameswaran. An RMS for temporal reasoning with abstraction. Lecture Notes in Computer Science, 497: 733–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wille:1991:IIS

[207] G. Wille. The information industry scenario in year 2000, trends and change of focus, for the IT professional and for the end-user. Lecture Notes in Computer Science, 498:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chung:1991:RUN

[208] L. Chung. Representation and utilization of non-functional requirements for information system design. Lecture Notes in Computer Science, 498:5–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cauvet:1991:AES

[209] C. Cauvet, C. Proix, and C. Rolland. ALECSI: An expert system for requirements engineering. Lecture Notes in Computer Science, 498:31–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fugini:1991:RRT

[210] M. Fugini, M. Guggino, and B. Pernici. Reusing requirements through a modeling and composition support tool. Lecture Notes in Computer Science, 498:50–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahlsen:1991:KKB

[211] M. Ahlsen, A. D’Atri, P. Johannesson, E. Laenens, N. Leone, P. Rullo, P. Rossi, F. Staes, L. Tarantino, L. Van Beirendonck, et al. The KIWIS knowledge base management system. Lecture Notes in Computer Science, 498:79–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlowska:1991:FMD

[212] M. E. Orlowska and K. G. Jeffery. A functional method of data processing based on relational algebra. Lecture Notes in Computer Science, 498:98–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiederhold:1991:DGT

[213] G. Wiederhold, S. Jajodia, and W. Litwin. Dealing with granularity of time in temporal databases. Lecture Notes in Computer Science, 498:

124–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olaisen:1991:ISR

[214] J. Olaisen. Information as a strategic resource: A question of communication. Lecture Notes in Computer Science, 498:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolander:1991:MFG

[215] K. Smolander, K. Lyytinen, V.-P. Tahvanainen, and P. Marttiin. MetaEdit — A flexible graphical environment for methodology modelling. Lecture Notes in Computer Science, 498:168– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gulla:1991:PIC

[216] J. A. Gulla, O. I. Lindland, and G. Willumsen. PPP — an integrated CASE environment. Lecture Notes in Computer Science, 498:194–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seppanen:1991:STC

[217] V. Seppanen, M. Heikkinen, and R. Lintulampi. SPADE — towards CASE tools that can guide design. Lecture Notes in Computer Science, 498: 222–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bitton:1991:DET

[218] D. Bitton, J. C. Millman, and S. Torgersen. DBE: An expert tool for database design. Lecture Notes in Computer Science, 498:240–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adreit:1991:EGE

[219] F. Adreit and M. Bonjour. Ecrins-Design — A graphical editor for semantic structures. Lecture Notes in Computer Science, 498:264–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hillston:1991:CSU

[220] J. Hillston, A. L. Opdahl, and R. Pooley. A case study using the IMSE experimentation tool. Lecture Notes in Computer Science, 498:284–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McBrien:1991:RLC

[221] P. McBrien, M. Niezette, D. Pantazis, A. H. Seltveit, U. Sundin, B. Theodoulidis, G. Tziallas, and R. Wohed. A rule language to capture and model business policy specifications. Lecture Notes in Computer Science, 498:307–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krogstie:1991:ISD

[222] J. Krogstie, P. McBrien, R. Owens, and A. H. Seltveit. Information systems development using a combination of process and rule based approaches. Lecture Notes in Computer Science, 498:319–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarzabek:1991:TAS

[223] S. Jarzabek and K. Tham. Towards automating software maintenance. Lecture Notes in Computer Science, 498: 336–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1991:SPP

[224] C. Liu. Software process planning and execution: Coupling vs. integration. Lecture Notes in Computer Science, 498:356–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1991:CTW

[225] R. Conradi and C. C. Malm. Cooperating transactions and workspaces in EPOS: Design and preliminary implementation. Lecture Notes in Computer Science, 498:375–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalman:1991:ICA

[226] K. Kalman. Implementation and critique of an algorithm which maps a relational database to a conceptual model. Lecture Notes in Computer Science, 498:393–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wangler:1991:AA

[227] B. Wangler. Attribute abstraction. Lecture Notes in Computer Science, 498: 416–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1991:ESM

[228] J. L. G. Dietz. The essential system model. Lecture Notes in Computer Science, 498:441–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papahristos:1991:FCE

[229] S. Papahristos and W. A. Gray. Federated CASE environment. Lecture Notes in Computer Science, 498:461– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kieback:1991:OIS

[230] A. Kieback and J. Niemeier. An organizational information systems enhancement advice unit (OISEAU). Lecture Notes in Computer Science, 498:479– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verhoef:1991:SMK

[231] T. F. Verhoef, A. H. M. ter Hofstede, and G. M. Wijers. Structuring modelling knowledge for CASE shells. Lecture Notes in Computer Science, 498:502–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuwongse:1991:OOA

[232] V. Wuwongse and J. Ma. An objectoriented approach to model management. Lecture Notes in Computer Science, 498:525–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oelmann:1991:RSS

[233] A. Oelmann. Representing a system specification with a temporal dimension in an object-oriented language. Lecture Notes in Computer Science, 498:540–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeTroyer:1991:OBR

[234] O. De Troyer. The OO-Binary relationship model: A truly object oriented conceptual model. Lecture Notes in Computer Science, 498:561–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1991:OIN

[235] B. Nilsson. Organisation’s IT needs in the next decade — the Swedish vision. Lecture Notes in Computer Science, 498:579–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curtis:1991:MTS

[236] M. C. Curtis and V. J. D. Sivess. Modelling the temporal summation of neural membranes using the Ada language. Lecture Notes in Computer Science, 499:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elrad:1991:UAR

[237] T. Elrad and V. Winans. The use of Ada in reactive systems: A 3dimensional model. Lecture Notes in Computer Science, 499:20–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maymir-Ducharme:1991:ART

[238] F. A. Maymir-Ducharme. Ada 9X realtime scheduling alternatives. Lecture Notes in Computer Science, 499:37–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjornson:1991:IES

[239] S. Bjornson. Information engineering for systems in Ada. Lecture Notes in Computer Science, 499:50–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsalidis:1991:AMT

[240] C. T. Tsalidis and A. E. Hatzimanikatis. AFADA, a measurement tool for Ada. Lecture Notes in Computer Science, 499:65–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curley:1991:ABA

[241] T. Curley. An approach to benchmarking Ada compilation systems. Lecture Notes in Computer Science, 499:87–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonGudenberg:1991:MST

[242] J. Wolff von Gudenberg. Modelling SIMD — type parallel arithmetic operations in Ada. Lecture Notes in Computer Science, 499:110–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1991:AAT

[243] J. Cheng and K. Ushijima. Analyzing Ada tasking deadlocks and livelocks using extended Petri nets. Lecture Notes in Computer Science, 499:

125–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lundberg:1991:CHP

[244] L. Lundberg. A coprocessor for high performance multiprocessor Ada tasking. Lecture Notes in Computer Science, 499:147–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutzmann:1991:SPA

[245] K. M. Gutzmann, D. L. Remkes, and J. L. Ragsdale. Software project activity network for managing the development and testing process. Lecture Notes in Computer Science, 499:168– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welz:1991:ASC

[246] A. Welz. Ada in safety critical applications. Lecture Notes in Computer Science, 499:193–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willain:1991:RTE

[247] J. Willain, P. Baesens, J. Sanzot, and S. van Vlierberghe. Real-time electronic funds transfer and Ada. Lecture Notes in Computer Science, 499: 199–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larre:1991:IHS

[248] J. R. Larre, N. Alfaro, J. J. Galan, and J. Garbajosa. Introducing HOOD into software process modelling based environments. Lecture Notes in Computer Science, 499:212–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delis:1991:DAA

[249] A. Delis and W. M. Thomas. Design assessment of Ada systems using static analysis. Lecture Notes in Computer Science, 499:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soupos:1991:UAI

[250] P. Soupos and S. Goutas. Using Ada to implement the transaction mechanism of a distributed object-oriented DBMS. Lecture Notes in Computer Science, 499:252–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgarten:1991:DSA

[251] U. Baumgarten. Distributed systems and Ada — current projects and approaches comparative study’s results. Lecture Notes in Computer Science, 499:260–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1991:RAD

[252] D. C. Levy, M. C. Randelhoff, and J. L. Tokar. Reconfigurable Ada distributed control system software. Lecture Notes in Computer Science, 499: 279–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badaro:1991:RAM

[253] N. Badaro and Th. Moineau. ROSE-Ada: a method and a tool to help reuse of Ada codes. Lecture Notes in Computer Science, 499:304–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Syms:1991:SRC

[254] T. Syms and C. L. Braun. Software reuse: Customer vs. contractor point-counterpoint. Lecture Notes in Computer Science, 499:326–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosskopf:1991:RIO

[255] A. Rosskopf. Reusable input/output packages for Ada avionic applications. Lecture Notes in Computer Science, 499:338–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1991:CPL

[256] J. L. Fernandez and J. A. de la Puente. Constructing a pilot library of components for avionic systems. Lecture Notes in Computer Science, 499:362– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guaspari:1991:FSL

[257] D. Guaspari. Formally specifying the logic of an automatic guidance controller. Lecture Notes in Computer Science, 499:372–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colombini:1991:AHL

[258] C. Colombini, A. della Torre, and R. Palotti. Ada as high level language for real-time systems exploiting RDBMS techniques. Lecture Notes in Computer Science, 499:384–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilbert:1991:EDT

[259] B. Gilbert, M. Taylor, and G. Bek. Experience developing two Ada applications for embedded real-time systems using different software processes. Lecture Notes in Computer Science, 499: 394–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lippens:1991:TIL

[260] J.-M. Lippens and J. Arnol. Testing and integrating a large embedded real-time system. Lecture Notes in Computer Science, 499:404–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Held:1991:CGP

[261] M. Held. On the computational geometry of pocket machining. Lecture Notes in Computer Science, 500:3–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1991:ASSa

[262] M. Bidoit, H.-J. Kreowski, P. Lescanne, F. Orejas, and D. Sannella. Algebraic system specification and development: A survey and annotated bibliography. Lecture Notes in Computer Science, 501:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1991:ASSb

[263] M. Bidoit, H.-J. Kreowski, P. Lescanne, F. Orejas, and D. Sannella. Algebraic system specification and development: Basic foundations. Lecture Notes in Computer Science, 501:7–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1991:ASSc

[264] M. Bidoit, H.-J. Kreowski, P. Lescanne, F. Orejas, and D. Sannella. Algebraic system specification and development models. Lecture Notes in Computer Science, 501:15–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1991:ASSd

[265] M. Bidoit, H.-J. Kreowski, P. Lescanne, F. Orejas, and D. Sannella. Algebraic system specification and development concepts. Lecture Notes in Computer Science, 501:25–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1991:ASSe

[266] M. Bidoit, H.-J. Kreowski, P. Lescanne, F. Orejas, and D. Sannella. Algebraic system specification and development: Support tools. Lecture Notes in Computer Science, 501:35–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1991:ASSf

[267] M. Bidoit, H.-J. Kreowski, P. Lescanne, F. Orejas, and D. Sannella. Algebraic system specification and development: Applications. Lecture Notes in Computer Science, 501:49–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mints:1991:PSP

[268] G. Mints and E. Tyugu. The programming system PRIZ. Lecture Notes in Computer Science, 502:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mints:1991:TTS

[269] G. Mints, J. M. Smith, and E. Tyugu. Type-theoretical semantics of some declarative languages. Lecture Notes in Computer Science, 502:18–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tammet:1991:URD

[270] T. Tammet. Using resolution for deciding solvable classes and building finite models. Lecture Notes in Computer Science, 502:33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meriste:1991:TKB

[271] M. Meriste and J. Penjam. Toward knowledge-based specifications of languages. Lecture Notes in Computer Science, 502:65–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1991:IIRa

[272] R. Freivalds. Inductive inference of recursive functions: Qualitative theory. Lecture Notes in Computer Science, 502:77–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1991:IIRb

[273] R. Freivalds, J. Barzdins, and K. Podnieks. Inductive inference of recursive functions: Complexity bounds. Lecture Notes in Computer Science, 502: 111–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazma:1991:ISD

[274] A. Brazma. Inductive synthesis of dot expressions. Lecture Notes in Computer Science, 502:156–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinber:1991:SMI

[275] E. Kinber. Some models of inductive syntactical synthesis from sample computations. Lecture Notes in Computer Science, 502:213–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barzdins:1991:IST

[276] G. Barzdins. Inductive synthesis of term rewriting systems. Lecture Notes in Computer Science, 502:253– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auzins:1991:ACT

[277] A. Auzins, J. Barzdins, J. Bicevskis, K. Cerans, and A. Kalnins. Automatic construction of test sets: Theoretical approach. Lecture Notes in Computer Science, 502:286–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borzovs:1991:ACT

[278] J. Borzovs, A. Kalnins, and I. Medvedis. Automatic construction of test sets: Practical approach. Lecture Notes in Computer Science, 502:360– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pranevitchius:1991:AAS

[279] H. Pranevitchius. Aggregate approach for specification, validation, simulation and implementation of computer network protocols. Lecture Notes in Computer Science, 502:433–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gecas:1991:CPS

[280] K. Gecas. A compositional proof system for distributed programs. Lecture Notes in Computer Science, 502:478– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1991:IFC

[281] R. Pliuskevicius. Investigation of finitary calculus for a discrete linear time logic by means of infinitary calculus. Lecture Notes in Computer Science, 502:504–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auguston:1991:RPL

[282] M. Auguston. RIGAL — a programming language for compiler writing. Lecture Notes in Computer Science, 502:529–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1991:CPV

[283] R. Freivalds. Complexity of probabilistic versus deterministic automata. Lecture Notes in Computer Science, 502: 565–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bal:1991:LPP

[284] Henri Bal. Languages for parallel programming. Lecture Notes in Computer Science, 503:3–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bratbergsengen:1991:RAO

[285] Kjell Bratbergsengen. Relational algebra operations. Lecture Notes in Computer Science, 503:24–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cacace:1991:OPS

[286] Filippo Cacace, Stefano Ceri, and Maurice Houtsma. An overview of parallel strategies for transitive closure on algebraic machines. Lecture Notes in Computer Science, 503:44–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia-Molina:1991:DMM

[287] Hector Garcia-Molina, Robert Abbott, Christopher Clifton, Carl Staelin, and Kenneth Salem. Data management with massive memory: A summary. Lecture Notes in Computer Science, 503:63–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harder:1991:AID

[288] Theo Harder. An approach to implement dynamically defined complex objects. Lecture Notes in Computer Science, 503:71–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hey:1991:EMM

[289] Anthony Hey. Experience with MIMD message-passing systems: Towards general purpose parallel computing. Lecture Notes in Computer Science, 503:99–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavington:1991:PAP

[290] Simon Lavington. Parallel associative processing for knowledge bases. Lecture Notes in Computer Science, 503: 112–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1991:PEJ

[291] Suresh Patel. Performance estimates of a join. Lecture Notes in Computer Science, 503:124–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1991:EPR

[292] Paul Watson and Paul Townsend. The EDS parallel relational database system. Lecture Notes in Computer Science, 503:149–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apers:1991:PPE

[293] Peter Apers, Bob Hertzberger, Ben Hulshof, Hans Oerlemans, and Martin Kersten. PRISMA, a platform for experiments with parallelism. Lecture Notes in Computer Science, 503:169– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apers:1991:PE

[294] Peter Apers, Bob Hertzberger, and Ben Hulshof. The PRISMA experiment. Lecture Notes in Computer Science, 503:181–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beemster:1991:BEA

[295] Marcel Beemster. Back-end aspects of a portable POOL-X implementation. Lecture Notes in Computer Science, 503:193–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vandenBerg:1991:LRP

[296] Carel van den Berg and Martin Kersten. Logging and recovery in PRISMA. Lecture Notes in Computer Science, 503:229–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grefen:1991:PHI

[297] Paul Grefen, Jan Flokstra, and Peter Apers. Parallel handling of integrity constraints. Lecture Notes in Computer Science, 503:242–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kersten:1991:PAD

[298] Martin Kersten, Satish Shair-Ali, and Carel van den Berg. Performance analysis of a dynamic query processing scheme. Lecture Notes in Computer Science, 503:259–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1991:ECA

[299] Henk Muller. Evaluation of a communication architecture by means of simulation. Lecture Notes in Computer Science, 503:275–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spee:1991:PIE

[300] Jan Willem Spee. A parallel implementation of the EQUIP expert system. Lecture Notes in Computer Science, 503:294–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderSpek:1991:PXI

[301] Juul van der Spek. POOL-X and its implementation. Lecture Notes in Computer Science, 503:309–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sun:1991:ARP

[302] Chengzheng Sun and Bob Hertzberger. Automatic recovery of a parallel stable file system. Lecture Notes in Computer Science, 503:345–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verbrugge:1991:DTN

[303] Nanda Verbrugge. Duplicates and translation of nested SQL queries into XRA. Lecture Notes in Computer Science, 503:354–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlot:1991:PA

[304] Marnix Vlot. The POOMA architecture. Lecture Notes in Computer Science, 503:365–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wester:1991:POS

[305] Rogier Wester and Ben Hulshof. The POOMA operating system. Lecture Notes in Computer Science, 503:396– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilschut:1991:PQE

[306] Annita Wilschut, Peter Apers, and Jan Flokstra. Parallel query execution in PRISMA/DB. Lecture Notes in Computer Science, 503:424–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manthey:1991:DLP

[307] R. Manthey. Declarative languages — paradigm of the past or challenge of the future? Lecture Notes in Computer Science, 504:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ait-Kaci:1991:GP

[308] H. Ait-Kaci. A glimpse of paradise. Lecture Notes in Computer Science, 504:17–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formica:1991:OOD

[309] A. Formica, M. Missikoff, and S. Vazzana. An object-oriented data model for artificial intelligence applications. Lecture Notes in Computer Science, 504:26–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ait-Kaci:1991:OL

[310] H. Ait-Kaci. An overview of LIFE. Lecture Notes in Computer Science, 504: 42–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hull:1991:SRD

[311] R. Hull and D. Jacobs. On the semantics of rules in database programming languages. Lecture Notes in Computer Science, 504:59–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lellahi:1991:TCD

[312] S. K. Lellahi and N. Spyratos. Towards a categorical data model supporting structured objects and inheritance. Lecture Notes in Computer Science, 504:86–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthes:1991:TSO

[313] F. Matthes, A. Ohori, and J. W. Schmidt. Typing schemes for objects with locality. Lecture Notes in Computer Science, 504:106–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceri:1991:OOL

[314] S. Ceri, F. Cacace, and L. Tanca. Object orientation and logic programming for databases: A season’s flirt or longterm marriage? Lecture Notes in Computer Science, 504:124–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottlob:1991:SOO

[315] G. Gottlob, G. Kappel, and M. Schrefl. Semantics of object-oriented data models — the evolving algebra approach. Lecture Notes in Computer Science, 504:144–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thalheim:1991:EER

[316] B. Thalheim. Extending the entityrelationship model for a high-level, theory-based database design. Lecture Notes in Computer Science, 504:161– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthes:1991:TDA

[317] F. Matthes and J. W. Schmidt. Towards database application systems: Types, kinds and other open invitations. Lecture Notes in Computer Science, 504:185–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zlatuska:1991:MIS

[318] J. Zlatuska. Modelling inheritance in a strongly typed functional model. Lecture Notes in Computer Science, 504: 212–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkinson:1991:BTL

[319] M. Atkinson, P. Richard, and P. Trinder. Bulk types for large scale programming. Lecture Notes in Computer Science, 504:228–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pirotte:1991:CGI

[320] A. Pirotte, D. Roelants, and E. Zimanyi. Controlled generation of intensional answers. Lecture Notes in Computer Science, 504:251–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eder:1991:DCR

[321] J. Eder, A. Rudloff, F. Matthes, and J. W. Schmidt. Data construction with recursive set expressions. Lecture Notes in Computer Science, 504:271– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atzeni:1991:LML

[322] P. Atzeni and L. Tanca. The LOGI-DATA+ model and language. Lecture Notes in Computer Science, 504:294– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stemple:1991:RBD

[323] D. Stemple and T. Sheard. A recursive base for database programming primitives. Lecture Notes in Computer Science, 504:311–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alagic:1991:TMD

[324] S. Alagic. Toward multiparadigm database interfaces. Lecture Notes in Computer Science, 504:333–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atzeni:1991:UDD

[325] P. Atzeni and R. Torlone. Updating datalog databases. Lecture Notes in Computer Science, 504:347–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalinichenko:1991:STQ

[326] L. Kalinichenko. SYNTHESIS: Towards a query facility for generalized information resources. Lecture Notes in Computer Science, 504:363–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zamulin:1991:DPL

[327] A. V. Zamulin, V. A. Kositov, G. A. Kucherov, E. V. Pak, and V. D. Ryshkov. The database programming language atlant: Principal features and implementation techniques. Lecture Notes in Computer Science, 504:385– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subieta:1991:LOO

[328] K. Subieta. LOQIS: The objectoriented database programming system. Lecture Notes in Computer Science, 504:403–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burdonov:1991:CPT

[329] I. Burdonov, V. Ivannikov, G. Kopytov, A. Kosachev, and S. Kuznetsov. The CLOS project: Towards an objectoriented environment for application development. Lecture Notes in Computer Science, 504:422–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirkova:1991:HPI

[330] R. Kirkova, M. Ikonomova, and E. Dencheva. The HYTECH project: IP tools based on hypertext technology. Lecture Notes in Computer Science, 504:428–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolinko:1991:VVI

[331] V. Kolinko, A. Mendelev, and S. Shevchenko. VISICOM: A visual interface to databases. Lecture Notes in Computer Science, 504:435–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Redko:1991:APA

[332] V. N. Red’ko, I. A. Basarab, and A. I. Solodovnic. The ADIS project on ADaptive Information Systems: Concepts and language technology. Lecture Notes in Computer Science, 504: 437–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zinovjev:1991:IIS

[333] E. Zinovjev, A. Strecalov, and O. Rogova. ISTSPN: Integrated software tool supporting Petri nets. Lecture Notes in Computer Science, 504:440–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zubenko:1991:FPT

[334] V. Zubenko and A. Protasov. The FOLI project: Towards an artificial intelligence language. Lecture Notes in Computer Science, 504:441–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chereshkin:1991:IRS

[335] D. Chereshkin and M. Tsalenko. Informatization and restructuring of Soviet society. Lecture Notes in Computer Science, 504:443–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnsson:1991:PEF

[336] Th. Johnsson. Parallel evaluation of functional programs: The ¡v, G¿- machine approach. Lecture Notes in Computer Science, 505:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

delaTorre:1991:TSM

[337] Pilar de la Torre and C. P. Kruskal. Towards a single model of efficient computation in real parallel machines. Lecture Notes in Computer Science, 505:6–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treleaven:1991:NCG

[338] P. C. Treleaven. Neural computing and the GALATEA project. Lecture Notes in Computer Science, 505:25–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hum:1991:NHS

[339] H. H. J. Hum and G. R. Gao. A novel high-speed memory organization for fine-grain multi-thread computing. Lecture Notes in Computer Science, 505:34–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langendoen:1991:EFH

[340] K. G. Langendoen, H. L. Muller, and L. O. Hertzberger. Evaluation of Futurebus hierarchical caching. Lecture Notes in Computer Science, 505:52–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filloque:1991:EGC

[341] J. M. Filloque, E. Gautrin, and B. Pottier. Efficient global computations on a processor network with programmable logic. Lecture Notes in Computer Science, 505:69–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogvorst:1991:PHD

[342] P. Hoogvorst, R. Keryell, Ph. Matherat, and N. Paris. POMP or how to design a massively parallel machine with small developments. Lecture Notes in Computer Science, 505:83–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasell:1991:FPA

[343] Jonas Vasell and Jesper Vasell. The function processor: An architecture for efficient execution of recursive functions. Lecture Notes in Computer Science, 505:101–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milikowski:1991:GLD

[344] R. Milikowski and W. G. Vree. The Gline: A distributed processor for graph reduction. Lecture Notes in Computer Science, 505:119–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tel:1991:DDT

[345] G. Tel and F. Mattern. The derivation of distributed termination detection algorithms from garbage collection schemes. Lecture Notes in Computer Science, 505:137–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piquer:1991:IRC

[346] J. M. Piquer. Indirect reference counting: A distributed garbage collection algorithm. Lecture Notes in Computer Science, 505:150–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korst:1991:PMS

[347] J. H. M. Korst, E. H. L. Aarts, J. K. Lenstra, and J. Wessels. Periodic multiprocessor scheduling. Lecture Notes in Computer Science, 505:166–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumslag:1991:ESL

[348] M. Baumslag, M. C. Heydemann, J. Opatrny, and D. Sotteau. Embeddings of shuffle-like graphs in hypercubes. Lecture Notes in Computer Science, 505:179–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDongen:1991:MUR

[349] V. Van Dongen. Mapping uniform recurrences onto small size arrays. Lecture Notes in Computer Science, 505: 191–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferianc:1991:ACM

[350] P. Ferianc and O. Sykora. Area complexity of multilective merging. Lecture Notes in Computer Science, 505: 209–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1991:DFE

[351] X. Zhong and S. Rajopadhye. Deriving fully efficient systolic arrays by quasi-linear allocation functions. Lecture Notes in Computer Science, 505: 219–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mongenet:1991:ATS

[352] C. Mongenet. Affine timings for systems of affine recurrence equations. Lecture Notes in Computer Science, 505:236–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parberry:1991:CCO

[353] I. Parberry. On the computational complexity of optimal sorting network verification. Lecture Notes in Computer Science, 505:252–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1991:MPH

[354] Sajal K. Das and Horng Wen-Bing. Managing a parallel heap efficiently. Lecture Notes in Computer Science, 505:270–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alvarez:1991:PCD

[355] C. Alvarez, J. L. Balcazar, J. Gabarro, and M. Santha. Parallel complexity in the design and analysis of concurrent systems. Lecture Notes in Computer Science, 505:288–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1991:FHL

[356] T. Hagerup, A. Schmitt, and H. Seidl. FORK: A high-level language for PRAMs. Lecture Notes in Computer Science, 505:304–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurson:1991:NNB

[357] A. R. Hurson, B. Jin, and S. H. Pakzad. Neural network-based decision making for large incomplete databases. Lecture Notes in Computer Science, 505:321–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Louri:1991:OCA

[358] A. Louri. An optical contentaddressable parallel processor for fast searching and retrieving. Lecture Notes in Computer Science, 505:338– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1991:TEH

[359] G. R. Gao, H. H. J. Hum, and J.- M. Monti. Towards an efficient hybrid dataflow architecture model. Lecture Notes in Computer Science, 505:355– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Govindarajan:1991:DFI

[360] R. Govindarajan and Yu Sheng. Data flow implementation of generalized guarded commands. Lecture Notes in Computer Science, 505:372–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duato:1991:DDF

[361] J. Duato. On the design of deadlockfree adaptive routing algorithms for multicomputers: Design methodologies. Lecture Notes in Computer Science, 505:390–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ilmberger:1991:TDP

[362] H. Ilmberger and S. Thurmel. A toolkit for debugging parallel Lisp programs. Lecture Notes in Computer Science, 505:406–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Misra:1991:LCP

[363] J. Misra. Loosely-coupled processes. Lecture Notes in Computer Science, 506:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBakker:1991:RMS

[364] J. W. de Bakker and E. P. de Vink. Rendezvous with metric semantics. Lecture Notes in Computer Science, 506:27–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janicki:1991:IPC

[365] R. Janicki and M. Koutny. Invariants and paradigms of concurrency theory. Lecture Notes in Computer Science, 506:59–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1991:AAF

[366] L. M. Alonso and R. Pena. Acceptance automata: A framework for specifying and verifying TCSP parallel systems. Lecture Notes in Computer Science, 506:75–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fanchon:1991:MDP

[367] J. Fanchon and D. Millot. Models for dynamically placed concurrent processes. Lecture Notes in Computer Science, 506:92–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sami:1991:FBA

[368] Y. Sami and G. Vidal-Naquet. Formalisation of the behavior of actors by colored Petri nets and some applications. Lecture Notes in Computer Science, 506:110–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1991:PRF

[369] Ambuj K. Singh. Program refinement in fair transition systems. Lecture Notes in Computer Science, 506:128– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yantchev:1991:CAR

[370] J. T. Yantchev. Communication abstraction and refinement. Lecture Notes in Computer Science, 506:148– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouge:1991:SLM

[371] L. Bouge. On the semantics of languages for massively parallel SIMD architectures. Lecture Notes in Computer Science, 506:166–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooman:1991:DRT

[372] J. Hooman. A denotational real-time semantics for shared processors. Lecture Notes in Computer Science, 506: 184–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nocker:1991:CC

[373] E. G. J. M. H. Nocker, J. E. W. Smetsers, M. C. J. D. van Eekelen, and M. J. Plasmeijer. Concurrent clean. Lecture Notes in Computer Science, 506:202–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanDelft:1991:SPL

[374] A. van Delft. The scriptic programming language. Lecture Notes in Computer Science, 506:220–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haridi:1991:SOS

[375] S. Haridi and C. Palamidessi. Structural operational semantics for kernel Andorra Prolog. Lecture Notes in Computer Science, 506:238–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagannathan:1991:CFC

[376] S. Jagannathan. Customization of firstclass tuple-spaces in a higher-order language. Lecture Notes in Computer Science, 506:254–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Factor:1991:FSP

[377] M. Factor. A formal specification of the process trellis. Lecture Notes in Computer Science, 506:277–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Autant:1991:SBN

[378] C. Autant, Z. Belmesk, and Ph. Schnoebelen. Strong bisimilarity on nets revisited. Lecture Notes in Computer Science, 506:295–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magee:1991:CAP

[379] J. Magee and N. Dulay. A configuration approach to parallel programming. Lecture Notes in Computer Science, 506:313–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaudiot:1991:CLS

[380] J.-L. Gaudiot and Lin Chih-Ming. Chaotic linear system solvers in a variable-grain data-driven multiprocessor system. Lecture Notes in Computer Science, 506:331–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waite:1991:PAC

[381] M. Waite, B. Giddings, and S. Lavington. Parallel associative combinator evaluation. Lecture Notes in Computer Science, 506:349–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hankin:1991:SAT

[382] C. Hankin. Static analysis of term graph rewriting systems. Lecture Notes in Computer Science, 506:367– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briat:1991:RDM

[383] J. Briat, M. Favre, C. Geyer, and J. Chassin de Kergommeaux. Reconfigurable, distributed-memory multiprocessor. Lecture Notes in Computer Science, 506:385–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaumont:1991:FSP

[384] A. Beaumont, S. Muthu Raman, P. Szeredi, and D. H. D. Warren. Flexible scheduling of OR-Parallelism in aurora: The bristol scheduler. Lecture Notes in Computer Science, 506:403–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veron:1991:VMS

[385] A. Veron, Xu Jiyang, S. A. Delgado-Rannauro, and K. Schuerman. Virtual memory support for OR-Parallel logic programming systems. Lecture Notes in Computer Science, 506:421– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szeredi:1991:IES

[386] P. Szeredi, R. Yang, and M. Carlsson. Interfacing engines and schedulers in OR-Parallel Prolog systems. Lecture Notes in Computer Science, 506:439– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1991:RCS

[387] Zhiyi Hwang, Shouren Hu, Chengzheng Sun, and Yaoqing Gao. Reduction of code space in parallel logic programming systems. Lecture Notes in Computer Science, 506:454–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1991:SLP

[388] S. Fujita, M. Yamashita, and T. Ae. Search level parallel processing of production systems. Lecture Notes in Computer Science, 506:471–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manning:1991:CCG

[389] Joseph Manning. Computational complexity of geometric symmetry detection in graphs. Lecture Notes in Computer Science, 507:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bagga:1991:LDM

[390] Kunwarjit S. Bagga and Frank W. Owens. Low dimensional middle cubes are Hamiltonian. Lecture Notes in Computer Science, 507:8–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kubicka:1991:AAC

[391] Ewa Kubicka, Grzegorz Kubicki, and Dionisios Kountanis. Approximation algorithms for the chromatic sum. Lecture Notes in Computer Science, 507: 15–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sager:1991:CEA

[392] Thomas J. Sager and Lin Shi-Jen. A color-exchange algorithm for exact graph coloring. Lecture Notes in Computer Science, 507:22–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johns:1991:DT

[393] Garry L. Johns and Tai-Chi Lee. Sdistance in trees. Lecture Notes in Computer Science, 507:29–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhugou:1991:ABC

[394] Mo Zhugou and Kenneth Williams. Algorithms on block-complete graphs. Lecture Notes in Computer Science, 507:34–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rubin:1991:HBL

[395] Stuart H. Rubin. Heuristic-based learning. Lecture Notes in Computer Science, 507:41–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinkowski:1991:TBA

[396] Ben Pinkowski. A template-based approach for recognition of intermittent sounds. Lecture Notes in Computer Science, 507:51–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chun-Hung:1991:MMU

[397] Tzeng Chun-Hung. A mathematical model of uncertain information. Lecture Notes in Computer Science, 507: 58–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gold:1991:WPT

[398] Gary Gold. A word to phoneme translator. Lecture Notes in Computer Science, 507:65–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgakis:1991:PEM

[399] D. Christine Georgakis, Martha Evens, Frank Naeymi-Rad, and David A. Trace. Performance evaluation of medical expert systems. Lecture Notes in Computer Science, 507:70–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoon:1991:SCI

[400] Hee Lee Yoon, Martha Evens, Joel A. Michael, and Allen A. Rovick. Spelling correction for an intelligent tutoring system. Lecture Notes in Computer Science, 507:77–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchheit:1991:INN

[401] Paul Buchheit. An informational network for a natural talking system. Lecture Notes in Computer Science, 507: 84–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatia:1991:FCD

[402] Sanjiv K. Bhatia, Jitender S. Deogun, and Vijay V. Raghavan. Formation of categories in document classification systems. Lecture Notes in Computer Science, 507:91–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ongwisesphaiboon:1991:ESC

[403] Vipa Ongwisesphaiboon, Ilene Burnstein, Martha Evens, Anjali Puri, and Martin E. Ginn. An expert system for creativity management. Lecture Notes in Computer Science, 507:98–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowski:1991:MFM

[404] Greg Makowski. A more flexible method for recognizing signals using back propagation: Piecewise linear regression vectors. Lecture Notes in Computer Science, 507:105–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gordon:1991:CAD

[405] Gerald Leonard Gordon. A cooperative algorithm for the dynamic stereo problem. Lecture Notes in Computer Science, 507:111–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haskell:1991:DHC

[406] Richard E. Haskell and Ali Noui-Mehidi. Design of hierarchical classifiers. Lecture Notes in Computer Science, 507:118–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1991:DIP

[407] Roy F. Keller, R. Mark Meyer, and Thomas L. Seevers. Designing and implementing parallel constructs. Lecture Notes in Computer Science, 507:125– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samal:1991:PPC

[408] Ashok Samal. Performance of parallel consistency algorithms. Lecture Notes in Computer Science, 507:132– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aboelaze:1991:DAC

[409] Mokhtar A. Aboelaze and Catherine E. Houstis. Delay analysis of the N-cube network. Lecture Notes in Computer Science, 507:139–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahmed:1991:NCP

[410] Kazem U. Ahmed and Yeh Der-Yun. A note on Conway’s parallel sorting algorithm. Lecture Notes in Computer Science, 507:147–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rover:1991:VPP

[411] Diane T. Rover, G. M. Prabhu, and Charles T. Wright. Visualization of program performance on concurrent computers. Lecture Notes in Computer Science, 507:154–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawes:1991:MON

[412] Robin W. Dawes. Minimum odd neighbourhood covers for trees. Lecture Notes in Computer Science, 507:161– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perl:1991:CLA

[413] Yehoshua Perl and Ashish Mehta. Cascading LZW algorithm with Huffman coding: A variable to variable length compression algorithm. Lecture Notes in Computer Science, 507:170– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McAlpin:1991:DIP

[414] John McAlpin and Christos Nikolopoulos. Dendrograms and irreducible prefix codes. Lecture Notes in Computer Science, 507:179–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wainer:1991:MPI

[415] Michael S. Wainer. Massively parallel implementations of adaptively subdividing fractal generating algorithms with parameter extensions. Lecture Notes in Computer Science, 507:185– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Si-Qing:1991:SPR

[416] Zheng Si-Qing. A simple and powerful representation of binary search trees. Lecture Notes in Computer Science, 507:192–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magliveras:1991:RPL

[417] Spyros S. Magliveras and Nasir D. Memon. Random permutations from logarithmic signatures. Lecture Notes in Computer Science, 507:199–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raja:1991:NDT

[418] Paruvachi V. R. Raja. Novel design techniques for RNS systolic VLSI arrays. Lecture Notes in Computer Science, 507:206–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunawardena:1991:CDT

[419] K. L. D. Gunawardena. A compound decision theory approach to digital signal reconstruction. Lecture Notes in Computer Science, 507:213–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vishnubhotla:1991:RTP

[420] Sarma R. Vishnubhotla. Real-time parameter control in space vehicles: A parallel architecture and the protocols. Lecture Notes in Computer Science, 507:220–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vishnubhotla:1991:MBP

[421] Sarma R. Vishnubhotla. A microinstruction based procedure to diagnose hardware failures in industrial robots. Lecture Notes in Computer Science, 507:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaur:1991:SVT

[422] Harbans Kaur, James Brown Grier, and Graham Campbell. A specification and verification tool for communication protocols. Lecture Notes in Computer Science, 507:234–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richards:1991:SLD

[423] Phil G. Richards and Chao Meng-Chun. Structured on-line directory and file organization. Lecture Notes in Computer Science, 507:242–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fotouhi:1991:GAS

[424] Farshad Fotouhi and Carlos E. Galarce. Genetic algorithms and the search for optimal database index selection. Lecture Notes in Computer Science, 507: 249–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saiedian:1991:OBA

[425] Hossein Saiedian. An object-based approach to the specification of office entities. Lecture Notes in Computer Science, 507:256–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sumali:1991:LAL

[426] Pin-Ngern Sumali, Robert Strutz, and Martha Evens. Lexical acquisition for lexical databases. Lecture Notes in Computer Science, 507:264–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lyness:1991:SGL

[427] James N. Lyness and W. Newman. A search for good lattice rules based on the reciprocal lattice generator matrix. Lecture Notes in Computer Science, 507:271–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Genz:1991:ANI

[428] Alan Genz. An adaptive numerical integration algorithm for simplices. Lecture Notes in Computer Science, 507: 279–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ewerbring:1991:HSV

[429] L. Magnus Ewerbring and Franklin T. Luk. The HK singular value decomposition of rank deficient matrix triplets. Lecture Notes in Computer Science, 507:286–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carino:1991:AIU

[430] Ricolindo Carino, Elise de Doncker, and Ian Robinson. Approximate integration using iterated Levin transformations. Lecture Notes in Computer Science, 507:293–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sprague:1991:RAS

[431] Thomas Sprague. Recent advances in shape preserving piecewise cubic interpolation. Lecture Notes in Computer Science, 507:300–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ancona:1991:SGM

[432] M. Ancona, K. S. Bagga, E. Bruzzone, L. De Floriani, and J. S. Deogun. Structured graph models: An efficient tool for VLSI design. Lecture Notes in Computer Science, 507:307–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saab:1991:SAC

[433] Youssef Saab and Vasant Rao. A stochastic algorithm for circuit bipartitioning. Lecture Notes in Computer Science, 507:313–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parks:1991:OCR

[434] Dee Parks. An optimal channelrouting algorithm for a restricted class of multi-terminal nets. Lecture Notes in Computer Science, 507:322–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewis:1991:CPE

[435] F. D. Lewis and N. Van Cleave. Correct and provably efficient methods for rectilinear Steiner spanning tree generation. Lecture Notes in Computer Science, 507:329–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alijani:1991:TPT

[436] Ghasem S. Alijani and Horst F. Wedde. A three-phase task scheduling scheme in A hard real-time distributed environment. Lecture Notes in Computer Science, 507:336–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramakrishnan:1991:BPF

[437] Sub Ramakrishnan. A broadcast protocol: Functional properties and specification. Lecture Notes in Computer Science, 507:343–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Ruby:1991:LED

[438] Mohamed El-Ruby, James Kenevan, Robert Carlson, and Khalid Khalil. Leader election in distributed computing systems. Lecture Notes in Computer Science, 507:350–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zimmerman:1991:NAS

[439] Guy W. Zimmerman and Abdol-Hossein Esfahanian. A new approach to system-level fault-tolerance in messagepassing MultiComputers. Lecture Notes in Computer Science, 507:357– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaven:1991:FLC

[440] Mike Beaven, Ryan Stansifer, and Dan Wetklow. A functional language with classes. Lecture Notes in Computer Science, 507:364–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapin:1991:UMC

[441] Steve J. Chapin and Ryan Stansifer. Using ML as a command language. Lecture Notes in Computer Science, 507: 371–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huei-Ning:1991:ISG

[442] Natasha Ma Huei-Ning, Martha Evens, David A. Trace, and Frank Naeymi-Rad. An interactive system for generating hospital progress notes. Lecture Notes in Computer Science, 507: 378–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rizki:1991:BEP

[443] Mateen M. Rizki, Louis A. Tamburino, and Michael A. Zmuda. Biological evolution as a paradigm for performance driven design processes. Lecture Notes in Computer Science, 507: 385–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elrod:1991:PCR

[444] David W. Elrod, Gerry M. Maggiora, and Robert G. Trenary. Predicting chemical reactions with a neural network. Lecture Notes in Computer Science, 507:392–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rizki:1991:MRE

[445] Mateen M. Rizki and Chen Jong. Mutation and recombination effects on the adaptability of sexual and asexual organisms. Lecture Notes in Computer Science, 507:399–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coleman:1991:LSN

[446] William P. Coleman, David P. Sanford, Andrea De Gaetano, and Fred Geisler. Logical structure of neurobiological information processing. Lecture Notes in Computer Science, 507:406–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogel:1991:DES

[447] Gretchen Vogel, Diane Rover, Charles Wright, and Gurpur Prabhu. Development of an experimental setup for studying parallel processing. Lecture Notes in Computer Science, 507:412– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yim:1991:CCS

[448] Roger Yim and Hu Gongzhu. Compiler course with a semi-challenging project. Lecture Notes in Computer Science, 507:419–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swafford:1991:RPB

[449] Bill E. Swafford. Report on a PC based system designed to enhance the teaching of IBM 370 assembly language. Lecture Notes in Computer Science, 507:426–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanders:1991:TSC

[450] Dean Sanders. Teaching the spirit of computational science. Lecture Notes in Computer Science, 507:433– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1991:CRW

[451] G. D. Cohen. Covering radius and writing on memories. Lecture Notes in Computer Science, 508:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heintz:1991:GPS

[452] J. Heintz, T. Krick, M.-F. Roy, and P. Solerno. Geometric problems solvable in single exponential time. Lecture Notes in Computer Science, 508:11–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simonis:1991:DC

[453] J. Simonis. A description of the [16, 7, 6] codes. Lecture Notes in Computer Science, 508:24–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamaguchi:1991:PSA

[454] K. Yamaguchi and H. Imai. Periodic sequences for absolute type shaft encoders. Lecture Notes in Computer Science, 508:36–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sinaki:1991:ECA

[455] G. Sinaki. Error-coded algorithms for on-line arithmetic. Lecture Notes in Computer Science, 508:46–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saitoh:1991:CCC

[456] Y. Saitoh and H. Imai. Constructions of codes correcting burst asymmetric errors. Lecture Notes in Computer Science, 508:59–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahn:1991:CMA

[457] T. N. Ahn, K. Sakaniwa, and T. R. N. Rao. A construction method for m-ary unidirectional error control codes. Lecture Notes in Computer Science, 508: 71–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peizhong:1991:FCG

[458] L. Peizhong and S. Guowen. Feasible calculation of the generator for combined LFSR sequences. Lecture Notes in Computer Science, 508:86–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vajda:1991:SCA

[459] I. Vajda and T. Nemetz. Substitution of characters in q-ary msequences. Lecture Notes in Computer Science, 508:96–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suehiro:1991:PPO

[460] N. Suehiro. Pseudo-polyphase orthogonal sequence sets with good crosscorrelation property. Lecture Notes in Computer Science, 508:106–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsufuji:1991:RVB

[461] S. Matsufuji and K. Imamura. Realvalued bent function and its application to the design of balanced quadriphase sequences with optimal correlation properties. Lecture Notes in Computer Science, 508:113–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnabl:1991:CMG

[462] G. Schnabl and M. Bossert. Coded modulation with generalized multiple concatenation of block codes. Lecture Notes in Computer Science, 508:122– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohno:1991:TCM

[463] R. Kohno, Ch. Yim, and H. Imai. Trellis coded modulation based on timevarying mapping and encoders for utilizing a channel intersymbol interference. Lecture Notes in Computer Science, 508:132–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikityuk:1991:UAC

[464] N. M. Nikityuk. Use of the algebraic coding theory in nuclear electronics. Lecture Notes in Computer Science, 508:144–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abbott:1991:SIA

[465] J. Abbott. Some ideas about faulttolerant Chinese remaindering. Lecture Notes in Computer Science, 508:155– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coma:1991:CIB

[466] J. R. Coma. On a categorial isomorphism between a class of completely regular codes and a class of distance regular graphs. Lecture Notes in Computer Science, 508:164–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heintz:1991:SEP

[467] J. Heintz, M.-F. Roy, and P. Solerno. Single exponential path finding in semialgebraic sets. part I: The case of a regular bounded hypersurface. Lecture Notes in Computer Science, 508:180–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1991:CAP

[468] M. E. Alonso, T. Mora, and M. Raimondo. On the complexity of algebraic power series. Lecture Notes in Computer Science, 508:197–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1991:LDA

[469] M. E. Alonso, T. Mora, and M. Raimondo. Local decomposition algorithms. Lecture Notes in Computer Science, 508:208–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langemyr:1991:AFP

[470] L. Langemyr. An asymptotically fast probabilistic algorithm for computing polynomial GCD’s over an algebraic number field. Lecture Notes in Computer Science, 508:222–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mourrain:1991:SAF

[471] B. Mourrain. Some algebra with formal matrices. Lecture Notes in Computer Science, 508:234–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalkbrener:1991:IRP

[472] M. Kalkbrener. Implicitization of rational parametric curves and surfaces. Lecture Notes in Computer Science, 508:249–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glesser:1991:IAI

[473] Ph. Glesser and M. Mignotte. An inequality about irreducible factors of integer polynomials (II). Lecture Notes in Computer Science, 508:260–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beneteau:1991:STM

[474] Pr. L. Beneteau. The symplectic trilinear mappings: An algorithmic approach of the classification; case of the field GF(3)^{†}. Lecture Notes in Computer Science, 508:267–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakata:1991:GBM

[475] S. Sakata. A Gröbner basis and a minimal polynomial set of a finite nD array. Lecture Notes in Computer Science, 508:280–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robbiano:1991:BDN

[476] L. Robbiano. Bounds for degrees and number of elements in Gröbner bases. Lecture Notes in Computer Science, 508:292–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ollivier:1991:SBD

[477] F. Ollivier. Standard bases of differential ideals. Lecture Notes in Computer Science, 508:304–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giusti:1991:CSB

[478] M. Giusti. Complexity of standard bases in projective dimension zero II. Lecture Notes in Computer Science, 508:322–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diab:1991:SAM

[479] M. Diab. Systolic architectures for multiplication over finite field GF(2^{m}). Lecture Notes in Computer Science, 508:329–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuechlin:1991:PPS

[480] W. Kuechlin. PARSAC-2: A parallel SAC-2 based on threads. Lecture Notes in Computer Science, 508:341– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morii:1991:EFF

[481] M. Morii and Y. Takamatsu. Exponentiation in finite fields using dual basis multiplier. Lecture Notes in Computer Science, 508:354–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooperman:1991:ACG

[482] G. Cooperman, L. Finkelstein, and N. Sarawagi. Applications of Cayley graphs. Lecture Notes in Computer Science, 508:367–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1991:DBT

[483] Y. Zheng, T. Matsumoto, and H. Imai. Duality between two cryptographic primitives. Lecture Notes in Computer Science, 508:379–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obbink:1991:SEE

[484] Henk Obbink. Systems engineering environments of ATMOSPHERE. Lecture Notes in Computer Science, 509:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gohner:1991:BIC

[485] Peter Gohner. Building IPSE’s by combining heterogeneous CASE tools. Lecture Notes in Computer Science, 509: 18–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merbeth:1991:FAI

[486] Gunter Merbeth. On the functional and architectural integration of CASE systems. Lecture Notes in Computer Science, 509:31–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavi:1991:DMD

[487] Jonah Z. Lavi and Michael Winokur. Development of a method driven CAS^{2}E tool. Lecture Notes in Computer Science, 509:44–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pocock:1991:VRO

[488] John N. Pocock. VSF and its relationship to open systems and standard repositories. Lecture Notes in Computer Science, 509:53–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliver:1991:ACI

[489] Huw Oliver. Adding control integration to PCTE. Lecture Notes in Computer Science, 509:69–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alderson:1991:MCT

[490] Albert Alderson. Meta-CASE technology. Lecture Notes in Computer Science, 509:81–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burchett:1991:MIS

[491] Rainer Burchett. Management issues in software development. Lecture Notes in Computer Science, 509:92–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brohl:1991:SSD

[492] Adolf Peter Brohl. Standardization of software development environments. Lecture Notes in Computer Science, 509:99–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohler:1991:SDE

[493] Lothar Kohler. A software development environment for CIM applications. Lecture Notes in Computer Science, 509:109–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinbauer:1991:ROT

[494] Dieter Steinbauer. A repository and other tools in a commercial development center. Lecture Notes in Computer Science, 509:114–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1991:LLD

[495] Johann Wagner. Lessons learned from DOMINO and GRAPES. Lecture Notes in Computer Science, 509:130– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hruschka:1991:SCA

[496] Peter Hruschka. Should CASE be application dependent? Lecture Notes in Computer Science, 509:143–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henhapl:1991:UFG

[497] Wolfgang Henhapl, Stefan Kaes, and Gregor Snelting. Utilizing fifth generation technology in software development tools. Lecture Notes in Computer Science, 509:153–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clement:1991:ISC

[498] Dominique Clement, Vincent Prunet, and Francis Montagnac. Integrated software components: A paradigm for control integration. Lecture Notes in Computer Science, 509:167–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjorner:1991:FMS

[499] Dines Bjorner and Soren Prehn. Formal methods in software development: Requirements for a CASE. Lecture Notes in Computer Science, 509:178– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagl:1991:MSA

[500] Manfred Nagl. Modeling of software architectures: Importance, notions, experiences. Lecture Notes in Computer Science, 509:211–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1991:CFM

[501] Jeff Kramer and Anthony Finkelstein. A configurable framework for method and tool integration. Lecture Notes in Computer Science, 509:233–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Humphrey:1991:RFS

[502] Watts S. Humphrey. Recent findings in software process maturity. Lecture Notes in Computer Science, 509:258– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruhn:1991:VVS

[503] Volker Gruhn. Validation and verification of software process models. Lecture Notes in Computer Science, 509: 271–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbrielli:1991:SLP

[504] M. Gabbrielli and G. Levi. On the semantics of logic programs (invited lecture). Lecture Notes in Computer Science, 510:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1991:LPR

[505] Chen Hong and Hsiang Jieh. Logic programming with recurrence domains. Lecture Notes in Computer Science, 510:20–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bucciarelli:1991:EES

[506] A. Bucciarelli and T. Ehrhard. Extensional embedding of a strongly stable model of PCF. Lecture Notes in Computer Science, 510:35–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ernoult:1991:UIS

[507] C. Ernoult and A. Mycroft. Uniform ideals and strictness analysis. Lecture Notes in Computer Science, 510:47–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breazu-Tannen:1991:LCA

[508] V. Breazu-Tannen and R. Subrahmanyam. Logical and computational aspects of programming with sets/ bags/ lists. Lecture Notes in Computer Science, 510:60–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1991:SBT

[509] A. Bouajjani, J. C. Fernandez, S. Graf, C. Rodriguez, and J. Sifakis. Safety for branching time semantics. Lecture Notes in Computer Science, 510:76–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fix:1991:PCM

[510] L. Fix, N. Francez, and O. Grumberg. Program composition and modular verification. Lecture Notes in Computer Science, 510:93–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1991:MCP

[511] R. Alur, C. Courcoubetis, and D. Dill. Model-checking for probabilistic realtime systems. Lecture Notes in Computer Science, 510:115–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1991:CBR

[512] R. Cleaveland and B. Steffen. Computing behavioural relations, logically. Lecture Notes in Computer Science, 510:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Asher:1991:PR

[513] Y. Ben-Asher, D. Peleg, R. Ramaswami, and A. Schuster. The power of reconfiguration. Lecture Notes in Computer Science, 510:139–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fouks:1991:GRT

[514] J.-D. Fouks. General resolution of tseitin formulas is hard. Lecture Notes in Computer Science, 510:151– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kannan:1991:PCP

[515] S. Kannan and A. Yao. Program checkers for probability generation. Lecture Notes in Computer Science, 510:163– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaneps:1991:RTR

[516] J. Kaneps and R. Freivalds. Running time to recognize nonregular languages by 2-Way probabilistic automata. Lecture Notes in Computer Science, 510: 174–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casas:1991:SRT

[517] R. Casas, J. Diaz, and C. Martinez. Statistics on random trees (invited lecture). Lecture Notes in Computer Science, 510:186–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1991:EPI

[518] K. G. Larsen. The expressive power of implicit specifications. Lecture Notes in Computer Science, 510:204–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1991:CIM

[519] Yi Wang. CCS+Time = an interleaving model for real time systems. Lecture Notes in Computer Science, 510:217–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1991:CSC

[520] V. Diekert, E. Ochmanski, and K. Reinhardt. On confluent semicommutations-decidability and complexity results. Lecture Notes in Computer Science, 510:229–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duchamp:1991:LFF

[521] G. Duchamp and D. Krob. Lazard’s factorizations of free partially commutative monoids. Lecture Notes in Computer Science, 510:242–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gastin:1991:KTI

[522] P. Gastin, A. Petit, and W. Zielonka. A Kleene theorem for infinite trace languages. Lecture Notes in Computer Science, 510:254–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1991:CSH

[523] N. Dershowitz. Canonical sets of Horn clauses. Lecture Notes in Computer Science, 510:267–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madlener:1991:SCP

[524] K. Madlener, P. Narendran, and F. Otto. A specialized completion procedure for monadic string-rewriting systems presenting groups. Lecture Notes in Computer Science, 510:279– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curien:1991:CRL

[525] P.-L. Curien and R. Di Cosmo. A confluent reduction for the lambdacalculus with surjective pairing and terminal object. Lecture Notes in Computer Science, 510:291–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernando:1991:PRP

[526] T. Fernando. Provably recursive programs and program extraction. Lecture Notes in Computer Science, 510: 303–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lengauer:1991:EAP

[527] T. Lengauer and D. Theune. Efficient algorithms for path problems with general cost criteria. Lecture Notes in Computer Science, 510:314–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1991:CSP

[528] H. N. Djidjev, G. E. Pantziou, and C. D. Zaroliagis. Computing shortest paths and distances in planar graphs. Lecture Notes in Computer Science, 510:327–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galil:1991:MBC

[529] Z. Galil and G. F. Italiano. Maintaining biconnected components of dynamic planar graphs. Lecture Notes in Computer Science, 510:339–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loebl:1991:EMC

[530] M. Loebl. Efficient maximal cubic graph cuts. Lecture Notes in Computer Science, 510:351–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vishkin:1991:SPA

[531] U. Vishkin. Structural parallel algorithmics (invited lecture). Lecture Notes in Computer Science, 510:363– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranjan:1991:IKS

[532] D. Ranjan, S. Chari, and P. Rohatgi. Improving known solutions is hard. Lecture Notes in Computer Science, 510:381–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemachandra:1991:CDS

[533] L. A. Hemachandra and A. Hoene. Collapsing degrees via strong computation. Lecture Notes in Computer Science, 510:393–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1991:FPG

[534] T. Hagerup. Fast parallel generation of random permutations. Lecture Notes in Computer Science, 510:405–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:1991:PAT

[535] H. Jung, M. Serna, and P. Spirakis. A parallel algorithm for two processors precedence constraint scheduling. Lecture Notes in Computer Science, 510: 417–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furer:1991:ENA

[536] M. Furer and B. Raghavachari. An efficient NC algorithm for finding Hamiltonian cycles in dense directed graphs. Lecture Notes in Computer Science, 510:429–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1991:LTA

[537] W. Thomas. On logics, tilings, and automata (invited lecture). Lecture Notes in Computer Science, 510:441– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jouannaud:1991:SSO

[538] J.-P. Jouannaud and M. Okada. Satisfiability of systems of ordinal notations with the subterm property is decidable. Lecture Notes in Computer Science, 510:455–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comon:1991:CAS

[539] H. Comon. Complete axiomatizations of some quotient term algebras. Lecture Notes in Computer Science, 510: 469–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bol:1991:MNP

[540] R. Bol and J. F. Groote. The meaning of negative premises in transition system specifications. Lecture Notes in Computer Science, 510:481–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1991:DHP

[541] W. Vogler. Deciding history preserving bisimilarity. Lecture Notes in Computer Science, 510:495–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1991:AAR

[542] L. Aceto and M. Hennessy. Adding action refinement to a finite process algebra. Lecture Notes in Computer Science, 510:506–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bini:1991:IPC

[543] D. Bini, L. Gemignani, and V. Pan. Improved parallel computations with matrices and polynomials. Lecture Notes in Computer Science, 510:520– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lagergren:1991:FMF

[544] J. Lagergren and S. Arnborg. Finding minimal forbidden minors using a finite congruence. Lecture Notes in Computer Science, 510:532–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1991:BAP

[545] H. L. Bodlaender and T. Kloks. Better algorithms for the pathwidth and treewidth of graphs. Lecture Notes in Computer Science, 510:544–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cucker:1991:TPC

[546] F. Cucker and A. Torrecillas. Two P-complete problems in the theory of the reals. Lecture Notes in Computer Science, 510:556–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honkala:1991:MBD

[547] J. Honkala and A. Salomaa. L morphisms: Bounded delay and regularity of ambiguity. Lecture Notes in Computer Science, 510:566–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruyere:1991:DDV

[548] V. Bruyere and C. de Felice. Degree et decomposability of variable-length codes. Lecture Notes in Computer Science, 510:575–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1991:ETI

[549] T. Wilke. An EILENBERG theorem for inf-languages. Lecture Notes in Computer Science, 510:588–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1991:BOC

[550] K. Culik and S. Dube. Balancing order and chaos in image generation. Lecture Notes in Computer Science, 510:600–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gurevich:1991:ACC

[551] Y. Gurevich. Average case complexity (invited lecture). Lecture Notes in Computer Science, 510:615–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tao:1991:MNP

[552] Jiang Tao and B. Ravikumar. Minimal NFA problems are hard. Lecture Notes in Computer Science, 510:629– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hashiguchi:1991:ADS

[553] K. Hashiguchi. Algorithms for determining the smallest number of nonterminals (states) sufficient for generating (accepting) a regular language. Lecture Notes in Computer Science, 510: 641–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhattacharya:1991:CST

[554] B. Bhattacharya and G. Toussaint. Computing shortest transversals. Lecture Notes in Computer Science, 510: 649–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazelle:1991:RSP

[555] B. Chazelle, H. Edelsbrunner, M. Grigni, L. Guibas, J. Hershberger, M. Sharir, and J. Snoeyink. Ray shooting in polygons using geodesic triangulations. Lecture Notes in Computer Science, 510:661–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bern:1991:EED

[556] M. Bern, D. Eppstein, and F. Yao. The expected extremes in a Delaunay triangulation. Lecture Notes in Computer Science, 510:674–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazelle:1991:CGG

[557] B. Chazelle. Computational geometry for the gourmet: Old fare and new dishes (invited lecture). Lecture Notes in Computer Science, 510:686– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duris:1991:PMR

[558] P. Duris and Z. Galil. On the power of multiple reads in a chip. Lecture Notes in Computer Science, 510:697– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groger:1991:LDT

[559] H. D. Groger and G. Turan. On linear decision trees computing Boolean functions. Lecture Notes in Computer Science, 510:707–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galil:1991:ALT

[560] Z. Galil and O. Margalit. An almost linear-time algorithm for the dense subset-sum problem. Lecture Notes in Computer Science, 510:719–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuller:1991:LAW

[561] S. Khuller, S. G. Mitchell, and V. V. Vazirani. On-line algorithms for weighted bipartite matching and stable marriages. Lecture Notes in Computer Science, 510:728–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naor:1991:SMP

[562] M. Naor. String matching with preprocessing of text and pattern. Lecture Notes in Computer Science, 510: 739–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravi:1991:OPA

[563] R. Ravi, A. Agrawal, and P. Klein. Ordering problems approximated: Singleprocessor scheduling and interval graph completion. Lecture Notes in Computer Science, 510:751–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saffer:1991:SPD

[564] J. R. Saffer, H. H. Barrett, H. B. Barber, and J. M. Woolfenden. Surgical probe design for a coincidence imaging system without a collimator. Lecture Notes in Computer Science, 511:8–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zubal:1991:VBM

[565] G. Zubal and C. Harrell. Voxel based Monte Carlo calculations of nuclear medicine images and applied variance reduction techniques. Lecture Notes in Computer Science, 511:23–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meikle:1991:SSC

[566] S. R. Meikle, B. F. Hutton, D. L. Bailey, R. R. Fulton, and K. Schindhelm. SPECT scatter correction in nonhomogeneous media. Lecture Notes in Computer Science, 511:34–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clack:1991:CBS

[567] R. Clack, G. L. Zeng, Y. Weng, P. E. Christian, and G. T. Gullberg. Cone beam single photon emission computed tomography using two orbits. Lecture Notes in Computer Science, 511:45–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guedon:1991:SBR

[568] J-P. V. Guedon and Y. J. Bizais. Spline-based regularization for discrete FBP reconstruction. Lecture Notes in Computer Science, 511:55–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henri:1991:RDB

[569] C. J. Henri, D. L. Collins, and T. M. Peters. Reconstruction of 3-D branching structures. Lecture Notes in Computer Science, 511:65–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Llacer:1991:PEU

[570] J. Llacer, E. Veklerov, and J. Nunez. Preliminary examination of the use of case specific medical information as “prior” in Bayesian reconstruction. Lecture Notes in Computer Science, 511:81–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1991:RSP

[571] Z. Liang, R. Jaszczak, and R. Coleman. On reconstruction and segmentation of piecewise continuous images. Lecture Notes in Computer Science, 511:94–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leahy:1991:IAM

[572] R. Leahy and X. Yan. Incorporation of anatomical MR data for improved functional imaging with PET. Lecture Notes in Computer Science, 511:105– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gindi:1991:BRF

[573] G. Gindi, M. Lee, A. Rangarajan, and I. G. Zubal. Bayesian reconstruction of functional images using registered anatomical images as priors. Lecture Notes in Computer Science, 511:121– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pelizzari:1991:IDP

[574] C. A. Pelizzari, K. K. Tan, D. N. Levin, G. T. Y. Chen, and J. Balter. Interactive 3-D patient-image registration. Lecture Notes in Computer Science, 511:132–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vandenElsen:1991:MGR

[575] P. A. van den Elsen and M. A. Viergever. Marker guided registration of electromagnetic dipole data with tomographic images. Lecture Notes in Computer Science, 511:142–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lemoine:1991:ABR

[576] D. Lemoine, C. Barillot, B. Gibaud, and E. Pasqualini. An anatomicalbased 3D registration system of multimodality and atlas data in neurosurgery. Lecture Notes in Computer Science, 511:154–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oghabian:1991:RBI

[577] M. A. Oghabian and A. Todd-Pokropek. Registration of brain images by a multi-resolution sequential method. Lecture Notes in Computer Science, 511:165–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerig:1991:ASD

[578] G. Gerig, J. Martin, R. Kikinis, O. Kubler, M. Shenton, and F. A. Jolesz. Automating segmentation of dual-echo MR head data. Lecture Notes in Computer Science, 511:175– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brummer:1991:ADB

[579] M. E. Brummer, R. M. Mersereau, R. L. Eisner, and R. R. J. Lewine. Automatic detection of brain contours in MRI data sets. Lecture Notes in Computer Science, 511:188–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bister:1991:TAA

[580] M. Bister, J. Cornelis, and Y. Taeymans. Towards automated analysis in 3D cardiac MR imaging. Lecture Notes in Computer Science, 511:205– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snyder:1991:SMR

[581] W. Snyder, A. Logenthiran, P. Santago, K. Link, G. Bilbro, and S. Rajala. Segmentation of magnetic resonance images using mean field annealing. Lecture Notes in Computer Science, 511:218–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karssemeijer:1991:SMA

[582] N. Karssemeijer. A stochastic model for automated detection of calcifications in digital mammograms. Lecture Notes in Computer Science, 511:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romeny:1991:SSN

[583] B. M. ter Haar Romeny, L. M. J. Florack, J. J. Koenderink, and M. A. Viergever. Scale-space: its natural operators and differential invariants. Lecture Notes in Computer Science, 511:239–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffin:1991:SSG

[584] L. D. Griffin, A. C. F. Colchester, and G. P. Robinson. Scale and segmentation of grey-level images using maximum gradient paths. Lecture Notes in Computer Science, 511:256–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monga:1991:UUL

[585] O. Monga, N. Ayache, and P. Sander. Using uncertainty to link 3D edge detection and local surface modelling. Lecture Notes in Computer Science, 511:273–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1991:BEU

[586] W. J. Lin, S. M. Pizer, and V. E. Johnson. Boundary estimation in ultrasound images. Lecture Notes in Computer Science, 511:285–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malandain:1991:TCD

[587] G. Malandain, G. Bertrand, and N. Ayache. Topological classification in digital space. Lecture Notes in Computer Science, 511:300–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herman:1991:SBI

[588] G. T. Herman and C. A. Bucholtz. Shape-based interpolation using a chamfer distance. Lecture Notes in Computer Science, 511:314–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bookstein:1991:TPS

[589] F. L. Bookstein. Thin-plate splines and the atlas problem for biomedical images. Lecture Notes in Computer Science, 511:326–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amini:1991:NRM

[590] A. A. Amini, R. L. Owen, P. Anandan, and J. S. Duncan. Non-rigid motion models for tracking the left ventricular wall. Lecture Notes in Computer Science, 511:343–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozma:1991:MBR

[591] H. I. Bozma and J. S. Duncan. Model-based recognition of multiple deformable objects using a gametheoretic framework. Lecture Notes in Computer Science, 511:358–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barber:1991:EBD

[592] D. C. Barber and W. B. Tindale. Extraction of background distributions from abnormal data sets: application to radiolabelled leucocyte imaging. Lecture Notes in Computer Science, 511:373–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Houston:1991:QCC

[593] A. S. Houston and K. S. Nijran. A quantitative comparison of current methods of factor analysis of dynamic structures (FADS) in renal dynamic studies. Lecture Notes in Computer Science, 511:386–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samal:1991:CAF

[594] M. Samal, M. Karny, and D. Zahalka. Confirmatory aspects in factor analysis of image sequences. Lecture Notes in Computer Science, 511:397–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldock:1991:TMI

[595] R. A. Baldock. Trainable models for the interpretation of echocardiogram images. Lecture Notes in Computer Science, 511:408–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ovalle:1991:KDE

[596] A. Ovalle and C. Garbay. KIDS: A distributed expert system for biomedical image interpretation. Lecture Notes in Computer Science, 511:419–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binaghi:1991:LUCa

[597] E. Binaghi and A. Rampini. Learning of uncertain classification rules in biomedical images: the case of colposcopic images. Lecture Notes in Computer Science, 511:434–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dellepiane:1991:FMP

[598] S. Dellepiane, G. Venturi, and G. Vernazza. A fuzzy model for the processing and recognition of MR pathological images. Lecture Notes in Computer Science, 511:444–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrett:1991:LDI

[599] H. H. Barrett, T. Gooley, K. Girodias, J. Rolland, T. White, and J. Yao. Linear discriminants and image quality. Lecture Notes in Computer Science, 511:458–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cromartie:1991:EAC

[600] R. Cromartie and S. M. Pizer. Edgeaffected context for adaptive contrast enhancement. Lecture Notes in Computer Science, 511:474–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohbuchi:1991:IVR

[601] R. Ohbuchi and H. Fuchs. Incremental volume rendering algorithm for interactive 3D ultrasound imaging. Lecture Notes in Computer Science, 511: 486–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gudmundsson:1991:CSV

[602] B. Gudmundsson and M. Randen. Compression of sequences of 3Dvolume surface projections. Lecture Notes in Computer Science, 511:501– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snyder:1991:MCO

[603] Alan Snyder. Modeling the C++ object model: An application of an abstract object model. Lecture Notes in Computer Science, 512:1– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0512/05120001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0512/05120001.
pdf.

Holzle:1991:ODT

[604] Urs Hölzle, Craig Chambers, and David Ungar. Optimizing dynamicallytyped object-oriented languages with polymorphic inline caches. Lecture Notes in Computer Science, 512:21– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0512/05120021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0512/05120021.
pdf.

Morzenti:1991:OOL

[605] Angelo Morzenti and Pierluigi San Pietro. An object-oriented logic language for modular system specification. Lecture Notes in Computer Science, 512:39–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0512/05120039.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0512/05120039.
pdf.

Wills:1991:CTF

[606] Alan Wills. Capsules and types in Fresco: Program verification in Smalltalk. Lecture Notes in Computer Science, 512:59–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0512/05120059.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0512/05120059.pdf.

Barbedette:1991:SML

[607] Gilles Barbedette. Schema modifications in the LISPO2 persistent object-oriented language. Lecture Notes in Computer Science, 512:77– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0512/05120077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0512/05120077.
pdf.

Delcourt:1991:DIC

[608] Christine Delcourt and Roberto Zicari. The design of an integrity consistency checker (ICC) for an object oriented database system. Lecture Notes in Computer Science, 512:97– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0512/05120097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0512/05120097.
pdf.

Neusius:1991:SA

[609] Christian Neusius. Synchronizing actions. Lecture Notes in Computer Science, 512:118–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0512/05120118.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0512/05120118.pdf.

Honda:1991:OCA

[610] Kohei Honda and Mario Tokoro. An object calculus for asynchronous communication. Lecture Notes in Computer Science, 512:133–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/0512/05120133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0512/05120133.
                                        pdf.

Reghizzi:1991:DRC

[611] Stefano Crespi Reghizzi, Guido Galli de Paratesi, and Stefano Genolini. Definition of reusable concurrent software components. Lecture Notes in Computer Science, 512:148–166, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0512/05120148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0512/05120148.
                                        pdf.

Cusack:1991:IOO

[612] Elspeth Cusack. Inheritance in object oriented Z. Lecture Notes in Computer Science, 512:167–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0512/05120167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0512/05120167.
                                        pdf.

Alencar:1991:OOO

[613] Antonio J. Alencar and Joseph A. Goguen. OOZE: An object oriented Z environment. Lecture Notes in Computer Science, 512:180–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/0512/05120180.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0512/05120180.
pdf.

Almarode:1991:IDI

[614] Jay Almarode. Issues in the design and implementation of a schema designer for an OODBMS. Lecture Notes in Computer Science, 512:200– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0512/05120200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0512/05120200.
pdf.

Medeiros:1991:OIU

[615] Claudia Bauzer Medeiros and Patrick Pfeffer. Object integrity using rules. Lecture Notes in Computer Science, 512:219–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0512/05120219.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0512/05120219.
pdf.

Matsuoka:1991:HGR

[616] Satoshi Matsuoka, Takuo Watanabe, and Akinori Yonezawa. Hybrid group reflective architecture for object-oriented concurrent reflective programming. Lecture Notes in Computer Science, 512:231–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/0512/05120231.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0512/05120231.
                                        pdf.

Rao:1991:IRS

[617] Ramana Rao. Implementational reflection in Silica. Lecture Notes in Computer Science, 512:251–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0512/05120251.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0512/05120251.
                                        pdf.

Lacourte:1991:EGO

[618] Serge Lacourte. Exceptions in guide, an object-oriented language for distributed applications. Lecture Notes in Computer Science, 512:268–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0512/05120268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0512/05120268.
                                        pdf.

Wolinski:1991:RCO

[619] Francis Wolinski and Jean-François Perrot. Representation of complex objects: Multiple facets with partwhole hierarchies. Lecture Notes in Computer Science, 512:288–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0512/05120288.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/0512/05120288.
pdf.

Mugridge:1991:MMS

[620] Warwick B. Mugridge, John Hamer, and John G. Hosking. Multimethods in a statically-typed programming language. Lecture Notes in Computer Science, 512:307–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0512/05120307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0512/05120307.
pdf.

Palsberg:1991:WTS

[621] Jens Palsberg and Michael I. Schwartzbach. What is type-safe code reuse? Lecture Notes in Computer Science, 512:325–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0512/05120325.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0512/05120325.
pdf.

Lippe:1991:ITI

[622] Ernst Lippe and Gert Florijn. Implementation techniques for integral version management. Lecture Notes in Computer Science, 512:342–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0512/05120342.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/0512/05120342.
                                        pdf.

Champeaux:1991:OOA

[623] Dennis de Champeaux. Objectoriented analysis and top-down software development. Lecture Notes in Computer Science, 512:360–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0512/05120360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0512/05120360.
                                        pdf.

deChampeaux:1991:OOA

[624] Dennis de Champeaux. Objectoriented analysis and top-down software development. Lecture Notes in Computer Science, 512:360–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstein:1991:ICD

[625] Paul L. Bergstein and Karl J. Lieberherr. Incremental class dictionary learning and optimization. Lecture Notes in Computer Science, 512:377– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0512/05120377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0512/05120377.
                                        pdf.

Mattos:1991:AKB

[626] N. M. Mattos. An approach to knowledge base management. Lecture Notes in Computer Science, 513:1–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ericson:1991:CSC

[627] Th. Ericson and V. A. Zinoviev. Concatenated spherical codes. Lecture Notes in Computer Science, 514:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elia:1991:NAG

[628] M. Elia and G. Taricco. A note on automorphism groups of codes and symbol error probability computation. Lecture Notes in Computer Science, 514:6–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1991:DPA

[629] T. Berger. A direct proof for the automorphism group of Reed–Solomon codes. Lecture Notes in Computer Science, 514:21–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beth:1991:FBC

[630] T. Beth, D. E. Lazic, and V. Senk. A family of binary codes with asymptotically good distance distribution. Lecture Notes in Computer Science, 514: 30–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlet:1991:TBF

[631] C. Carlet. A transformation on Boolean functions, its consequences on some problems related to Reed–Muller codes. Lecture Notes in Computer Science, 514:42–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langevin:1991:CRR

[632] P. Langevin. Covering radius of RM (1,9) in RM (3,9). Lecture Notes in Computer Science, 514:51–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodier:1991:WDB

[633] F. Rodier. The weights of the duals of binary BCH codes of designed distance delta = 9. Lecture Notes in Computer Science, 514:60–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Augot:1991:MDS

[634] D. Augot, P. Charpin, and N. Sendrier. The minimum distance of some binary codes via the Newton’s identities. Lecture Notes in Computer Science, 514:65–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burger:1991:MCB

[635] J. Burger, H. Chabanne, and M. Girault. Minimum-change binary blockcodes which are well balanced. Lecture Notes in Computer Science, 514:76–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mabogunje:1991:CUE

[636] A. O. Mabogunje and P. G. Farrell. Construction of unequal error protection codes. Lecture Notes in Computer Science, 514:87–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1991:UED

[637] G. D. Cohen, L. Gargano, and U. Vaccaro. Unidirectional error-detecting codes. Lecture Notes in Computer Science, 514:94–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montpetit:1991:CPC

[638] A. Montpetit. Coherent partitions and codes. Lecture Notes in Computer Science, 514:106–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astie-Vidal:1991:LRL

[639] A. Astie-Vidal and V. Dugat. (1/ lambda)-regular and (1/lambda)- homogeneous tournaments. Lecture Notes in Computer Science, 514:114– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeBrigand:1991:DCH

[640] D. Le Brigand. Decoding of codes on hyperelliptic curves. Lecture Notes in Computer Science, 514:126–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rotillon:1991:DCK

[641] D. Rotillon and J. A. Thiong Ly. Decoding of codes on the Klein quartic. Lecture Notes in Computer Science, 514:135–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voss:1991:AGF

[642] C. Voss. Asymptotically good families of geometric Goppa codes and the Gilbert–Varshamov bound. Lecture Notes in Computer Science, 514: 150–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perret:1991:MCS

[643] M. Perret. Multiplicative character sums and non linear geometric codes. Lecture Notes in Computer Science, 514:158–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girault:1991:SIS

[644] M. Girault. A survey of identification schemes. Lecture Notes in Computer Science, 514:168–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harari:1991:CCS

[645] S. Harari. A correlation cryptographic scheme. Lecture Notes in Computer Science, 514:180–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patarin:1991:PPB

[646] J. Patarin. Pseudorandom permutations based on the DES scheme. Lecture Notes in Computer Science, 514: 193–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fell:1991:LCT

[647] H. J. Fell. Linear complexity of transformed sequences. Lecture Notes in Computer Science, 514:205–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chasse:1991:SRL

[648] G. Chasse. Some remarks on a LFSR “disturbed” by other sequences. Lecture Notes in Computer Science, 514: 215–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Creutzburg:1991:PCF

[649] R. Creutzburg. Parameters for complex FFTs infinite residue class rings. Lecture Notes in Computer Science, 514: 222–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domingo-Ferrer:1991:CTP

[650] J. Domingo-Ferrer and L. Huguet-Rotger. A cryptographic tool for programs protection. Lecture Notes in Computer Science, 514:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haccoun:1991:DTC

[651] D. Haccoun. Decoding techniques for convolutional codes. Lecture Notes in Computer Science, 514:242–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sfez:1991:WOV

[652] R. Sfez and G. Battail. A weightedoutput variant of the Viterbi algorithm for concatenated schemes using a convolutional inner code. Lecture Notes in Computer Science, 514:259–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filho:1991:CMC

[653] R. Baldini Filho and P. G. Farrell. Coded modulation with convolutional codes over rings. Lecture Notes in Computer Science, 514:271–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sgarro:1991:STC

[654] A. Sgarro. A Shannon-theoretic coding theorem in authentication theory. Lecture Notes in Computer Science, 514: 282–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fioretto:1991:JFE

[655] A. Fioretto and A. Sgarro. Joint fractional entropy. Lecture Notes in Computer Science, 514:292–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battail:1991:ASC

[656] G. Battail and M. Guazzo. On the adaptive source coding. Lecture Notes in Computer Science, 514:298– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capocelli:1991:MCL

[657] R. M. Capocelli and A. De Santis. Minimum codeword length and redundancy of Huffman codes. Lecture Notes in Computer Science, 514:309–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calderbank:1991:BCC

[658] A. R. Calderbank. Binary covering codes and high speed data transmission. Lecture Notes in Computer Science, 514:320–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battail:1991:CMG

[659] G. Battail, H. Magalhaes De Oliviera, and Z. Weidong. Coding and modulation for the Gaussian channel, in the absence or in the presence of fluctuations. Lecture Notes in Computer Science, 514:337–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sfez:1991:CTM

[660] R. Sfez, J. C. Belfiore, K. Leeuwin, and A. Fihel. Comparison of two modulation-coding schemes for lowrate digital land mobile radio communication. Lecture Notes in Computer Science, 514:350–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leeuwin:1991:CBT

[661] K. Leeuwin, J. C. Belfiore, and G. Kawas Kaleh. Chernoff bound of trellis coded modulation over correlated Rayleigh channel. Lecture Notes in Computer Science, 514:364– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darmon:1991:HFA

[662] M. M. Darmon and P. R. Sadot. A hybrid FEC-ARQ communication system using sequential decoding. Lecture Notes in Computer Science, 514:378– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Politano:1991:MRS

[663] J.-L. Politano and D. Deprey. A 30 mbits/s (255,223) Reed–Solomon decoder. Lecture Notes in Computer Science, 514:385–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dressler:1991:PPS

[664] Oskar Dressler and Adam Farquhar. Putting the problem solver back in the driver’s seat: Contextual control of the AMTS. Lecture Notes in Computer Science, 515:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Junker:1991:VBT

[665] Ulrich Junker. Variations on backtracking for TMS. Lecture Notes in Computer Science, 515:17–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inoue:1991:APC

[666] Katsumi Inoue. An abductive procedure for the CMS/ATMS. Lecture Notes in Computer Science, 515:34–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakas:1991:KAA

[667] A. C. Kakas and P. Mancarella. Knowledge assimilation and abduction. Lecture Notes in Computer Science, 515: 54–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordano:1991:TMS

[668] Laura Giordano and Alberto Martelli. Truth maintenance systems and belief revision. Lecture Notes in Computer Science, 515:71–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1991:PAB

[669] Didier Dubois, Jerome Lang, and Henri Prade. A possibilistic assumptionbased truth maintenance system with uncertain justifications, and its application to belief revision. Lecture Notes in Computer Science, 515:87–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tayrac:1991:AEA

[670] Pierre Tayrac. ARC: An extended ATMS based on directed CAT-Correct resolution. Lecture Notes in Computer Science, 515:107–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujiwara:1991:LFA

[671] Yasushi Fujiwara and Shinichi Honiden. On logical foundations of the ATMS. Lecture Notes in Computer Science, 515:125–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1991:SST

[672] Cees Witteveen. A skeptical semantics for truth maintenance. Lecture Notes in Computer Science, 515:136–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1991:SAB

[673] Peter Jackson and John Pais. Semantic accounts of belief revision. Lecture Notes in Computer Science, 515:155– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bronsard:1991:CRF

[674] Francois Bronsard and Uday S. Reddy. Conditional rewriting in focus. Lecture Notes in Computer Science, 516:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1991:MLU

[675] Nachum Dershowitz. A maximal-literal unit strategy for Horn clauses. Lecture Notes in Computer Science, 516:14–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klop:1991:ETR

[676] Jan Willem Klop and Roel de Vrijer. Extended term rewriting systems (invited paper). Lecture Notes in Computer Science, 516:26–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kounalis:1991:PSC

[677] Emmanuel Kounalis and Michael Rusinowitch. A proof system for conditional algebraic specifications. Lecture Notes in Computer Science, 516:51–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meseguer:1991:CRL

[678] Jose Meseguer. Conditional rewriting logic: Deduction, models and concurrency (invited paper). Lecture Notes in Computer Science, 516:64–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohan:1991:ERP

[679] Chilukuri K. Mohan. Equivalences of rewrite programs (*). Lecture Notes in Computer Science, 516:92–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1991:FRI

[680] Chen Hong, Hsiang Jieh, and Kong Hwa-Chung. On finite representations of infinite sequences of terms. Lecture Notes in Computer Science, 516:100– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yiyun:1991:ITI

[681] Chen Yiyun and Michael J. O’Donnell. Infinite terms and infinite rewritings. Lecture Notes in Computer Science, 516:115–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yiyun:1991:TCN

[682] Chen Yiyun and Michael J. O’Donnell. Testing confluence of nonterminating rewriting systems. Lecture Notes in Computer Science, 516:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hodgson:1991:SOI

[683] Bernard R. Hodgson and Clement F. Kent. A survey of ordinal interpretations of type ϵ_{0} for termination of rewriting systems (*). Lecture Notes in Computer Science, 516:137–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirchner:1991:MRS

[684] Helene Kirchner and Miki Hermann. Meta-rule synthesis from crossed rewrite systems. Lecture Notes in Computer Science, 516:143–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anantharaman:1991:AAE

[685] Siva Anantharaman and Maria Paola Bonacina. An application of automated equational reasoning to manyvalued logic (*). Lecture Notes in Computer Science, 516:156–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1991:CFO

[686] Leo Bachmair and Harald Ganzinger. Completion of first-order clauses with equality by strict superposition (invited paper). Lecture Notes in Computer Science, 516:162–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertling:1991:KBC

[687] Hubert Bertling. Knuth–Bendix completion of Horn clause programs for restricted linear resolution and paramodulation. Lecture Notes in Computer Science, 516:181–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bevers:1991:PCC

[688] Eddy Bevers and Johan Lewi. Proof by consistency in conditional equational theories. Lecture Notes in Computer Science, 516:194–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonacina:1991:CPS

[689] Maria Paola Bonacina and Hsiang Jieh. Completion procedures as semidecision procedures (invited paper). Lecture Notes in Computer Science, 516:206–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devie:1991:LC

[690] Herve Devie. Linear completion. Lecture Notes in Computer Science, 516: 233–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieuwenhuis:1991:CR

[691] Robert Nieuwenhuis and Fernando Orejas. Clausal rewriting. Lecture Notes in Computer Science, 516:246– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbanera:1991:AAR

[692] Franco Barbanera. Adding algebraic rewriting to the calculus of constructions: Strong normalization preserved. Lecture Notes in Computer Science, 516:262–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bousdira:1991:SCC

[693] Wadoud Bousdira and Jean-Luc Remy. On sufficient completeness of conditional specifications. Lecture Notes in Computer Science, 516:272–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lichtenstein:1991:FFP

[694] N. Lichtenstein and S. Kaplan. FPL: Functional plus logic programming. an integration of the FP and Prolog languages. Lecture Notes in Computer Science, 516:287–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Middeldorp:1991:CDU

[695] Aart Middeldorp. Confluence of the disjoint union of conditional term rewriting systems. Lecture Notes in Computer Science, 516:295–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plump:1991:ITR

[696] Detlef Plump. Implementing term rewriting by graph reduction: Termination of combined systems. Lecture Notes in Computer Science, 516:307– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aida:1991:CCR

[697] Hitoshi Aida, Joseph Goguen, and Jose Meseguer. Compiling concurrent rewriting onto the rewrite rule machine. Lecture Notes in Computer Science, 516:320–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoy:1991:DSR

[698] Sergio Antoy. Design strategies for rewrite rules. Lecture Notes in Computer Science, 516:333–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galbiati:1991:SUL

[699] Louis Galbiati and Carolyn Talcott. A simplifier for untyped lambda expressions. Lecture Notes in Computer Science, 516:342–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanEekelen:1991:PGR

[700] M. C. J. D. van Eekelen, M. J. Plasmeijer, and J. E. W. Smetsers. Parallel graph rewriting on loosely coupled machine architectures. Lecture Notes in Computer Science, 516:354–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1991:TET

[701] Roberto M. Amadio. Typed equivalence, type assignment and type containment. Lecture Notes in Computer Science, 516:372–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emele:1991:FPS

[702] Martin Emele and Remi Zajac. A fixedpoint semantics for feature type systems (*). Lecture Notes in Computer Science, 516:383–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stell:1991:USO

[703] John G. Stell. Unique-sort order-sorted theories: A description as monad morphisms. Lecture Notes in Computer Science, 516:389–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yong:1991:ELB

[704] Sun Yong. Equational logics (Birkhoff’s method revisited) (*). Lecture Notes in Computer Science, 516:401–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waldmann:1991:COS

[705] Uwe Waldmann. Compatibility of order-sorted rewrite rules. Lecture Notes in Computer Science, 516:407– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguzzi:1991:UTC

[706] G. Aguzzi, U. Modigliani, and M. C. Verri. A universal termination condition for solving goals in equational languages (*). Lecture Notes in Computer Science, 516:418–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darlington:1991:CED

[707] John Darlington and Guo Yike. Constrained equational deduction. Lecture Notes in Computer Science, 516:424– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1991:HOU

[708] Tobias Nipkow. Higher-order unification, polymorphism, and subsorts. Lecture Notes in Computer Science, 516: 436–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhenyu:1991:SOU

[709] Qian Zhenyu. Second-order unification in the presence of linear shallow algebraic equations (*). Lecture Notes in Computer Science, 516:448–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snyder:1991:ISH

[710] Wayne Snyder and Christopher Lynch. An inference system for Horn clause logic with equality: A foundation for conditional E-unification and for logic programming in the presence of equality. Lecture Notes in Computer Science, 516:454–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nokel:1991:TDS

[711] K. Nokel. Temporally distributed symptoms in technical diagnosis. Lecture Notes in Computer Science, 517: 1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1991:ITF

[712] J. G. Williams. Instantiation theory: On the foundations of automated deduction. Lecture Notes in Computer Science, 518:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunto:1991:CSC

[713] W. Cunto, J. I. Munro, and P. V. Poblete. A case study in comparison based complexity: Finding the nearest value(s). Lecture Notes in Computer Science, 519:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aronov:1991:ZSH

[714] B. Aronov and M. Sharir. On the zone of a surface in a hyperplane arrangement. Lecture Notes in Computer Science, 519:13–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pellegrini:1991:RSI

[715] M. Pellegrini. Ray-shooting and isotopy classes of lines in 3-dimensional space. Lecture Notes in Computer Science, 519:20–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1991:FLA

[716] P. F. Dietz. Finding level-ancestors in dynamic trees. Lecture Notes in Computer Science, 519:32–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sundaram:1991:TCA

[717] R. Sundaram, K. S. Singh, and C. Pandu Rangan. Treewidth of circular-arc graphs. Lecture Notes in Computer Science, 519:41–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devillers:1991:FDD

[718] O. Devillers, S. Meiser, and M. Teillaud. Fully dynamic Delaunay triangulation in logarithmic expected time per operation. Lecture Notes in Computer Science, 519:42–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1991:CVD

[719] H. Djidjev and A. Lingas. On computing the Voronoi diagram for restricted planar figures. Lecture Notes in Computer Science, 519:54–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1991:MP

[720] J. Diaz, A. M. Gibbons, M. S. Paterson, and J. Toran. The minsumcut problem. Lecture Notes in Computer Science, 519:65–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katoh:1991:EAM

[721] N. Katoh and K. Iwano. Efficient algorithms for the minimum range cut problems. Lecture Notes in Computer Science, 519:80–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akl:1991:MAM

[722] S. G. Akl. Memory access in models of parallel computation: From folklore to synergy and beyond. Lecture Notes in Computer Science, 519:92–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1991:FNM

[723] P. K. Agarwal, J. Matousek, and S. Suri. Farthest neighbors, maximum spanning trees and related problems in higher dimensions. Lecture Notes in Computer Science, 519:105–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salowe:1991:SIS

[724] J. S. Salowe. Shallow interdistance selection and interdistance enumeration. Lecture Notes in Computer Science, 519:117–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilar:1991:SMA

[725] O. R. Aguilar, A. Kumar Datta, and S. Ghosh. Sharing memory in asynchronous message passing systems. Lecture Notes in Computer Science, 519:129–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1991:LTS

[726] L. Yu and D. J. Rosenkrantz. A linear-time scheme for version reconstruction. Lecture Notes in Computer Science, 519:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanson:1991:ISL

[727] E. N. Hanson. The interval skip list: A data structure for finding all intervals that overlap a point. Lecture Notes in Computer Science, 519:153–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arkin:1991:GKP

[728] E. M. Arkin, S. Khuller, and J. S. B. Mitchell. Geometric knapsack problems. Lecture Notes in Computer Science, 519:165–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1991:FDS

[729] Y. Han. A fast derandomization scheme and its applications. Lecture Notes in Computer Science, 519:177– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lengauer:1991:UPP

[730] T. Lengauer and D. Theune. Unstructured path problems and the making of semirings. Lecture Notes in Computer Science, 519:189–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1991:NGG

[731] F. Yao. Neighborhood graphs and geometric embedding. Lecture Notes in Computer Science, 519:201–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1991:FOB

[732] J. S. B. Mitchell and E. L. Wynters. Finding optimal bipartitions of points and polygons. Lecture Notes in Computer Science, 519:202–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czyzowicz:1991:IP

[733] J. Czyzowicz, I. Stojmenovic, and J. Urrutia. Immobilizing a polytope. Lecture Notes in Computer Science, 519:214–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquet:1991:WCW

[734] P. Jacquet and W. Szpankowski. What can we learn about suffix trees from independent tries? Lecture Notes in Computer Science, 519:228–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

dAmore:1991:CAW

[735] F. d’Amore, A. Marchetti-Spaccamela, and U. Nanni. Competitive algorithms for the weighted list update problem. Lecture Notes in Computer Science, 519:240–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1991:OAR

[736] B. J. Nilsson and S. Schuierer. An optimal algorithm for the rectilinear link center of a rectilinear polygon. Lecture Notes in Computer Science, 519: 249–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1991:GSL

[737] G. Das and G. Narasimhan. Geometric searching and link distance. Lecture Notes in Computer Science, 519: 261–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naor:1991:REE

[738] D. Naor and V. V. Vazirani. Representing and enumerating edge connectivity cuts in RNC. Lecture Notes in Computer Science, 519:273–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kant:1991:PGA

[739] G. Kant and H. L. Bodlaender. Planar graph augmentation problems. Lecture Notes in Computer Science, 519:

286–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frederickson:1991:PSL

[740] G. N. Frederickson. Parametric search and locating supply centers in trees. Lecture Notes in Computer Science, 519:299–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1991:BLR

[741] C. D. Yang, D. T. Lee, and C. K. Wong. On bends and lengths of rectilinear paths: A graph-theoretic approach. Lecture Notes in Computer Science, 519:320–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hershberger:1991:CML

[742] J. Hershberger and J. Snoeyink. Computing minimum length paths of a given homotopy class. Lecture Notes in Computer Science, 519:331–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bar-Ilan:1991:AAS

[743] J. Bar-Ilan and D. Peleg. Approximation algorithms for selecting network centers. Lecture Notes in Computer Science, 519:343–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravi:1991:FDP

[744] S. S. Ravi, D. J. Rosenkrantz, and G. K. Tayi. Facility dispersion problems: Heuristics and special cases. Lecture Notes in Computer Science, 519: 355–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1991:OGC

[745] S. Carlsson, B. J. Nilsson, and S. Ntafos. Optimum guard covers and m-watchmen routes for restricted polygons. Lecture Notes in Computer Science, 519:367–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1991:ANS

[746] P. K. Agarwal and M. U. Sharir. Applications of a new space partitioning technique. Lecture Notes in Computer Science, 519:379–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eppstein:1991:OAD

[747] D. Eppstein. Offline algorithms for dynamic minimum spanning tree problems. Lecture Notes in Computer Science, 519:392–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moret:1991:EAA

[748] B. M. E. Moret and H. D. Shapiro. An empirical analysis of algorithms for constructing a minimum spanning tree. Lecture Notes in Computer Science, 519:400–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhattacharya:1991:LTA

[749] B. K. Bhattacharya, A. Mukhopadhyay, and G. T. Toussaint. A linear time algorithm for computing the shortest line segment from which a polygon is weakly externally visible.

Lecture Notes in Computer Science, 519:412–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vegter:1991:DMV

[750] G. Vegter. Dynamically maintaining the visibility graph. Lecture Notes in Computer Science, 519:425–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heffernan:1991:OAC

[751] P. J. Heffernan and J. S. B. Mitchell. An optimal algorithm for computing visibility in the plane. Lecture Notes in Computer Science, 519:437–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Italiano:1991:FPD

[752] G. F. Italiano and N. Sarnak. Fully persistent data structures for disjoint set union problems. Lecture Notes in Computer Science, 519:449–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapoor:1991:AGA

[753] S. Kapoor and H. Ramesh. Algorithms for generating all spanning trees of undirected, directed and weighted graphs. Lecture Notes in Computer Science, 519:461–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munro:1991:SMV

[754] J. I. Munro and V. Raman. Sorting multisets and vectors in-place. Lecture Notes in Computer Science, 519:473– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abrahamson:1991:PLE

[755] K. Abrahamson, A. Adler, L. Higham, and D. Kirkpatrick. Probabilistic leader election on rings of known size. Lecture Notes in Computer Science, 519:481–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lassez:1991:ENT

[756] J.-L. Lassez, M. Maher, and K. Marriott. Elimination of negation in term algebras. Lecture Notes in Computer Science, 520:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lescanne:1991:ROT

[757] P. Lescanne. Rewrite orderings and termination of rewrite systems. Lecture Notes in Computer Science, 520:17–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1991:FFM

[758] Z. Manna and A. Pnueli. On the faithfulness of formal models. Lecture Notes in Computer Science, 520:28–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1991:MC

[759] M. Nielsen. Models for concurrency. Lecture Notes in Computer Science, 520:43–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salwicki:1991:HFT

[760] A. Salwicki. On a hierarchy of file types and a tower of their theories. Lecture Notes in Computer Science, 520:47–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alessi:1991:SCI

[761] F. Alessi and F. Barbanera. Strong conjunction and intersection types. Lecture Notes in Computer Science, 520:64–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astesiano:1991:PHO

[762] E. Astesiano and M. Cerioli. Partial higher-order specifications. Lecture Notes in Computer Science, 520:74–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonnier:1991:UIS

[763] S. Bonnier. Unification in incompletely specified theories: a case study. Lecture Notes in Computer Science, 520:84–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudol:1991:OL

[764] G. Boudol, I. Castellani, M. Hennessy, and A. Kiehn. Observing localities. Lecture Notes in Computer Science, 520:93–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costa:1991:ADD

[765] G. Costa and G. Reggio. Abstract dynamic data types: a temporal logic approach. Lecture Notes in Computer Science, 520:103–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerny:1991:GWC

[766] A. Cerny. Generating words by cellular automata. Lecture Notes in Computer Science, 520:113–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degano:1991:ARP

[767] P. Degano and R. Gorrieri. Atomic refinement in process description languages. Lecture Notes in Computer Science, 520:121–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1991:RCT

[768] V. Diekert, P. Gastin, and A. Petit. Recognizable complex trace languages. Lecture Notes in Computer Science, 520:131–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domenjoud:1991:SSL

[769] E. Domenjoud. Solving systems of linear Diophantine equations: an algebraic approach. Lecture Notes in Computer Science, 520:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowek:1991:SOP

[770] G. Dowek. A second-order pattern matching algorithm for the cube of typed lambda-calculi. Lecture Notes in Computer Science, 520:151–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egidi:1991:LCV

[771] L. Egidi, F. Honsell, and S. Ronchi della Rocca. The lazy call-by-value lambda-calculus. Lecture Notes in Computer Science, 520:161–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorrieri:1991:LSB

[772] R. Gorrieri and C. Laneve. The limit of split _{n}-bisimulations for CCS agents. Lecture Notes in Computer Science, 520:170–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansel:1991:SAL

[773] G. Hansel, D. Krob, and Ch. Michaux. Stochastic automata and length distributions of rational languages. Lecture Notes in Computer Science, 520:181– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilken:1991:TCS

[774] B. P. Hilken and D. E. Rydeheard. Towards a categorical semantics type classes. Lecture Notes in Computer Science, 520:191–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howell:1991:SPP

[775] R. R. Howell, P. Jancar, and L. E. Rosier. Single-path Petri nets. Lecture Notes in Computer Science, 520: 202–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1991:BPG

[776] J. Hromkovic and B. Monien. The bisection problem for graphs of degree 4 (configuring Transputer systems). Lecture Notes in Computer Science, 520:211–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1991:SRC

[777] O. H. Ibarra, Jiang Tao, and Wang Hui. Some results concerning 2-D on-line tessellation acceptors and 2-D alternating finite automata. Lecture Notes in Computer Science, 520:221–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inverardi:1991:INF

[778] P. Inverardi and M. Nesi. Infinite normal forms for non-linear term rewriting systems. Lecture Notes in Computer Science, 520:231–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jokinen:1991:TAA

[779] P. Jokinen and E. Ukkonen. Two algorithms for approximate string matching in static texts. Lecture Notes in Computer Science, 520:240–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karhumaki:1991:ECT

[780] J. Karhumaki, W. Rytter, and S. Jarominek. Efficient constructions of test sets for regular and context-free languages. Lecture Notes in Computer Science, 520:249–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kavadias:1991:CRC

[781] D. Kavadias, L. M. Kirousis, and P. Spirakis. The complexity of the reliable connectivity problem. Lecture Notes in Computer Science, 520:259– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesner:1991:PMO

[782] D. Kesner. Pattern matching in ordersorted languages. Lecture Notes in Computer Science, 520:267–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konikowska:1991:TTT

[783] B. Konikowska. Two over three: a two-valued logic for software specification and validation over a three-valued predicate calculus. Lecture Notes in Computer Science, 520:277–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kounalis:1991:SCP

[784] E. Kounalis, D. Lugiez, and L. Pottier. A solution of the complement problem in associative-commutative theories. Lecture Notes in Computer Science, 520:287–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnan:1991:MRT

[785] Padmanabhan Krishnan. A model for real-time systems. Lecture Notes in Computer Science, 520:298–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1991:SC

[786] Huong Lam Nguyen and Long Van Do. On strict codes. Lecture Notes in Computer Science, 520:308–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leib:1991:DCS

[787] H. Leib and F. Henglein. A decidable case of the semi-unification problem. Lecture Notes in Computer Science, 520:318–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leopold:1991:MDH

[788] C. Leopold. Maintaining dictionaries in a hierarchical memory. Lecture Notes in Computer Science, 520:328– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinel:1991:ULB

[789] Ch. Meinel and S. Waack. Upper and lower bounds for certain graphaccessibility problems on bounded alternating omega-branching programs. Lecture Notes in Computer Science, 520:337–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1991:CDB

[790] U. Montanari and V. Sassone. CCS dynamic bisimulation is progressing. Lecture Notes in Computer Science, 520: 346–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdallah:1991:SSM

[791] M. A. Nait Abdallah. Syntax and semantics of a monotonic framework for non-monotonic reasoning. Lecture Notes in Computer Science, 520:357– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niwinski:1991:CSI

[792] D. Niwinski. On the cardinality of sets of infinite trees recognizable by finite automata. Lecture Notes in Computer Science, 520:367–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paech:1991:ETL

[793] B. Paech. Extending temporal logic by explicit concurrency. Lecture Notes in Computer Science, 520:377–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez:1991:EPC

[794] R. Pino Perez. An extensional partial combinatory algebra based on lambdaterms. Lecture Notes in Computer Science, 520:387–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poigne:1991:OMO

[795] A. Poigne. Once more on order-sorted algebras. Lecture Notes in Computer Science, 520:397–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roos:1991:CTS

[796] Y. Roos and P. A. Wacrenier. Composition of two semi commutations. Lecture Notes in Computer Science, 520: 406–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruzicka:1991:EDA

[797] P. Ruzicka. An efficient decision algorithm for the uniform semi-unification problem. Lecture Notes in Computer Science, 520:415–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shvachko:1991:DMP

[798] K. V. Shvachko. Different modifications of pointer machines and their computational power. Lecture Notes in Computer Science, 520:426–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shenoy:1991:STE

[799] P. P. Shenoy. On Spohn’s theory of epistemic beliefs. Lecture Notes in Computer Science, 521:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennes:1991:FAD

[800] R. Kennes and P. Smets. Fast algorithms for Dempster-Shafer theory. Lecture Notes in Computer Science, 521:14–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kruse:1991:CIS

[801] R. Kruse and E. Schwecke. On the combination of information sources. Lecture Notes in Computer Science, 521: 24–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spies:1991:AAQ

[802] M. Spies. Application aspects of qualitative conditional independence. Lecture Notes in Computer Science, 521: 31–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deutsch-McLeish:1991:SPB

[803] M. Deutsch-McLeish. A study of probabilities and belief functions under conflicting evidence: Comparisons and new methods. Lecture Notes in Computer Science, 521:41–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohlas:1991:PBF

[804] J. Kohlas and P.-A. Monney. Propagating belief functions through constraints systems. Lecture Notes in Computer Science, 521:50–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moral:1991:UUI

[805] S. Moral and L. M. de Campos. Updating uncertain information. Lecture Notes in Computer Science, 521:58–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yen-Teh:1991:AMB

[806] Hsia Yen-Teh. Assessing multiple beliefs according to one body of evidence — why it may be necessary, and how we might do it correctly. Lecture Notes in Computer Science, 521:68–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paass:1991:PDRa

[807] G. Paass. Probabilistic default reasoning. Lecture Notes in Computer Science, 521:76–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramson:1991:KRB

[808] B. Abramson. On knowledge representation in belief networks. Lecture Notes in Computer Science, 521:86–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luo:1991:SSS

[809] Z. Luo and A. Gammerman. STOSS — A stochastic simulation system for Bayesian belief networks. Lecture Notes in Computer Science, 521:97–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coletti:1991:CEV

[810] G. Coletti, A. Gilio, and R. Scozzafava. Conditional events with vague information in expert systems. Lecture Notes in Computer Science, 521:106–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OLeary:1991:RSR

[811] D. E. O’Leary. On representation of source reliability in weight of evidence. Lecture Notes in Computer Science, 521:115–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1991:AMH

[812] T. P. Martin and J. F. Baldwin. An abstract mechanism for handling uncertainty. Lecture Notes in Computer Science, 521:126–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bobrowicz:1991:MBM

[813] O. Bobrowicz, C. Choulet, A. Haurat, F. Sandoz, and M. Tebaa. A method to build membership functions. Lecture Notes in Computer Science, 521:136–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jian-Ming:1991:AAF

[814] Gao Jian-Ming and A. Nakamura. Algebraic analysis of fuzzy indiscernibility. Lecture Notes in Computer Science, 521:143–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ovchinnikov:1991:MFP

[815] S. Ovchinnikov. On modelling fuzzy preference relations. Lecture Notes in Computer Science, 521:154–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilar-Martin:1991:CCA

[816] J. Aguilar-Martin, M. Martin, and N. Piera. Conceptual connectivity analysis by means of fuzzy partitions. Lecture Notes in Computer Science, 521:165–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

diNola:1991:TSR

[817] A. di Nola, W. Kolodziejczyk, and S. Sessa. Transitive solutions of relational equations on finite sets and linear lattices. Lecture Notes in Computer Science, 521:173–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wygralak:1991:GCN

[818] M. Wygralak. Generalized cardinal numbers and their ordering. Lecture Notes in Computer Science, 521:183– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1991:IBA

[819] A. Gonzalez and M.-A. Vila. An interval-based approach for working with fuzzy numbers. Lecture Notes in Computer Science, 521:193–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kyburg:1991:BS

[820] H. E. Kyburg. Beyond specificity. Lecture Notes in Computer Science, 521: 204–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregoire:1991:ALI

[821] E. Gregoire. About the logical interpretation of ambiguous inheritance hierarchies. Lecture Notes in Computer Science, 521:213–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1991:NMR

[822] P. Balbiani. Non-monotonic reasoning and modal logic, from negation as failure to default logic. Lecture Notes in Computer Science, 521:223–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernasconi:1991:NUB

[823] C. Bernasconi, S. Rivoira, and S. Termini. On the notion of uncertain belief revision systems. Lecture Notes in Computer Science, 521:232–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cori:1991:NMR

[824] M. Cori. Non-monotonic reasoning in a semantic network. Lecture Notes in Computer Science, 521:239–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1991:IPH

[825] D. Dubois and H. Prade. Inference in possibilistic hypergraphs. Lecture Notes in Computer Science, 521:250– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1991:SEP

[826] J. Lang. Semantic evaluation in possibilistic logic, application to min-max discrete optimisation problems. Lecture Notes in Computer Science, 521: 260–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agusti-Cullell:1991:FMV

[827] J. Agusti-Cullell, F. Esteva, P. Garcia, and L. Godo. Formalizing multiplevalued logics as institutions. Lecture Notes in Computer Science, 521:269– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bottoni:1991:EPR

[828] P. Bottoni, L. Mari, and P. Mussio. Empirical plausible reasoning by multiple-valued logic. Lecture Notes in Computer Science, 521:279–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anger:1991:TTR

[829] F. D. Anger and R. V. Rodriguez. Time, tense and relativity revisited. Lecture Notes in Computer Science, 521:286–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Longo:1991:IMB

[830] G. Longo. Information and the mindbody problem. Lecture Notes in Computer Science, 521:298–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benvenuti:1991:GIF

[831] P. Benvenuti, D. Vivona, and M. Divari. A general information for fuzzy sets. Lecture Notes in Computer Science, 521:307–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramer:1991:ITB

[832] A. Ramer. Information theory based on fuzzy (possibilistic) rules. Lecture Notes in Computer Science, 521:317– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morrissey:1991:MUG

[833] J. M. Morrissey. Measuring uncertainty given imprecise attribute values. Lecture Notes in Computer Science, 521:327–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zighed:1991:DOL

[834] A. Zighed, D. Tounissoux, J. P. Auray, and C. Largeron. Discrimination by optimizing a local consistency criterion. Lecture Notes in Computer Science, 521:337–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forte:1991:MLI

[835] B. Forte and V. Kolbas. Minimum loss of information and image segmentation. Lecture Notes in Computer Science, 521:349–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1991:TGT

[836] J. F. Baldwin. Towards a general theory of evidential reasoning. Lecture Notes in Computer Science, 521:360– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Longo:1991:PWM

[837] G. Longo and A. Sgarro. A pragmatic way out of the maze of uncertainty measures. Lecture Notes in Computer Science, 521:370–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderLubbe:1991:MRM

[838] J. C. A. van der Lubbe, E. Backer, and W. Krijgsman. Models for reasoning with multitype uncertainty in expert systems. Lecture Notes in Computer Science, 521:377–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saffiotti:1991:HBS

[839] A. Saffiotti. A hybrid belief system for doubtful agents. Lecture Notes in Computer Science, 521:393–402, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roos:1991:HRU

[840] N. Roos. How to reason with uncertain knowledge. Lecture Notes in Computer Science, 521:403–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ligomenides:1991:CUR

[841] P. A. Ligomenides. Computation and uncertainty in regulated synergetic machines. Lecture Notes in Computer Science, 521:413–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacprzyk:1991:ILI

[842] J. Kacprzyk and C. Iwanski. Inductive learning from incomplete and imprecise examples. Lecture Notes in Computer Science, 521:424–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosc:1991:SAE

[843] P. Bosc and O. Pivert. Some algorithms for evaluating fuzzy relational queries. Lecture Notes in Computer Science, 521:431–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1991:TID

[844] M. H. Williams and Q. Kong. Time and incompleteness in a deductive database. Lecture Notes in Computer Science, 521:443–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yager:1991:ALS

[845] R. R. Yager, K. M. Ford, and A. J. Canas. An approach to the linguistic summarization of data. Lecture Notes in Computer Science, 521:456– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terenziani:1991:MUA

[846] P. Terenziani, L. Lesmo, and E. Gerbino. Management of uncertainty in the attachment problem in natural language processing. Lecture Notes in Computer Science, 521:469–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacas:1991:TAS

[847] J. Jacas and J. Recasens. A topological approach to some cluster methods. Lecture Notes in Computer Science, 521:479–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binaghi:1991:FKB

[848] E. Binaghi, A. della Ventura, A. Rampini, and R. Schettini. A fuzzy knowledge-based system for biomedical image interpretation. Lecture Notes in Computer Science, 521:488–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernard-Weil:1991:MCS

[849] E. Bernard-Weil. Management of chaotic systems with the model for the regulation of agonistic antagonistic couples. Lecture Notes in Computer Science, 521:498–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearman:1991:FMT

[850] A. D. Pearman, J. Montero, and J. Tejada. Fuzzy multicriteria techniques: An application to transport planning. Lecture Notes in Computer Science, 521:510–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salski:1991:FLA

[851] A. Salski and C. Sperlbaum. Fuzzy logic approach to modelling in ecosystem research. Lecture Notes in Computer Science, 521:520–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kershenbaum:1991:AFM

[852] A. Kershenbaum and T. Rubinson. An application of fuzzy multi-criteria decision making for topological design of large networks. Lecture Notes in Computer Science, 521:528–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esposito:1991:DMD

[853] F. Esposito, D. Malerba, and G. Semeraro. A distance measure for decision making in uncertain domains. Lecture Notes in Computer Science, 521: 538–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerf:1991:DLP

[854] M. Cerf, S. Poitrenaud, J.-F. Richard, M. Sebillotte, and C. A. Tijus. Decisions and lack of precision in crop management: The role of processing both objects and procedures through semantic networks. Lecture Notes in Computer Science, 521:548–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallant:1991:NNE

[855] S. I. Gallant and Y. Hayashi. A neural network expert system with confidence measurements. Lecture Notes in Computer Science, 521:562–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Console:1991:DUD

[856] L. Console, C. Borlo, A. Casale, and P. Torasso. Dealing with uncertainty in a distributed expert system architecture. Lecture Notes in Computer Science, 521:568–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Machado:1991:CNN

[857] R. J. Machado and A. F. da Rocha. The combinatorial neural network: A connectionist model for knowledge based systems. Lecture Notes in Computer Science, 521:578–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1991:MDA

[858] M. E. Cohen and D. L. Hudson. A medical decision aid based on a neural network model. Lecture Notes in Computer Science, 521:588–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHagan:1991:FNB

[859] M. O’Hagan. A fuzzy neuron based upon maximum entropy ordered weighted averaging. Lecture Notes in Computer Science, 521:598–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aylett:1991:HHE

[860] Ruth Aylett, Alan Fish, and Simon Bartrum. HELP — A hierarchical execution-led planner for robotic domains. Lecture Notes in Computer Science, 522:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Downs:1991:ICR

[861] Joseph Downs and Han Reichgelt. Integrating classical and reactive planning within an architecture for autonomous agents. Lecture Notes in Computer Science, 522:13–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drabble:1991:AAP

[862] Brian Drabble and Richard Kirby. Associating AI planner entities with an underlying time point network. Lecture Notes in Computer Science, 522:27–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fronhofer:1991:DCM

[863] Bertram Fronhofer. Default connections in a modal planning framework. Lecture Notes in Computer Science, 522:39–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghassem-Sani:1991:RP

[864] Reza Ghassem-Sani and Sam Steel. Recursive plans. Lecture Notes in Computer Science, 522:53–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gobel:1991:LRP

[865] Richard Gobel and Robert Rodosek. A language for representing planning problems. Lecture Notes in Computer Science, 522:64–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosse:1991:TTS

[866] Gerd Große and Richard Waldinger. Towards a theory of simultaneous actions. Lecture Notes in Computer Science, 522:78–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Materne:1991:MCE

[867] Stefan Materne and Joachim Hertzberg. MTMM — correcting and extending time map management. Lecture Notes in Computer Science, 522: 88–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Regnier:1991:CDP

[868] Pierre Regnier and Bernard Fade. Complete determination of parallel actions and temporal optimization in linear plans of action. Lecture Notes in Computer Science, 522:100–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steel:1991:KSP

[869] Sam Steel. Knowledge subgoals in plans. Lecture Notes in Computer Science, 522:112–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1991:TCO

[870] Tobias Nipkow and Gregor Snelting. Type classes and overloading resolution via order-sorted unification. Lecture Notes in Computer Science, 523:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volpano:1991:CMT

[871] Dennis M. Volpano and Geoffrey S. Smith. On the complexity of ML typability with overloading. Lecture Notes in Computer Science, 523:15–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thatte:1991:CTI

[872] Satish R. Thatte. Coercive type isomorphism. Lecture Notes in Computer Science, 523:29–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schauser:1991:CCM

[873] Klaus Erik Schauser, David E. Culler, and Thorsten von Eicken. Compilercontrolled multithreading for lenient parallel languages. Lecture Notes in Computer Science, 523:50–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traub:1991:MTC

[874] Kenneth R. Traub. Multi-thread code generation for dataflow architectures from non-strict programs. Lecture Notes in Computer Science, 523:73–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maranget:1991:GPI

[875] Luc Maranget. GAML: a parallel implementation of lazy ML. Lecture Notes in Computer Science, 523:102– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meijer:1991:FPB

[876] Erik Meijer, Maarten Fokkinga, and Ross Paterson. Functional programming with bananas, lenses, envelopes and barbed wire. Lecture Notes in Computer Science, 523:124–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Launchbury:1991:STS

[877] John Launchbury. A strongly-typed self-applicable partial evaluator. Lecture Notes in Computer Science, 523: 145–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weise:1991:AOP

[878] Daniel Weise, Roland Conybeare, Erik Ruf, and Scott Seligman. Automatic online partial evaluation. Lecture Notes in Computer Science, 523: 165–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swarup:1991:AAL

[879] Vipin Swarup, Uday S. Reddy, and Evan Ireland. Assignments for applicative languages. Lecture Notes in Computer Science, 523:192–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wakeling:1991:LL

[880] David Wakeling and Colin Runciman. Linearity and laziness. Lecture Notes in Computer Science, 523:215–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fradet:1991:SDS

[881] Pascal Fradet. Syntactic detection of single-threading using continuations. Lecture Notes in Computer Science, 523:241–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lillie:1991:PMT

[882] David J. Lillie and Peter G. Harrison. A projection model of types. Lecture Notes in Computer Science, 523:259– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frandsen:1991:WEI

[883] Gudmund S. Frandsen and Carl Sturtivant. What is an efficient implementation of the lambda-calculus? Lecture Notes in Computer Science, 523:289–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mairson:1991:OPT

[884] Harry G. Mairson. Outline of a proof theory of parametricity. Lecture Notes in Computer Science, 523:313– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leung:1991:RAS

[885] Allen Leung and Prateek Mishra. Reasoning about simple and exhaustive demand in higher-order lazy languages. Lecture Notes in Computer Science, 523:328–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1991:SAL

[886] Thomas P. Jensen. Strictness analysis in logical form. Lecture Notes in Computer Science, 523:352–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraki:1991:NAI

[887] Gebreselassie Baraki. A note on abstract interpretation of polymorphic functions. Lecture Notes in Computer Science, 523:367–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aditya:1991:IP

[888] Shail Aditya and Rishiyur S. Nikhil. Incremental polymorphism. Lecture Notes in Computer Science, 523:379– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leroy:1991:DM

[889] Xavier Leroy and Michel Mauny. Dynamics in ML. Lecture Notes in Computer Science, 523:406–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiken:1991:IRT

[890] Alexander Aiken and Brian R. Murphy. Implementing regular tree expressions. Lecture Notes in Computer Science, 523:427–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henglein:1991:ETI

[891] Fritz Henglein. Efficient type inference for higher-order binding-time analysis. Lecture Notes in Computer Science, 523:448–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holst:1991:FA

[892] Carsten Kehler Holst. Finiteness analysis. Lecture Notes in Computer Science, 523:473–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Consel:1991:BSS

[893] Charles Consel and Olivier Danvy. For a better support of static data flow. Lecture Notes in Computer Science, 523:496–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiueh:1991:ATC

[894] Tzi cker Chiueh. An architectural technique for cache-level garbage collection. Lecture Notes in Computer Science, 523:520–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1991:MSE

[895] Paul S. Barth, Rishiyur S. Nikhil, and Arvind. M-structures: Extending a parallel, non-strict, functional language with state. Lecture Notes in Computer Science, 523:538–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heytens:1991:LCA

[896] Michael L. Heytens and Rishiyur S. Nikhil. List comprehensions in agna, A parallel persistent object system. Lecture Notes in Computer Science, 523: 569–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smetsers:1991:GEC

[897] Sjaak Smetsers, Eric Nocker, John van Groningen, and Rinus Plasmeijer. Generating efficient code for lazy functional languages. Lecture Notes in Computer Science, 523:592–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hannan:1991:MAM

[898] John Hannan. Making abstract machines less abstract. Lecture Notes in Computer Science, 523:618–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1991:UVF

[899] Simon L. Peyton Jones and John Launchbury. Unboxed values as first class citizens in a non-strict functional language. Lecture Notes in Computer Science, 523:636–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1991:DSE

[900] C. Andre. Delays in synchronized elementary net systems. Lecture Notes in Computer Science, 524:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradfield:1991:PTP

[901] J. C. Bradfield. Proving temporal properties of Petri nets. Lecture Notes in Computer Science, 524:29–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carstensen:1991:CTE

[902] H. Carstensen. The complexity of testing the equivalence of transition sequences. Lecture Notes in Computer Science, 524:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chehaibar:1991:URN

[903] G. Chehaibar. Use of reentrant nets in modular analysis of colored nets. Lecture Notes in Computer Science, 524: 58–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desel:1991:AN

[904] J. Desel. On abstractions of nets. Lecture Notes in Computer Science, 524: 78–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimitrovici:1991:SCN

[905] C. Dimitrovici, U. Hummert, and L. Petrucci. Semantics, composition and net properties of algebraic highlevel nets. Lecture Notes in Computer Science, 524:93–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1991:TSL

[906] J. Esparza and M. Silva. Top-down synthesis of live and bounded free choice nets. Lecture Notes in Computer Science, 524:118–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

diGiovanni:1991:HN

[907] R. di Giovanni. Hood nets. Lecture Notes in Computer Science, 524:140– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hopkins:1991:DN

[908] R. P. Hopkins. Distributable nets. Lecture Notes in Computer Science, 524: 161–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itter:1991:CSD

[909] F. Itter and C. Relewicz. Computer supported design of kanban controlled production. Lecture Notes in Computer Science, 524:188–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janicki:1991:OSN

[910] R. Janicki and M. Koutny. Optimal simulations, nets and reachability graphs. Lecture Notes in Computer Science, 524:205–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinci:1991:ISD

[911] V. O. Pinci and R. M. Shapiro. An integrated software development methodology based on hierarchical colored Petri nets. Lecture Notes in Computer Science, 524:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1991:NSG

[912] E. Smith. On net systems generated by process foldings. Lecture Notes in Computer Science, 524:253–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Souissi:1991:LPC

[913] Y. Souissi. On liveness preservation by composition of nets via a set of places. Lecture Notes in Computer Science, 524:277–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valette:1991:SIP

[914] R. Valette and B. Bako. Software implementation of Petri nets and compilation of rule-based systems. Lecture Notes in Computer Science, 524:296– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plunnecke:1991:BPN

[915] H. Plunnecke and W. Reisig. Bibliography on Petri nets 1990. Lecture Notes in Computer Science, 524:317– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ottmann:1991:GAT

[916] Th. Ottmann. Geometric algorithms and their complexity. Lecture Notes in Computer Science, 525:3–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriegel:1991:CSA

[917] H.-P. Kriegel, Th. Brinkhoff, and R. Schneider. The combination of spatial access methods and computational geometry in geographic database systems. Lecture Notes in Computer Science, 525:5–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheiney:1991:FQN

[918] J. P. Cheiney and A. Touir. FI-Quadtree: A new data structure for content-oriented retrieval and fuzzy search. Lecture Notes in Computer Science, 525:23–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radermacher:1991:IME

[919] F. J. Radermacher. The importance of metaknowledge for environmental information systems. Lecture Notes in Computer Science, 525:35–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1991:EOO

[920] A. Gupta, T. E. Weymouth, and R. Jain. An extended object-oriented data model for large image bases. Lecture Notes in Computer Science, 525: 45–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1991:IGR

[921] Q. Zhou and B. J. Garner. On the integration for GIS and remotely sensed data: Towards an integrated system to handle the large volume of spatial data. Lecture Notes in Computer Science, 525:63–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voisard:1991:TTG

[922] A. Voisard. Towards a toolbox for geographic user interfaces. Lecture Notes in Computer Science, 525:75–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calcinelli:1991:MAG

[923] D. Calcinelli and M. Mainguenaud. The management of the ambiguities in a graphical query language for geographical information systems. Lecture Notes in Computer Science, 525:99–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Svensson:1991:GSQ

[924] P. Svensson and H. Zhexue. Geo-SAL: A query language for spatial data analysis. Lecture Notes in Computer Science, 525:119–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egenhofer:1991:RAB

[925] M. J. Egenhofer. Reasoning about binary topological relations. Lecture Notes in Computer Science, 525:143– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dutta:1991:TCR

[926] S. Dutta. Topological constraints: A representational framework for approximative spatial and temporal reasoning. Lecture Notes in Computer Science, 525:161–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lomet:1991:GPI

[927] D. B. Lomet. Grow and post index trees: Role, techniques and future potential. Lecture Notes in Computer Science, 525:183–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1991:SLS

[928] K.-J. Li and R. Laurini. The spatial locality and a spatial indexing method by dynamic clustering in hypermap systems. Lecture Notes in Computer Science, 525:207–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frank:1991:PGD

[929] A. U. Frank. Properties of geographic data: Requirements for spatial access methods. Lecture Notes in Computer Science, 525:225–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoel:1991:EPS

[930] E. G. Hoel and H. Samet. Efficient processing of spatial queries in line segment databases. Lecture Notes in Computer Science, 525:237–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriegel:1991:POD

[931] H.-P. Kriegel, H. Horn, and M. Schiwietz. The performance of object decomposition techniques for spatial query processing. Lecture Notes in Computer Science, 525:257–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seeger:1991:PCS

[932] B. Seeger. Performance comparison of segment access methods implemented on top of the buddy-tree. Lecture Notes in Computer Science, 525:277– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aref:1991:EDS

[933] W. G. Aref and H. Samet. Extending a DBMS with spatial operations. Lecture Notes in Computer Science, 525:299–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lorie:1991:UCO

[934] R. A. Lorie. The use of a complex object language in geographic data management. Lecture Notes in Computer Science, 525:319–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shekhar:1991:MGD

[935] S. Shekhar and T. A. Yang. Motion in a geographical database system. Lecture Notes in Computer Science, 525:339–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vanecek:1991:SIC

[936] G. Vanecek and V. Ferrucci. A spatial index for convex simplical complexes in d dimensions. Lecture Notes in Computer Science, 525:361–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orenstein:1991:ACO

[937] J. Orenstein. An algorithm for computing the overlay of k-dimensional spaces. Lecture Notes in Computer Science, 525:381–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deJong:1991:EMP

[938] W. de Jong, A. Schijf, and P. Scheuermann. Encoding and manipulating pictorial data with S^{+}-trees. Lecture Notes in Computer Science, 525:401– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haas:1991:EED

[939] L. M. Haas and W. F. Cody. Exploiting extensible DBMS in integrated geographic information systems. Lecture Notes in Computer Science, 525:423– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1991:SMG

[940] H. Lu, B. C. Ooi, A. D’Souza, and C. C. Low. Storage management in geographic information systems. Lecture Notes in Computer Science, 525: 451–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samet:1991:PWD

[941] H. Samet. Panel: Why does GIS industry ignore university research? Lecture Notes in Computer Science, 525: 471–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plotkin:1991:STC

[942] Gordon Plotkin. A semantics for type checking. Lecture Notes in Computer Science, 526:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giannini:1991:TIP

[943] Paola Giannini and Simona Ronchi Della Rocca. Type inference in polymorphic type discipline. Lecture Notes in Computer Science, 526:18–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tatsuta:1991:MRD

[944] Makoto Tatsuta. Monotone recursive definition of predicates and its realizability interpretation. Lecture Notes in Computer Science, 526:38–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1991:APO

[945] Masahiko Sato. Adding proof objects and inductive definition mechanisms to Frege structures. Lecture Notes in Computer Science, 526:53–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phoa:1991:TMD

[946] Wesley Phoa. From term models to domains. Lecture Notes in Computer Science, 526:88–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunter:1991:AIM

[947] Carl A. Gunter, Elsa L. Gunter, and David B. MacQueen. An abstract interpretation for ML equality kinds. Lecture Notes in Computer Science, 526: 112–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jim:1991:FAC

[948] Trevor Jim and Albert R. Meyer. Full abstraction and the context lemma. Lecture Notes in Computer Science, 526:131–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arun-Kumar:1991:EPP

[949] S. Arun-Kumar and M. Hennessy. An efficiency preorder for processes. Lecture Notes in Computer Science, 526: 152–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabinovich:1991:NAM

[950] Alexander Rabinovich and Boris A. Trakhtenbrot. On nets, algebras and modularity. Lecture Notes in Computer Science, 526:176–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1991:TCH

[951] Bengt Jonsson and Joost N. Kok. Towards a complete hierarchy of compositional dataflow models. Lecture Notes in Computer Science, 526:204– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Constable:1991:TTF

[952] Robert L. Constable. Type theory as a foundation for computer science. Lecture Notes in Computer Science, 526:226–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pnueli:1991:WSS

[953] A. Pnueli and M. Shalev. What is in a step: On the semantics of statecharts. Lecture Notes in Computer Science, 526:244–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apt:1991:PTG

[954] Krzysztof R. Apt and Dino Pedreschi. Proving termination of general Prolog programs. Lecture Notes in Computer Science, 526:265–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1991:AEP

[955] John C. Mitchell. On abstraction and the expressive power of programming languages. Lecture Notes in Computer Science, 526:290–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchi:1991:RLP

[956] Kazuhiro Fuchi and Koichi Furukawa. Role of logic programming in the FGCS project. Lecture Notes in Computer Science, 526:311–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1991:ADS

[957] M. Abadi, M. Burrows, C. Kaufman, and B. Lampson. Authentication and delegation with smart-cards. Lecture Notes in Computer Science, 526:326– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1991:DFA

[958] Bernhard Steffen. Data flow analysis as model checking. Lecture Notes in Computer Science, 526:346–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfeiffer:1991:ADB

[959] Phil Pfeiffer and Rebecca Parsons Selke. On the adequacy of dependencebased representations for programs with heaps. Lecture Notes in Computer Science, 526:365–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagiya:1991:PEP

[960] Masami Hagiya. From programmingby-example to proving-by-example. Lecture Notes in Computer Science, 526:387–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lassez:1991:LLP

[961] Jean-Louis Lassez. From LP to LP: Programming with constraints. Lecture Notes in Computer Science, 526: 420–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armbruster:1991:PRA

[962] Dieter Armbruster. Polynomial recursion analysis in Pascal like programs. Lecture Notes in Computer Science, 526:447–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hungar:1991:CPPa

[963] Hardi Hungar. Complexity of proving program correctness. Lecture Notes in Computer Science, 526:459–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1991:SNP

[964] David A. Basin and Douglas J. Howe. Some normalization properties of Martin-Lof’s type theory, and applications. Lecture Notes in Computer Science, 526:475–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1991:PEC

[965] Ryu Hasegawa. Parametricity of extensionally collapsed term models of polymorphism and their categorical properties. Lecture Notes in Computer Science, 526:495–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishizaki:1991:PCL

[966] Shin ya Nishizaki. Programs with continuations and linear logic. Lecture Notes in Computer Science, 526:513– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walker:1991:PCS

[967] David Walker. pi-calculus semantics of object-oriented programming languages. Lecture Notes in Computer Science, 526:532–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hense:1991:WSO

[968] Andreas V. Hense. Wrapper semantics of an object-oriented programming language with state. Lecture Notes in Computer Science, 526:548–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryan:1991:SAA

[969] Mark Ryan, Jose Fiadeiro, and Tom Maibaum. Sharing actions and attributes in modal action logic. Lecture Notes in Computer Science, 526: 569–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treinen:1991:FOD

[970] Ralf Treinen. First order data types and first order logic. Lecture Notes in Computer Science, 526:594–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanovich:1991:EPS

[971] Max I. Kanovich. Efficient program synthesis: Semantics, logic, complexity. Lecture Notes in Computer Science, 526:615–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirokawa:1991:PTS

[972] Sachio Hirokawa. Principal typeschemes of BCI-Lambda-Terms. Lecture Notes in Computer Science, 526: 633–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbanera:1991:IUT

[973] Franco Barbanera and Mariangiola Dezani-Ciancaglini. Intersection and union types. Lecture Notes in Computer Science, 526:651–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynolds:1991:CLI

[974] John C. Reynolds. The coherence of languages with intersection types. Lecture Notes in Computer Science, 526: 675–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1991:SUI

[975] Susumu Hayashi. Singleton, union and intersection types for program extraction. Lecture Notes in Computer Science, 526:701–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curien:1991:SEC

[976] Pierre-Louis Curien and Giorgio Ghelli. Subtyping + extensionality: Confluence of betaeta top reduction in flsim. Lecture Notes in Computer Science, 526:731–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardelli:1991:ESF

[977] Luca Cardelli, Simone Martini, John C. Mitchell, and Andre Scedrov. An extension of system F with subtyping. Lecture Notes in Computer Science, 526:750–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scott:1991:WLR

[978] Dana S. Scott. Will logicians be replaced by machines? (abstract). Lecture Notes in Computer Science, 526: 771–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

America:1991:FTP

[979] P. America. Formal techniques for parallel object-oriented languages. Lecture Notes in Computer Science, 527:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouali:1991:CMR

[980] A. Bouali and R. de Simone. Causal models for rational algebraic processes. Lecture Notes in Computer Science, 527:18–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1991:ASB

[981] R. De Nicola. Action and state-based logics for process algebras. Lecture Notes in Computer Science, 527:20–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1991:TSD

[982] J.-C. Fernandez and L. Mounier. A tool set for deciding behavioral equivalences. Lecture Notes in Computer Science, 527:23–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goltz:1991:CBM

[983] U. Goltz. Causality based models for the design of concurrent systems. Lecture Notes in Computer Science, 527:

43–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milner:1991:MLM

[984] R. Milner, J. Parrow, and D. Walker. Modal logics for mobile processes. Lecture Notes in Computer Science, 527: 45–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olderog:1991:TDC

[985] E.-R. Olderog. Towards a design calculus for communicating programs. Lecture Notes in Computer Science, 527: 61–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1991:TTA

[986] L. Aceto and A. Ingolfsdottir. A theory of testing for ACP. Lecture Notes in Computer Science, 527:78–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1991:RSP

[987] J. C. M. Baeten and J. A. Bergstra. Real space process algebra. Lecture Notes in Computer Science, 527:96–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBoer:1991:FFP

[988] F. S. de Boer, J. N. Kok, C. Palamidessi, and J. J. M. M. Rutten. The failure of failures in a paradigm for asynchronous communication. Lecture Notes in Computer Science, 527:111– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBoer:1991:ETL

[989] F. S. de Boer and C. Palamidessi. Embedding as a tool for language comparison: On the CSP hierarchy. Lecture Notes in Computer Science, 527:127– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camilleri:1991:COC

[990] J. Camilleri. A conditional operator for CCS. Lecture Notes in Computer Science, 527:142–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drost:1991:AFT

[991] N. J. Drost. Algebraic formulations of trace theory. Lecture Notes in Computer Science, 527:157–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1991:CSL

[992] J. Esparza and M. Silva. Compositional synthesis of live and bounded free choice Petri nets. Lecture Notes in Computer Science, 527:172–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1991:OAS

[993] G. L. Ferrari and U. Montanari. The observation algebra of spatial pomsets. Lecture Notes in Computer Science, 527:188–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Francez:1991:SLT

[994] N. Francez and I. R. Forman. Synchrony loosening transformations for interacting processes. Lecture Notes in Computer Science, 527:203–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldman:1991:CML

[995] K. J. Goldman. A compositional model for layered distributed systems. Lecture Notes in Computer Science, 527: 220–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groote:1991:PAG

[996] J. F. Groote and A. Ponse. Process algebra with guards. Lecture Notes in Computer Science, 527:235–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumberg:1991:MCM

[997] O. Grumberg and D. E. Long. Model checking and modular verification. Lecture Notes in Computer Science, 527: 250–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunawardena:1991:GLC

[998] J. Gunawardena. Geometric logic, causality and event structures. Lecture Notes in Computer Science, 527: 266–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquet:1991:EHC

[999] J.-M. Jacquet and L. Monteiro. Extended Horn clauses: the framework and some semantics. Lecture Notes in Computer Science, 527:281–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssen:1991:ASA

[1000] W. Janssen, M. Poel, and J. Zwiers.

Action systems and action refinement in the development of parallel systems. Lecture Notes in Computer Science, 527:298–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janicki:1991:ISN

[1001] R. Janicki and M. Koutny. Invariant semantics of nets with inhibitor arcs. Lecture Notes in Computer Science, 527:317–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffrey:1991:ATO

[1002] A. Jeffrey. Abstract timed observation and process algebra. Lecture Notes in Computer Science, 527:332–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1991:SBS

[1003] B. Jonsson. Simulations between specifications of distributed systems. Lecture Notes in Computer Science, 527: 346–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaivola:1991:UTP

[1004] R. Kaivola and A. Valmari. Using truth-preserving reductions to improve the clarity of Kripke-models. Lecture Notes in Computer Science, 527:361– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klusener:1991:CRT

[1005] A. S. Klusener. Completeness in real time process algebra. Lecture Notes in Computer Science, 527:376–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnan:1991:DC

[1006] P. Krishnan. Distributed CCS. Lecture Notes in Computer Science, 527: 393–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merritt:1991:TCA

[1007] M. Merritt, F. Modugno, and M. R. Tuttle. Time-constrained automata. Lecture Notes in Computer Science, 527:408–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:1991:RPR

[1008] F. Moller and C. Tofts. Relating processes with respect to speed. Lecture Notes in Computer Science, 527:424– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murphy:1991:TBT

[1009] D. Murphy. Testing, betting and timed true concurrency. Lecture Notes in Computer Science, 527:439–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolte:1991:FMT

[1010] D. Nolte and L. Priese. Fairness in models with true concurrency. Lecture Notes in Computer Science, 527:455– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qin:1991:EVD

[1011] H. Qin. Efficient verification of determinate processes. Lecture Notes in Computer Science, 527:470–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnoebelen:1991:EPB

[1012] Ph. Schnoebelen. Experiments on processes with backtracking. Lecture Notes in Computer Science, 527:480– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spruit:1991:SFG

[1013] P. Spruit and R. Wieringa. Some finite-graph models for process algebra. Lecture Notes in Computer Science, 527:495–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stolen:1991:MDT

[1014] K. Stolen. A method for the development of totally correct shared-state parallel programs. Lecture Notes in Computer Science, 527:510–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valmari:1991:RLT

[1015] A. Valmari and M. Clegg. Reduced labelled transition systems save verification effort. Lecture Notes in Computer Science, 527:526–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appel:1991:ILS

[1016] A. W. Appel and D. B. MacQueen. Invited lecture: Standard ML of new jersey. Lecture Notes in Computer Science, 528:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naish:1991:AEN

[1017] L. Naish. Adding equations to NU-Prolog. Lecture Notes in Computer Science, 528:15–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamasaki:1991:EFL

[1018] S. Yamasaki. Extraction of functional from logic program. Lecture Notes in Computer Science, 528:27–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kredel:1991:MSC

[1019] H. Kredel. The Mas specification component. Lecture Notes in Computer Science, 528:39–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1991:DIC

[1020] T. C. N. Graham and G. Kock. Domesticating imperative constructs so that they can live in a functional world. Lecture Notes in Computer Science, 528:51–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poetzsch-Heffter:1991:LBS

[1021] A. Poetzsch-Heffter. Logic-based specification of visibility rules. Lecture Notes in Computer Science, 528:63–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ertl:1991:OIS

[1022] M. A. Ertl and A. Krall. Optimal instruction scheduling using constraint logic programming. Lecture Notes in Computer Science, 528:75–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baiardi:1991:AMP

[1023] F. Baiardi and D. M. Di Bella. An architectural model for OR-Parallellism on distributed memory systems. Lecture Notes in Computer Science, 528: 87–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langendoen:1991:FPR

[1024] K. G. Langendoen and W. G. Vree.

FRATS: A parallel reduction strategy for shared memory. Lecture Notes in Computer Science, 528:99–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpuente:1991:NIC

[1025] M. Alpuente and M. Falaschi. Narrowing as an incremental constraint satisfaction algorithm. Lecture Notes in Computer Science, 528:111–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chakravarty:1991:ILN

[1026] M. M. T. Chakravarty and H. C. R.

Lock. The implementation of lazy narrowing. Lecture Notes in Computer Science, 528:123–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Consel:1991:SDG

[1027] C. Consel and Cheng Khoo Siau. Semantics-directed generation of a Prolog compiler. Lecture Notes in Computer Science, 528:135–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boerger:1991:FSD

[1028] E. Boerger and B. Demoen. A framework to specify database update views for Prolog. Lecture Notes in Computer Science, 528:147–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferenczi:1991:CMD

[1029] S. Ferenczi. Concepts for a modular and distributed Prolog language. Lecture Notes in Computer Science, 528: 159–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davison:1991:PPT

[1030] A. Davison. From Parlog to polka in two easy steps. Lecture Notes in Computer Science, 528:171–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aasa:1991:PSI

[1031] A. Aasa. Precedences in specifications and implementations of programming languages. Lecture Notes in Computer Science, 528:183–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kebler:1991:RHA

[1032] C. W. Kebler, W. J. Paul, and T. Rauber. A randomized heuristic approach to register allocation. Lecture Notes in Computer Science, 528:195– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halbwachs:1991:GEC

[1033] N. Halbwachs, P. Raymond, and C. Ratel. Generating efficient code from data-flow programs. Lecture Notes in Computer Science, 528:207–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koster:1991:BBG

[1034] C. H. A. Koster and J. G. Beney. On the borderline between grammars and programs. Lecture Notes in Computer Science, 528:219–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogt:1991:EIE

[1035] H. Vogt, D. Swierstra, and M. Kuiper.

Efficient incremental evaluation of higher order attribute grammars. Lecture Notes in Computer Science, 528: 231–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paakki:1991:PSI

[1036] J. Paakki. PROFIT: A system integrating logic programming and attribute grammars. Lecture Notes in Computer Science, 528:243–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ait-Kaci:1991:ILT

[1037] H. Ait-Kaci and A. Podelski. Invited lecture: Towards a meaning of LIFE. Lecture Notes in Computer Science, 528:255–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gloess:1991:ULO

[1038] P. Y. Gloess. U-log, an ordered sorted logic with typed attributes. Lecture Notes in Computer Science, 528:275– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1991:CPA

[1039] W. Chen and D. S. Warren. Compilation of predicate abstractions in higher-order logic programming. Lecture Notes in Computer Science, 528: 287–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1991:AIK

[1040] U. Nilsson. Abstract interpretation: A kind of magic. Lecture Notes in Computer Science, 528:299–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

File:1991:AIT

[1041] G. File and P. Sottero. Abstract interpretation for type checking. Lecture Notes in Computer Science, 528:311– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giannotti:1991:TRI

[1042] F. Giannotti and M. Hermenegildo. A technique for recursive invariance detection and selective program specialization. Lecture Notes in Computer Science, 528:323–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loogen:1991:DDD

[1043] R. Loogen and S. Winkler. Dynamic detection of determinism in functional logic languages. Lecture Notes in Computer Science, 528:335–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Proietti:1991:UDF

[1044] M. Proietti and A. Pettorossi. Unfoldingdefinition-folding, in this order, for avoiding unnecessary variables in logic programs. Lecture Notes in Computer Science, 528:347–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheong:1991:EIS

[1045] P. H. Cheong and L. Fribourg. Efficient integration of simplification into Prolog. Lecture Notes in Computer Science, 528:359–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoy:1991:LEL

[1046] S. Antoy. Lazy evaluation in logic. Lecture Notes in Computer Science, 528: 371–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boye:1991:SRO

[1047] J. Boye. S-SLD-Resolution — an operational semantics for logic programs with external procedures. Lecture Notes in Computer Science, 528: 383–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHentenryck:1991:OSC

[1048] P. Van Hentenryck and Y. Deville. Operational semantics of constraint logic programming over finite domains. Lecture Notes in Computer Science, 528: 395–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1991:CSC

[1049] A. Brogi and M. Gabbrielli. Constraints for synchronizing coarsegrained sequential logic processes. Lecture Notes in Computer Science, 528: 407–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loyola:1991:EET

[1050] D. Loyola. Extending execution trees for debugging and animation in logic programming. Lecture Notes in Computer Science, 528:419–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraus:1991:NBT

[1051] U. Fraus and H. HuBmann. A narrowing-based theorem prover. Lecture Notes in Computer Science, 528: 421–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1991:AS

[1052] M. Hanus. The ALF system. Lecture Notes in Computer Science, 528:423– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vollmer:1991:EGE

[1053] J. Vollmer. Experiences with gentle:

Efficient compiler construction based on logic programming. Lecture Notes in Computer Science, 528:425–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baum:1991:SFF

[1054] B. Baum and P. Forbrig. The system FLR (fast laboratory for recomposition). Lecture Notes in Computer Science, 528:427–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attardi:1991:MI

[1055] G. Attardi and M. Gaspari. Multilanguage interoperability. Lecture Notes in Computer Science, 528:429–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Legeard:1991:SOC

[1056] B. Legeard and E. Legros. Short overview of the CLPS system. Lecture Notes in Computer Science, 528: 431–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allender:1991:SS

[1057] E. Allender and V. Gore. On strong separations from AC^{0}. Lecture Notes in Computer Science, 529:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchmann:1991:NTA

[1058] J. Buchmann. Number theoretic algorithms and cryptology. Lecture Notes in Computer Science, 529:16–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1991:CIG

[1059] J. y. Cai. Computations over infinite groups. Lecture Notes in Computer Science, 529:22–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinrich:1991:EMC

[1060] S. Heinrich. Efficiency of Monte Carlo algorithms in numerical analysis. Lecture Notes in Computer Science, 529: 33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karpinski:1991:AAC

[1061] M. Karpinski. Approximation algorithms for counting problems in finite fields. Lecture Notes in Computer Science, 529:45–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Razborov:1991:LBD

[1062] A. A. Razborov. Lower bounds for deterministic and nondeterministic branching programs. Lecture Notes in Computer Science, 529:47–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reischuk:1991:GTM

[1063] R. Reischuk. Graph theoretical methods for the design of parallel algorithms. Lecture Notes in Computer Science, 529:61–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnorr:1991:LBR

[1064] C. P. Schnorr and M. Euchner. Lattice basis reduction: Improved practical algorithms and solving subset sum problems. Lecture Notes in Computer Science, 529:68–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traub:1991:IBC

[1065] J. F. Traub. Information-based complexity: Recent results and open problems. Lecture Notes in Computer Science, 529:86–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turan:1991:SSA

[1066] G. Turan. A survey of some aspects of computational learning theory. Lecture Notes in Computer Science, 529:89–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baier:1991:CNI

[1067] C. Baier and M. E. Majster-Cederbaum.

The consistency of a noninterleaving and an interleaving model for full TCSP. Lecture Notes in Computer Science, 529:105–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bank:1991:GBI

[1068] B. Bank, T. Krick, R. Mandel, and P. Solerno. A geometrical bound for integer programming with polynomial constraints. Lecture Notes in Computer Science, 529:121–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barcucci:1991:CBS

[1069] E. Barcucci, R. Pinzani, E. Rodella, and R. Sprugnoli. A characterization of binary search networks. Lecture Notes in Computer Science, 529:126– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beauquier:1991:AEN

[1070] D. Beauquier, M. Nivat, and D. Niwinski. About the effect of the number of successful paths in an infinite tree on the recognizability by a finite automaton with Buchi conditions. Lecture Notes in Computer Science, 529:136–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breveglieri:1991:DDA

[1071] L. Breveglieri, C. Citrini, and S. Crespi-Reghizzi. Deterministic dequeue automata and LL(1) parsing of breadth-depth grammars. Lecture Notes in Computer Science, 529:146– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruschi:1991:CCM

[1072] D. Bruschi and G. Pighizzini. The complexity of computing maximal word functions. Lecture Notes in Computer Science, 529:157–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buntrock:1991:UFL

[1073] G. Buntrock, B. Jenner, K.-J. Lange, and P. Rossmanith. Unambiguity and fewness for logarithmic space. Lecture Notes in Computer Science, 529:168– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chardin:1991:DRS

[1074] M. Chardin. Differential resultants and subresultants. Lecture Notes in Computer Science, 529:180–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1991:UBS

[1075] B. S. Chlebus and I. Vrto. Unifying binary-search trees and permutations. Lecture Notes in Computer Science, 529:190–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dassow:1991:CCH

[1076] J. Dassow and K.-J. Lange. Computational complexity and hardest languages of automata with abstract storages. Lecture Notes in Computer Science, 529:200–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fachini:1991:STA

[1077] E. Fachini, A. Monti, M. Napoli, and D. Parente. Systolic Y-tree automata: Closure properties and decision problems. Lecture Notes in Computer Science, 529:210–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groger:1991:NPL

[1078] H. D. Groger. A new partition lemma for planar graphs and its application to circuit complexity. Lecture Notes in Computer Science, 529:220–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmeister:1991:SNT

[1079] T. Hofmeister, W. Hohberg, and S. Kohling. Some notes on threshold circuits, and multiplication in depth 4. Lecture Notes in Computer Science, 529:230–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1991:NLB

[1080] J. Hromkovic. Nonlinear lower bounds on the number of processors of circuits with sublinear separators. Lecture Notes in Computer Science, 529: 240–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1991:SBS

[1081] O. H. Ibarra and N. Q. Tran. On space-bounded synchronized alternating Turing machines. Lecture Notes in Computer Science, 529:248–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joux:1991:ICD

[1082] A. Joux and J. Stern. Improving the critical density of the Lagarias– Odlyzko attack against subset sum problems. Lecture Notes in Computer Science, 529:258–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jukna:1991:OVS

[1083] S. Jukna. Optimal versus stable in Boolean formulae. Lecture Notes in Computer Science, 529:265–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaib:1991:GLB

[1084] M. Kaib. The GauBeta lattice basis reduction algorithm succeeds with any given norm. Lecture Notes in Computer Science, 529:275–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaneps:1991:ROL

[1085] J. Kaneps. Regularity of one-letter languages acceptable by 2-Way finite probabilistic automata. Lecture Notes in Computer Science, 529:287– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knijnenburg:1991:SAS

[1086] P. Knijnenburg and J. N. Kok. On the semantics of atomized statements: The parallel-choice option. Lecture Notes in Computer Science, 529:297–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kounalis:1991:APM

[1087] E. Kounalis and M. Rusinowitch. Automatic proof methods for algebraic specifications. Lecture Notes in Computer Science, 529:307–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kratsch:1991:CGR

[1088] D. Kratsch and L. A. Hemachandra.

On the complexity of graph reconstruction. Lecture Notes in Computer Science, 529:318–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levcopoulos:1991:OAP

[1089] C. Levcopoulos and O. Petersson. An optimal adaptive in-place sorting algorithm. Lecture Notes in Computer Science, 529:329–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Louchard:1991:DSM

[1090] G. Louchard, C. Kenyon, and R. Schott. Data structures maxima. Lecture Notes in Computer Science, 529:339–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez:1991:ACA

[1091] C. Martinez. Average-case analysis of equality of binary trees under the BST probability model. Lecture Notes in Computer Science, 529:350–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neraud:1991:SRT

[1092] J. Neraud. On the subsets of rank two in a free monoid: A fast decision algorithm. Lecture Notes in Computer Science, 529:360–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plandowski:1991:EAT

[1093] W. Plandowski, W. Rytter, and T. Szymacha. Exact analysis of three tree contraction algorithms. Lecture Notes in Computer Science, 529:370– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomaa:1991:DNP

[1094] K. Salomaa and S. Yu. Degrees of nondeterminism for pushdown automata. Lecture Notes in Computer Science, 529:380–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroder:1991:OET

[1095] H. Schroder, O. Sykora, and I. Vrto.

Optimal embedding of a toroidal array in a linear array. Lecture Notes in Computer Science, 529:390–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uhlig:1991:BFL

[1096] D. Uhlig. Boolean functions with a large number of subfunctions and small complexity and depth. Lecture Notes in Computer Science, 529:395–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valiveti:1991:ALL

[1097] R. S. Valiveti, B. J. Oommen, and J. R. Zgierski. Adaptive linear list reorganisation for a system processing set queries. Lecture Notes in Computer Science, 529:405–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wanke:1991:DIS

[1098] E. Wanke. On the decidability of integer subgraph problems on contextfree graph languages. Lecture Notes in Computer Science, 529:415–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrhard:1991:SDS

[1099] T. Ehrhard and P. Malacaria. Stone duality for stable functions. Lecture Notes in Computer Science, 530:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1991:BDS

[1100] R. M. Amadio. Bifinite domains: Stable case. Lecture Notes in Computer Science, 530:16–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tennent:1991:LVN

[1101] R. D. Tennent. Local variables and non-interference in Algol-like languages. Lecture Notes in Computer Science, 530:34–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edalat:1991:CIS

[1102] A. Edalat and M. B. Smyth. Categories of information systems. Lecture Notes in Computer Science, 530:37–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoofman:1991:CGM

[1103] R. Hoofman and H. Schellinx. Collapsing graph models by preorders. Lecture Notes in Computer Science, 530:53–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHearn:1991:LLI

[1104] P. W. O’Hearn. Linear logic and interference control. Lecture Notes in Computer Science, 530:74–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burroni:1991:HDW

[1105] A. Burroni. Higher dimensional word problem. Lecture Notes in Computer Science, 530:94–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirokawa:1991:BFH

[1106] S. Hirokawa. BCK-Formulas having unique proofs. Lecture Notes in Computer Science, 530:106–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blute:1991:PNC

[1107] R. Blute. Proof nets and coherence theorems. Lecture Notes in Computer Science, 530:121–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moggi:1991:MAD

[1108] E. Moggi. A modular approach to denotational semantics. Lecture Notes in Computer Science, 530:138–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarzembski:1991:PPA

[1109] G. Jarzembski. Programs in partial algebras — a categorical approach. Lecture Notes in Computer Science, 530: 140–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jay:1991:TRU

[1110] C. B. Jay. Tail recursion from universal invariants. Lecture Notes in Computer Science, 530:151–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coquand:1991:DPI

[1111] T. Coquand. A direct proof of the intuitionistic Ramsey theorem. Lecture Notes in Computer Science, 530:164– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pavlovic:1991:CP

[1112] D. Pavlovic. Constructions and predicates. Lecture Notes in Computer Science, 530:173–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1991:RMI

[1113] B. Jacobs, E. Moggi, and T. Streicher. Relating models of impredicative type theories. Lecture Notes in Computer Science, 530:197–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phoa:1991:TRS

[1114] W. Phoa. Two results on set-theoretic polymorphism. Lecture Notes in Computer Science, 530:219–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1991:AGG

[1115] A. Corradini and U. Montanari. An algebra of graphs and graph rewriting. Lecture Notes in Computer Science, 530:236–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stark:1991:DNF

[1116] E. W. Stark. Dataflow networks are fibrations. Lecture Notes in Computer Science, 530:261–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degano:1991:ACT

[1117] P. Degano, S. Kasangian, and S. Vigna. Applications of the calculus of trees to process description languages. Lecture Notes in Computer Science, 530:282–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1991:TLM

[1118] Edmund M. Clarke. Temporal logic model checking: Two techniques for avoiding the state explosion problem. Lecture Notes in Computer Science, 531:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eveking:1991:AVE

[1119] Hans Eveking. Automatic verification of extensions of hardware descriptions. Lecture Notes in Computer Science, 531:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthelot:1991:PEA

[1120] G. Berthelot, C. Johnen, and Laure Petrucci. PAPETRI: Environment for the analysis of Petri nets. Lecture Notes in Computer Science, 531:13–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coudert:1991:VTP

[1121] Olivier Coudert, Jean Christophe Madre, and Christian Berthet. Verifying temporal properties of sequential machines without building their state diagrams. Lecture Notes in Computer Science, 531:23–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryant:1991:FVD

[1122] Randal E. Bryant and Carl-Johan Seger. Formal verification of digital circuits using symbolic ternary system models. Lecture Notes in Computer Science, 531:33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiraishi:1991:VMC

[1123] Hiromi Hiraishi, Shintaro Meki, and Kiyoharu Hamaguchi. Vectorized model checking for computation tree logic. Lecture Notes in Computer Science, 531:44–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pixley:1991:ICT

[1124] Carl Pixley. Introduction to a computational theory and implementation of sequential hardware equivalence. Lecture Notes in Computer Science, 531: 54–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roy:1991:AA

[1125] Valerie Roy and Robert de Simone. Auto/autograph. Lecture Notes in Computer Science, 531:65–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1991:DPV

[1126] Hiroshi Nakamura, Yuji Kukimoto, Masahiro Fujita, and Hidehiko Tanaka. A data path verifier for register transfer level using temporal logic language tokio. Lecture Notes in Computer Science, 531:76–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camurati:1991:UMC

[1127] Paolo Camurati, M. Gilli, P. Prinetto, and M. Sonza Reorda. The use of model checking in ATPG for sequential circuits. Lecture Notes in Computer Science, 531:86–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lloret:1991:CDV

[1128] Jean Christophe Lloret, Pierre Azema, and Francois Vernadat. Compositional design and verification of communication protocols, using labelled Petri nets. Lecture Notes in Computer Science, 531:96–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ness:1991:IAA

[1129] Linda Ness. Issues arising in the analysis of L.0. Lecture Notes in Computer Science, 531:106–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langevin:1991:ARV

[1130] Michel Langevin. Automated RTL verification based on predicate calculus. Lecture Notes in Computer Science, 531:116–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1991:UPV

[1131] Richard Lai, Ken R. Parker, and T. S.

Dillon. On using Protean to verify ISO FTAM protocol. Lecture Notes in Computer Science, 531:126–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1991:QTR

[1132] E. Allen Emerson, A. K. Mok, A. Prasad Sistla, and Jai Srinivasan. Quantitative temporal reasoning. Lecture Notes in Computer Science, 531: 136–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Probst:1991:UPO

[1133] David K. Probst and Hon F. Li. Using partial-order semantics to avoid the state explosion problem in asynchronous systems. Lecture Notes in Computer Science, 531:146–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valmari:1991:SAS

[1134] Antti Valmari. A stubborn attack on state explosion. Lecture Notes in Computer Science, 531:156–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janicki:1991:UOS

[1135] Ryszard Janicki and Maciej Koutny. Using optimal simulations to reduce reachability graphs. Lecture Notes in Computer Science, 531:166–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godefroid:1991:UPOa

[1136] Patrice Godefroid. Using partial orders to improve automatic verification methods. Lecture Notes in Computer Science, 531:176–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1991:CMF

[1137] Susanne Graf and Bernhard Steffen.

Compositional minimization of finite state systems. Lecture Notes in Computer Science, 531:186–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1991:MMG

[1138] A. Bouajjani, Jean-Claude Fernandez, and Nicholas Halbwachs. Minimal model generation. Lecture Notes in Computer Science, 531:197–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Josko:1991:CDE

[1139] Bernhard Josko. A context dependent equivalence relation between Kripke structures. Lecture Notes in Computer Science, 531:204–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shurek:1991:MFC

[1140] Gil Shurek and Orna Grumberg. The modular framework of computer-aided verification. Lecture Notes in Computer Science, 531:214–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burch:1991:VLP

[1141] Jerry R. Burch. Verifying liveness properties by verifying safety properties. Lecture Notes in Computer Science, 531:224–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcoubetis:1991:MEA

[1142] Costas Courcoubetis, Moshe Vardi, Pierre Wolper, and Mihalis Yannakakis. Memory efficient algorithms for the verification of temporal properties. Lecture Notes in Computer Science, 531:233–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuxu:1991:UAD

[1143] Peng Wuxu and S. Purushothaman. A unified approach to the deadlock detection problem in networks of communicating finite state machines. Lecture Notes in Computer Science, 531:243– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamaguchi:1991:BTR

[1144] Kiyoharu Hamaguchi, Hiromi Hiraishi, and Shuzo Yajima. Branching time regular temporal logic for model checking with linear time complexity. Lecture Notes in Computer Science, 531:253– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yodaiken:1991:AFP

[1145] Victor Yodaiken. The algebraic feedback product of automata. Lecture Notes in Computer Science, 531:263– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong-Toi:1991:SPS

[1146] Howard Wong-Toi and David L. Dill. Synthesizing processes and schedulers from temporal specifications. Lecture Notes in Computer Science, 531:272– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golaszewski:1991:TDS

[1147] Christian H. Golaszewski and Robert P.

Kurshan. Task-driven supervisory control of discrete event systems. Lecture Notes in Computer Science, 531:282– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buy:1991:PLB

[1148] Ugo Buy and Robert Moll. A proof lattice-based technique for analyzing liveness of resource controllers. Lecture Notes in Computer Science, 531: 292–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loewenstein:1991:VMC

[1149] Paul Loewenstein and David L. Dill.

Verification of a multiprocessor cache protocol using simulation relations and higher-order logic. Lecture Notes in Computer Science, 531:302–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrington:1991:CAP

[1150] David A. Carrington and K. A. Robinson. Computer assistance for program refinement. Lecture Notes in Computer Science, 531:312–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morris:1991:PVS

[1151] James M. Morris and Mark Howard.

Program verification by symbolic execution of hyperfinite ideal machines. Lecture Notes in Computer Science, 531:322–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbeau:1991:EKM

[1152] Michel Barbeau and Gregor V. Bochmann. Extension of the Karp and Miller procedure to Lotos specifications. Lecture Notes in Computer Science, 531:333–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Josephs:1991:ADI

[1153] Mark B. Josephs and Jan Tijmen Udding. An algebra for delay-insensitive circuits. Lecture Notes in Computer Science, 531:343–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madelaine:1991:FCS

[1154] Eric Madelaine and Didier Vergamini. Finiteness conditions and structural construction of automata for all process algebras. Lecture Notes in Computer Science, 531:353–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1991:AEB

[1155] Rance Cleaveland. On automatically explaining bisimulation inequivalence. Lecture Notes in Computer Science, 531:364–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drewes:1991:NHR

[1156] F. Drewes and H.-J. Kreowski. A note on hyperedge replacement. Lecture Notes in Computer Science, 532:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1991:GGB

[1157] J. Engelfriet and G. Rozenberg. Graph grammars based on node rewriting: An introduction to NLC graph grammars. Lecture Notes in Computer Science, 532:12–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1991:TIA

[1158] H. Ehrig, M. Korff, and M. Lowe. Tutorial introduction to the algebraic approach of graph grammars based on double and single pushouts. Lecture Notes in Computer Science, 532:24–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1991:LEG

[1159] B. Courcelle. The logical expression of graph properties (abstract). Lecture Notes in Computer Science, 532:38–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:PDU

[1160] Anonymous. Panel discussion: The use of graph grammars in applications. Lecture Notes in Computer Science, 532:41–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Himsolt:1991:ITD

[1161] M. Himsolt. Graphed: An interactive tool for developing graph grammars. Lecture Notes in Computer Science, 532:61–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schurr:1991:IEI

[1162] A. Schurr. IPSEN-Environment: An integrated and Incremental Project Support ENvironment. Lecture Notes in Computer Science, 532:66–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schurr:1991:PET

[1163] A. Schurr. PROGRESS-Editor: A text-oriented hybrid editor for PROgrammed Graph REwriting SyStems. Lecture Notes in Computer Science, 532:67–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wanke:1991:PTA

[1164] E. Wanke. PLEXUS: Tools for analyzing graph grammars. Lecture Notes in Computer Science, 532:68–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnborg:1991:ATG

[1165] S. Arnborg, B. Courcelle, A. Proskurowski, and D. Seese. An algebraic theory of graph reduction. Lecture Notes in Computer Science, 532:70–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailey:1991:PVL

[1166] D. A. Bailey, J. E. Cuny, and C. D. Fisher. Programming with very large graphs. Lecture Notes in Computer Science, 532:84–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthelmann:1991:DGO

[1167] K. Barthelmann. Describing Gottler’s operational graph grammars with pushouts. Lecture Notes in Computer Science, 532:98–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauderon:1991:GSS

[1168] M. Bauderon. General solution to a system of recursive equations on hypergraphs. Lecture Notes in Computer Science, 532:113–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBoer:1991:CMO

[1169] M. J. M. de Boer. Construction of map OL-Systems for developmental sequences of plant cell layers. Lecture Notes in Computer Science, 532:127– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1991:LGG

[1170] F. J. Brandenburg. Layout graph grammars: The placement approach. Lecture Notes in Computer Science, 532: 144–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1991:CCC

[1171] F. J. Brandenburg and M. P. Chytil.

Cycle chain code picture languages. Lecture Notes in Computer Science, 532:157–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bunke:1991:EIG

[1172] H. Bunke, T. Glauser, and T.-H. Tran. An efficient implementation of graph grammars based on the RETE matching algorithm. Lecture Notes in Computer Science, 532:174–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caucal:1991:AGG

[1173] D. Caucal. An application of graph grammars to the elimination of redundancy from functions defined by schemes. Lecture Notes in Computer Science, 532:190–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1991:GEC

[1174] T. C. Chen. Graphic equivalence and computer optimization. Lecture Notes in Computer Science, 532:207– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1991:GGL

[1175] A. Corradini, U. Montanari, F. Rossi, H. Ehrig, and M. Lowe. Graph grammars and logic programming. Lecture Notes in Computer Science, 532:221– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1991:GRS

[1176] B. Courcelle. Graphs as relational structures: An algebraic and logical approach. Lecture Notes in Computer Science, 532:238–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1991:CFH

[1177] B. Courcelle, J. Engelfriet, and G. Rozenberg. Context-free handlerewriting hypergraph grammars. Lecture Notes in Computer Science, 532: 253–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1991:GGH

[1178] H. Ehrig, A. Habel, H.-J. Kreowski, and F. Parisi-Presicce. From graph grammars to high level replacement systems. Lecture Notes in Computer Science, 532:269–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1991:ASG

[1179] H. Ehrig and F. Parisi-Presicce. Algebraic specification grammars: A junction between module specifications and graph grammars. Lecture Notes in Computer Science, 532:292–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1991:CCF

[1180] J. Engelfriet. A characterization of context-free NCE graph languages by monadic second-order logic on trees. Lecture Notes in Computer Science, 532:311–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1991:TGP

[1181] J. Engelfriet and L. Heyker. The term generating power of context-free hypergraph grammars. Lecture Notes in Computer Science, 532:328–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engels:1991:EAE

[1182] G. Engels. Elementary actions on an extended entity-relationship database. Lecture Notes in Computer Science, 532:344–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fracchia:1991:PBG

[1183] F. D. Fracchia and P. Prusinkiewicz. Physically-based graphical interpretation of marker cellwork L- systems. Lecture Notes in Computer Science, 532: 363–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glauert:1991:DEG

[1184] J. R. W. Glauert, J. R. Kennaway, and M. R. Sleep. Dactl: An experimental graph rewriting language. Lecture Notes in Computer Science, 532:378– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottler:1991:UGG

[1185] H. Gottler, J. Gunther, and G. Nieskens. Use graph grammars to design CAD-Systems! Lecture Notes in Computer Science, 532:396–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1991:CG

[1186] A. Habel and H.-J. Kreowski. Collage grammars. Lecture Notes in Computer Science, 532:411–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hess:1991:FMA

[1187] L. Hess and B. Mayoh. The four musicians: Analogies and expert systems — A graphic approach. Lecture Notes in Computer Science, 532:430–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssens:1991:STC

[1188] D. Janssens and G. Rozenberg. Structured transformations and computation graphs for actor grammars. Lecture Notes in Computer Science, 532: 446–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeltsch:1991:GIB

[1189] E. Jeltsch and H.-J. Kreowski. Grammatical inference based on hyperedge replacement. Lecture Notes in Computer Science, 532:461–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaplan:1991:SCL

[1190] S. M. Kaplan, J. P. Loyall, and S. K. Goering. Specifying concurrent languages and systems with deltagrammars. Lecture Notes in Computer Science, 532:475–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennaway:1991:GRS

[1191] R. Kennaway. Graph rewriting in some categories of partial morphisms. Lecture Notes in Computer Science, 532: 490–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korff:1991:AGG

[1192] M. Korff. Application of graph grammars to rule-based systems. Lecture Notes in Computer Science, 532:505– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lautemann:1991:TAT

[1193] C. Lautemann. Tree automata, tree decomposition and hyperedge replacement. Lecture Notes in Computer Science, 532:520–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lichtblau:1991:RRC

[1194] U. Lichtblau. Recognizing rooted context-free flowgraph languages in polynomial time. Lecture Notes in Computer Science, 532:538–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Litovsky:1991:CGR

[1195] I. Litovsky and Y. Metivier. Computing with graph relabelling systems with priorities. Lecture Notes in Computer Science, 532:549–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luck:1991:DWC

[1196] H. B. Luck and J. Luck. Doublewall cellwork systems for plant meristems. Lecture Notes in Computer Science, 532:564–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maggiolo-Schettini:1991:PDR

[1197] A. Maggiolo-Schettini and J. Winkowski. Programmed derivations of relational structures. Lecture Notes in Computer Science, 532:582–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagl:1991:SEG

[1198] M. Nagl and A. Schurr. A specification environment for graph grammars. Lecture Notes in Computer Science, 532:599–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paz:1991:TGS

[1199] A. Paz. The theory of graphoids: A survey. Lecture Notes in Computer Science, 532:610–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plump:1991:GRT

[1200] D. Plump. Graph-reducible term rewriting systems. Lecture Notes in Computer Science, 532:622–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosenfeld:1991:NGD

[1201] A. Rosenfeld. A note on graph decimation. Lecture Notes in Computer Science, 532:637–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schurr:1991:PVL

[1202] A. Schurr. PROGRESS: A VHLlanguage based on graph grammars. Lecture Notes in Computer Science, 532:641–??, 1991. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Taentzer:1991:MOC

[1203] G. Taentzer and H. Schween. Movement of objects in configuration spaces modelled by graph grammars. Lecture Notes in Computer Science, 532:660– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1991:RER

[1204] W. Vogler. Recognizing edge replacement graph languages in cubic time. Lecture Notes in Computer Science, 532:676–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1991:CGT

[1205] H. Ehrig and M. Lowe. Computing by graph transformation: Overall aims and new results. Lecture Notes in Computer Science, 532:688–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnborg:1991:MSO

[1206] Stefan Arnborg, Andrzej Proskurowski, and Detlef Seese. Monadic second order logic, tree automata and forbidden minors. Lecture Notes in Computer Science, 533:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blass:1991:RTA

[1207] Andreas Blass and Yuri Gurevich. On the reduction theory for average case complexity. Lecture Notes in Computer Science, 533:17–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borger:1991:PAT

[1208] Egon Borger and Dean Rosenzweig.

From Prolog algebras towards WAM – A mathematical study of implementation. Lecture Notes in Computer Science, 533:31–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borger:1991:FOS

[1209] Egon Borger and Peter H. Schmitt. A formal operational semantics for languages of type Prolog III. Lecture Notes in Computer Science, 533:67–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burgard:1991:ECG

[1210] Wolfram Burgard. Efficiency considerations on goal-directed forward chaining for logic programs. Lecture Notes in Computer Science, 533:80–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cantone:1991:DPT

[1211] Domenico Cantone, V. Cutello, and J. T. Schwartz. Decision problems for Tarski and Presburger arithmetics extended with sets. Lecture Notes in Computer Science, 533:95–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durdanovic:1991:FGC

[1212] Igor Durdanovic. A fast garbage collection algorithm for WAM-Based PRO-LOG. Lecture Notes in Computer Science, 533:110–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fermuller:1991:RVD

[1213] Christian Fermuller. A resolution variant deciding some classes of clause sets. Lecture Notes in Computer Science, 533:128–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flogel:1991:SQB

[1214] Andreas Flogel, Marek Karpinski, and Hans Kleine Buning. Subclasses of quantified Boolean formulas. Lecture Notes in Computer Science, 533:145– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1991:APD

[1215] Dov M. Gabbay. Algorithmic proof with diminishing resources part 1. Lecture Notes in Computer Science, 533: 156–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goerdt:1991:CPV

[1216] Andreas Goerdt. Cutting plane versus Frege proof systems. Lecture Notes in Computer Science, 533:174–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grandjean:1991:RCM

[1217] Etienne Grandjean and J. M. Robson. RAM with compact memory: A realistic and robust model of computation. Lecture Notes in Computer Science, 533:195–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habart:1991:RTR

[1218] Karol Habart. Randomness and Turing reducibility restraints. Lecture Notes in Computer Science, 533:234–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahnle:1991:TET

[1219] Reiner Hahnle. Towards an efficient tableau proof procedure for multiplevalued logics. Lecture Notes in Computer Science, 533:248–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hertrampf:1991:IPS

[1220] Ulrich Hertrampf and Klaus Wagner.

Interactive proof systems: Provers, rounds, and error bounds. Lecture Notes in Computer Science, 533:261– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhisheng:1991:LBD

[1221] Huang Zhisheng. Logics for belief dependence. Lecture Notes in Computer Science, 533:274–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaspars:1991:GSA

[1222] Jan Jaspars. A generalization of stability and its application to circumscription of positive introspective knowledge. Lecture Notes in Computer Science, 533:289–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundici:1991:CAE

[1223] Daniel Mundici. The complexity of adaptive error-correcting codes. Lecture Notes in Computer Science, 533: 300–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudlak:1991:RTB

[1224] Pavel Pudlak. Ramsey’s theorem in bounded arithmetic. Lecture Notes in Computer Science, 533:308–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranaivoson:1991:NLB

[1225] Solomampionona Ranaivoson. Nontrivial lower bounds for some NP-Problems on directed graphs. Lecture Notes in Computer Science, 533:318–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauszer:1991:EMA

[1226] Cecylia M. Rauszer. Expansions and models of autoepistemic theories. Lecture Notes in Computer Science, 533: 340–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stark:1991:EFM

[1227] Robert F. Stark. On the existence of fixpoints in Moore’s autoepistemic logic and the non-monotonic logic of McDermott and doyle. Lecture Notes in Computer Science, 533:354–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szabo:1991:TLA

[1228] Manfred E. Szabo. On the tracking of loops in automated deductions. Lecture Notes in Computer Science, 533:366–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vollmer:1991:GLT

[1229] Heribert Vollmer. The gap-language-technique revisited. Lecture Notes in Computer Science, 533:389–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barzdins:1991:AIC

[1230] G. Barzdins. ADT implementation and completion by induction from examples. Lecture Notes in Computer Science, 534:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettaz:1991:AAT

[1231] M. Bettaz. An association of algebraic term nets and abstract data types for specifying real communication protocols. Lecture Notes in Computer Science, 534:11–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clerici:1991:SLG

[1232] S. Clerici and F. Orejas. The specification language GSBL. Lecture Notes in Computer Science, 534:31–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimitrovici:1991:CAH

[1233] C. Dimitrovici and U. Hummert. Composition of algebraic high-level nets.

Lecture Notes in Computer Science, 534:52–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1991:MOR

[1234] H. Ehrig and F. Parisi-Presicce. A match operation for rule-based modular system design. Lecture Notes in Computer Science, 534:74–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GroBe-Rhode:1991:TOO

[1235] M. GroBe-Rhode. Towards objectoriented algebraic specifications. Lecture Notes in Computer Science, 534: 98–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1991:ICT

[1236] S. Lange and K. P. Jantke. Inductive completion for transformation of equational specifications. Lecture Notes in Computer Science, 534:117–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehmann:1991:NIS

[1237] T. Lehmann. A notion of implementation for the specification language OB-SCURE. Lecture Notes in Computer Science, 534:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mascari:1991:MTS

[1238] G. Mascari and A. Vincenzi. Modeltheoretic specifications and back-and-forth equivalences. Lecture Notes in Computer Science, 534:166–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinke:1991:UAH

[1239] K. Meinke. Universal algebra in higher types (extended abstract). Lecture Notes in Computer Science, 534:185– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieuwenhuis:1991:CRA

[1240] R. Nieuwenhuis and F. Orejas. Clausal rewriting: Applications and implementation. Lecture Notes in Computer Science, 534:204–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orejas:1991:CBS

[1241] F. Orejas and P. Nivela. Constraints for behavioural specifications. Lecture Notes in Computer Science, 534:220– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reggio:1991:EID

[1242] G. Reggio. Entities: An institution for dynamic systems. Lecture Notes in Computer Science, 534:246–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reichel:1991:CAC

[1243] H. Reichel. A 2-Category approach to critical pair completion. Lecture Notes in Computer Science, 534:266– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sannella:1991:KSF

[1244] D. Sannella and A. Tarlecki. A kernel specification formalism with higher-order parameterisation. Lecture Notes in Computer Science, 534:274– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sannella:1991:EMP

[1245] D. Sannella and A. Tarlecki. Extended ML: Past, present and future. Lecture Notes in Computer Science, 534:297–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Streicher:1991:DTC

[1246] T. Streicher and M. Wirsing. Dependent types considered necessary for specification languages. Lecture Notes in Computer Science, 534:323– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1991:GTL

[1247] E. G. Wagner. Generic types in a language for data directed design. Lecture Notes in Computer Science, 534: 341–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolz:1991:DCL

[1248] D. Wolz. Design of a compiler for lazy pattern driven narrowing. Lecture Notes in Computer Science, 534:362– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caferra:1991:UOT

[1249] R. Caferra, M. Herment, and N. Zabel.

User-oriented theorem proving with the ATINF graphic proof editor. Lecture Notes in Computer Science, 535:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

delCerro:1991:MAPa

[1250] L. Farinas del Cerro and A. Herzig.

A modal analysis of possibility theory. Lecture Notes in Computer Science, 535:11–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1991:MIR

[1251] D. Gabbay and A. Hunter. Making inconsistency respectable: a logical framework for inconsistency in reasoning. Lecture Notes in Computer Science, 535:19–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlowska:1991:RPS

[1252] E. Orlowska. Relational proof systems for some AI logics. Lecture Notes in Computer Science, 535:33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paun:1991:FGC

[1253] Gh. Paun. Formal grammars and cognitive architectures. Lecture Notes in Computer Science, 535:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiedermann:1991:ESN

[1254] J. Wiedermann. Efficient simulations of nondeterministic computations and their speed-up by the ring of cooperating machines. Lecture Notes in Computer Science, 535:59–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freund:1991:SCD

[1255] M. Freund. A semantic characterization of disjunctive relations. Lecture Notes in Computer Science, 535:72–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunter:1991:EDT

[1256] A. Hunter. Execution of defeasible temporal clauses for building preferred models. Lecture Notes in Computer Science, 535:84–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klopotek:1991:PFF

[1257] M. Klopotek. On the phenomenon of flattening “flexible prediction” concept hierarchy. Lecture Notes in Computer Science, 535:99–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1991:PLL

[1258] J. Lang. Possibilistic logic as a logical framework for min-max discrete optimisation problems and prioritized constraints. Lecture Notes in Computer Science, 535:112–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markov:1991:ADD

[1259] Z. Markov. An approach to data-driven learning. Lecture Notes in Computer Science, 535:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marquis:1991:EAP

[1260] P. Marquis. Extending abduction from propositional to first-order logic. Lecture Notes in Computer Science, 535: 141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petermann:1991:BET

[1261] U. Petermann. Building in equational theories into the connection method. Lecture Notes in Computer Science, 535:156–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfalzgraf:1991:LFP

[1262] J. Pfalzgraf. Logical fiberings and polycontextural systems. Lecture Notes in Computer Science, 535:170–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rusinowitch:1991:ADA

[1263] M. Rusinowitch and L. Vigneron.

Automated deduction with associative commutative operators. Lecture Notes in Computer Science, 535:185– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strasser:1991:TLK

[1264] A. Strasser. Towards a lattice of knowledge representation systems. Lecture Notes in Computer Science, 535:200– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sefranek:1991:IHN

[1265] J. Sefranek. Inconsistencies handling: nonmonotonic and paraconsistent reasoning. Lecture Notes in Computer Science, 535:214–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taucaite:1991:ASS

[1266] G. Taucaite and J. Laurinskas. An approach to structural synthesis of data processing programs. Lecture Notes in Computer Science, 535:220–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vauzeilles:1991:NFI

[1267] J. Vauzeilles. Negation as failure and intuitionistic three-valued logic. Lecture Notes in Computer Science, 535: 227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miola:1991:SCA

[1268] A. Miola. Symbolic computation and artificial intelligence. Lecture Notes in Computer Science, 535:244–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shafer:1991:KAS

[1269] Steve Shafer. Keynote address: Soft ware for hard physics. Lecture Notes in Computer Science, 536:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:FAS

[1270] Anonymous. A family album of software project courses. Lecture Notes in Computer Science, 536:3–4, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knoke:1991:MSP

[1271] Peter J. Knoke. Medium size project model: Variations on a theme. Lecture Notes in Computer Science, 536:5–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calliss:1991:CSM

[1272] Frank W. Calliss and Debra L.

Trantina. A controlled software maintenance project. Lecture Notes in Computer Science, 536:25–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shaw:1991:MUP

[1273] Mary Shaw and James E. Tomayko.

Models for undergraduate project courses in software engineering. Lecture Notes in Computer Science, 536:33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:SETa

[1274] Anonymous. Software engineering training in government and industry. Lecture Notes in Computer Science, 536:73–74, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelly:1991:EAS

[1275] R. D. Kelly, D. L. Lapay, and F. S.

Pitcairn. The establishment of an appropriate software engineering training program. Lecture Notes in Computer Science, 536:75–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mann:1991:ITS

[1276] P. Mann, A. Mason, and M. T. Norris. Industrial training for software engineers. Lecture Notes in Computer Science, 536:99–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mead:1991:SEG

[1277] Nancy R. Mead and Patricia K. Lawlis. Software engineering: Graduate-level courses for AFIT professional continuing education. Lecture Notes in Computer Science, 536:114–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:PCI

[1278] Anonymous. Panel on curriculum issues. Lecture Notes in Computer Science, 536:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnes:1991:CCI

[1279] Bruce H. Barnes and Jean B. Rogers. Computing curricula 1991 — its implications for software engineering education. Lecture Notes in Computer Science, 536:129–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavi:1991:CBS

[1280] Jonah Z. Lavi, Ashok Agrawala, Raymond Buhr, Ken Jackson, Michael Jackson, and Bernard Lang. Computer based systems engineering workshop. Lecture Notes in Computer Science, 536:149–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:SETb

[1281] Anonymous. Software engineering teaching styles. Lecture Notes in Computer Science, 536:165–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Budgen:1991:TAP

[1282] David Budgen and Chic Rattray.

Teaching about process issues in software engineering. Lecture Notes in Computer Science, 536:167–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bustard:1991:LAT

[1283] David Bustard. A layered approach to teaching software project management. Lecture Notes in Computer Science, 536:181–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:TD

[1284] Anonymous. Teaching design. Lecture Notes in Computer Science, 536:193– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquot:1991:SLT

[1285] J. P. Jacquot and J. Guyard. Seven lessons to teach design. Lecture Notes in Computer Science, 536:195– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Northrop:1991:DEI

[1286] Linda M. Northrop and William E.

Richardson. Design evolution: Implications for academia and industry. Lecture Notes in Computer Science, 536: 205–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierce:1991:TSD

[1287] Keith Pierce, Linda Deneen, and Gary Shute. Teaching software design in the freshman year. Lecture Notes in Computer Science, 536:219–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:STR

[1288] Anonymous. Special topics in real time and environments. Lecture Notes in Computer Science, 536:233–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchewka:1991:TSE

[1289] Conni Goodman Marchewka. Teaching software engineering for real-time design. Lecture Notes in Computer Science, 536:235–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werth:1991:IAC

[1290] Laurie Honour Werth. Industryacademia collaboration to provide CASE tools for software engineering classes. Lecture Notes in Computer Science, 536:245–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:DSE

[1291] Anonymous. Developing software engineering expertise. Lecture Notes in Computer Science, 536:257–258, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weitzenfeld:1991:DSE

[1292] Julian S. Weitzenfeld. Developing SE expertise. Lecture Notes in Computer Science, 536:259–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riedl:1991:WWL

[1293] Thomas R. Riedl, Julian S. Weitzenfeld, Jared T. Freeman, Gary A. Klein, and John Musa. What we have learned about software engineering expertise. Lecture Notes in Computer Science, 536:261–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeman:1991:ISE

[1294] Jared T. Freeman, Thomas R. Riedl, Julian S. Weitzenfeld, Gary A. Klein, and John Musa. Instruction for software engineering expertise. Lecture Notes in Computer Science, 536:271– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weitzenfeld:1991:KES

[1295] Julian S. Weitzenfeld, Thomas R.

Riedl, Jared T. Freeman, Gary A. Klein, and John Musa. Knowledge elicitation for software engineering expertise. Lecture Notes in Computer Science, 536:283–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biham:1991:DCL

[1296] Eli Biham and Adi Shamir. Differential cryptanalysis of DES-like cryptosystems (invited talk), (extended abstract). Lecture Notes in Computer Science, 537:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0537/05370002.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0537/05370002.pdf.

Gilbert:1991:SAF

[1297] Henry Gilbert and Guy Chassé. A statistical attack of the FEAL-8 cryptosystem. Lecture Notes in Computer Science, 537:22–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370022.
pdf.

Zeng:1991:ILS

[1298] Kencheng Zeng, C. H. Yang, and T. R. N. Rao. An improved linear syndrome algorithm in cryptanalysis with applications. Lecture Notes in Computer Science, 537:34– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370034.
pdf.

Brassard:1991:QBC

[1299] Gilles Brassard and Claude Crépeau. Quantum bit commitment and coin tossing protocols. Lecture Notes in Computer Science, 537:49–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370049.
pdf.

Beaver:1991:SLC

[1300] Donald Beaver, Joan Feigenbaum, Joe Kilian, and Phillip Rogaway. Security with low communication overhead (extended abstract). Lecture Notes in Computer Science, 537:62– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370062.
pdf.

Goldwasser:1991:FCG

[1301] Shafi Goldwasser and Leonid Levin.

Fair computation of general functions in presence of immoral majority. Lecture Notes in Computer Science, 537: 77–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370077.
pdf.

Brassard:1991:OWG

[1302] Gilles Brassard and Moti Yung.

One-way group actions. Lecture Notes in Computer Science, 537:94– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370094.
pdf.

LaMacchia:1991:SLS

[1303] B. A. LaMacchia and Andrew M. Odlyzko. Solving large sparse linear systems over finite fields. Lecture Notes in Computer Science, 537:109– 133, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370109.
pdf;   http://www.research.att.
com/~amo/doc/arch/sparse.linear.
eqs.pdf;   http://www.research.
att.com/~amo/doc/arch/sparse.linear.
eqs.ps;   http://www.research.att.
com/~amo/doc/arch/sparse.linear.
eqs.tex.

Buchmann:1991:CDL

[1304] Johannes A. Buchmann and Stephan Düllmann. On the computation of discrete logarithms in class groups (extended abstract). Lecture Notes in Computer Science, 537:134–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370134.
pdf.

Chuang:1991:MER

[1305] Chih-Chwen C. Chuang and James George Dunham. Matrix extensions of the RSA algorithm. Lecture Notes in Computer Science, 537:140–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370140.
pdf.

Koblitz:1991:CEC

[1306] Neal Koblitz. Constructing elliptic curve cryptosystems in characteristic 2. Lecture Notes in Computer Science, 537:156–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0537/05370156.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0537/05370156.
pdf.

Beth:1991:ITS

[1307] Thomas Beth and Yvo Desmedt. Identification tokens — or: Solving the chess grandmaster problem. Lecture Notes in Computer Science, 537:169– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370169.
pdf.

Desmedt:1991:AUS

[1308] Yvo Desmedt and Moti Yung. Arbitrated unconditionally secure authentication can be unconditionally protected against arbiter’s attacks (extended abstract). Lecture Notes in Computer Science, 537:177–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370177.
pdf.

Boyar:1991:CUS

[1309] Joan Boyar, David Chaum, Ivan Bjerre Damgård, and Torben Pryds Pedersen. Convertible undeniable signatures. Lecture Notes in Computer Science, 537: 189–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370189.
pdf.

Chaum:1991:USD

[1310] David Chaum and Sandra Roijakkers. Unconditionally secure digital signatures. Lecture Notes in Computer Science, 537:206–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370206.
pdf.

Simmons:1991:GSS

[1311] Gustavus J. Simmons. Geometric shared secret and/or shared control schemes (invited talk). Lecture Notes in Computer Science, 537:216– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370216.
pdf.

Brickell:1991:SIB

[1312] Ernest F. Brickell and Douglas R. Stinson. Some improved bounds on the information rate of perfect secret sharing schemes (extended abstract). Lecture Notes in Computer Science, 537:242– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370242.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370242.
pdf.

Micali:1991:CCT

[1313] Silvio Micali and Tal Rabin. Collective coin tossing without assumptions nor broadcasting. Lecture Notes in Computer Science, 537:253–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370253.
pdf.

Yacobi:1991:KDP

[1314] Yacov Yacobi. A key distribution “Paradox”. Lecture Notes in Computer Science, 537:268–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370268.
                                        pdf.

Fumy:1991:MAK

[1315] Walter Fumy and Michael Munzert. A modular approach to key distribution. Lecture Notes in Computer Science, 537:274–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370274.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370274.
                                        pdf.

Zheng:1991:SPO

[1316] Yuliang Zheng, Tsutomu Matsumoto, and Hideki Imai. Structural properties of one-way hash functions. Lecture Notes in Computer Science, 537:285– 302, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370285.
                                        pdf.

Rivest:1991:MMD

[1317] Ronald L. Rivest. The MD4 message digest algorithm. Lecture Notes in Computer Science, 537:303–311, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370303.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370303.
pdf.

Kilian:1991:AZK

[1318] Joe Kilian. Achieving zero-knowledge robustly. Lecture Notes in Computer Science, 537:313–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0537/05370313.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0537/05370313.pdf.

Beaver:1991:HIZ

[1319] Donald Beaver, Joan Feigenbaum, and Victor Shoup. Hiding instances in zero-knowledge proof systems (extended abstract). Lecture Notes in Computer Science, 537:326–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370326.
pdf.

Kurosawa:1991:MLZ

[1320] Kaoru Kurosawa and Shigeo Tsujii. Multi-language zero-knowledge interactive proof systems. Lecture Notes in Computer Science, 537:339– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370339.
pdf.

Lapidot:1991:PVN

[1321] Dror Lapidot and Adi Shamir. Publicly verifiable non-interactive zeroknowledge proofs. Lecture Notes in Computer Science, 537:353–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370353.
pdf.

DeSantis:1991:CAN

[1322] A. De Santis and M. Yung. Cryptographic applications of the noninteractive metaproof and many-prover systems. Lecture Notes in Computer Science, 537:366–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santis:1991:CAN

[1323] Alfredo De Santis and Moti Yung. Cryptographic applications of the noninteractive metaproof and many-prover systems (preliminary version). Lecture Notes in Computer Science, 537:366– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0537/05370366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370366.
pdf.

Kilian:1991:IPP

[1324] Joe Kilian. Interactive proofs with provable security against honest verifiers. Lecture Notes in Computer Science, 537:378–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0537/05370378.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0537/05370378.pdf.

Schrift:1991:UNB

[1325] A. W. Schrift and Adi Shamir. On the universality of the next bit test. Lecture Notes in Computer Science, 537: 394–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370394.
pdf.

Maurer:1991:UST

[1326] Ueli M. Maurer. A universal statistical test for random bit generators. Lecture Notes in Computer Science, 537: 409–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370409.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/0537/05370409.
                                        pdf.

McInnes:1991:IPK

[1327] James L. McInnes and Benny Pinkas. On the impossibility of private key cryptography with weakly random keys. Lecture Notes in Computer Science, 537:421–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0537/05370421.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0537/05370421.pdf.

Haber:1991:HTS

[1328] Stuart Haber and W. Scott Stornetta. How to time-stamp a digital document. Lecture Notes in Computer Science, 537:437–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370437.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370437.
                                        pdf.

Okamoto:1991:HUR

[1329] Tatsuaki Okamoto and Kazuo Ohta. How to utilize the randomness of zeroknowledge proofs (extended abstract). Lecture Notes in Computer Science, 537:456–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0537/05370456.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/0537/05370456.
pdf.

Merkle:1991:FSE

[1330] Ralph C. Merkle. Fast software encryption functions. Lecture Notes in Computer Science, 537:476–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370476.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370476.
pdf.

deWaleffe:1991:CSC

[1331] Dominique de Waleffe and Jean-Jacques Quisquater. CORSAIR: A smart card for public key cryptosystems. Lecture Notes in Computer Science, 537:502–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0537/05370502.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0537/05370502.pdf.

Waleffe:1991:CSC

[1332] Dominique de Waleffe and Jean-Jacques Quisquater. CORSAIR: A smart card for public key cryptosystems. Lecture Notes in Computer Science, 537:502–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0537/05370502.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0537/05370502.pdf.

Kompella:1991:FCC

[1333] Kireeti Kompella and Leonard M. Adleman. Fast checkers for cryptography. Lecture Notes in Computer Science, 537:515–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0537/05370515.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0537/05370515.pdf.

Cleve:1991:CTI

[1334] Richard Cleve. Complexity theoretic issues concerning block ciphers related to D.E.S. Lecture Notes in Computer Science, 537:530–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370530.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370530.
                                        pdf.

Cusick:1991:RIC

[1335] Thomas W. Cusick and Michael C. Wood. The REDOC II cryptosystem. Lecture Notes in Computer Science, 537:545–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0537/05370545.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370545.
                                        pdf.

Kim:1991:RCM

[1336] Kwangjo Kim, Tsutomu Matsumoto, and Hideki Imai. A recursive construction method of S-boxes satisfying strict avalanche criterion. Lecture Notes in Computer Science, 537:564– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370564.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370564.
pdf.

Oorschot:1991:CPP

[1337] Paul C. van Oorschot. A comparison of practical public-key cryptosystems based on integer factorization and discrete logarithms. Lecture Notes in Computer Science, 537:576– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370576.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370576.
pdf.

vanOorschot:1991:CPP

[1338] Paul C. van Oorschot. A comparison of practical public key cryptosystems based on integer factorization and discrete logarithms. Lecture Notes in Computer Science, 537:576– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370576.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370576.
pdf.

Koyama:1991:NPC

[1339] Kenji Koyama and Routo Terada. Nonlinear parity circuits and their cryptographic applications. Lecture Notes in Computer Science, 537:582– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370582.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370582.
                                        pdf.

Staffelbach:1991:CSC

[1340] Othmar Staffelbach and Willi Meier. Cryptographic significance of the carry for ciphers based on integer addition. Lecture Notes in Computer Science, 537:601–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370601.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370601.
                                        pdf.

LaMacchia:1991:CDL

[1341] B. A. LaMacchia and Andrew M. Odlyzko. Computation of discrete logarithms in prime fields (extended abstract). Lecture Notes in Computer Science, 537:616–618, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370616.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370616.
                                        pdf;   http://www.research.att.

com/~amo/doc/arch/prime.discrete.
logs.pdf;   http://www.research.
att.com/~amo/doc/arch/prime.discrete.
logs.ps;   http://www.research.
att.com/~amo/doc/arch/prime.discrete.
logs.tex.

Even:1991:SMM

[1342] Shimon Even. Systolic modular multiplication. Lecture Notes in Computer Science, 537:619–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370619.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370619.
pdf.

Rivest:1991:FFM

[1343] Ronald L. Rivest. Finding four million large random primes. Lecture Notes in Computer Science, 537:625– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0537/05370625.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0537/05370625.
pdf.

Miyaguchi:1991:FCF

[1344] Shoji Miyaguchi. The FEAL cipher family. Lecture Notes in Computer Science, 537:627–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0537/05370627.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        0537/05370627.pdf.

Yacobi:1991:DLC

[1345] Yacov Yacobi. Discrete-log with compressible exponents. Lecture Notes in Computer Science, 537:639–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0537/05370639.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0537/05370639.
                                        pdf.

Anonymous:1991:AIa

[1346] Anonymous. Author index. Lecture Notes in Computer Science, 537:644– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0537/0537ind.pdf.

Kojima:1991:UAI

[1347] M. Kojima, N. Megiddo, T. Noma, and A. Yoshise. A unified approach to interior point algorithms for linear complementarity problems. Lecture Notes in Computer Science, 538:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cellini:1991:ASS

[1348] P. Cellini, P. Gianni, and C. Traverso. Algorithms for the shape of semialgebraic sets. A new approach. Lecture Notes in Computer Science, 539:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janwa:1991:PAG

[1349] H. Janwa. On the parameters of algebraic geometric codes. Lecture Notes in Computer Science, 539:19–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaltofen:1991:WMS

[1350] E. Kaltofen and B. D. Saunders. On Wiedemann’s method of solving sparse linear systems. Lecture Notes in Computer Science, 539:29–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Litsyn:1991:FAD

[1351] S. N. Litsyn. Fast algorithms for decoding orthogonal and related codes. Lecture Notes in Computer Science, 539: 39–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasconcelos:1991:JMC

[1352] W. V. Vasconcelos. Jacobian matrices and constructions in algebra. Lecture Notes in Computer Science, 539:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1991:HPH

[1353] T. Becker. Homogenity, pseudohomogenity, and Gröbner basis computations. Lecture Notes in Computer Science, 539:65–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beth:1991:ANS

[1354] T. Beth and F. Schaefer. Arithmetic on non-supersingular elliptic curves. Lecture Notes in Computer Science, 539:

74–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1991:ISA

[1355] G. Butler. Implementing some algorithms of kantor. Lecture Notes in Computer Science, 539:82–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canny:1991:CRG

[1356] J. F. Canny. Computing roadmaps of general semi-algebraic sets. Lecture Notes in Computer Science, 539:94–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canny:1991:ISD

[1357] J. F. Canny. An improved sign determination algorithm. Lecture Notes in Computer Science, 539:108–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chung:1991:NGH

[1358] H. Chung. The 2-nd generalized Hamming weight of double-error correcting binary BCH codes and their dual codes. Lecture Notes in Computer Science, 539:118–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conti:1991:BAI

[1359] P. Conti and C. Traverso. Buchberger algorithm and integer programming. Lecture Notes in Computer Science, 539:130–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diab:1991:NSA

[1360] M. Diab. New systolic architectures for cyclic code encoding. Lecture Notes in Computer Science, 539:140–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinneen:1991:ACE

[1361] M. J. Dinneen, M. R. Fellows, and V. Faber. Algebraic constructions of efficient broadcast networks. Lecture Notes in Computer Science, 539:152– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanEijl:1991:ECW

[1362] C. van Eijl, G. Cohen, and G. Zemor. Error-correction for WIMs and WUMs. Lecture Notes in Computer Science, 539:159–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallo:1991:SCR

[1363] G. Gallo, B. Mishra, and F. Ollivier.

Some constructions in rings of differential polynomials. Lecture Notes in Computer Science, 539:171–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holmquist:1991:CED

[1364] L. P. Holmquist and L. L. Kinney. Concurrent error detection in sequential circuits using convolutional codes. Lecture Notes in Computer Science, 539: 183–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krick:1991:ACR

[1365] T. Krick and A. Logar. An algorithm for the computation of the radical of an ideal in the ring of polynomials. Lecture Notes in Computer Science, 539:195–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuechlin:1991:IMP

[1366] W. Kuechlin, D. Lutz, and N. Nevin. Integer multiplication in PARSAC-2 on stock microprocessors. Lecture Notes in Computer Science, 539:206–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lachaud:1991:PTC

[1367] G. Lachaud and J. Stern. Polynomialtime construction of spherical codes. Lecture Notes in Computer Science, 539:218–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langemyr:1991:AMA

[1368] L. Langemyr. Algorithms for a multiple algebraic extension II. Lecture Notes in Computer Science, 539:224–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langevin:1991:OCR

[1369] P. Langevin. On the orphans and covering radius of the Reed–Muller codes. Lecture Notes in Computer Science, 539:234–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1991:JAE

[1370] Y. Li and X. Wang. A joint authentication and encryption scheme based on algebraic coding theory. Lecture Notes in Computer Science, 539:241– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lobstein:1991:ACS

[1371] A. C. Lobstein and P. Sole. Arithmetic codes — survey, recent and new results. Lecture Notes in Computer Science, 539:246–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morelos-Zaragoza:1991:SRL

[1372] R. Morelos-Zaragoza and S. Lin.

Some results on linear unequal-error-protection codes specified by their generator matrix. Lecture Notes in Computer Science, 539:259–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Socias:1991:API

[1373] G. Moreno Socias. An ackermannian polynomial ideal. Lecture Notes in Computer Science, 539:269–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mostowski:1991:CCC

[1374] T. Mostowski and E. Rannou. Complexity of the computation of the canonical whitney stratification of an algebraic set in C^{n}. Lecture Notes in Computer Science, 539:281–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otto:1991:SUR

[1375] F. Otto. Some undecidability results for weakly confluent monadic stringrewriting systems. Lecture Notes in Computer Science, 539:292–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedersen:1991:CMS

[1376] P. Pedersen. Calculating multidimensional symmetric functions using Jacobi’s formula. Lecture Notes in Computer Science, 539:304–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedersen:1991:MST

[1377] P. Pedersen. Multivariate Sturm theory. Lecture Notes in Computer Science, 539:318–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perret:1991:BSG

[1378] M. Perret. Binary spherical geometric codes. Lecture Notes in Computer Science, 539:333–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poli:1991:ACG

[1379] A. Poli and M. Belkasmi. An algebraic construction of generalized Beenker’s codes. Lecture Notes in Computer Science, 539:340–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rifa:1991:ITC

[1380] J. Rifa and J. Borrell. Improving the time complexity of the computation of irreducible and primitive polynomials in finite fields. Lecture Notes in Computer Science, 539:352–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rifa:1991:CTC

[1381] J. Rifa and J. Pujol. Completely transitive codes and distance transitive graphs. Lecture Notes in Computer Science, 539:360–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Risler:1991:PCP

[1382] J.-J. Risler. Placement of curved polygons. Lecture Notes in Computer Science, 539:368–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodier:1991:WED

[1383] F. Rodier. On the weights of the elements of the duals of binary BCH codes. Lecture Notes in Computer Science, 539:384–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossi:1991:COS

[1384] F. Rossi and W. Spangher. Computation of the openness of some loci of modules. Lecture Notes in Computer Science, 539:390–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saitoh:1991:RBE

[1385] Y. Saitoh and H. Imai. Random and byte error correcting codes for asymmetric or unidirectional error control. Lecture Notes in Computer Science, 539:403–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakata:1991:FMP

[1386] S. Sakata. Finding a minimal polynomial vector set of a vector of nD arrays. Lecture Notes in Computer Science, 539:414–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sole:1991:CCC

[1387] P. Sole. Covering codes and combinatorial optimization. Lecture Notes in Computer Science, 539:426–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ucan:1991:DQP

[1388] O. N. Ucan, U. Aygolu, and E. Panayirci. Decoding of quadrature partial response-trellis coded signals (QPR-TCM) in the presence of intersymbol interference and noise. Lecture Notes in Computer Science, 539:434– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ulmer:1991:ASL

[1389] F. Ulmer. On algebraic solutions of linear differential equations with primitive unimodular Galois group. Lecture Notes in Computer Science, 539:446– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vainstein:1991:EDC

[1390] F. S. Vainstein. Error detection and correction in numerical computations by algebraic methods. Lecture Notes in Computer Science, 539:456–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeng:1991:FSD

[1391] K. Zeng, D.-Y. Wei, and T. R. N.

Rao. d-functions in V_{k}(F_{2}) and selfdecimation of m-sequences. Lecture Notes in Computer Science, 539:465– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1991:MMC

[1392] L. Zhang and B. Vuketic. Multilevel modulation codes for Rayleigh fading channels. Lecture Notes in Computer Science, 539:477–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filgueiras:1991:SLC

[1393] M. Filgueiras and A. P. Tomas. Solving linear constraints on finite domains through parsing. Lecture Notes in Computer Science, 541:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menezes:1991:CSF

[1394] F. Menezes and P. Barahona. Constraint solving in finite domains under user control. Lecture Notes in Computer Science, 541:17–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomas:1991:NMS

[1395] A. P. Tomas and M. Filgueiras. A new method for solving linear constraints on the natural numbers. Lecture Notes in Computer Science, 541:30–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lano:1991:CBF

[1396] K. Lano. A constraint-based fuzzy inference system. Lecture Notes in Computer Science, 541:45–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mamede:1991:CBL

[1397] M. Mamede and L. Monteiro. A constraint-based language for querying taxonomic systems. Lecture Notes in Computer Science, 541:60–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mendes:1991:HPS

[1398] J. A. Mendes and P. R. Henriques. Heuristic parsing and search space pruning. Lecture Notes in Computer Science, 541:76–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelson:1991:WSM

[1399] P. C. Nelson and A. A. Toptsis. Waveshaping in multiprocessor bidirectional heuristic state space search. Lecture Notes in Computer Science, 541:92–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1991:ESM

[1400] L. Moniz Pereira, J. J. Alferes, and J. N. Aparicio. The extended stable models of contradiction removal semantics. Lecture Notes in Computer Science, 541:105–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1991:MRC

[1401] D. Kumar and S. C. Shapiro. Modelling a rational cognitive agent in SNePS. Lecture Notes in Computer Science, 541:120–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1991:SPI

[1402] G. David and A. Porto. Semantics of property inheritance in a hierarchic system with explicit negation. Lecture Notes in Computer Science, 541:135– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardoso:1991:TCD

[1403] A. Cardoso and E. Costa. Time in confluences: Dealing with delays for consistency-checking. Lecture Notes in Computer Science, 541:151–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rutten:1991:TRI

[1404] E. Rutten. A temporal representation for imperatively structured plans of actions. Lecture Notes in Computer Science, 541:165–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ribeiro:1991:MIA

[1405] C. Ribeiro and A. Porto. Maximal intervals: An approach to temporal reasoning. Lecture Notes in Computer Science, 541:180–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decker:1991:CDP

[1406] M. Decker, G. Moerkotte, H. Muller, and J. Posegga. Consistency driven planning. Lecture Notes in Computer Science, 541:195–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramos:1991:EAP

[1407] C. Ramos and E. Oliveira. An efficient approach to planning in assembly tasks. Lecture Notes in Computer Science, 541:210–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Friedrich:1991:TTR

[1408] G. Friedrich, G. Gottlob, and W. Nejdl. Towards a theory of the repair process. Lecture Notes in Computer Science, 541:222–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calejo:1991:DSD

[1409] M. Calejo and L. Moniz Pereira. Declarative source debugging. Lecture Notes in Computer Science, 541:237– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kratzer:1991:NAD

[1410] K. P. Kratzer. A neural approach to data compression and classification. Lecture Notes in Computer Science, 541:250–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vorch:1991:GPC

[1411] R. Vorc’h. Generalization for a propositional calculus: A constraints-based approach. Lecture Notes in Computer Science, 541:264–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winkelbauer:1991:ALU

[1412] L. Winkelbauer and C. Stary. Adaptive learning using a qualitative feedback loop. Lecture Notes in Computer Science, 541:278–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biermann:1991:APV

[1413] A. W. Biermann and R. W. Smith. An architecture for pragmatic voice interactive systems. Lecture Notes in Computer Science, 542:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doyle:1991:LRD

[1414] J. Doyle, Y. Shoham, and M. P. Wellman. A logic of relative desire. Lecture Notes in Computer Science, 542:16–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiederhold:1991:RAI

[1415] G. Wiederhold. The roles of artificial intelligence in information systems. Lecture Notes in Computer Science, 542:38–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eick:1991:IVO

[1416] C. F. Eick. Integrating variables and operations into rule-based forward chaining systems. Lecture Notes in Computer Science, 542:52–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mutchler:1991:MPV

[1417] D. Mutchler. The multi-player version of minimax displays game-tree pathology. Lecture Notes in Computer Science, 542:62–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portinale:1991:GHD

[1418] L. Portinale. Generalization handling in a dynamic case memory. Lecture Notes in Computer Science, 542:72–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

An:1991:KBS

[1419] H. An and L. J. Henschen. Knowledge based semantic query optimization. Lecture Notes in Computer Science, 542:82–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baral:1991:CKB

[1420] C. Baral, S. Kraus, J. Minker, and V. S. Subrahmanian. Combining knowledge bases consisting of first order theories. Lecture Notes in Computer Science, 542:92–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatia:1991:UPI

[1421] S. K. Bhatia, J. S. Deogun, and V. V. Raghavan. User profiles for information retrieval. Lecture Notes in Computer Science, 542:102–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1991:ADO

[1422] I-M. A. Chen and R.-C. Lee. An approach to deriving object hierarchies from database schema and contents. Lecture Notes in Computer Science, 542:112–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desai:1991:MID

[1423] B. C. Desai and L. Zhang. Multilevel interface to a distributed database system. Lecture Notes in Computer Science, 542:122–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1991:EEL

[1424] J. Han and Q. Wang. Efficient evaluation of linear recursions with function symbols in deductive databases. Lecture Notes in Computer Science, 542: 132–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1991:SEA

[1425] J. Y j. Hsu. Synthesizing efficient agents from partial programs. Lecture Notes in Computer Science, 542:142– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufman:1991:KED

[1426] K. A. Kaufman, R. S. Michalski, and L. Kerschberg. Knowledge extraction from databases: design principles of the INLEN system. Lecture Notes in Computer Science, 542:152–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ng:1991:SMS

[1427] R. Ng and V. S. Subrahmanian. Stable model semantics for probabilistic deductive databases. Lecture Notes in Computer Science, 542:162–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1991:TIV

[1428] S. K. M. Wong, P. Lingras, and Y. Y. Yao. Towards implementing valuation based systems with relational databases. Lecture Notes in Computer Science, 542:172–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1991:RCC

[1429] B.-T. B. Chu. On the representation of continuous causal relationships. Lecture Notes in Computer Science, 542:183–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malec:1991:PTN

[1430] J. Malec. Process transition networks: a formal graphical knowledge representation tool. Lecture Notes in Computer Science, 542:193–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munoz:1991:UND

[1431] R. A. Munoz and C.-C. Yang. Using normal deduction graphs in default reasoning. Lecture Notes in Computer Science, 542:203–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ras:1991:MIS

[1432] Z. W. Ras and J. Xiao. A model of information sharing for fault-tolerant flexible manufacturing systems. Lecture Notes in Computer Science, 542: 213–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richards:1991:IBT

[1433] B. Richards, Y. Jiang, and H. Choi.

On interval-based temporal planning — an IQ strategy. Lecture Notes in Computer Science, 542:226–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sobolewski:1991:PCT

[1434] M. Sobolewski. Percept conceptualizations and their knowledge representation schemes. Lecture Notes in Computer Science, 542:236–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terenziani:1991:CTP

[1435] P. Terenziani. Combining time points and time intervals in a hybrid knowledge representation formalism. Lecture Notes in Computer Science, 542:246– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thirunarayan:1991:APF

[1436] K. Thirunarayan. An analysis of property-flow view vs. individual-flow view of inheritance. Lecture Notes in Computer Science, 542:256–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wendlandt:1991:SET

[1437] E. B. Wendlandt and J. R. Driscoll. Semantic extensions to text retrieval. Lecture Notes in Computer Science, 542: 266–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winslett:1991:UEB

[1438] M. Winslett and T. S.-C. Chou. Updates with equality: beyond the Herbrand universe assumption. Lecture Notes in Computer Science, 542:276– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bala:1991:UGA

[1439] J. Bala, K. DeJong, and P. Pachowicz. Using genetic algorithms to improve the performance of classification rules produced by symbolic inductive method. Lecture Notes in Computer Science, 542:286–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botta:1991:LQF

[1440] M. Botta and A. Giordana. Learning quantitative features in a symbolic environment. Lecture Notes in Computer Science, 542:296–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coates-Stephens:1991:AAP

[1441] S. Coates-Stephens. Automatic acquisition of proper noun meanings. Lecture Notes in Computer Science, 542: 306–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Console:1991:IRB

[1442] L. Console, A. Giordana, and L. Saitta.

Investigating the relationships between abduction and inverse resolution in propositional calculus. Lecture Notes in Computer Science, 542:316–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delisle:1991:EBL

[1443] S. Delisle, S. Matwin, J. Wang, and L. Zupan. Explanation-based learning helps acquire knowledge from natural language texts. Lecture Notes in Computer Science, 542:326–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glover:1991:HPR

[1444] C. W. Glover, N. S. V. Rao, and E. M.

Oblow. Hybrid pattern recognition system capable of self-modification. Lecture Notes in Computer Science, 542: 338–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goker:1991:TAI

[1445] A. Goker and T. L. McCluskey. Towards an adaptive information retrieval system. Lecture Notes in Computer Science, 542:348–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gordon:1991:ICA

[1446] D. F. Gordon. Improving the comprehensibility, accuracy, and generality of reactive plans. Lecture Notes in Computer Science, 542:358–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grzymala-Busse:1991:UAV

[1447] J. W. Grzymala-Busse. On the unknown attribute values in learning from examples. Lecture Notes in Computer Science, 542:368–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janikow:1991:NSI

[1448] C. Z. Janikow. A new system for inductive learning in attribute-based spaces. Lecture Notes in Computer Science, 542:378–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michalewicz:1991:GCR

[1449] M. Michalewicz and Z. Michalewicz. Generation of classification rules. Lecture Notes in Computer Science, 542: 389–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salzberg:1991:DMI

[1450] S. Salzberg. Distance metrics for instance-based learning. Lecture Notes in Computer Science, 542:399–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spears:1991:SCO

[1451] W. M. Spears and V. Anand. A study of crossover operators in genetic programming. Lecture Notes in Computer Science, 542:409–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tecuci:1991:IUB

[1452] G. Tecuci and R. S. Michalski. Input understanding as a basis for multistrategy task-adaptive learning. Lecture Notes in Computer Science, 542: 419–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zembowicz:1991:ADE

[1453] R. Zembowicz and J. M. Zytkow. Automated discovery of empirical equations from data. Lecture Notes in Computer Science, 542:429–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zytkow:1991:CMH

[1454] J. M. Zytkow and P. J. Fischer.

Constructing models of hidden structure. Lecture Notes in Computer Science, 542:441–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akama:1991:ALP

[1455] Seiki Akama. Amalgamated logic programming and non-monotonic reasoning. Lecture Notes in Computer Science, 542:450–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baral:1991:WSN

[1456] C. Baral, J. Lobo, and J. Minker.

WF^{3}: a semantics for negation in normal disjunctive logic programs. Lecture Notes in Computer Science, 542: 459–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calmet:1991:RMK

[1457] J. Calmet and I. A. Tjandra. Representation of mathematical knowledge. Lecture Notes in Computer Science, 542:

469–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatalic:1991:GLF

[1458] P. Chatalic and C. Froidevaux. Graded logics: a framework for uncertain and defeasible knowledge. Lecture Notes in Computer Science, 542:479–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1991:SSL

[1459] J. Chen and S. Kundu. The strong semantics for logic programs. Lecture Notes in Computer Science, 542:490– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conrad:1991:SPA

[1460] J. M. Conrad, D. Bahler, and J. Bowen. Static parallel arc consistency in constraint satisfaction. Lecture Notes in Computer Science, 542:500–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flach:1991:TTI

[1461] P. A. Flach. Towards a theory of inductive logic programming. Lecture Notes in Computer Science, 542:510– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1991:AVA

[1462] Y. J. Jiang. Another view of autoepistemic logic and truth maintenance system. Lecture Notes in Computer Science, 542:520–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuo:1991:DAL

[1463] K. Kuo. Dynamic autoepistemic logic. Lecture Notes in Computer Science, 542:530–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lorenz:1991:NTR

[1464] S. Lorenz. Nonmonotonic temporal reasoning: persistence, justified causation and event minimization. Lecture Notes in Computer Science, 542:540– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marquis:1991:NR

[1465] P. Marquis. Novelty revisited. Lecture Notes in Computer Science, 542: 550–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1991:TCR

[1466] M. J. Miller and D. Perlis. Typicality constants and range defaults: some pros and cons of a cognitive model of default reasoning. Lecture Notes in Computer Science, 542:560–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murray:1991:RPD

[1467] N. V. Murray and E. Rosenthal. Resolution and path dissolution in multiplevalued logics. Lecture Notes in Computer Science, 542:570–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nie:1991:HWN

[1468] X. Nie. How well are non-Horn clauses handled? Lecture Notes in Computer Science, 542:580–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nirkhe:1991:FDC

[1469] M. Nirkhe, S. Kraus, and D. Perlis. Fully deadline-coupled planning: one step at a time. Lecture Notes in Computer Science, 542:589–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaza:1991:CPL

[1470] J. A. Plaza. Completeness for propositional logic programs with negation. Lecture Notes in Computer Science, 542:600–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajasekar:1991:DLP

[1471] A. Rajasekar. Disjunctive logic programs and constraint processing. Lecture Notes in Computer Science, 542: 610–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stachniak:1991:NRC

[1472] Z. Stachniak. Note on resolution circuits. Lecture Notes in Computer Science, 542:620–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomason:1991:SAM

[1473] R. H. Thomason. A semantic analysis of monotonic inheritance with roles and relations. Lecture Notes in Computer Science, 542:630–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brewka:1991:CIR

[1474] Gerhard Brewka, David Makinson, and Karl Schlechta. Cumulative inference relations for JTMS and logic programming. Lecture Notes in Computer Science, 543:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dix:1991:CRS

[1475] Jurgen Dix. Cumulativity and rationality in semantics of normal logic programs. Lecture Notes in Computer Science, 543:13–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herre:1991:NRL

[1476] Heinrich Herre. Nonmonotonic reasoning and logic programs. Lecture Notes in Computer Science, 543:38–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freund:1991:SIO

[1477] Michael Freund. Supracompact inference operations. Lecture Notes in Computer Science, 543:59–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jager:1991:NND

[1478] Gerhard Jager. Notions of nonmonotonic derivability. Lecture Notes in Computer Science, 543:74–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marek:1991:RSM

[1479] Wiktor Marek, Grigori F. Schwarz, and Miroslaw Truszczynski. Ranges of strong modal nonmonotonic logics. Lecture Notes in Computer Science, 543:85–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thiele:1991:GCI

[1480] Helmut Thiele. On generation of cumulative inference operators by default deduction rules. Lecture Notes in Computer Science, 543:100–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1991:QMR

[1481] Emil Weydert. Qualitative magnitude reasoning. Lecture Notes in Computer Science, 543:138–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jantke:1991:MNI

[1482] Klaus P. Jantke. Monotonic and nonmonotonic inductive inference of functions and patterns. Lecture Notes in Computer Science, 543:161–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1991:NPT

[1483] Steffen Lange. A note on polynomialtime inference of k-variable pattern languages. Lecture Notes in Computer Science, 543:178–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiehagen:1991:TII

[1484] Rolf Wiehagen. A thesis in inductive inference. Lecture Notes in Computer Science, 543:184–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeugmann:1991:IIO

[1485] Thomas Zeugmann. Inductive inference of optimal programs: A survey and open problems. Lecture Notes in Computer Science, 543:208–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalinski:1991:AEI

[1486] Jurgen Kalinski. Autoepistemic expansions with incomplete belief introspection. Lecture Notes in Computer Science, 543:223–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morreau:1991:IIK

[1487] Michael Morreau and Hans Rott. Is it impossible to keep up to date? Lecture Notes in Computer Science, 543: 233–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesse:1991:TMA

[1488] W. Hesse. Two metamodels for application system development — conventional vs. object-oriented approach. Lecture Notes in Computer Science, 544:3–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krieg-Bruckner:1991:TMP

[1489] B. Krieg-Bruckner. Transformational meta program development. Lecture Notes in Computer Science, 544:19–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Partsch:1991:ACS

[1490] H. A. Partsch and N. Volker. Another case study on reusability of transformational developments pattern matching according to knuth, morris, and pratt. Lecture Notes in Computer Science, 544:35–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennicker:1991:FMS

[1491] R. Hennicker and M. Wirsing. A formal method for the systematic reuse of specification components. Lecture Notes in Computer Science, 544:49–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1991:DPD

[1492] M. Broy. Deductive program development: Evaluation in reverse Polish notation as an example. Lecture Notes in Computer Science, 544:79–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pepper:1991:LPD

[1493] P. Pepper. Literate program derivation: A case study. Lecture Notes in Computer Science, 544:101–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinbruggen:1991:PVS

[1494] R. Steinbruggen. Programs viewed as SKOLEM functions. Lecture Notes in Computer Science, 544:125–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1991:CGC

[1495] U. Berger, W. Meixner, and B. Moller.

Calculating a garbage collector. Lecture Notes in Computer Science, 544: 137–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berghammer:1991:UEF

[1496] R. Berghammer and H. Ehler. On the use of elements of functional programming in program development by transformations. Lecture Notes in Computer Science, 544:193–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloos:1991:TDC

[1497] C. Delgado Kloos and W. Dosch.

Transformational development of circuit descriptions for binary adders. Lecture Notes in Computer Science, 544: 217–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alblas:1991:IAG

[1498] Henk Alblas. Introduction to attribute grammars. Lecture Notes in Computer Science, 545:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kastens:1991:AGS

[1499] Uwe Kastens. Attribute grammars as a specification method. Lecture Notes in Computer Science, 545:16–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alblas:1991:AEM

[1500] Henk Alblas. Attribute evaluation methods. Lecture Notes in Computer Science, 545:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kastens:1991:IVO

[1501] Uwe Kastens. Implementation of visitoriented attribute evaluators. Lecture Notes in Computer Science, 545:114– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

opdenAkker:1991:SAA

[1502] Rieks op den Akker and Erik Sluiman. Storage allocation for attribute evaluators using stacks and queues. Lecture Notes in Computer Science, 545:140– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moncke:1991:GFA

[1503] Ulrich Moncke and Reinhard Wilhelm. Grammar flow analysis. Lecture Notes in Computer Science, 545: 151–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

opdenAkker:1991:AEP

[1504] Rieks op den Akker, Borivoj Melichar, and Jorma Tarhio. Attribute evaluation and parsing. Lecture Notes in Computer Science, 545:187–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alblas:1991:IAE

[1505] Henk Alblas. Incremental attribute evaluation. Lecture Notes in Computer Science, 545:215–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourdan:1991:SPA

[1506] Martin Jourdan. A survey of parallel attribute evaluation methods. Lecture Notes in Computer Science, 545:234– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swierstra:1991:HOA

[1507] Doaitse Swierstra and Harald Vogt.

Higher order attribute grammars. Lecture Notes in Computer Science, 545: 256–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koskimies:1991:OOA

[1508] Kai Koskimies. Object-orientation in attribute grammars. Lecture Notes in Computer Science, 545:297–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maluszynski:1991:AGL

[1509] Jan Maluszynski. Attribute grammars and logic programs: A comparison of concepts. Lecture Notes in Computer Science, 545:330–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koster:1991:AGP

[1510] Kees Koster. Affix grammars for programming languages. Lecture Notes in Computer Science, 545:358–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hedin:1991:ISS

[1511] Gorel Hedin. Incremental staticsemantic analysis for object-oriented languages using door attribute grammars. Lecture Notes in Computer Science, 545:374–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kastens:1991:AGC

[1512] Uwe Kastens. Attribute grammars in a compiler construction environment. Lecture Notes in Computer Science, 545:380–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riedewald:1991:PUA

[1513] Gunter Riedewald. Prototyping by using an attribute grammar as a logic program. Lecture Notes in Computer Science, 545:401–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forbrig:1991:UGA

[1514] Peter Forbrig. Using the generative aspect of attribute grammars in a knowledge based way. Lecture Notes in Computer Science, 545:438–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gyimothy:1991:NLI

[1515] Tibor Gyimothy. Natural language interface construction using attribute grammars. Lecture Notes in Computer Science, 545:460–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koster:1991:AGN

[1516] Kees Koster. Affix grammars for natural languages. Lecture Notes in Computer Science, 545:469–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourdan:1991:IEF

[1517] Martin Jourdan and Didier Parigot.

Internals and externals of the FNC-2 attribute grammar system. Lecture Notes in Computer Science, 545:485– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipps:1991:OOS

[1518] Peter Lipps, Ulrich Moncke, and Reinhard Wilhelm. An overview of the OPTRAN system. Lecture Notes in Computer Science, 545:505–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilhelm:1991:ARO

[1519] Reinhard Wilhelm. Attribute reevaluation in OPTRAN. Lecture Notes in Computer Science, 545:507–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lammel:1991:TWS

[1520] Ralf Lammel. The translator writing system RUEGEN — KS. Lecture Notes in Computer Science, 545:508– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rollinger:1991:IL

[1521] Claus-Rainer Rollinger and Otthein Herzog. Introducing LILOG. Lecture Notes in Computer Science, 546:3–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauer:1991:TUC

[1522] Wilfried Brauer, Peter Lockemann, and Helmut Schnelle. Text understanding — the challenges to come. Lecture Notes in Computer Science, 546:14–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorre:1991:FNL

[1523] Jochen Dorre and Roland Seiffert. A formalism for natural language — STUF. Lecture Notes in Computer Science, 546:33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorre:1991:LS

[1524] Jochen Dorre. The language of STUF. Lecture Notes in Computer Science, 546:39–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seiffert:1991:CPS

[1525] Roland Seiffert. Chart-parsing of STUF grammars. Lecture Notes in Computer Science, 546:51–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorre:1991:SW

[1526] Jochen Dorre and Ingo Raasch. The STUF workbench. Lecture Notes in Computer Science, 546:55–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seiffert:1991:UIL

[1527] Roland Seiffert. Unification-ID/LP grammars: Formalization and parsing. Lecture Notes in Computer Science, 546:63–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erbach:1991:FPL

[1528] Gregor Erbach. A flexible parser for a linguistic development environment. Lecture Notes in Computer Science, 546:74–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konig:1991:GHM

[1529] Esther Konig. GAP-Handling mechanisms in categorial grammars. Lecture Notes in Computer Science, 546:88–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanngiesser:1991:OLL

[1530] Siegfried Kanngießer. Outlines of the LEU/2 lexicology. Lecture Notes in Computer Science, 546:105–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiller:1991:MPT

[1531] Anne Schiller and Petra Steffens. Morphological processing in the two-level paradigm. Lecture Notes in Computer Science, 546:112–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gust:1991:RWM

[1532] Helmar Gust. Representing word meanings. Lecture Notes in Computer Science, 546:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rickheit:1991:SIL

[1533] Mechthild Rickheit. Sortal information in lexical concepts. Lecture Notes in Computer Science, 546:143–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludewig:1991:IVE

[1534] Petra Ludewig. Incremental vocabulary extensions in text understanding systems. Lecture Notes in Computer Science, 546:153–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emde:1991:MLK

[1535] Werner Emde. Managing lexical knowledge in LEU/2. Lecture Notes in Computer Science, 546:167–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiss:1991:GL

[1536] Tibor Kiss. The grammars of LILOG. Lecture Notes in Computer Science, 546:183–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wesche:1991:APS

[1537] Birgit Wesche. An alternative phrase structure account of symmetric coordination. Lecture Notes in Computer Science, 546:200–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiss:1991:VOH

[1538] Tibor Kiss and Birgit Wesche. Verb order and head movement. Lecture Notes in Computer Science, 546:216– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosch:1991:BTN

[1539] Peter Bosch. The bermuda triangle: Natural language semantics between linguistics, knowledge representation, and knowledge processing. Lecture Notes in Computer Science, 546: 243–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderSandt:1991:PAL

[1540] Rob A. van der Sandt and Bart Geurts.

Presupposition, anaphora, and lexical content. Lecture Notes in Computer Science, 546:259–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayer:1991:ADR

[1541] Rolf Mayer. Anaphora and domain restriction. Lecture Notes in Computer Science, 546:297–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberle:1991:RTS

[1542] Kurt Eberle. On representing the temporal structure of texts. Lecture Notes in Computer Science, 546:313– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Link:1991:TPL

[1543] Godehard Link and Hinrich Schutze. The treatment of plurality in lsub LILOG. Lecture Notes in Computer Science, 546:342–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pletat:1991:KRL

[1544] Udo Pletat. The knowledge representation language lsub LILOG. Lecture Notes in Computer Science, 546:357– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wachsmuth:1991:KPK

[1545] Ipke Wachsmuth and Barbara Gangler. Knowledge packets and knowledge packet structures. Lecture Notes in Computer Science, 546:380–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitt:1991:DAT

[1546] Peter H. Schmitt. Deductive aspects of three-valued logic. Lecture Notes in Computer Science, 546:394–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bollinger:1991:LIE

[1547] Toni Bollinger, Sven Lorenz, and Udo Pletat. The LILOG inference engine. Lecture Notes in Computer Science, 546:402–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blasius:1991:KBC

[1548] Karl Hans Blasius. Knowledge based control of the LILOG inference engine: Kinds of metaknowledge. Lecture Notes in Computer Science, 546: 428–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nebel:1991:ADF

[1549] Bernhard Nebel and Gert Smolka.

Attributive description formalisms … and the rest of the world. Lecture Notes in Computer Science, 546:439– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klose:1991:BKL

[1550] Gudrun Klose and Kai von Luck. The background knowledge of the LILOG system. Lecture Notes in Computer Science, 546:455–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1991:LOL

[1551] Ewald Lang. The LILOG ontology from a linguistic point of view. Lecture Notes in Computer Science, 546: 464–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borkel:1991:KEE

[1552] Manfred Borkel and Peter Gerstl. A knowledge engineering environment for LILOG. Lecture Notes in Computer Science, 546:482–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Angele:1991:KEC

[1553] Jurgen Angele, Dieter Fensel, Dieter Landes, Susanne Neubert, and Rudi Studer. Knowledge engineering in the context of related fields of research. Lecture Notes in Computer Science, 546:490–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benzschawel:1991:LDD

[1554] Stefan Benzschawel, Erich Gehlen, Michael Ley, Thomas Ludwig, Albert Maier, and Bernd Walter. LILOG-DB: Database support for knowledge based systems. Lecture Notes in Computer Science, 546:501–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1991:PSE

[1555] Christopher Habel. Processing of spatial expressions in LILOG. Lecture Notes in Computer Science, 546:598– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pribbenow:1991:PL

[1556] Simone Pribbenow. Phenomena of localization. Lecture Notes in Computer Science, 546:609–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maienborn:1991:VMP

[1557] Claudia Maienborn. Verbs of motion and position: On the optionality of the local argument. Lecture Notes in Computer Science, 546:621–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carstensen:1991:WHC

[1558] Kai-Uwe Carstensen and Geoff Simmons. Why a hill can’t be a valley: Representing gestalt and position properties of objects with object schemata. Lecture Notes in Computer Science, 546:632–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khenkhar:1991:OOR

[1559] Mohammed N. Khenkhar. Objectoriented representation of depictions on the basis of cell matrices. Lecture Notes in Computer Science, 546:645– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Novak:1991:IGC

[1560] Hans-Joachim Novak. Integrating a generation component into a natural language understanding system. Lecture Notes in Computer Science, 546: 659–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobes:1991:KST

[1561] Zuzana Dobes and Hans-Joachim Novak. From knowledge structures to text structures. Lecture Notes in Computer Science, 546:670–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wendholt:1991:F

[1562] Birgit Wendholt. The formulator. Lecture Notes in Computer Science, 546: 685–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonLuck:1991:CCL

[1563] Kai von Luck and Thomas Pirlein. Constructing a context for LEU/2. Lecture Notes in Computer Science, 546: 703–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilms:1991:TUS

[1564] Jan Wilms. The text understanding system LEU/2. Lecture Notes in Computer Science, 546:711–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arning:1991:TBL

[1565] Andreas Arning. The trace of building a large AI system. Lecture Notes in Computer Science, 546:719–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biham:1991:DCF

[1566] Eli Biham and Adi Shamir. Differential cryptanalysis of Feal and N-hash. Lecture Notes in Computer Science, 547: 1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470001.
                                        pdf.

Lai:1991:MCD

[1567] Xuejia Lai, James L. Massey, and Sean Murphy. Markov ciphers and differential cryptanalysis. Lecture Notes in Computer Science, 547:17– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470017.
pdf.

Camion:1991:KHF

[1568] Paul Camion and Jacques Patarin. The knapsack hash function proposed at Crypto ’89 can be broken. Lecture Notes in Computer Science, 547:39– 53, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470039.
pdf.

Coster:1991:ILD

[1569] Matthijs J. Coster, B. A. LaMacchia, Andrew M. Odlyzko, and Claus P. Schnorr. An improved low-density subset sum algorithm. Lecture Notes in Computer Science, 547:54– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470054.
pdf.

Korzhik:1991:CMP

[1570] Valery I. Korzhik and Andrey I.

Turkin. Cryptanalysis of McEliece’s public-key cryptosystem. Lecture Notes in Computer Science, 547:68– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470068.
                                        pdf.

deRooij:1991:SSS

[1571] Peter de Rooij. On the security of the Schnorr scheme using preprocessing. Lecture Notes in Computer Science, 547:71–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470071.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470071.
                                        pdf.

Burmester:1991:BIP

[1572] Mike V. D. Burmester and Yvo Desmedt. Broadcast interactive proofs. Lecture Notes in Computer Science, 547:81–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0547/05470081.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470081.
                                        pdf.

Okamoto:1991:DZK

[1573] Tatsuaki Okamoto, David Chaum, and Kazuo Ohta. Direct zero knowledge proofs of computational power in five rounds. Lecture Notes in Computer Science, 547:96– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470096.
pdf.

Crepeau:1991:ROT

[1574] Claude Crépeau and Miklós Sántha.

On the reversibility of oblivious transfer. Lecture Notes in Computer Science, 547:106–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0547/05470106.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0547/05470106.pdf.

Mund:1991:ZLC

[1575] Sibylle Mund. Ziv–Lempel complexity for periodic sequences and its cryptographic application. Lecture Notes in Computer Science, 547:114– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470114.
pdf.

Habutsu:1991:SKC

[1576] Toshiki Habutsu, Yoshifumi Nishio, Iwao Sasase, and Shinsaku Mori. A secret key cryptosystem by iterating a chaotic map. Lecture Notes in Computer Science, 547:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470127.
                                        pdf.

Preneel:1991:BFS

[1577] Bart Preneel, René Govaerts, and Joos Vandewalle. Boolean functions satisfying higher order propagation criteria. Lecture Notes in Computer Science, 547:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470141.
                                        pdf.

Jansen:1991:MOC

[1578] Cees J. A. Jansen. The maximum order complexity of sequence ensembles. Lecture Notes in Computer Science, 547: 153–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470153.
                                        pdf.

Golic:1991:NOS

[1579] Jovan Dj. Golic. The number of output sequences of a binary sequence generator. Lecture Notes in Computer Science, 547:160–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/0547/05470160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470160.
pdf.

Dai:1991:LCP

[1580] Zong-Duo D. Dai and Jun-Hui H. Yang.

Linear complexity of periodically repeated random sequences. Lecture Notes in Computer Science, 547:168– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470168.
pdf.

Chepyzhov:1991:FCA

[1581] Vladimir Chepyzhov and Bernard Smeets. On a fast correlation attack on certain stream ciphers. Lecture Notes in Computer Science, 547:176– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470176.
pdf.

Meier:1991:APR

[1582] Willi Meier and Othmar Staffelbach.

Analysis of pseudo random sequences generated by cellular automata. Lecture Notes in Computer Science, 547: 186–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470186.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470186.
                                        pdf.

Chambers:1991:BSR

[1583] William G. Chambers and Zong-Duo D. Dai. On binary sequences from recursions “modulo 2^{e}” made non-linear by the bit-by-bit “XOR” function. Lecture Notes in Computer Science, 547:200–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470200.
                                        pdf.

Desmedt:1991:WUS

[1584] Yvo Desmedt and Moti Yung. Weaknesses of undeniable signature schemes (extended abstract). Lecture Notes in Computer Science, 547:205–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470205.
                                        pdf.

Pedersen:1991:DPA

[1585] Torben Pryds Pedersen. Distributed provers with applications to undeniable signatures. Lecture Notes in Computer Science, 547:221–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/0547/05470221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470221.
pdf.

Fujioka:1991:IBP

[1586] Atsushi Fujioka, Tatsuaki Okamoto, and Kazuo Ohta. Interactive biproof systems and undeniable signature schemes. Lecture Notes in Computer Science, 547:243–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470243.
pdf.

Chaum:1991:GS

[1587] David Chaum and Eugène van Heyst.

Group signatures. Lecture Notes in Computer Science, 547:257–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470257.
pdf.

Boyd:1991:ESD

[1588] Colin Boyd. Enhancing secrecy by data compression: Theoretical and practical aspects. Lecture Notes in Computer Science, 547:266–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470266.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470266.
                                        pdf.

Schnorr:1991:FIC

[1589] Claus P. Schnorr. Factoring integers and computing discrete logarithms via Diophantine approximation. Lecture Notes in Computer Science, 547:281– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470281.
                                        pdf.

Huber:1991:SCC

[1590] Klaus Huber. Some considerations concerning the selection of RSA moduli. Lecture Notes in Computer Science, 547:294–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0547/05470294.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470294.
                                        pdf.

Portz:1991:UIN

[1591] Michael Portz. On the use of interconnection networks in cryptography. Lecture Notes in Computer Science, 547: 302–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470302.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470302.
                                        pdf.

Beth:1991:NSE

[1592] Thomas Beth and F. Schaefer. Non supersingular elliptic curves for public key cryptosystems. Lecture Notes in Computer Science, 547:316–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470316.
pdf.

Morain:1991:BCE

[1593] François Morain. Building cyclic elliptic curves modulo large primes. Lecture Notes in Computer Science, 547: 328–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470328.
pdf.

Shizuya:1991:CHD

[1594] Hiroki Shizuya, Toshiya Itoh, and Kouichi Sakurai. On the complexity of hyperelliptic discrete logarithm problem. Lecture Notes in Computer Science, 547:337–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470337.
pdf.

Dawson:1991:ESB

[1595] M. H. Dawson and Stafford E. Tavares. An expanded set of S-box design criteria based on information theory and its relation to differential-like attacks. Lecture Notes in Computer Science, 547: 352–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470352.
                                        pdf.

OConnor:1991:ENP

[1596] Luke O’Connor. Enumerating nondegenerate permutations. Lecture Notes in Computer Science, 547:368– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470368.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470368.
                                        pdf.

Nyberg:1991:PNB

[1597] Kaisa Nyberg. Perfect nonlinear Sboxes. Lecture Notes in Computer Science, 547:378–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0547/05470378.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0547/05470378.pdf.

Rueppel:1991:FAS

[1598] Rainer A. Rueppel. A formal approach to security architectures. Lecture Notes in Computer Science, 547: 387–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470387.
pdf.

Horster:1991:DLB

[1599] Patrick Horster and Hans-Joachim Knobloch. Discrete logarithm based protocols. Lecture Notes in Computer Science, 547:399–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0547/05470399.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0547/05470399.pdf.

Matsumoto:1991:HIT

[1600] Tsutomu Matsumoto and Hideki Imai. Human identification through insecure channel. Lecture Notes in Computer Science, 547:409–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470409.
pdf.

Goldburg:1991:ACA

[1601] B. Goldburg, E. Dawson, and S. Sridharan. The automated cryptanalysis of analog speech scramblers. Lecture Notes in Computer Science, 547:422– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470422.
                                        pdf.

Sadeghiyan:1991:COW

[1602] Babak Sadeghiyan and Józef P. Pieprzyk. A construction for one way hash functions and pseudorandom bit generators. Lecture Notes in Computer Science, 547:431–445, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470431.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470431.
                                        pdf.

Fujioka:1991:EED

[1603] Atsushi Fujioka, Tatsuaki Okamoto, and Shoji Miyaguchi. ESIGN: An efficient digital signature implementation for smart cards. Lecture Notes in Computer Science, 547:446– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470446.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470446.
                                        pdf.

Maurer:1991:NAD

[1604] Ueli M. Maurer. New approaches to the design of self-synchronizing stream ciphers. Lecture Notes in Computer Science, 547:458–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470458.
pdf.

Pieprzyk:1991:RAS

[1605] Józef P. Pieprzyk and Reihaneh Safavi-Naini. Randomized authentication systems. Lecture Notes in Computer Science, 547:472–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470472.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470472.
pdf.

Gabidulin:1991:INC

[1606] E. M. Gabidulin, A. V. Paramonov, and O. V. Tretjakov. Ideals over a non-commutative ring and their application in cryptology. Lecture Notes in Computer Science, 547:482– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470482.
pdf.

Girault:1991:SCP

[1607] Marc Girault. Self-certified public keys.

Lecture Notes in Computer Science, 547:490–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0547/05470490.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470490.
                                        pdf.

Maurer:1991:NIP

[1608] Ueli M. Maurer and Yacov Yacobi. Non-interactive public-key cryptography. Lecture Notes in Computer Science, 547:498–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0547/05470498.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0547/05470498.pdf.

Zemor:1991:HFG

[1609] Gilles Zémor. Hash functions and graphs with large girths. Lecture Notes in Computer Science, 547:508– 511, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470508.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470508.
                                        pdf.

Muller:1991:DPP

[1610] Winfried B. Müller and Alan Oswald. Dickson pseudoprimes and primality testing. Lecture Notes in Computer Science, 547:512–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470512.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/0547/05470512.
pdf.

Gibson:1991:EGC

[1611] J. Keith Gibson. Equivalent Goppa codes and trapdoors to McEliece’s public key cryptosystem. Lecture Notes in Computer Science, 547:517– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470517.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470517.
pdf.

Pedersen:1991:TCT

[1612] Torben Pryds Pedersen. A threshold cryptosystem without a trusted party (extended abstract). Lecture Notes in Computer Science, 547:522– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470522.
pdf.

Mihaljevic:1991:CCP

[1613] Miodrag J. Mihaljevic and Jovan Dj.

Golic. A comparison of cryptanalytic principles based on iterative error-correction. Lecture Notes in Computer Science, 547:527–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470527.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470527.
                                        pdf.

Biham:1991:CCM

[1614] Eli Biham. Cryptanalysis of the chaotic-map cryptosystem suggested at EUROCRYPT ’91. Lecture Notes in Computer Science, 547:532–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470532.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470532.
                                        pdf.

Berkovits:1991:HBS

[1615] Shimshon Berkovits. How to broadcast a secret. Lecture Notes in Computer Science, 547:535–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470535.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470535.
                                        pdf.

Pieprzyk:1991:PAE

[1616] Józef P. Pieprzyk. Probabilistic analysis of elementary randomizers. Lecture Notes in Computer Science, 547:542– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0547/05470542.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0547/05470542.
                                        pdf.

Preneel:1991:RIP

[1617] Bart Preneel, David Chaum, Walter Fumy, Cees J. A. Jansen, Peter Landrock, and G. Roelofsen. Race integrity primitives evaluation (RIPE): A status report. Lecture Notes in Computer Science, 547:547– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470547.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470547.
pdf.

Brynielsson:1991:ILT

[1618] Lennart Brynielsson. The information leakage through a randomly generated function. Lecture Notes in Computer Science, 547:552–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0547/05470552.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0547/05470552.
pdf.

Chaum:1991:SWW

[1619] David Chaum. Some weaknesses of “weaknesses of undeniable signatures”. Lecture Notes in Computer Science, 547:554–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0547/05470554.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0547/05470554.
pdf.

Gabbay:1991:ALD

[1620] D. M. Gabbay. Abduction in labelled deductive systems — A conceptual abstract. Lecture Notes in Computer Science, 548:3–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardenfors:1991:NIE

[1621] P. Gardenfors. Nonmonotonic inference, expectations, and neural networks. Lecture Notes in Computer Science, 548:12–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:Ib

[1622] Anonymous. Introduction. Lecture Notes in Computer Science, 548:31–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amarger:1991:IQC

[1623] S. Amarger, D. Dubois, and H. Prade. Imprecise quantifiers and conditional probabilities. Lecture Notes in Computer Science, 548:33–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besnard:1991:DL

[1624] P. Besnard. Default logics. Lecture Notes in Computer Science, 548:38–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cano:1991:PUD

[1625] J. Cano, M. Delgado, and S. Moral. Propagation of uncertainty in dependence graphs. Lecture Notes in Computer Science, 548:42–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1991:EAB

[1626] M. Clarke and N. Wilson. Efficient algorithms for belief functions based on the relationship between belief and probability. Lecture Notes in Computer Science, 548:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1991:BOP

[1627] D. Dubois, J. Lang, and H. Prade. A brief overview of possibilistic logic. Lecture Notes in Computer Science, 548: 53–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

delCerro:1991:MAPb

[1628] L. Farinas del Cerro and A. Herzig.

A modal analysis of possibility theory. Lecture Notes in Computer Science, 548:58–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1991:ELL

[1629] J. Fox, P. Krause, and M. Dohnal. An extended logic language for representing belief. Lecture Notes in Computer Science, 548:63–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froidevaux:1991:GDL

[1630] Ch. Froidevaux, P. Chatalic, and J. Mengin. Graded default logics. Lecture Notes in Computer Science, 548: 70–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godo:1991:LEU

[1631] L. Godo and R. Lopez de Mantaras. Linguistically expressed uncertainty: Its elicitation and use in modular expert systems. Lecture Notes in Computer Science, 548:76–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kruse:1991:RMD

[1632] R. Kruse, J. Gebhardt, and F. Klawonn. Reasoning with mass distributions and the context model. Lecture Notes in Computer Science, 548:81–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1991:AP

[1633] S. Parsons. Advance prototyping. Lecture Notes in Computer Science, 548: 86–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smets:1991:TBM

[1634] P. Smets, Y-T Hsia, A. Saffiotti, R. Kennes, H. Xu, and E. Umkehrer. The transferable belief model. Lecture Notes in Computer Science, 548:91–98, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acid:1991:LC

[1635] S. Acid, L. M. de Campos, A. Gonzalez, and R. Molina. Learning with CASTLE. Lecture Notes in Computer Science, 548:99–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1991:NAI

[1636] J. F. Baldwin. A new approach to inference under uncertainty for knowledge based systems. Lecture Notes in Computer Science, 548:107–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binaghi:1991:LUCb

[1637] E. Binaghi. Learning of uncertain classification rules in medical diagnosis. Lecture Notes in Computer Science, 548:115–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brewka:1991:ADT

[1638] G. Brewka. Assertional default theories. Lecture Notes in Computer Science, 548:120–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardona:1991:RRU

[1639] L. Cardona, J. Kohlas, and P. A.

Monney. The reliability of reasoning with unreliable rules and propositions. Lecture Notes in Computer Science, 548:125–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chateauneuf:1991:UVR

[1640] A. Chateauneuf, R. Kast, and A. Lapied. Uncertainty in the valuation of risky assets. Lecture Notes in Computer Science, 548:130–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coletti:1991:AQJ

[1641] G. Coletti, A. Gilio, and R. Scozzafava. Assessment of qualitative judgements for conditional events in expert systems. Lecture Notes in Computer Science, 548:135–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CorreadaSilva:1991:ARA

[1642] F. S. Correa da Silva, D. Robertson, and P. Chung. Automated reasoning about an uncertain domain. Lecture Notes in Computer Science, 548:141– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cussens:1991:UDL

[1643] J. Cussens and A. Hunter. Using defeasible logic for a window on a probabilistic database: Some preliminary notes. Lecture Notes in Computer Science, 548:146–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diday:1991:DAU

[1644] E. Diday. From data analysis to uncertainty knowledge analysis. Lecture Notes in Computer Science, 548:153– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiNola:1991:DFR

[1645] A. Di Nola, W. Pedrycz, and S. Sessa. Difference fuzzy relation equations: Studies in dynamical systems. Lecture Notes in Computer Science, 548:161– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Driankov:1991:TLF

[1646] D. Driankov and H. Hellendoorn. Towards a logic for a fuzzy logic controller. Lecture Notes in Computer Science, 548:166–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Etzion:1991:HAD

[1647] O. Etzion. Handling active databases with partial inconsistencies. Lecture Notes in Computer Science, 548:171– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gacogne:1991:EPT

[1648] L. Gacogne. An extension of the possibility theory in view of the formalization of approximate reasoning. Lecture Notes in Computer Science, 548: 176–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goodwin:1991:PRP

[1649] S. D. Goodwin, E. Neufeld, and A. Trudel. Probabilistic regions of persistence. Lecture Notes in Computer Science, 548:182–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregoire:1991:FPL

[1650] E. Gregoire. Formalizing pertinence links in inheritance reasoning: Preliminary report. Lecture Notes in Computer Science, 548:190–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinsohn:1991:HAM

[1651] J. Heinsohn. A hybrid approach for modeling uncertainty in terminological logics. Lecture Notes in Computer Science, 548:198–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hellendoorn:1991:FCR

[1652] H. Hellendoorn and M. Reinfrank. Fuzzy control research at Siemens Corporate R and D. Lecture Notes in Computer Science, 548:206–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Junker:1991:HPO

[1653] U. Junker and G. Brewka. Handling partially ordered defaults in TMS. Lecture Notes in Computer Science, 548: 211–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1991:CED

[1654] F. Levy. Computing extensions of default theories. Lecture Notes in Computer Science, 548:219–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lohmann:1991:ERA

[1655] G. Lohmann. An evidential reasoning approach to the classification of satellite images. Lecture Notes in Computer Science, 548:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luo:1991:PPR

[1656] Z. Luo and A. Gammerman. PRESS — A probabilistic reasoning expert system shell. Lecture Notes in Computer Science, 548:232–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ma:1991:IUR

[1657] Y. Ma and D. C. Wilkins. Induction of uncertain rules and the sociopathicity property in Dempster-Shafer theory. Lecture Notes in Computer Science, 548:238–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacNish:1991:HDL

[1658] C. MacNish. Hierarchical default logic. Lecture Notes in Computer Science, 548:246–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1991:LIM

[1659] A. Nakamura. A logic of imprecise monadic predicates and its relation to the S5-Modal fuzzy logic. Lecture Notes in Computer Science, 548:254– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oppel:1991:ECS

[1660] U. G. Oppel. Every complex system can be determined by a causal probabilistic network without cycles and every such network determines a Markov field. Lecture Notes in Computer Science, 548:262–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paass:1991:PDRb

[1661] G. Paass. Probabilistic default reasoning involving continuous variables. Lecture Notes in Computer Science, 548: 267–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papini:1991:RPC

[1662] O. Papini. Revision in propositional calculus. Lecture Notes in Computer Science, 548:272–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pontet:1991:CBA

[1663] T. Pontet. A constraint-based approach to uncertain and imprecise reasoning. application to expert systems. Lecture Notes in Computer Science, 548:277–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quinio:1991:RCS

[1664] P. Quinio and T. Matsuyama. Random closed sets: A unified approach to the representation of imprecision and uncertainty. Lecture Notes in Computer Science, 548:282–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauzy:1991:KET

[1665] A. Rauzy. Knowledge extraction in trivalued propositional logic. Lecture Notes in Computer Science, 548:287– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rhodes:1991:UME

[1666] P. C. Rhodes and G. R. Garside. Using maximum entropy to identify unsafe assumptions in probabilistic expert systems. Lecture Notes in Computer Science, 548:292–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruspini:1991:TU

[1667] E. H. Ruspini. On truth and utility. Lecture Notes in Computer Science, 548:297–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaub:1991:CCD

[1668] T. Schaub. On commitment and cumulativity in default logics. Lecture Notes in Computer Science, 548:305– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwind:1991:TBC

[1669] C. B. Schwind and V. Risch. A tableau-based characterisation for default logic. Lecture Notes in Computer Science, 548:310–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snow:1991:RPW

[1670] P. Snow. Restraining the proliferation of worlds in probabilistic logic entailments. Lecture Notes in Computer Science, 548:318–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spies:1991:MUE

[1671] M. Spies. Managing uncertainty in environmental analysis: An application to measurement data interpretation. Lecture Notes in Computer Science, 548:323–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sucar:1991:HUK

[1672] L. E. Sucar and D. F. Gillies. Handling uncertainty in knowledge-based computer vision. Lecture Notes in Computer Science, 548:328–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thone:1991:PRF

[1673] H. Thone, U. Guntzer, and W. Kiebling. Probabilistic reasoning with facts and rules in deductive databases. Lecture Notes in Computer Science, 548: 333–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandenberghe:1991:ERA

[1674] R. M. Vandenberghe and R. M. de Caluwe. An entity-relationship approach to the modelling of vagueness in databases. Lecture Notes in Computer Science, 548:338–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voorbraak:1991:PMS

[1675] F. Voorbraak. A preferential model semantics for default logic. Lecture Notes in Computer Science, 548:344– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1991:EHN

[1676] E. Weydert. Elementary hyperentailment — nonmonotonic reasoning about defaults. Lecture Notes in Computer Science, 548:352–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wahlster:1991:KBM

[1677] W. Wahlster, E. Andre, W. Graf, and T. Rist. Knowledge-based media coordination in intelligent user interfaces. Lecture Notes in Computer Science, 549:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amati:1991:SNC

[1678] G. Amati. Some notes on cumulative reasoning. Lecture Notes in Computer Science, 549:18–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonatti:1991:FTV

[1679] P. A. Bonatti. A family of threevalued autoepistemic logics. Lecture Notes in Computer Science, 549:28–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simi:1991:VSB

[1680] M. Simi. Viewpoints subsume beliefs, truths and situations. Lecture Notes in Computer Science, 549:38–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1991:AMT

[1681] D. Basin, F. Giunchigia, and P. Traverso. Automating meta-theory creation and system extension. Lecture Notes in Computer Science, 549:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fronhofer:1991:IEI

[1682] B. Fronhofer. Implementing exceptions in inheritance by concept transforming actions. Lecture Notes in Computer Science, 549:58–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cadoli:1991:AE

[1683] M. Cadoli and M. Schaerf. Approximate entailment. Lecture Notes in Computer Science, 549:68–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordano:1991:NFC

[1684] L. Giordano, M. Martelli, and M. L. Sapino. Negation as failure and constraints through abduction and defaults. Lecture Notes in Computer Science, 549:78–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donni:1991:HSD

[1685] F. M. Donni, M. Lenzerini, D. Nardi, and A. Schaerf. A hybrid system with datalog and concept languages. Lecture Notes in Computer Science, 549:88–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellman:1991:OSC

[1686] J. Ellman and G. Mezzanatto. Optimal search for conjunctive goals using constraints. Lecture Notes in Computer Science, 549:98–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen-Winkeln:1991:FWI

[1687] R. M. Jansen-Winkeln, A. Ndiaye, and N. Reithinger. FSS-WASTL: interactive knowledge acquisition for a semantic lexicon. Lecture Notes in Computer Science, 549:108–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slack:1991:HEC

[1688] J.-M. Slack. Hybrid encoding: constraints on addressing structure. Lecture Notes in Computer Science, 549: 117–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonarini:1991:UCD

[1689] A. Bonarini, E. Cappelletti, and A. Corrao. Uncertainty and cyclic dependencies: a proposal and a network implementation. Lecture Notes in Computer Science, 549:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1991:ARM

[1690] A. Brogi, E. Lamma, P. Mancarelli, and P. Mello. Abductive reasoning in a multi-theory framework. Lecture Notes in Computer Science, 549:137– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evertsz:1991:AIH

[1691] R. Evertsz and E. Motta. The abstract interpretation of hybrid rule/ frame-based systems. Lecture Notes in Computer Science, 549:147–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bagnasco:1991:TRC

[1692] C. Bagnasco, P. Petrin, and L. Spampinato. Taxonomic reasoning in configuration tasks. Lecture Notes in Computer Science, 549:157–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brajnik:1991:RUT

[1693] G. Brajnik, L. Chittaro, G. Guida, C. Tasso, and E. Toppano. Representation and use of teleological knowledge in the multi-modeling approach. Lecture Notes in Computer Science, 549: 167–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Console:1991:TID

[1694] L. Console, D. Theseider-Dupre, and P. Torasso. Towards the integration of different knowledge sources in modelbased diagnosis. Lecture Notes in Computer Science, 549:177–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cermignani:1991:ISN

[1695] S. Cermignani and G. Tornielli. Integrating statistics, numerical analysis and dependency-recording in modelbased diagnosis. Lecture Notes in Computer Science, 549:187–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpineto:1991:VSI

[1696] C. Carpineto. Version-space induction with multiple concept languages. Lecture Notes in Computer Science, 549: 198–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerbone:1991:KCS

[1697] G. Cerbone and T. G. Dietterich. Knowledge compilation to speed up numerical optimization. Lecture Notes in Computer Science, 549:208–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorigo:1991:NPA

[1698] M. Dorigo. New perspectives about default hierarchies formation in learning classifier systems. Lecture Notes in Computer Science, 549:218–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botta:1991:UCM

[1699] M. Botta, F. Brancadori, S. Ravotto, S. Sperotto, and L. Saitta. Use of causal model to learn diagnostic knowledge in a real domain. Lecture Notes in Computer Science, 549:228–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordana:1991:EIR

[1700] A. Giordana, L. Saitta, D. Roverso, R. Finelli, and M. Paderni. Extending inverse resolution to build up abstractions. Lecture Notes in Computer Science, 549:238–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basili:1991:UWA

[1701] R. Basili, M. T. Pazienza, and P. Velardi. Using word association for syntactic disambiguation. Lecture Notes in Computer Science, 549:250–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fum:1991:CMT

[1702] D. Fum, C. Tasso, L. Tiepolo, and A. Tramontini. A computational model of tense selection and its experimentation within an intelligent tutor. Lecture Notes in Computer Science, 549: 261–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piastra:1991:ECF

[1703] M. Piastra and R. Bolognesi. An efficient context-free parsing algorithm with semantic actions. Lecture Notes in Computer Science, 549:271–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strapparava:1991:SIS

[1704] C. Strapparava. From scopings to interpretation: the semantic interpretation within the AIFresco system. Lecture Notes in Computer Science, 549: 281–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lombardo:1991:PDG

[1705] V. Lombardo. Parsing dependency grammars. Lecture Notes in Computer Science, 549:291–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardizzone:1991:SBN

[1706] E. Ardizzone, A. Chella, R. Pirrone, and F. Sorbello. A system based on neural architectures for the reconstruction of 3-D shapes from images. Lecture Notes in Computer Science, 549:302–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burbello:1991:RHL

[1707] F. Burbello, S. B. Pollard, J. Porrill, and J. E. W. Mayhew. Retrieval of high-level data structures from stereo vision data. Lecture Notes in Computer Science, 549:312–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adorni:1991:HLL

[1708] G. Adorni, A. Broggi, G. Conte, V. D’Andrea, and C. Sansoe. Highlevel and low-level computer vision: towards an integrated approach. Lecture Notes in Computer Science, 549:322– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardizzone:1991:ALG

[1709] E. Ardizzone, F. Callari, A. Chella, and F. Frixione. An associative link from geometric to symbolic representations in artificial vision. Lecture Notes in Computer Science, 549:332–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vercelli:1991:TSB

[1710] G. Vercelli, R. Zaccaria, and P. Morasso. A theory of sensor-based robot navigation using local information. Lecture Notes in Computer Science, 549:342–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1991:ICG

[1711] M. Klein and S. C.-Y. Lu. Insights into cooperative group design: experience with the LAN designer system. Lecture Notes in Computer Science, 549:354–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abbruzzese:1991:DIA

[1712] F. Abbruzzese and E. Minicozzi. Direct interaction among active data structures: a tool for building AI systems. Lecture Notes in Computer Science, 549:365–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artale:1991:IKR

[1713] A. Artale, F. Cesarini, and G. Soda. Introducing knowledge representation techniques in database models. Lecture Notes in Computer Science, 549: 375–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gioiello:1991:SOM

[1714] M. Gioiello, G. Vassallo, A. Chella, and F. Sorbello. Self-organizing maps: a new digital architecture. Lecture Notes in Computer Science, 549:385– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertin:1991:EEM

[1715] A. Bertin, F. Buciol, G. Dondossola, and C. Lanza. Electrical equipment maintenance training and its application in industrial environment. Lecture Notes in Computer Science, 549:399– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kantardzic:1991:TSD

[1716] M. Kantardzic, V. Okanovic, S. Rizvic, A. Filipovic, and H. Glavic. Temporal and spatial data and default assumptions in DIPSY-E system. Lecture Notes in Computer Science, 549:

410–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milanese:1991:VIA

[1717] R. Milanese, J.-M. Bost, and T. Pun.

Visual indexing with an attentive system. Lecture Notes in Computer Science, 549:415–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

dePinto:1991:NNP

[1718] P. de Pinto and M. Sette. On neural network programming. Lecture Notes in Computer Science, 549:420– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abbati:1991:LRS

[1719] M. Abbati and S. Odoardi. A logic for the representation of spatial knowledge. Lecture Notes in Computer Science, 549:425–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fringuelli:1991:TMA

[1720] B. Fringuelli, S. Marcugini, A. Milani, and S. Rivora. Truth maintenance in approximate reasoning. Lecture Notes in Computer Science, 549: 430–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McIntyre:1991:SDD

[1721] A. McIntyre. A strategy for design and development of complex knowledgesystems. Lecture Notes in Computer Science, 549:435–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attardi:1991:ECF

[1722] G. Attardi, M. Gaspari, and P. Iglio. Efficient compilation of first order predicates. Lecture Notes in Computer Science, 549:440–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brajnik:1991:FTA

[1723] G. Brajnik, C. Tasso, and A. Vaccher. A flexible tool for assumptionbased user modeling. Lecture Notes in Computer Science, 549:445–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frasconi:1991:KNM

[1724] P. Frasconi, M. Gori, M. Maggini, and G. Soda. KL: a neural model for capturing structure in speech. Lecture Notes in Computer Science, 549:450– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavelli:1991:LDW

[1725] G. Lavelli and B. Magnini. Lexical discrimination within a multilevel semantics approach. Lecture Notes in Computer Science, 549:455–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costantini:1991:MRA

[1726] S. Costantini and G. A. Lanzarone. Metalevel representation of analogical inference. Lecture Notes in Computer Science, 549:460–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardissono:1991:SIC

[1727] L. Ardissono and P. Terenziani. Semantic interpretation of copulative sentences. Lecture Notes in Computer Science, 549:465–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badaloni:1991:TPR

[1728] S. Badaloni and A. Zanardo. Typicality for plausible reasoning. Lecture Notes in Computer Science, 549:470– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roberto:1991:DIU

[1729] V. Roberto. A distributed imageunderstanding system. Lecture Notes in Computer Science, 549:475–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldman:1991:SCM

[1730] Stuart I. Feldman. Software configuration management: Past uses and future challenges. Lecture Notes in Computer Science, 550:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coutaz:1991:ADU

[1731] Joelle Coutaz. Architectural design for user interfaces. Lecture Notes in Computer Science, 550:7–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernstrom:1991:ESF

[1732] Christer Fernstrom. The eureka software factory: Concepts and accomplishments. Lecture Notes in Computer Science, 550:23–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dick:1991:ISF

[1733] J. Dick and J. Loubersac. Integrating structured and formal methods: A visual approach to VDM. Lecture Notes in Computer Science, 550:37–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattel:1991:RDD

[1734] T. Cattel. Rational design of distributed applications. Lecture Notes in Computer Science, 550:60–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dauchy:1991:TDS

[1735] P. Dauchy and B. Marre. Test data selection from algebraic specifications: Application to an automatic subway module. Lecture Notes in Computer Science, 550:80–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderLinden:1991:SCC

[1736] F. J. van der Linden. Specification in COLD-1 of a CAD-Package for drawing shadow masks. Lecture Notes in Computer Science, 550:101–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghezzi:1991:AAL

[1737] C. Ghezzi and R. A. Kemmerer. AS-TRAL: An assertion language for specifying real-time systems. Lecture Notes in Computer Science, 550:122– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Creusot:1991:EEE

[1738] D. Creusot, P. Lemoine, O. Roux, Y. Trinquet, A. Kung, O. Marbach, and C. Serrano-Morales. Execution environment for ELECTRE applications. Lecture Notes in Computer Science, 550:147–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopetz:1991:EAT

[1739] H. Kopetz, R. Zainlinger, G. Fohler, H. Kantz, P. Puschner, and W. Schutz. An engineering approach towards hard real-time system design. Lecture Notes in Computer Science, 550:166– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honiden:1991:AAI

[1740] S. Honiden, N. Uchihira, and K. Itoh.

An application of artificial intelligence to prototyping process in performance design for real-time systems. Lecture Notes in Computer Science, 550:189– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amador:1991:DRS

[1741] J. Amador, B. de Vicente, and A. Alonso. Dynamically replaceable software: A design method. Lecture Notes in Computer Science, 550:210– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berzins:1991:SMM

[1742] V. Berzins. Software merge: Models and methods for combining changes to programs. Lecture Notes in Computer Science, 550:229–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sijtsma:1991:TSD

[1743] B. A. Sijtsma and J. W. Mager. A theory for software design extraction. Lecture Notes in Computer Science, 550: 251–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coen-Porisini:1991:SES

[1744] A. Coen-Porisini and F. De Paoli. SESADA: An environment supporting software specialization. Lecture Notes in Computer Science, 550:266– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Selby:1991:MDC

[1745] R. W. Selby and R. K. Madsen. Metricdriven classification analysis. Lecture Notes in Computer Science, 550:290– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voas:1991:DFM

[1746] J. Voas. A dynamic failure model for predicting the impact that a program location has on the program. Lecture Notes in Computer Science, 550:308– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rozman:1991:RBS

[1747] I. Rozman, J. Gyorkos, and T. Dogsa.

Relation between source code metrics and structure analysis metrics. Lecture Notes in Computer Science, 550: 332–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shepperd:1991:AVS

[1748] M. Shepperd and D. Ince. Algebraic validation of software metrics. Lecture Notes in Computer Science, 550:343– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Presicce:1991:AVI

[1749] F. Parisi Presicce and A. Pierantonio. An algebraic view of inheritance and subtyping in object oriented programming. Lecture Notes in Computer Science, 550:364–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barghouti:1991:SRB

[1750] N. S. Barghouti and G. E. Kaiser.

Scaling up rule-based software development environments. Lecture Notes in Computer Science, 550:380–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snelting:1991:IBS

[1751] G. Snelting, F. J. Grosch, and U. Schroeder. Inference-based support for programming in the large. Lecture Notes in Computer Science, 550:396– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1991:TOO

[1752] T. Collins, K. Ewert, C. Gerety, J. Gustafson, and I. Thomas. TICKLE: Object-oriented description and composition services for software engineering environments. Lecture Notes in Computer Science, 550:409–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tattersall:1991:IPS

[1753] C. Tattersall. Integrated project support environments, text generation and technical writing. Lecture Notes in Computer Science, 550:424–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schefstrom:1991:AE

[1754] D. Schefstrom. The arcs experience. Lecture Notes in Computer Science, 550:443–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuggetta:1991:PP

[1755] A. Fuggetta. Panel presentation. Lecture Notes in Computer Science, 550: 465–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frasca:1991:PSF

[1756] V. Frasca. The production of software in the FINSIEL group. Lecture Notes in Computer Science, 550:468–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1991:ISE

[1757] C. Jackson. The impact of software engineering researches on industrial practice — A personal view. Lecture Notes in Computer Science, 550:472– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Troy:1991:IMP

[1758] R. Troy. Impact of methods on productivity and quality. Lecture Notes in Computer Science, 550:480–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feather:1991:REG

[1759] M. S. Feather. Requirements engineering: Getting right from wrong. Lecture Notes in Computer Science, 550: 485–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finkelstein:1991:NAR

[1760] A. Finkelstein. A (neat) alphabet of requirements engineering issues. Lecture Notes in Computer Science, 550: 489–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenspan:1991:SSR

[1761] Sol Greenspan. The scruffy side of requirements engineering. Lecture Notes in Computer Science, 550:492– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Potts:1991:EAT

[1762] C. Potts. Expediency and appropriate technology: An agenda for requirements engineering research in the 1990s. Lecture Notes in Computer Science, 550:495–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hruschka:1991:CSS

[1763] P. Hruschka. CASE support for the software process. Lecture Notes in Computer Science, 550:497–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1991:CSS

[1764] J. Kramer. CASE support for the software process: a research viewpoint. Lecture Notes in Computer Science, 550:499–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1991:CSL

[1765] D. Robinson. CASE support for large systems. Lecture Notes in Computer Science, 550:504–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1991:CSB

[1766] W. E. Fischer. CASE seen from both sides of the fence. Lecture Notes in Computer Science, 550:509–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1991:CSS

[1767] B. Lang. CASE support for the software process: Advances and problems. Lecture Notes in Computer Science, 550:512–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1991:DOB

[1768] Michael Jackson. Description is our business. Lecture Notes in Computer Science, 551:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milner:1991:CPO

[1769] Robin Milner. Concurrent processes as objects. Lecture Notes in Computer Science, 551:9–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guttag:1991:LAS

[1770] John V. Guttag. The larch approach to specification. Lecture Notes in Computer Science, 551:10–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Penny:1991:FSM

[1771] David A. Penny, Richard C. Holt, and Michael W. Godfrey. Formal specification in metamorphic programming. Lecture Notes in Computer Science, 551:11–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garlan:1991:FDS

[1772] David Garlan and David Notkin. Formalizing design spaces: Implicit invocation mechanisms. Lecture Notes in Computer Science, 551:31–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damm:1991:TCV

[1773] Flemming Damm, Bo Stig Hansen, and Hans Bruun. On type checking in VDM and related consistency issues. Lecture Notes in Computer Science, 551:45–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clement:1991:CTP

[1774] T. Clement. Combining transformation and posit-and-prove in a VDM development. Lecture Notes in Computer Science, 551:63–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plat:1991:CSA

[1775] Nico Plat, Jan van Katwijk, and Kees Pronk. A case for structured analysis/formal design. Lecture Notes in Computer Science, 551:81–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bril:1991:MOM

[1776] Reinder J. Bril. A model-oriented method for algebraic specifications using COLD-1 as notation. Lecture Notes in Computer Science, 551:106– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldschlag:1991:MFS

[1777] David M. Goldschlag. A mechanical formalization of several fairness notions. Lecture Notes in Computer Science, 551:125–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossig:1991:SSD

[1778] Stephan Rossig and Michael Schenke. Specification and stepwise development of communicating systems. Lecture Notes in Computer Science, 551:149– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benveniste:1991:WOS

[1779] Marc Benveniste. Writing operational semantics in Z: A structural approach. Lecture Notes in Computer Science, 551:164–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doma:1991:ESA

[1780] Veronika Doma and Robin Nicholl. EZ:

A system for automatic prototyping of Z specifications. Lecture Notes in Computer Science, 551:189–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanHee:1991:ZHL

[1781] K. M. van Hee, L. J. Somers, and M. Voorhoeve. Z and high level Petri nets. Lecture Notes in Computer Science, 551:204–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruun:1991:ASS

[1782] Hans Bruun, Bo Stig Hansen, and Flemming Damm. An approach to the static semantics of VDM-SL. Lecture Notes in Computer Science, 551:220– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1991:BEC

[1783] Michael J. Butler. Behavioural extension for CSP. Lecture Notes in Computer Science, 551:254–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bednarczyk:1991:CDF

[1784] Marek A. Bednarczyk and Andrzej M. Borzyszkowski. Cpo’s do not form a Cpo, and yet recursion works. Lecture Notes in Computer Science, 551:268– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Middelburg:1991:LML

[1785] C. A. Middelburg and G. R. Renardel de Lavalette. LPF and MPLomega — A logical comparison of VDM SL and COLD-K. Lecture Notes in Computer Science, 551:279–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattel:1991:TTD

[1786] Thierry Cattel. Tactical tools for distributing VDM specifications. Lecture Notes in Computer Science, 551:309– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stolen:1991:ARA

[1787] Ketil Stolen. An attempt to reason about shared-state concurrency in the style of VDM. Lecture Notes in Computer Science, 551:324–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elvang-Goransson:1991:RAV

[1788] Morten Elvang-Goransson. Reasoning about VDM specifications. Lecture Notes in Computer Science, 551:343– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arthan:1991:FSP

[1789] R. D. Arthan. Formal specification of a proof tool. Lecture Notes in Computer Science, 551:356–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bicarregui:1991:RAV

[1790] J. C. Bicarregui and B. Ritchie. Reasoning about VDM developments using the VDM support tool in mural. Lecture Notes in Computer Science, 551:371–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Craigen:1991:EO

[1791] Dan Craigen, Sentot Kromodimoeljo, Irwin Meisels, Bill Pase, and Mark Saaltink. EVES: An overview. Lecture Notes in Computer Science, 551: 389–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1991:DTV

[1792] Matthias Weber. Deriving transitivity of VDM-Reification in DEVA. Lecture Notes in Computer Science, 551:406– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonkers:1991:UPP

[1793] H. B. M. Jonkers. Upgrading the preand postcondition technique. Lecture Notes in Computer Science, 551:428– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1991:FDS

[1794] Paul Smith and Richard Keighley. The formal development of a secure transaction mechanism. Lecture Notes in Computer Science, 551:457–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deLavalette:1991:FDS

[1795] Gerard R. Renardel de Lavalette. Formal development of a serial copy management system. Lecture Notes in Computer Science, 551:477–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schewe:1991:SRI

[1796] Klaus-Dieter Schewe, Joachim W. Schmidt, and Ingrid Wetzel. Specification and refinement in an integrated database application environment. Lecture Notes in Computer Science, 551:496–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zave:1991:TPS

[1797] Pamela Zave and Michael Jackson. Techniques for partial specification and specification of switching systems. Lecture Notes in Computer Science, 551: 511–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parkin:1991:SMS

[1798] G. I. Parkin and G. O’Neill. Specification of the MAA standard in VDM. Lecture Notes in Computer Science, 551:526–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wing:1991:UWI

[1799] Jeannette M. Wing and Amy Moormann Zaremski. Unintrusive ways to integrate formal specifications in practice. Lecture Notes in Computer Science, 551:545–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bear:1991:OHS

[1800] Stephen Bear. An overview of HP-SL. Lecture Notes in Computer Science, 551:571–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Houston:1991:CPR

[1801] Iain Houston and Steve King. CICS project report: Experiences and results from the use of Z in IBM. Lecture Notes in Computer Science, 551:588– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinnaes:1991:DMI

[1802] D. Kinnaes and K. De Vlaminck. A debugger for a Meta-IV-like metalanguage. Lecture Notes in Computer Science, 551:597–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1991:ESM

[1803] Peter Gorm Larsen and Paul Bogh Lassen. An executable subset of Meta-IV with loose specification. Lecture Notes in Computer Science, 551:604– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marshall:1991:UVW

[1804] Lynn S. Marshall and Linda Simon. Using VDM within an object-oriented framework. Lecture Notes in Computer Science, 551:619–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puccetti:1991:ISD

[1805] A. Puccetti. The integrated software development and verification system ates. Lecture Notes in Computer Science, 551:629–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalmers:1991:URF

[1806] D. L. Chalmers, B. Dandanell, J. Gortz, J. S. Pedersen, and E. Zierau. Using RAISE — first impressions from the LaCoS applications. Lecture Notes in Computer Science, 551:645– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McParland:1991:SCP

[1807] P. McParland and P. Kilpatrick. A specification of a complex programming language statement. Lecture Notes in Computer Science, 551:658– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffman:1991:PS

[1808] Berthold Hoffman and Bernd Krieg-Bruckner. The PROSPECTRA system. Lecture Notes in Computer Science, 551:668–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dick:1991:VPV

[1809] Jeremy Dick and Jerome Loubersac.

The visual presentation of VDM specifications. Lecture Notes in Computer Science, 551:670–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1991:MS

[1810] Richard Moore and Peter Froome. mural and SpecBox. Lecture Notes in Computer Science, 551:673–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1991:VDC

[1811] Uwe Schmidt and Hans-Martin Horcher. The VDM domain compiler: A VDM class library generator. Lecture Notes in Computer Science, 551:675– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plat:1991:DVS

[1812] Nico Plat, Kees Pronk, and Marcel Verhoef. The delft VDM-SL front end. Lecture Notes in Computer Science, 551:677–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heping:1991:PTV

[1813] H. Heping and H. Zedan. Prototyping with temporal VDM: A status report. Lecture Notes in Computer Science, 551:681–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanHee:1991:ET

[1814] K. M. van Hee, L. J. Somers, and M. Voorhoeve. The ExSpect tool. Lecture Notes in Computer Science, 551: 683–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jordan:1991:CCA

[1815] David Jordan. Cadiz — computer aided design in Z. Lecture Notes in Computer Science, 551:685–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dollin:1991:HST

[1816] Chris Dollin. The HP-ST toolset. Lecture Notes in Computer Science, 551: 687–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Group:1991:RT

[1817] CRI RAISE Tools Group. The RAISE toolset. Lecture Notes in Computer Science, 551:689–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Houston:1991:IZT

[1818] Iain Houston. The IBM Z tool. Lecture Notes in Computer Science, 551: 691–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damm:1991:VSE

[1819] Flemming M. Damm, Hans Bruun, and Bo Stig Hansen. The VDM-SL editor and consistency checker. Lecture Notes in Computer Science, 551:693– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1991:BT

[1820] Matthew Lee and Ib. H. Sorensen.

B-tool. Lecture Notes in Computer Science, 551:695–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaue:1991:VSC

[1821] Christoph Blaue. A VDM subset compiler. Lecture Notes in Computer Science, 551:697–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:Pa

[1822] Anonymous. Preface. Lecture Notes in Computer Science, 551:v–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:TC

[1823] Anonymous. Table of contents. Lecture Notes in Computer Science, 551:ix–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guttag:1991:TLL

[1824] John V. Guttag and James J. Horning. A tutorial on Larch and LCL, A Larch/C interface language. Lecture Notes in Computer Science, 552:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woodcock:1991:TRC

[1825] J. C. P. Woodcock. A tutorial on the refinement calculus. Lecture Notes in Computer Science, 552:79–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacanAirchinnigh:1991:TLN

[1826] Micheal Mac an Airchinnigh. Tutorial lecture notes on the Irish school of the VDM. Lecture Notes in Computer Science, 552:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

George:1991:RSL

[1827] Chris George. The RAISE specification language: A tutorial. Lecture Notes in Computer Science, 552:238–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahl:1991:FDA

[1828] Ole-Johan Dahl and Olaf Owe. Formal development with ABEL. Lecture Notes in Computer Science, 552:320– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krieg-Bruckner:1991:PMS

[1829] Bernd Krieg-Bruckner, Einar W. Karlsen, Junbo Liu, and Owen Traynor. The PROSPECTRA methodology and system: Uniform transformational (meta-) development. Lecture Notes in Computer Science, 552:363– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abrial:1991:BML

[1830] J. R. Abrial. The B method for large software. specification, design and coding (abstract). Lecture Notes in Computer Science, 552:398–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Good:1991:MMD

[1831] Don I. Good and William D. Young.

Mathematical methods for digital systems development. Lecture Notes in Computer Science, 552:406–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:C

[1832] Anonymous. Contents. Lecture Notes in Computer Science, 552:ix–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:Pb

[1833] Anonymous. Preface. Lecture Notes in Computer Science, 552:v–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aurenhammer:1991:POP

[1834] Franz Aurenhammer, Gerd Stockl, and Emo Welzl. The post office problem for fuzzy point sets. Lecture Notes in Computer Science, 553:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1991:OAA

[1835] Bruno Becker, Paolo Giulio Franciosa, Stephan Gschwind, Thomas Ohler, Gerald Thiemt, and Peter Widmayer. An optimal algorithm for approximating a set of rectangles by two minimum area rectangles. Lecture Notes in Computer Science, 553:13–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bieri:1991:LAC

[1836] Hanspeter Bieri and Peter-Michael Schmidt. An on-line algorithm for constructing sweep planes in regular position. Lecture Notes in Computer Science, 553:27–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dominguez:1991:PAT

[1837] Salvador Dominguez and Oliver Gunther. Performance analysis of three curve presentation schemes. Lecture Notes in Computer Science, 553:37–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durst:1991:PDC

[1838] Martin J. Durst. Preclassification and delayed classification of boundary entities in arbitrary dimensions. Lecture Notes in Computer Science, 553:57–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engel:1991:MP

[1839] Peter Engel. On the morphology of polytopes in R^{d}. Lecture Notes in Computer Science, 553:71–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shiaofen:1991:RGM

[1840] Fang Shiaofen and Beat Bruderlin. Robustness in geometric modeling — tolerance-based methods. Lecture Notes in Computer Science, 553:85–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ihler:1991:SNC

[1841] Edmund Ihler, Gabriele Reich, and Peter Widmayer. On shortest networks for classes of points in the plane. Lecture Notes in Computer Science, 553:103–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1991:DSP

[1842] Xiao-Yi Jiang and Horst Bunke. Determination of the symmetries of polyhedra and an application to object recognition. Lecture Notes in Computer Science, 553:113–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1991:MAS

[1843] Rolf Klein. Moving along a street. Lecture Notes in Computer Science, 553: 123–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monney:1991:PGR

[1844] Paul-Andre Monney. Planar geometric reasoning with the theory of hints. Lecture Notes in Computer Science, 553:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1991:SAS

[1845] Heinrich Muller and Matthias Otte.

Solving algebraic systems in Bernstein-Bezier representation. Lecture Notes in Computer Science, 553:161–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nievergelt:1991:XPE

[1846] Jurg Nievergelt, Peter Schorn, Michele De Lorenzi, Christoph Ammann, and Adrian Brungger. XYZ: A project in experimental geometric computation. Lecture Notes in Computer Science, 553:171–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schorn:1991:IXG

[1847] Peter Schorn. Implementing the XYZ GeoBench: A programming environment for geometric algorithms. Lecture Notes in Computer Science, 553: 187–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1991:CRL

[1848] Bengt J. Nilsson and Sven Schuierer. Computing the rectilinear link diameter of a polygon. Lecture Notes in Computer Science, 553:203–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noltemeier:1991:LFM

[1849] Hartmut Noltemeier. Layout of flexible manufacturing systems — selected problems. Lecture Notes in Computer Science, 553:217–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roos:1991:DVD

[1850] Thomas Roos and Hartmut Noltemeier. Dynamic Voronoi diagrams in motion planning. Lecture Notes in Computer Science, 553:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schipper:1991:GTM

[1851] Haijo Schipper. Generating triangulations of 2-Manifolds. Lecture Notes in Computer Science, 553:237–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1991:TTN

[1852] Ralf Schneider and Hans-Peter Kriegel.

The TR^{∗}-tree: A new representation of polygonal objects supporting spatial queries and operations. Lecture Notes in Computer Science, 553:249– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1991:VDB

[1853] Thomas Schreiber. A Voronoi diagram based adaptive K-means-type clustering algorithm for multidimensional weighted data. Lecture Notes in Computer Science, 553:265–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuierer:1991:GSV

[1854] Sven Schuierer, Gregory J. E. Rawlins, and Derick Wood. A generalization of staircase visibility. Lecture Notes in Computer Science, 553:277–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1991:NSL

[1855] Klaus Simon. A new simple linear algorithm to recognize interval graphs. Lecture Notes in Computer Science, 553:289–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stifter:1991:PAC

[1856] Sabine Stifter. Predictions about collision free paths from intersection tests. Lecture Notes in Computer Science, 553:309–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grahne:1991:PII

[1857] G. Grahne. The problem of incomplete information in relational databases. Lecture Notes in Computer Science, 554:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1991:WAF

[1858] G. Barth. What is all the fuss about neural networks? Lecture Notes in Computer Science, 555:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauer:1991:NPI

[1859] W. Brauer. The new paradigm of informatics. Lecture Notes in Computer Science, 555:15–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlson:1991:VTN

[1860] P. A. Carlson. Virtual text and new habits of mind. Lecture Notes in Computer Science, 555:25–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cremers:1991:ACP

[1861] A. B. Cremers and Th. N. Hibbard. Axioms for concurrent processes. Lecture Notes in Computer Science, 555:54–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1991:NMI

[1862] K. Culik and S. Dube. New methods for image generation and compression. Lecture Notes in Computer Science, 555:69–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domik:1991:RVU

[1863] G. Domik. The role of visualization in understanding data. Lecture Notes in Computer Science, 555:91–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edelsbrunner:1991:ZTH

[1864] H. Edelsbrunner, R. Seidel, and M. Sharir. On the zone theorem for hyperplane arrangements. Lecture Notes in Computer Science, 555:108– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estivill-Castro:1991:SMD

[1865] V. Estivill-Castro and D. Wood. Sorting, measures of disorder, and worstcase performance. Lecture Notes in Computer Science, 555:124–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fellner:1991:OOP

[1866] D. W. Fellner. Object-oriented programming — does it help in computer graphics? Lecture Notes in Computer Science, 555:132–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hambrusch:1991:NAA

[1867] S. E. Hambrusch and H.-Y. Tu.

New algorithms and approaches for 1dimensional layout compaction. Lecture Notes in Computer Science, 555: 152–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriegel:1991:QPS

[1868] H.-P. Kriegel. Query processing in spatial database systems. Lecture Notes in Computer Science, 555:172–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moret:1991:HFM

[1869] B. M. E. Moret and H. D. Shapiro. How to find a minimum spanning tree in practice. Lecture Notes in Computer Science, 555:192–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1991:CTD

[1870] G. Muller. Current trends in distributed systems. Lecture Notes in Computer Science, 555:204–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norrie:1991:OOD

[1871] D. H. Norrie and A. D. Kwok. Objectoriented distributed artificial intelligence. Lecture Notes in Computer Science, 555:225–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ottmann:1991:TPV

[1872] T. Ottmann. Trees — a personal view. Lecture Notes in Computer Science, 555:243–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preece:1991:SUT

[1873] J. Preece. Supporting user testing in human-computer interaction design. Lecture Notes in Computer Science, 555:256–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rozenberg:1991:LS

[1874] G. Rozenberg. Labeled 2-Structures.

Lecture Notes in Computer Science, 555:268–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomaa:1991:VRS

[1875] A. Salomaa. Verifying and recasting secret ballots in computer networks. Lecture Notes in Computer Science, 555:283–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaer:1991:FTV

[1876] J. Schaer and M. G. Stone. Face traverses and a volume algorithm for polyhedra. Lecture Notes in Computer Science, 555:290–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwald:1991:AAP

[1877] A. Schwald. Ada — on the acceptance problems of a planned innovation. Lecture Notes in Computer Science, 555:298–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Six:1991:UID

[1878] H.-W. Six and J. Voss. User interface development: Problems and experiences. Lecture Notes in Computer Science, 555:306–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomek:1991:HPF

[1879] I. Tomek and H. Maurer. Hypermedia — from the past to the future. Lecture Notes in Computer Science, 555:320–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegner:1991:LFD

[1880] L. M. Wegner. Let the fingers do the walking: Object manipulation in an NF^{2} database editor. Lecture Notes in Computer Science, 555:337–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welzl:1991:SED

[1881] E. Welzl. Smallest enclosing disks (balls and ellipsoids). Lecture Notes in Computer Science, 555:359–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1991:FC

[1882] M. R. Williams. The first computer. Lecture Notes in Computer Science, 555:371–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witten:1991:EPF

[1883] I. H. Witten and D. L. Maulsby. Evaluating programs formed by example: an informational heuristic. Lecture Notes in Computer Science, 555:388– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadimitriou:1991:DMI

[1884] C. H. Papadimitriou. Decision-making with incomplete information. Lecture Notes in Computer Science, 557:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1991:MIS

[1885] D. T. Lee and Majid Sarrafzadeh. Maximum independent set of a permutation graph in k tracks. Lecture Notes in Computer Science, 557:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1991:ASR

[1886] Yaw-Ling Lin and Steven S. Skiena.

Algorithms for square roots of graphs. Lecture Notes in Computer Science, 557:12–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1991:DME

[1887] Satoshi Fujita, Masafumi Yamashita, and Tadashi Ae. Distributed k-mutual exclusion problem and k-coteries. Lecture Notes in Computer Science, 557: 22–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raghunathan:1991:SEB

[1888] Arvind Raghunathan and Huzur Saran.

Is the shuffle-exchange better than the butterfly? Lecture Notes in Computer Science, 557:32–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1991:WRA

[1889] A. C. Yao. Weighted random assignments with application to hashing. Lecture Notes in Computer Science, 557: 42, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakano:1991:SFT

[1890] Shin-Ichi Nakano and Takao Nishizeki. Scheduling file transfers under port and channel constraints. Lecture Notes in Computer Science, 557:43–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wen-Lian:1991:SDC

[1891] Hsu Wen-Lian and Ma Tze-Heng. Substitution decomposition on chordal graphs and applications. Lecture Notes in Computer Science, 557:52–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1991:MSP

[1892] Atsushi Takahashi, Ueno Shuichi, and Kajitani Yoji. Mixed-searching and proper-path-width. Lecture Notes in Computer Science, 557:61–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1991:SWR

[1893] Frank Wagner and Barbara Wolfers. Short wire routing in convex grids. Lecture Notes in Computer Science, 557: 72–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1991:NAK

[1894] Dorothea Wagner. A new approach to knock-knee channel routing. Lecture Notes in Computer Science, 557:83–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1991:CPA

[1895] Tetsuo Asano and Takeshi Tokuyama.

Circuit partitioning algorithms: Graph model versus geometry model. Lecture Notes in Computer Science, 557:94–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boros:1991:IMP

[1896] Endre Boros, Peter L. Hammer, Toshihide Ibaraki, and Kazuhiko Kawakami. Identifying 2-Monotonic positive Boolean functions in polynomial time. Lecture Notes in Computer Science, 557:104–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1991:ACA

[1897] T. H. Hu, C. Y. Tang, and R. C. T.

Lee. An average case analysis of Monien and Speckenmeyer’s mechanical theorem proving algorithm. Lecture Notes in Computer Science, 557: 116–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwong-Fai:1991:LAN

[1898] Chan Kwong-Fai and Wah Lam Tak.

An on-line algorithm for navigating in unknown terrain. Lecture Notes in Computer Science, 557:127–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janardan:1991:MWD

[1899] Ravi Janardan. On maintaining the width and diameter of a planar pointset online. Lecture Notes in Computer Science, 557:137–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edelsbrunner:1991:OTR

[1900] Herbert Edelsbrunner. Optimal triangulations by retriangulating. Lecture Notes in Computer Science, 557:150– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guibas:1991:APS

[1901] Leonidas J. Guibas, John E. Hershberger, Joseph S. B. Mitchell, and Jack Scott Snoeyink. Approximating polygons and subdivisions with minimum link paths. Lecture Notes in Computer Science, 557:151–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xue-Hou:1991:IAC

[1902] Tan Xue-Hou, Tomio Hirata, and Yasuyoshi Inagaki. An incremental algorithm for constructing shortest watchman routes. Lecture Notes in Computer Science, 557:163–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1991:HGP

[1903] H. S. Lee and R. C. Chang. On hitting grid points in a convex polygon with straight lines. Lecture Notes in Computer Science, 557:176–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkouar:1991:CSH

[1904] A. Benkouar, Y. G. Manoussakis, V. Th. Paschos, and R. Saad. On the complexity of some Hamiltonian and Eulerian problems in edge-colored complete graphs. Lecture Notes in Computer Science, 557:190–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1991:DPI

[1905] Takao Asano. Dynamic programming on intervals. Lecture Notes in Computer Science, 557:199–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szpankowski:1991:COT

[1906] Wojciech Szpankowski. Combinatorial optimization through order statistics. Lecture Notes in Computer Science, 557:208–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guibas:1991:CAG

[1907] Leonidas J. Guibas. Combinatorics and algorithms of geometric arrangements. Lecture Notes in Computer Science, 557:218–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1991:ARS

[1908] Susumu Hasegawa, Hiroshi Imai, and Koji Hakata. An analysis of randomized shear sort on the mesh computer. Lecture Notes in Computer Science, 557:219–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

I-Chen:1991:EPD

[1909] Wu I-Chen. Efficient parallel divideand-conquer for a class of interconnection topologies. Lecture Notes in Computer Science, 557:229–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1991:OSR

[1910] M. Y. Chan, F. Y. L. Chin, and C. K. Poon. Optimal specified root embedding of full binary trees in faulty hypercubes. Lecture Notes in Computer Science, 557:241–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleischer:1991:TLB

[1911] Rudolf Fleischer. A tight lower bound for the worst case of bottom-up Heapsort. Lecture Notes in Computer Science, 557:251–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moffat:1991:HSS

[1912] Alistair Moffat and Ola Petersson. Historical searching and sorting. Lecture Notes in Computer Science, 557:263– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1991:CEW

[1913] Arne Andersson and Tony W. Lai. Comparison-efficient and write-optimal searching and sorting. Lecture Notes in Computer Science, 557:273–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1991:NNR

[1914] F. Yao. Nearest neighbors revisited. Lecture Notes in Computer Science, 557:283–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estivill-Castro:1991:CRT

[1915] Vladimir Estivill-Castro and Murray Sherk. Competitiveness and response time in on-line algorithms. Lecture Notes in Computer Science, 557:284– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jan-Ming:1991:LTO

[1916] Ho Jan-Ming. A linear time optimal via assignment algorithm for threedimensional channel routing. Lecture Notes in Computer Science, 557:294– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Longpre:1991:SIO

[1917] Luc Longpre and Sarah Mocas. Symmetry of information and one-way functions. Lecture Notes in Computer Science, 557:308–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lih-Hsing:1991:LTA

[1918] Hsu Lih-Hsing, J. Y. Hwang, T. Y.

Ho, and C. H. Tsai. A linear time algorithm to recognize the double Euler trail for series-parallel networks. Lecture Notes in Computer Science, 557: 316–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsan-Sheng:1991:FSA

[1919] Hsu Tsan-Sheng and Vijaya Ramachandran. On finding a smallest augmentation to biconnect a graph. Lecture Notes in Computer Science, 557:326–??, 1991. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1991:FAE

[1920] Michael Kaufmann and Gerhard Klar. A faster algorithm for edge-disjoint paths in planar graphs. Lecture Notes in Computer Science, 557:336– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenhof:1991:OCM

[1921] Hans-Peter Lenhof and Michiel Smid. An optimal construction method for generalized convex layers. Lecture Notes in Computer Science, 557:349– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smid:1991:RPL

[1922] Michiel Smid. Rectangular point location and the dynamic closest pair problem. Lecture Notes in Computer Science, 557:364–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ip-Wang:1991:PAS

[1923] Chan Ip-Wang and Donald K. Friesen. Parallel algorithms for some dominance problems based on a CREW PRAM. Lecture Notes in Computer Science, 557:375–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diks:1991:PAF

[1924] Krzysztof Diks, Oscar Garrido, and Andrzej Lingas. Parallel algorithms for finding maximal k-dependent sets and maximal f-matchings. Lecture Notes in Computer Science, 557:385–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooman:1991:SCV

[1925] J. Hooman. Specification and compositional verification of real-time systems. Lecture Notes in Computer Science, 558:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1991:FAP

[1926] G. Butler. Fundamental algorithms for permutation groups. Lecture Notes in Computer Science, 559:xii + 238, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1991:PC

[1927] Manuel Blum. Program checking. Lecture Notes in Computer Science, 560: 1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blass:1991:RRS

[1928] Andreas Blass and Yuri Gurevich. Randomizing reductions of search problems. Lecture Notes in Computer Science, 560:10–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sands:1991:TAC

[1929] David Sands. Time analysis, cost equivalence, and program refinement. Lecture Notes in Computer Science, 560:25–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitra:1991:AES

[1930] Subrata Mitra and G. Sivakumar. AC-Equation solving. Lecture Notes in Computer Science, 560:40–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1991:OIC

[1931] Qing-Ping Tan and Huo-Wang Chen. On the operational interpretation of complex types. Lecture Notes in Computer Science, 560:57–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lodaya:1991:TLL

[1932] Kamal Lodaya and R. Ramanujam. Tense logics for local reasoning in distributed systems. Lecture Notes in Computer Science, 560:71–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1991:FSS

[1933] Luca Aceto and Uffe Engberg. Failures semantics for a simple process language with refinement. Lecture Notes in Computer Science, 560:89–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hungar:1991:CPPb

[1934] Hardi Hungar. Correctness of programs over poor signatures. Lecture Notes in Computer Science, 560:109–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1991:CIV

[1935] John Case, Sanjay Jain, and Arun Sharma. Complexity issues for vacillatory function identification. Lecture Notes in Computer Science, 560:121– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeSaec:1991:PAP

[1936] Bertrand Le Saec, Jean-Eric Pin, and Pascal Weil. A purely algebraic proof of McNaughton’s theorem on infinite words. Lecture Notes in Computer Science, 560:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tao:1991:SCM

[1937] Jiang Tao, Edward McDowell, and B. Ravikumar. The structure and complexity of minimal NFAs over a unary alphabet. Lecture Notes in Computer Science, 560:152–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahajan:1991:RCA

[1938] Meena Mahajan and Kamala Krithivasan. Relativised cellular automata and complexity classes. Lecture Notes in Computer Science, 560: 172–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1991:COL

[1939] Sam Kim and Robert McNaughton.

Computing the order of a locally testable automaton. Lecture Notes in Computer Science, 560:186–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldsmith:1991:SCI

[1940] Judy Goldsmith, Lane A. Hemachandra, and Kenneth Kunen. On the structure and complexity of infinite sets with minimal perfect hash functions. Lecture Notes in Computer Science, 560:212–223, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1991:NHS

[1941] Manindra Agrawal. NP-hard sets and creativeness over constant time languages. Lecture Notes in Computer Science, 560:224–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1991:CPI

[1942] Iain A. Stewart. Complete problems involving Boolean labelled structures and projection translations. Lecture Notes in Computer Science, 560:242– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vinay:1991:BCP

[1943] V. Vinay. Is BP crplus P a probabilistic class? Lecture Notes in Computer Science, 560:261–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munro:1991:FSP

[1944] J. Ian Munro and Venkatesh Raman. Fast stable in-place sorting with O(n) data moves. Lecture Notes in Computer Science, 560:266–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paschos:1991:TAS

[1945] V. Th. Paschos. A theorem on the approximation of set cover and vertex cover. Lecture Notes in Computer Science, 560:278–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patkar:1991:FAP

[1946] Sachin B. Patkar and H. Narayanan.

A fast algorithm for the principal partition of a graph. Lecture Notes in Computer Science, 560:288–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niepel:1991:UCE

[1947] Inga Niepel and Peter Rossmanith.

Uniform circuits and exclusive read PRAMs. Lecture Notes in Computer Science, 560:307–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furer:1991:CPG

[1948] Martin Furer and Balaji Raghavachari.

Contracting planar graphs efficiently in parallel. Lecture Notes in Computer Science, 560:319–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krizanc:1991:FDS

[1949] Danny Krizanc, Lata Narayanan, and Rajeev Raman. Fast deterministic selection on mesh-connected processor arrays. Lecture Notes in Computer Science, 560:336–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mansour:1991:IST

[1950] Yishay Mansour, James K. Park, Baruch Schieber, and Sandeep Sen. Improved selection in totally monotone arrays. Lecture Notes in Computer Science, 560:347–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadimitriou:1991:DSC

[1951] Christos H. Papadimitriou, Venkat Rangan, and Martha Sideri. Designing secure communication protocols from trust specifications. Lecture Notes in Computer Science, 560:360–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosh:1991:CSP

[1952] Subir Kumar Ghosh, Anil Maheshwari, Sudebkumar Prasant Pal, Sanjeev Saluja, and C. E. Veni Madhavan. Computing the shortest path tree in a weak visibility polygon. Lecture Notes in Computer Science, 560:369– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhattacharya:1991:UAS

[1953] Binay K. Bhattacharya. Usefulness of angle-sweep over line-sweep. Lecture Notes in Computer Science, 560:390– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1991:STS

[1954] C. Ding, G. Xiao, and W. Shan. The stability theory of stream ciphers. Lecture Notes in Computer Science, 561: 1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breu:1991:AST

[1955] R. Breu. Algebraic specification techniques in object oriented programming environments. Lecture Notes in Computer Science, 562:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Satyanarayanan:1991:ARL

[1956] M. Satyanarayanan. An agenda for research in large-scale distributed data repositories. Lecture Notes in Computer Science, 563:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saltzer:1991:FSI

[1957] J. H. Saltzer. File system indexing and backup. Lecture Notes in Computer Science, 563:13–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gifford:1991:IFS

[1958] D. W. Gifford and J. W. O’Toole. Intelligent file systems for object repositories. Lecture Notes in Computer Science, 563:20–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloman:1991:UDS

[1959] M. Sloman. Using domains for specifying management policy. Lecture Notes in Computer Science, 563:25–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:DSa

[1960] Anonymous. Discussion summary. Lecture Notes in Computer Science, 563: 31–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joy:1991:FOS

[1961] B. Joy. Future operating systems: An environmental scenario and consequential challenges. Lecture Notes in Computer Science, 563:34–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lazowska:1991:OSS

[1962] E. D. Lazowska. Operating system support for high-performance architectures. Lecture Notes in Computer Science, 563:40–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlichtiger:1991:CCS

[1963] P. Schlichtiger. Closely coupled systems. Lecture Notes in Computer Science, 563:44–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karshmer:1991:OSR

[1964] A. I. Karshmer and J. N. Thomas. Are operating systems at RISC? Lecture Notes in Computer Science, 563:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zahorjan:1991:OSB

[1965] J. Zahorjan. Operating systems of the 90s and beyond (with emphasis on multiprocessing). Lecture Notes in Computer Science, 563:53–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennett:1991:MDS

[1966] J. K. Bennett, J. B. Carter, and W. Zwaenepoel. Munin: Distributed shared memory using multi-protocol release consistency. Lecture Notes in Computer Science, 563:56–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:DSb

[1967] Anonymous. Discussion summary. Lecture Notes in Computer Science, 563: 61–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmutz:1991:AHC

[1968] H. Schmutz. Autonomous heterogeneous computing — some open problems. Lecture Notes in Computer Science, 563:63–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Randell:1991:CP

[1969] B. Randell. Coping with the past. Lecture Notes in Computer Science, 563: 72–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nehmer:1991:IOS

[1970] J. Nehmer. The immortality of operating systems, or: Is research in operating systems still justified? Lecture Notes in Computer Science, 563:77–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:DSc

[1971] Anonymous. Discussion summary. Lecture Notes in Computer Science, 563: 84–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopetz:1991:ETV

[1972] H. Kopetz. Event-triggered versus time-triggered real-time systems. Lecture Notes in Computer Science, 563: 87–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tripathi:1991:PSS

[1973] S. K. Tripathi and V. Nirkhe. Prescheduling for synchronization in hard real-time systems. Lecture Notes in Computer Science, 563:102–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeLann:1991:NFO

[1974] G. Le Lann. Notes on future operating systems for real-time dependable distributed computing. Lecture Notes in Computer Science, 563:109–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:DSd

[1975] Anonymous. Discussion summary. Lecture Notes in Computer Science, 563: 116–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cristian:1991:BCI

[1976] F. Cristian. Basic concepts and issues in fault-tolerant distributed systems. Lecture Notes in Computer Science, 563:119–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlichting:1991:OSF

[1977] F. D. Schlichting. Operating systems and fault-tolerance. Lecture Notes in Computer Science, 563:150–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birman:1991:FTS

[1978] K. Birman. Fault-tolerance in sixthgeneration operating systems. Lecture Notes in Computer Science, 563:154– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babaoglu:1991:FTS

[1979] O. Babaoglu. Fault tolerance support in future operating systems. Lecture Notes in Computer Science, 563:162– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banatre:1991:FSS

[1980] J. P. Banatre and M. Banatre. Fast stable storage as a basis for faulttolerant architectures. Lecture Notes in Computer Science, 563:167–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horn:1991:SMA

[1981] C. Horn, B. Coghlan, N. Harris, and J. Jones. Stable memory — another look. Lecture Notes in Computer Science, 563:171–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroger:1991:SLS

[1982] R. Kroger and E. Nett. System level support for dependable distributed applications. Lecture Notes in Computer Science, 563:178–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:DSf

[1983] Anonymous. Discussion summary. Lecture Notes in Computer Science, 563: 182–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neuman:1991:PSI

[1984] B. C. Neuman. Protection and security issues for future systems. Lecture Notes in Computer Science, 563:184– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cole:1991:SIS

[1985] R. Cole. Some issues in security arising from distribution. Lecture Notes in Computer Science, 563:202–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stiegler:1991:SSD

[1986] H. G. Stiegler. Security: System development in a new context. Lecture Notes in Computer Science, 563:207– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartig:1991:ISF

[1987] H. Hartig, W. Kuhnhauser, and J. Liedtke. Issues in security and fault tolerance. Lecture Notes in Computer Science, 563:212–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:DSg

[1988] Anonymous. Discussion summary. Lecture Notes in Computer Science, 563: 217–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Needham:1991:WNS

[1989] R. Needham. What next? some speculations. Lecture Notes in Computer Science, 563:220–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spector:1991:STS

[1990] A. Z. Spector. Some thoughts on systems challenges for the 1990s. Lecture Notes in Computer Science, 563:223– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gien:1991:NGO

[1991] M. Gien. Next generation operating systems architecture. Lecture Notes in Computer Science, 563:227–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokoro:1991:TCS

[1992] M. Tokoro. Toward computing systems for the 2000s. Lecture Notes in Computer Science, 563:233–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1991:EIF

[1993] F. Hofmann, P. Schlenk, and T. Eirich.

Encapsulation and interaction in future operating systems. Lecture Notes in Computer Science, 563:241–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wettstein:1991:EOS

[1994] H. D. Wettstein. The explainable operating system. Lecture Notes in Computer Science, 563:247–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1991:SOO

[1995] M. Shapiro. Soul: An object-oriented OS framework for object support. Lecture Notes in Computer Science, 563: 251–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shrivastava:1991:OSS

[1996] S. Shrivastava and D. McCue. Operating system support for object-oriented distributed systems. Lecture Notes in Computer Science, 563:256–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1991:DSh

[1997] Anonymous. Discussion summary. Lecture Notes in Computer Science, 563: 258–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamming:1991:THM

[1998] M. G. Lamming. Towards a human memory prosthesis. Lecture Notes in Computer Science, 563:262–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mullender:1991:SND

[1999] S. J. Mullender. Systems for the nineties — distributed multimedia systems. Lecture Notes in Computer Science, 563:273–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrtwich:1991:RPS

[2000] R. G. Herrtwich. The role of performance, scheduling and resource reservation in multimedia systems. Lecture Notes in Computer Science, 563:279– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molzberger:1991:PC

[2001] P. Molzberger. On parallel consciousness. Lecture Notes in Computer Science, 565:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucadou:1991:CNL

[2002] W. v. Lucadou. Complementarity and non-locality in complex systems. Lecture Notes in Computer Science, 565: 21–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caianiello:1991:SU

[2003] E. Caianiello. Systems and uncertainty. Lecture Notes in Computer Science, 565:56–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quadranti:1991:CSL

[2004] P. Quadranti. The construction of space and the logics of quantum mechanics. Lecture Notes in Computer Science, 565:67–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gernert:1991:CAC

[2005] D. Gernert. Cellular automata and the concept of space. Lecture Notes in Computer Science, 565:94–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fritsch:1991:MSP

[2006] G. Fritsch. Multipurpose and special purpose computers. Lecture Notes in Computer Science, 565:103–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knecht:1991:IDC

[2007] R. Knecht. Implementation of divideand-conquer algorithms on multiprocessors. Lecture Notes in Computer Science, 565:121–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vollmar:1991:SRS

[2008] R. Vollmar. Some remarks on synchronization problems. Lecture Notes in Computer Science, 565:137–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makhaniok:1991:EID

[2009] M. Makhaniok, V. Cherniavsky, R. Manner, and O. Stucky. Effective implementation of distributed arbitration in multiprocessor systems. Lecture Notes in Computer Science, 565:148– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pensini:1991:FTS

[2010] M. P. Pensini, G. Mauri, and F. Gardin. Flowshop and travelling salesman problem. Lecture Notes in Computer Science, 565:157–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauser:1991:ACN

[2011] R. Hauser, H. Horner, R. Manner, and M. Makhaniok. Architectural considerations for NERV — a general purpose neural network simulation system. Lecture Notes in Computer Science, 565:183–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobson:1991:ANN

[2012] V. G. Dobson. Associative neural networks in analog VLSI: Advantages of decrementing algorithms. Lecture Notes in Computer Science, 565:196– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eikmeyer:1991:CMU

[2013] H.-J. Eikmeyer. Connectionist models of utterance production. Lecture Notes in Computer Science, 565:219– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalenoort:1991:SOI

[2014] G. J. Dalenoort. Self-organization of informational systems. Lecture Notes in Computer Science, 565:231–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rieger:1991:DSR

[2015] B. B. Rieger. Distributed semantic representations of word meanings. Lecture Notes in Computer Science, 565:243– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wermus:1991:LCR

[2016] H. Wermus. Logic of cognitive representations and their evolution. Lecture Notes in Computer Science, 565:274– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atmanspacher:1991:MPC

[2017] H. Atmanspacher, F. R. Krueger, and H. Scheingraber. A modal propositional calculus for quantum facts and dynamical theories. Lecture Notes in Computer Science, 565:304–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundemann:1991:SOC

[2018] F. W. Mundemann. Self-organization in computational systems. Lecture Notes in Computer Science, 565:315– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1991:AEL

[2019] J. D. Becker. Adaptation and extension: Learning in rigid systems. Lecture Notes in Computer Science, 565: 326–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katsikas:1991:CDG

[2020] A. A. Katsikas and J. S. Nicolis. Chaotic dynamics of generating Markov partitions, and linguistic sequences mimicking Zipf’s law. Lecture Notes in Computer Science, 565:335– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuster:1991:OCM

[2021] P. Schuster. Optimization and complexity in molecular biology and physics. Lecture Notes in Computer Science, 565:364–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwefel:1991:UEC

[2022] H. P. Schwefel. Understanding evolution as a collective strategy for groping in the dark. Lecture Notes in Computer Science, 565:388–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muhlenbein:1991:PGA

[2023] H. Muhlenbein. Parallel genetic algorithms, population genetics, and combinatorial optimization. Lecture Notes in Computer Science, 565:398– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorges-Schleuter:1991:APG

[2024] M. Gorges-Schleuter. ASPARAGOS, A parallel genetic algorithm and population genetics. Lecture Notes in Computer Science, 565:407–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1991:ETI

[2025] P. M. Allen and M. J. Lesser. Evolution: Travelling in an imaginary landscape. Lecture Notes in Computer Science, 565:419–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banzhaf:1991:FGM

[2026] W. Banzhaf. Finding the global minimum of a low-dimensional spin-glass model. Lecture Notes in Computer Science, 565:442–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ablay:1991:TTR

[2027] P. Ablay. Ten theses regarding the design of controlled evolutionary strategies. Lecture Notes in Computer Science, 565:457–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atmanspacher:1991:ILD

[2028] H. Atmanspacher. Incommensurability of Liouvillean dynamics and information dynamics. Lecture Notes in Computer Science, 565:482–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lohmann:1991:SES

[2029] R. Lohmann. Selforganization by evolutionary strategies in visual systems. Lecture Notes in Computer Science, 565:500–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesser:1991:STG

[2030] J. Hesser, R. Manner, and O. Stucky. On Steiner trees and genetic algorithms. Lecture Notes in Computer Science, 565:509–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lanzelotte:1991:ONQ

[2031] R. Lanzelotte, P. Valduriez, M. Ziane, and J.-P. Cheiney. Optimization of nonrecursive queries in OODBs. Lecture Notes in Computer Science, 566: 1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiao:1991:OMN

[2032] Z. Jiao and P. M. D. Gray. Optimization of methods in a navigational query language. Lecture Notes in Computer Science, 566:22–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Straube:1991:EPG

[2033] D. D. Straube and M. Tamer Ozsu. Execution plan generation for an objectoriented data model. Lecture Notes in Computer Science, 566:43–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Low:1991:EAM

[2034] C. C. Low, H. Lu, B. C. Ooi, and J. Han. Efficient access methods in deductive and object-oriented databases. Lecture Notes in Computer Science, 566:68–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przymusinski:1991:SDL

[2035] T. C. Przymusinski. Semantics of disjunctive logic programs and deductive databases. Lecture Notes in Computer Science, 566:85–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paredaens:1991:GLD

[2036] J. Paredaens, P. Peelman, and L. Tanca. G-log: A declarative graphical query language. Lecture Notes in Computer Science, 566:108–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giannotti:1991:NDD

[2037] F. Giannotti, D. Pedreschi, D. Sacca, and C. Zaniolo. Nondeterminism in deductive databases. Lecture Notes in Computer Science, 566:129–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1991:DSE

[2038] W. Chen. Declarative specification and evaluation of database updates. Lecture Notes in Computer Science, 566: 147–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waller:1991:SUC

[2039] E. Waller. Schema updates and consistency. Lecture Notes in Computer Science, 566:167–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholl:1991:UVO

[2040] M. H. Scholl, C. Laasch, and M. Tresch. Updatable views in object-oriented databases. Lecture Notes in Computer Science, 566:189–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wood:1991:SCB

[2041] P. T. Wood. Syntactic characterizations of 1-Bounded datalog programs. Lecture Notes in Computer Science, 566:208–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBussche:1991:EOC

[2042] J. Van den Bussche. Evaluation and optimization of complex object selections. Lecture Notes in Computer Science, 566:226–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sagiv:1991:TEC

[2043] Y. Sagiv. On testing effective computability of magic programs. Lecture Notes in Computer Science, 566:244– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ullman:1991:CDO

[2044] J. D. Ullman. A comparison of deductive and object-oriented database systems. Lecture Notes in Computer Science, 566:263–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torlone:1991:UDD

[2045] R. Torlone and P. Atzeni. Updating deductive databases with functional dependencies. Lecture Notes in Computer Science, 566:278–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caseau:1991:COO

[2046] Y. Caseau. Constraints in an objectoriented deductive database. Lecture Notes in Computer Science, 566:292– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coburn:1991:PCG

[2047] N. Coburn and G. E. Weddell. Path constraints for graph-based data models: Towards a unified theory of typing constraints, equations, and functional dependencies. Lecture Notes in Computer Science, 566:312–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minohara:1991:PDT

[2048] T. Minohara and M. Tokoro. Providing dynamic type abstractions and specifications for persistent information. Lecture Notes in Computer Science, 566:332–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sciore:1991:MVO

[2049] E. Sciore. Multidimensional versioning for object-oriented databases. Lecture Notes in Computer Science, 566:355– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ngu:1991:CNC

[2050] A. Ngu, L. Wong, and S. Widjojo. On canonical and non-canonical classifications. Lecture Notes in Computer Science, 566:371–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsukamoto:1991:DTR

[2051] M. Tsukamoto, S. Nishio, and M. Fujio. DOT: A term representation using DOT algebra for knowledgebases. Lecture Notes in Computer Science, 566:391–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brass:1991:SIL

[2052] S. Brass and U. W. Lipeck. Semantics of inheritance in logical object specifications. Lecture Notes in Computer Science, 566:411–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wieringa:1991:FOU

[2053] R. J. Wieringa. A formalization of objects using equational dynamic logic. Lecture Notes in Computer Science, 566:431–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkinson:1991:VPS

[2054] M. Atkinson. A vision of persistent systems. Lecture Notes in Computer Science, 566:453–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeusfeld:1991:ROO

[2055] M. Jeusfeld and M. Jarke. From relational to object-oriented integrity simplification. Lecture Notes in Computer Science, 566:460–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuchenhoff:1991:ECD

[2056] V. Kuchenhoff. On the efficient computation of the difference between consecutive database states. Lecture Notes in Computer Science, 566:478–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheiney:1991:SOP

[2057] J.-P. Cheiney and Y.-N. Huang. Setoriented propagation of updates into transitively closed relations. Lecture Notes in Computer Science, 566:503– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imielinski:1991:QDP

[2058] T. Imielinski, S. Naqui, and K. Vadaparty. Querying design and planning databases. Lecture Notes in Computer Science, 566:524–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalinichenko:1991:GIR

[2059] L. Kalinichenko and V. Zadorozhny. A generalized information resource query language and basic query evaluation technique. Lecture Notes in Computer Science, 566:546–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boucelma:1991:EFQ

[2060] O. Boucelma and J. Le Maitre. An extensible functional query language for an object-oriented database system. Lecture Notes in Computer Science, 566:567–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaniolo:1991:EPD

[2061] C. Zaniolo. Efficient processing of declarative rule-based languages for databases. Lecture Notes in Computer Science, 567:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorochevsky:1991:DHP

[2062] M. Dorochevsky, J. Noye, and O. Thibault. Has dedicated hardware for Prolog a future? Lecture Notes in Computer Science, 567:17–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moss:1991:CAL

[2063] C. Moss. Commercial applications of large Prolog knowledge bases. Lecture Notes in Computer Science, 567:32–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1991:CCG

[2064] G. Ellis. Compiling conceptual graphs.

Lecture Notes in Computer Science, 567:41–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willems:1991:SKG

[2065] M. Willems. Subsumption in knowledge graphs. Lecture Notes in Computer Science, 567:56–??, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1991:TKR

[2066] F. Baader and B. Hollunder. A terminological knowledge representation system with complete inference algorithms. Lecture Notes in Computer Science, 567:67–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaeger:1991:IDC

[2067] H. Jaeger. An introduction to dynamic concept systems. Lecture Notes in Computer Science, 567:87–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenzerini:1991:QCB

[2068] M. Lenzerini and A. Schaerf. Querying concept-based knowledge bases. Lecture Notes in Computer Science, 567: 107–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artale:1991:SCO

[2069] A. Artale, F. Cesarini, and G. Soda. Subsumption computation in an object-oriented data model. Lecture Notes in Computer Science, 567:124– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guarino:1991:CPI

[2070] N. Guarino. A concise presentation of ITL. Lecture Notes in Computer Science, 567:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Backofen:1991:DDL

[2071] R. Backofen, L. Euler, and G. Gorz. Distributed disjunctions for LIFE. Lecture Notes in Computer Science, 567: 161–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ribeiro:1991:RMT

[2072] C. Ribeiro and A. Porto. Reasoning with maximal time intervals. Lecture Notes in Computer Science, 567:171– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriauciukas:1991:NML

[2073] V. Kriauciukas. Nonclassical models for logic programs. Lecture Notes in Computer Science, 567:179–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borger:1991:LOS

[2074] E. Borger and E. Riccobene. Logical operational semantics of Parlog — part I: And-parallellism. Lecture Notes in Computer Science, 567:191–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markov:1991:TBC

[2075] Z. Markov. A tool for building connectionist-like networks based on term unification. Lecture Notes in Computer Science, 567:199–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riekert:1991:PDA

[2076] W.-F. Riekert. Providing declarative access to a processing system for satellite image data. Lecture Notes in Computer Science, 567:214–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treur:1991:DFD

[2077] J. Treur. Declarative functionality descriptions of interactive reasoning modules. Lecture Notes in Computer Science, 567:221–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caseau:1991:RAC

[2078] Y. Caseau. Rule-aided constraint resolution in LAURE. Lecture Notes in Computer Science, 567:237–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mantsivoda:1991:FFL

[2079] A. V. Mantsivoda. Flang: A functional logic language. Lecture Notes in Computer Science, 567:257–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonzon:1991:PFD

[2080] P. Bonzon. Processing functional definitions as declarative knowledge: A reduced bytecode implementation of a functional logic machine. Lecture Notes in Computer Science, 567:271– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1991:RSO

[2081] A. King and P. Soper. Reducing scheduling overheads for concurrent logic programs. Lecture Notes in Computer Science, 567:279–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kautz:1991:GFK

[2082] H. Kautz and B. Selman. A general framework for knowledge compilation. Lecture Notes in Computer Science, 567:287–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosmadopoulos:1991:DDT

[2083] Y. Cosmadopoulos, M. Sergot, and R. W. Southwick. Data-driven transformation of meta-interpreters: A sketch. Lecture Notes in Computer Science, 567:301–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruynooghe:1991:IEC

[2084] M. Bruynooghe, V. Dumortier, and G. Janssens. Improving the efficiency of constraint logic programming languages by deriving specialized versions. Lecture Notes in Computer Science, 567:309–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yaoqing:1991:PPS

[2085] G. Yaoqing, W. Dingxing, Q. Xiaolin, H. Zhiyi, and H. Shouren. Parallelizing Prolog on shared-memory multiprocessors. Lecture Notes in Computer Science, 567:318–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamma:1991:PAR

[2086] E. Lamma and P. Mello. Processing abductive reasoning via contextual logic programming. Lecture Notes in Computer Science, 567:336–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1991:EIN

[2087] M. Hanus. Efficient implementation of narrowing and rewriting. Lecture Notes in Computer Science, 567:344– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1991:TTI

[2088] P. Balbiani, A. Herzig, and M. Lima Marques. TIM: The toulouse inference machine for non-classical logic programming. Lecture Notes in Computer Science, 567:366–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boley:1991:PDP

[2089] H. Boley, M. Meier, C. Moss, M. M. Richter, and A. A. Voronkov. PANEL: Declarative and procedural paradigms — do they really compete? Lecture Notes in Computer Science, 567:383– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aronsson:1991:GIP

[2090] M. Aronsson and L.-H. Eriksson. The GCLA II programming language. Lecture Notes in Computer Science, 567: 399–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beierle:1991:BDP

[2091] C. Beierle, G. Meyer, and H. Semle. A brief description of the PROTOS-L system. Lecture Notes in Computer Science, 567:402–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fricke:1991:PPC

[2092] S. Fricke and M. Hein. PCPL — PRO-LOG constraint processing library, version 2.0. Lecture Notes in Computer Science, 567:405–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guarino:1991:MDI

[2093] N. Guarino. A mini-description of the ITL system. Lecture Notes in Computer Science, 567:408–??, 1991.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanschke:1991:TCL

[2094] P. Hanschke, A. Abecker, and D. Drollinger. TAXON: A concept language with concrete domains. Lecture Notes in Computer Science, 567:411– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1991:ASE

[2095] M. Hanus. The ALF system: An efficient implementation of a functional logic language. Lecture Notes in Computer Science, 567:414–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinkelmann:1991:THC

[2096] K. Hinkelmann, M. Harm, and T. Labisch. Transforming Horn clauses for forward reasoning. Lecture Notes in Computer Science, 567:417–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mantsivoda:1991:IFL

[2097] A. Mantsivoda and V. Petukhin. Implementation of the functional logic language flang. Lecture Notes in Computer Science, 567:420–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meier:1991:SEP

[2098] M. Meier and J. Schimpf. SEPIA 3.0 — an extensible Prolog system. Lecture Notes in Computer Science, 567: 422–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1991:FEF

[2099] M. Meyer, J. Muller, and S. Schrodl. FIDO: Exploring finite domain consistency techniques in logic programming. Lecture Notes in Computer Science, 567:425–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burckert:1991:RPL

[2100] H.-J. Burckert. A resolution principle for a logic with restricted quantifiers. Lecture Notes in Computer Science, 568:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gittins:1991:DPS

[2101] M. Gittins. Debugging parallel Strand programs. Lecture Notes in Computer Science, 569:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorochevsky:1991:CHG

[2102] M. Dorochevsky, K. Schuerman, A. Veron, and J. Xu. Constraint handling, garbage collection and execution model issues in ElipSys. Lecture Notes in Computer Science, 569:17–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kefalas:1991:CSM

[2103] P. Kefalas and T. J. Reynolds. Controlling search with meta-brave. Lecture Notes in Computer Science, 569:29–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szeredi:1991:SOP

[2104] P. Szeredi. Solving optimization problems in the aurora Or-parallel Prolog system. Lecture Notes in Computer Science, 569:39–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moolenaar:1991:KAP

[2105] R. Moolenaar, H. Van Acker, and B. Demoen. καππα: A kernel Andorra Prolog parallel architecture design. Lecture Notes in Computer Science, 569:54–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deCastroDutra:1991:FSA

[2106] I. de Castro Dutra. A flexible scheduler in the Andorra-I system. Lecture Notes in Computer Science, 569:70–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bahgat:1991:PAM

[2107] R. Bahgat. The Pandora Abstract Machine: An extension of JAM. Lecture Notes in Computer Science, 569:83–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ali:1991:PMB

[2108] K. Ali, R. Karlsson, and S. Mudambi.

Performance of MUSE on the BBN Butterfly TC2000. Lecture Notes in Computer Science, 569:104–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaumont:1991:SSS

[2109] A. Beaumont. Scheduling strategies and speculative work. Lecture Notes in Computer Science, 569:120–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ertel:1991:PCP

[2110] W. Ertel. Performance of competitive Or-parallelism. Lecture Notes in Computer Science, 569:132–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1991:APC

[2111] G. Gupta and M. Hermenegildo. ACE: And/Or-parallel copying-based execution of logic programs. Lecture Notes in Computer Science, 569:146–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBosschere:1991:BCP

[2112] K. O. M. De Bosschere. Blackboard communication in Prolog. Lecture Notes in Computer Science, 569:159– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Succi:1991:DPL

[2113] G. Succi and G. Marino. Data parallelism in logic programming. Lecture Notes in Computer Science, 569:173– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourzoufi:1991:EBM

[2114] H. Bourzoufi, G. Goncalves, P. Lecouffe, and B. Toursel. An efficient binding management in Or-parallel model. Lecture Notes in Computer Science, 569:185–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goswami:1991:IIL

[2115] A. Goswami, M. Bell, and M. Joseph.

ISL: An interval logic for the specification of real-time programs. Lecture Notes in Computer Science, 571:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1991:DSS

[2116] Chaochen Zhou, M. R. Hansen, A. P.

Ravn, and H. Rischel. Duration specifications for shared processors. Lecture Notes in Computer Science, 571:21–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coenen:1991:CSF

[2117] J. Coenen and J. Hooman. A compositional semantics for fault-tolerant real-time systems. Lecture Notes in Computer Science, 571:33–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniels:1991:MRT

[2118] Mats Daniels. Modelling real-time behaviour with an interval time calculus. Lecture Notes in Computer Science, 571:53–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millet:1991:MRL

[2119] Odile Millet. Multicycles and RTL logic satisfiability. Lecture Notes in Computer Science, 571:73–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moitra:1991:VPT

[2120] Abha Moitra. Voluntary preemption:

a tool in the design of hard real-time systems. Lecture Notes in Computer Science, 571:87–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narayana:1991:OTP

[2121] K. T. Narayana. Observing task preemption in Ada 9X. Lecture Notes in Computer Science, 571:107–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breveglieri:1991:RTS

[2122] L. Breveglieri, A. Cherubini, and S. Crespi-Reghizzi. Real-time scheduling by queue automata. Lecture Notes in Computer Science, 571:131– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1991:BCR

[2123] J. Davies, D. Jackson, and S. Schneider. Broadcast communication for real-time processes. Lecture Notes in Computer Science, 571:149–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deLemos:1991:ATR

[2124] R. de Lemos, A. Saeed, and T. Anderson. Analysis of timeliness requirements in safety-critical systems. Lecture Notes in Computer Science, 571: 171–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yodaiken:1991:VRN

[2125] V. Yodaiken and K. Ramamritham. Verification of a reliable net protocol. Lecture Notes in Computer Science, 571:193–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shankar:1991:MVG

[2126] N. Shankar. Mechanical verification of a generalized protocol for Byzantine fault tolerant clock synchronization. Lecture Notes in Computer Science, 571:217–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rushby:1991:FSV

[2127] J. Rushby. Formal specification and verification of a fault-masking and transient-recovery model for digital flight-control systems. Lecture Notes in Computer Science, 571:237–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deylon:1991:FTS

[2128] B. Deylon and O. Maler. On faulttolerant symbolic computations. Lecture Notes in Computer Science, 571: 259–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heidtmann:1991:TLA

[2129] K. D. Heidtmann. Temporal logic applied to reliability modelling of faulttolerant systems. Lecture Notes in Computer Science, 571:271–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnan:1991:SAT

[2130] P. Krishnan and P. D. Mosses. Specifying asynchronous transfer of control. Lecture Notes in Computer Science, 571:291–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssen:1991:PDL

[2131] W. Janssen and J. Zwiers. Protocol design by layered decomposition: A compositional approach. Lecture Notes in Computer Science, 571:307–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurki-Suonio:1991:SRT

[2132] R. Kurki-Suonio, K. Systa, and J. Vain. Scheduling in real-time models. Lecture Notes in Computer Science, 571: 327–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naik:1991:TAR

[2133] Y. Naik. A temporal approach to requirement specification of real-time systems. Lecture Notes in Computer Science, 571:341–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaice:1991:RGR

[2134] J. A. Plaice. RLucid, a general real-time dataflow language. Lecture Notes in Computer Science, 571:363– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardell-Oliver:1991:MTV

[2135] R. Cardell-Oliver. A mechanized theory for the verification of realtime program code uUing higher order logic. Lecture Notes in Computer Science, 571:375–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidge:1991:SVR

[2136] C. J. Fidge. Specification and verification of real-time behaviour using Z and RTL. Lecture Notes in Computer Science, 571:393–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholefield:1991:TFF

[2137] D. J. Scholefield and H. S. M. Zedan.

TAM: A formal framework for the development of distributed real-time systems. Lecture Notes in Computer Science, 571:411–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richard:1991:ACA

[2138] M. Richard and O. Roux. An attempt to confront asynchronous reality to synchronous modelization in the ESTEREL language. Lecture Notes in Computer Science, 571:429–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBoer:1991:RTB

[2139] F. de Boer and J. Hooman. The realtime behaviour of asynchronously communicating processes. Lecture Notes in Computer Science, 571:451–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1991:ACR

[2140] J. C. M. Baeten and J. A. Bergstra.

Asynchronous communication in real space process algebra. Lecture Notes in Computer Science, 571:473–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffrey:1991:TTP

[2141] A. Jeffrey. Translating timed process algebra into prioritized process algebra. Lecture Notes in Computer Science, 571:493–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ortega-Mallen:1991:OST

[2142] Y. Ortega-Mallen. Operational semantics for timed observations. Lecture Notes in Computer Science, 571:507– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murphy:1991:RTC

[2143] D. Murphy and D. Pitt. Real-timed concurrent refineable behaviours. Lecture Notes in Computer Science, 571: 529–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toetenel:1991:SDM

[2144] H. Toetenel and J. van Katwijk. Stepwise development of model-oriented real-time specifications from action/ event models. Lecture Notes in Computer Science, 571:547–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coesmans:1991:FSF

[2145] P. Coesmans and M. J. Wieczorek. Formal specification of fault-tolerant real-time systems using minimal 3-Sorted modal logic. Lecture Notes in Computer Science, 571:571–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesten:1991:THS

[2146] Y. Kesten and A. Pnueli. Timed and hybrid statecharts and their textual representation. Lecture Notes in Computer Science, 571:591–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stirling:1991:TIS

[2147] C. Stirling. Taming infinite state spaces. Lecture Notes in Computer Science, 575:1–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huttel:1991:SGB

[2148] H. Huttel. Silence is golden: Branching bisimilarity is decidable for context– free processes. Lecture Notes in Computer Science, 575:2–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korver:1991:CDF

[2149] H. Korver. Computing distinguishing formulas for branching bisimulation. Lecture Notes in Computer Science, 575:13–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1991:CCS

[2150] H. Andersen and G. Winskel. Compositional checking of satisfaction. Lecture Notes in Computer Science, 575:24–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1991:ABF

[2151] R. De Nicola, A. Fantechi, S. Gnesi, and G. Ristori. An action based framework for verifying logical and behavioural properties of concurrent systems. Lecture Notes in Computer Science, 575:37–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1991:LTM

[2152] R. Cleaveland and B. Steffen. A linear-time model-checking algorithm for the alternation-free modal mucalculus. Lecture Notes in Computer Science, 575:48–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sistla:1991:ATV

[2153] A. P. Sistla and L. D. Zuck. Automatic temporal verification of buffer systems. Lecture Notes in Computer Science, 575:59–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bevier:1991:MCP

[2154] W. Bevier and J. Søgaard-Andersen. Mechanically checked proofs of kernel specifications. Lecture Notes in Computer Science, 575:70–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gjessing:1991:TAF

[2155] S. Gjessing, S. Krogdahl, and E. Munthe-Kaas. A top down approach to the formal specification of SCI cache coherence. Lecture Notes in Computer Science, 575:83–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avrunin:1991:IPA

[2156] G. S. Avrunin, J. C. Corbett, and U. A.

Buy. Integer programming in the analysis of concurrent systems. Lecture Notes in Computer Science, 575:92–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbeau:1991:LMF

[2157] M. Barbeau and G. v. Bochmann. The Lotos model of a fault protected system and its verification using a Petri net based approach. Lecture Notes in Computer Science, 575:103–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rasse:1991:EDF

[2158] A. Rasse. Error diagnosis in finite communicating systems. Lecture Notes in Computer Science, 575:114–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vemuri:1991:TPV

[2159] R. Vemuri and A. Sridhar. Temporal precondition verification of design transformations. Lecture Notes in Computer Science, 575:125–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1991:PPA

[2160] H. Lin. Pam: A process algebra manipulator. Lecture Notes in Computer Science, 575:136–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1991:CWP

[2161] C. T. Jensen. The concurrency workbench with priorities. Lecture Notes in Computer Science, 575:147–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mauw:1991:PAP

[2162] S. Mauw and G. Veltink. A proof assistant for PSF. Lecture Notes in Computer Science, 575:158–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finkel:1991:ASE

[2163] A. Finkel and L. Petrucci. Avoiding state explosion by composition of minimal covering graphs. Lecture Notes in Computer Science, 575:169–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1991:FVB

[2164] J.-C. Fernandez and L. Mounier. “on the fly” verification of behavioural equivalences and preorders. Lecture Notes in Computer Science, 575:181– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jard:1991:BMA

[2165] C. Jard and Th. Jeron. Boundedmemory algorithms for verification onthe-fly. Lecture Notes in Computer Science, 575:192–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Enders:1991:GBS

[2166] R. Enders, T. Filkorn, and D. Taubner. Generating BDDs for symbolic model checking in CCS. Lecture Notes in Computer Science, 575:203–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiraishi:1991:VSM

[2167] H. Hiraishi, K. Hamaguchi, H. Ochi, and S. Yajima. Vectorized symbolic model checking of computation tree logic for sequential machine verification. Lecture Notes in Computer Science, 575:214–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filkorn:1991:FES

[2168] T. Filkorn. Functional extension of symbolic model checking. Lecture Notes in Computer Science, 575:225– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mao:1991:APT

[2169] W. Mao and G. Milne. An automated proof technique for finite–state machine equivalence. Lecture Notes in Computer Science, 575:233–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brinksma:1991:DSP

[2170] E. Brinksma. From data structures to process structure. Lecture Notes in Computer Science, 575:244–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dill:1991:CLI

[2171] D. L. Dill, A. J. Hu, and H. Wong-Toi.

Checking for language inclusion using simulation relations. Lecture Notes in Computer Science, 575:255–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deFrancesco:1991:SDM

[2172] N. de Francesco and P. Inverardi. A semantic driven method to check the finiteness of CCS processes. Lecture Notes in Computer Science, 575:266– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mutz:1991:UHP

[2173] M. Mutz. Using the HOL prove assistant for proving the correctness of term rewriting rules reducing terms of sequential behaviour. Lecture Notes in Computer Science, 575:277–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nesi:1991:MPI

[2174] M. Nesi. Mechanizing a proof by induction of process algebra specifications in higher order logic. Lecture Notes in Computer Science, 575:288–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seger:1991:TLF

[2175] C.-J. Seger and J. J. Joyce. A two-level formal verification methodology using HOL and COSMOS. Lecture Notes in Computer Science, 575:299–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christoff:1991:EAV

[2176] L. Christoff and I. Christoff. Efficient algorithms for verification of equivalences for probabilistic processes. Lecture Notes in Computer Science, 575: 310–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Probst:1991:POM

[2177] D. K. Probst and H. F. Li. Partialorder model checking: A guide for the perplexed. Lecture Notes in Computer Science, 575:322–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godefroid:1991:UPOb

[2178] P. Godefroid and P. Wolper. Using partial orders for the efficient verification of deadlock freedom and safety properties. Lecture Notes in Computer Science, 575:332–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feigenbaum:1991:CRP

[2179] J. Feigenbaum, C. Lund, and J. A.

Kahn. Complexity results for POM-SET languages. Lecture Notes in Computer Science, 575:343–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldschlag:1991:MVS

[2180] D. M. Goldschlag. Mechanically verifying safety and liveness properties of delay insensitive circuits. Lecture Notes in Computer Science, 575:354–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1991:AMP

[2181] K. Schneider, R. Kumar, and T. Kropf.

Automating most parts of hardware tools in HOL. Lecture Notes in Computer Science, 575:365–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sifakis:1991:OST

[2182] J. Sifakis. An overview and synthesis on timed process algebras. Lecture Notes in Computer Science, 575:376– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcoubetis:1991:MMD

[2183] C. Courcoubetis and M. Yannakakis. Minimum and maximum delay problems in real-time systems. Lecture Notes in Computer Science, 575:399– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamaguchi:1991:FVS

[2184] K. Hamaguchi, H. Hiraishi, and S. Yajima. Formal verification of speeddependent asynchronous circuits using symbolic model checking of branching time regular temporal logic. Lecture Notes in Computer Science, 575:410– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabrielian:1991:VPH

[2185] A. Gabrielian and R. Iyer. Verifying properties of HMS machine specifications of real-time systems. Lecture Notes in Computer Science, 575:421– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffrey:1991:LTP

[2186] A. Jeffrey. A linear time process algebra. Lecture Notes in Computer Science, 575:432–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holmer:1991:DPR

[2187] U. Holmer, W. Yi, and K. Larsen. Deciding properties of regular real timed processes. Lecture Notes in Computer Science, 575:443–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corcoubetis:1991:ABP

[2188] C. Corcoubetis, S. Graf, and J. Sifakis.

An algebra of Boolean processes. Lecture Notes in Computer Science, 575: 454–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langevin:1991:CGS

[2189] M. Langevin and E. Cerny. Comparing generic state machines. Lecture Notes in Computer Science, 575:466– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deJong:1991:ATA

[2190] G. G. de Jong. An automata theoretic approach to temporal logic. Lecture Notes in Computer Science, 575: 477–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koyama:1991:NPK

[2191] Kenji Koyama, Ueli M. Maurer, Tatsuaki Okamoto, and Scott A. Vanstone. New public-key schemes based on elliptic curves over the ring Z_{n}. Lecture Notes in Computer Science, 576:252– 266, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/0576/05760252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760252.
pdf.

Abadi:1991:CAC

[2192] M. Abadi, M. Burrows, B. Lampson, and G. Plotkin. A calculus for access control in distributed systems (invited). Lecture Notes in Computer Science, 576:1– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760001.
pdf.

Toussaint:1991:DCK

[2193] Marie-Jeanne Toussaint. Deriving the complete knowledge of participants in cryptographic protocols (extended abstract). Lecture Notes in Computer Science, 576:24–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760024.
pdf.

Bird:1991:SDT

[2194] Ray Bird, Inder Gopal, Amir Herzberg, Phil Janson, Shay Kutten, Refik Molva, and Moti Yung. Systematic design of two-party authentication protocols. Lecture Notes in Computer Science, 576:44–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760044.
pdf.

Stinson:1991:CCA

[2195] Douglas R. Stinson. Combinatorial characterizations of authentication codes. Lecture Notes in Computer Science, 576:62–73, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760062.
pdf.

Stinson:1991:UHA

[2196] Douglas R. Stinson. Universal hashing and authentication codes. Lecture Notes in Computer Science, 576:74– 85, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760074.
pdf.

Camion:1991:CIF

[2197] Paul Camion, Claude Carlet, P. Charpin, and N. Sendrier. On correlationimmune functions. Lecture Notes in Computer Science, 576:86–100, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760086.
                                        pdf.

Capocelli:1991:SSS

[2198] R. M. Capocelli, A. De Santis, L. Gargano, and U. Vaccaro. On the size of shares for secret sharing schemes. Lecture Notes in Computer Science, 576:101–113, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760101.
                                        pdf.

Dwork:1991:VSS

[2199] Cynthia Dwork. On verification in secret sharing. Lecture Notes in Computer Science, 576:114–128, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760114.
                                        pdf.

Pedersen:1991:NII

[2200] Torben Pryds Pedersen. Noninteractive and information-theoretic secure verifiable secret sharing. Lecture Notes in Computer Science, 576:129– 140, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/0576/05760129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760129.
pdf.

Fischer:1991:MSK

[2201] Michael J. Fischer and Rebecca N.

Wright. Multiparty secret key exchange using a random deal of cards (extended abstract). Lecture Notes in Computer Science, 576:141–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760141.
pdf.

Biham:1991:DCS

[2202] Eli Biham and Adi Shamir. Differential cryptanalysis of Snefru, Khafre, REDOC-II, LOKI and Lucifer (extended abstract). Lecture Notes in Computer Science, 576:156–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760156.
pdf.

Tardy-Corfdir:1991:KPA

[2203] Anne Tardy-Corfdir and Henry Gilbert.

A known plaintext attack of FEAL-4 and FEAL-6. Lecture Notes in Computer Science, 576:172–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760172.
                                        pdf.

Morita:1991:SCT

[2204] Hikaru Morita, Kazuo Ohta, and Shoji Miyaguchi. A switching closure test to analyze cryptosystems (extended abstract). Lecture Notes in Computer Science, 576:183–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760183.
                                        pdf.

Boer:1991:ALT

[2205] Bert den Boer and Antoon Bosselaers. An attack on the last two rounds of MD4. Lecture Notes in Computer Science, 576:194–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760194.
                                        pdf.

denBoer:1991:ALT

[2206] Bert den Boer and Antoon Bosselaers. An attack on the last two rounds of MD4. Lecture Notes in Computer Science, 576:194–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0576/05760194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760194.
pdf.

Chee:1991:CNP

[2207] Yeow Meng Chee, Antoine Joux, and Jacques Stern. The cryptanalysis of a new public-key cryptosystem based on modular knapsacks. Lecture Notes in Computer Science, 576:204– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760204.
pdf.

Lapidot:1991:ORT

[2208] Dror Lapidot and Adi Shamir. A one-round, two-prover, zero-knowledge protocol for NP. Lecture Notes in Computer Science, 576:213–224, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760213.
pdf.

Kilian:1991:IPS

[2209] Joe Kilian and Ronitt Rubinfeld. Interactive proofs with space bounded provers. Lecture Notes in Computer Science, 576:225–231, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/0576/05760225.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0576/05760225.pdf.

Teng:1991:FIC

[2210] Shang-Hua Teng. Functional inversion and communication complexity. Lecture Notes in Computer Science, 576: 232–241, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0576/05760232.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760232.
                                        pdf.

Rudich:1991:UIP

[2211] Steven Rudich. The use of interaction in public cryptosystems (extended abstract). Lecture Notes in Computer Science, 576:242–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760242.
                                        pdf.

Okamoto:1991:EAC

[2212] Tatsuaki Okamoto and Kouichi Sakurai. Efficient algorithms for the construction of hyperelliptic cryptosystems. Lecture Notes in Computer Science, 576:267–278, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0576/05760267.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
0576/05760267.pdf.

Koblitz:1991:CCG

[2213] Neal Koblitz. CM-curves with good cryptographic properties. Lecture Notes in Computer Science, 576:279– 287, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760279.
pdf.

Tsujii:1991:NIB

[2214] Shigeo Tsujii and Jinhui Chao. A new ID-based key sharing system. Lecture Notes in Computer Science, 576:288– 299, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760288.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760288.
pdf.

Luby:1991:PRG

[2215] Michael Luby. Pseudo-random generators from one-way functions (invited abstract). Lecture Notes in Computer Science, 576:300–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760300.
pdf.

Patarin:1991:NRP

[2216] Jacques Patarin. New results on pseudorandom permutation generators based on the DES scheme. Lecture Notes in Computer Science, 576:301– 213, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760301.
                                        pdf.

Walter:1991:FMM

[2217] Colin D. Walter. Faster modular multiplication by operand scaling. Lecture Notes in Computer Science, 576:313– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760313.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760313.
                                        pdf.

Okamoto:1991:UEC

[2218] Tatsuaki Okamoto and Kazuo Ohta. Universal electronic cash. Lecture Notes in Computer Science, 576:324– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760324.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760324.
                                        pdf.

Pfitzmann:1991:HBR

[2219] Birgit Pfitzmann and Michael Waidner. How to break and repair a “provably secure” untraceable payment system (extended abstract). Lecture Notes in Computer Science, 576:338– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760338.
pdf.

Bennett:1991:PQO

[2220] Charles H. Bennett, Gilles Brassard, Claude Crépeau, and Marie-Hélène Skubiszewska. Practical quantum oblivious transfer. Lecture Notes in Computer Science, 576:351–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760351.
pdf.

Broscius:1991:EPH

[2221] Albert G. Broscius and Jonathan M.

Smith. Exploiting parallelism in hardware implementation of the DES. Lecture Notes in Computer Science, 576: 367–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760367.
pdf.

Beaver:1991:FSI

[2222] Donald Beaver. Foundations of secure interactive computing. Lecture Notes in Computer Science, 576:377– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760377.
                                        pdf.

Micali:1991:SCA

[2223] Silvio Micali and Phillip Rogaway. Secure computation (abstract). Lecture Notes in Computer Science, 576:392– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760392.
                                        pdf.

Iversen:1991:CSC

[2224] Kenneth R. Iversen. A cryptographic scheme for computerized general elections. Lecture Notes in Computer Science, 576:405–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760405.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760405.
                                        pdf.

Beaver:1991:EMP

[2225] Donald Beaver. Efficient multiparty protocols using circuit randomization. Lecture Notes in Computer Science, 576:420–??, 1991. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0576/05760420.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0576/05760420.
pdf.

Rackoff:1991:NIZ

[2226] Charles Rackoff and Daniel R. Simon. Non-interactive zero-knowledge proof of knowledge and chosen ciphertext attack. Lecture Notes in Computer Science, 576:433–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760433.
pdf.

Damgaard:1991:TPP

[2227] Ivan Bjerre Damgård. Towards practical public key systems secure against chosen ciphertext attacks. Lecture Notes in Computer Science, 576:445– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0576/05760445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0576/05760445.
pdf.

Desmedt:1991:SGA

[2228] Yvo Desmedt and Yair Frankel. Shared generation of authenticators and signatures (extended abstract). Lecture Notes in Computer Science, 576:457– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760457.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760457.
                                        pdf.

Chaum:1991:CSU

[2229] David Chaum, Eugène van Heijst, and Birgit Pfitzmann. Cryptographically strong undeniable signatures, unconditionally secure for the signer (extended abstract). Lecture Notes in Computer Science, 576:470–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0576/05760470.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0576/05760470.
                                        pdf.

Anonymous:1991:AIb

[2230] Anonymous. Author index. Lecture Notes in Computer Science, 576:485– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0576/0576ind.pdf.

Jacquet:1991:CMA

[2231] J.-M. (Jean-Marie) Jacquet. Conclog: a methodological approach to concurrent logic programming, volume 556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1991. CODEN LNCSD9. ISBN 3-540-54938-2 (Berlin), 0-387-54938-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 781 pp. LCCN QA76.63 .J33 1991.

Herman:1992:UPG

[2232] I. (Ivan) Herman. The use of projective geometry in computer graphics, volume 564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1992. CODEN LNCSD9. ISBN 3-540-55075-5 (Berlin), 0-387-55075-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 146 pp. LCCN T385 .H47 1991. Outgrowth of author’s thesis (Ph. D.)–University of Leiden, Netherlands,